A Bibliography of Papers in Lecture Notes in Computer Science (1996), Part 2 of 2

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

10 February 2022 Version 1.30

Title word cross-reference

(n − 1) [2526]. (t, n) [2439]. 1/p [1327]. 2 [775, 1963, 1726, 1029, 2713, 2783, 931, 993, 807, 1622, 1907, 813, 2163]. 3 [2532, 1805, 1912, 1026, 445, 1309, 1182, 320, 904, 1029, 1296, 1246, 1257, 1256, 2521, 807, 657, 1796, 1287, 1704, 816, 1791]. 4 [1337, 2534, 791]. 81/2 [945, 950]. ^{t} [356]. ax + b mod p [353, 354]. d [887, 2572]. E [1116]. ℓ [827]. f [1712]. h [1174]. k [1490, 2604, 1465, 2573, 1874, 2562]. λ [1443]. λ(?) [65]. log_{2} n/ log log n [2561]. log log n [560]. log n [2561]. m [750]. N [2740, 738, 776, 966, 2123, 1872, 1646]. O [834]. O(m + n log n) [753]. p [827, 830, 1176]. Q [832, 1481]. Q(t) [832]. ρ [2152]. × [2175]. Z [1179, 771].

-adic [830, 1176]. -ary [887, 1337, 1646, 738, 2123]. -Body [2740]. -Bounded [750]. -Calculus [1443, 65]. -Categorical [1174]. -Center [1465]. -Connected [1490]. -Cube [776, 966]. -D [1726, 1029, 2713, 1029, 993, 1246, 1257, 2521, 813, 2163, 1791]. -Diagnosable [1963]. -Dimensional [775, 904, 2783, 931, 1622, 2572]. -Error [2604]. -Free [771]. -gram [1872]. -Grams [1481]. -Isogenies [827]. -Lenses [1712]. -level [1912]. -manifold [2526]. -Oriented [265]. -Pairs [2573]. -piecewise [1874]. -Primes [1182]. -restricted [560]. -Root [356]. -standard [2152]. -strong [1574]. -Torsion [827]. -Trees [1490, 2562].

-Unification [1116].

/Inclusive [1735].

00 [2086]. 0I0T [1168].

1 [2086, 1737, 405]. 1-Solutions [2086]. 1.0 [172]. 1.5 [145]. 13 [163]. 15 [686]. 15-Puzzle [686]. 16th [2803]. 1996 [1790].

2 [2685, 2096, 832, 872, 943, 2777, 1071]. 2-Torsion [832]. 20n [2252, 2604]. 21st [217].

3 [2397, 1546]. 3-Satisfiability [1546]. 3T0AP [183].

56 [2397]. 56-3 [2397].

68 [2641].

737 [2397].

860 [1165].

9001 [1920]. ’95 [109].

= [1651].

A-Fields [2248]. A-NETL [263]. A/D [1164]. Abductive [1130, 604]. Abelian [822, 833, 190, 117]. ABL [2281]. Abnormal [1510]. Abnormalities [1288]. Aboutness [642]. Absfol [161]. Absolutely [1136]. Abstract [1813, 2284, 59, 2707, 1223, 1834, 14, 1215, 2665, 1570, 2055, 541, 2792, 1833, 148, 1584, 190, 562, 568, 186, 578]. Abstracting [1970, 81]. Abstraction [2363, 1498, 2317, 240, 113, 161]. Abstractions [2410, 2387, 1571]. AC [149, 116, 212]. AC-Complete [116]. AC-theories [212]. Acceleration [1765]. Acceptance [1741]. acceptors [1868]. Access [2494, 2700, 1074, 1390, 2605, 1377, 2435, 1841, 598, 1010, 552, 2422]. Account [2689]. Accurate [1264, 1981, 1248, 1054]. achieve [176]. achievement [197]. Achieving [2271]. Acitve [1379]. ACL2 [2326]. Acoustic [785]. Acquired [1988]. Acquiring [648, 583]. Acquisition [641, 1992, 1539, 2719, 804, 1991]. Across [2089, 1666, 408]. Action [1992, 2287, 1510, 619, 1120, 46, 537, 2434]. Actions [1496, 621, 26, 1510, 2411, 622, 1527]. Activation [597]. Active [1352, 2029, 1354, 486, 2030, 516, 1917, 1380, 2512]. Activity [404, 2349, 474, 1946, 2777]. Activity-Based [2349]. ActNet [275]. Actor [275, 1973]. Actors [2652]. Actual [338]. Adapt [483]. Adaptable [1939, 2663, 467]. Adaptable-Size [467]. Adaptation [2365, 1653, 2367, 2775, 2373, 2628, 2385]. Adapting [684, 504]. Adaption [1665, 1660]. Adaptive [1488, 2156, 439, 1356, 700, 1739, 423, 1656, 1783, 867, 286, 935, 271, 2080, 686, 2379, 292, 1008, 657, 2243, 616, 883, 1785, 1699, 1658, 1725, 668]. Adaptivity [894]. Adder [1029, 991, 1721]. Adding [2654, 1507, 2765, 1190]. Addition [687, 524, 990]. Address [1947]. Addresses [2454]. Addressing [2698]. Adequacy [1089]. adic [830, 1176]. Admissibility [2152]. Admission [1485, 1965]. Advanced [1303, 2695, 2047, 201, 2774, 39]. Advances [2477]. Advection [2694]. Advisor [284]. Advisory [437]. Aerodynamic [1711]. Affine [1032, 1314]. after [1300]. Again [882]. Against [1368, 365, 2580]. Agenda [1910]. Agent [1929, 1615, 630, 1184, 2027, 628, 290, 291, 1666, 2154, 278, 2374, 627]. Agents [2155, 1816, 1844, 2667, 1783, 2477, 1617, 1375, 1930, 2169, 1528, 1551, 2175, 2801]. Aggregate [2040, 1069, 1389]. aggregation [87]. Agnostic [359]. Agrarian [316]. Agreement [1967, 2250, 2263, 2444].

Ahead [732, 114]. AI [2179, 2172]. Aided [239, 2000, 321, 1268, 216, 811, 1197, 1727, 231, 244]. Air [2364, 2694, 1704]. Aircraft [725, 1002]. AL [2283]. Alarm [2506]. ALDY [2081]. ALF [1086]. Algebra [992, 1567, 1168, 2698, 1195, 2706, 1193, 2729, 26, 1188, 1187, 194, 2752, 1401]. Algebraic [1575, 1568, 2285, 825, 35, 1220, 41, 1212, 58, 2330, 837, 2293, 1176, 42, 12, 1014, 1225, 1227, 1187, 134, 617, 1197, 393, 2299, 32, 1326, 1325, 1393, 855, 1238, 1240, 2640, 2191, 64, 545]. algebraically [558]. Algebras [1553, 2654, 1582, 1232, 47]. Algol [378, 2641]. Algol-like [378]. Algorithm [2156, 2692, 965, 2697, 497, 2013, 1270, 1767, 1684, 1182, 982, 1126, 2166, 906, 671, 1711, 1114, 1655, 2010, 1769, 958, 836, 1662, 1903, 902, 1715, 644, 2722, 2724, 1099, 1750, 1656, 1248, 1997, 979, 794, 2731, 2732, 1664, 1513, 529, 1892, 2162, 972, 601, 2158, 2141, 1083, 1262, 752, 634, 2642, 2740, 2742, 1665, 776, 966, 487, 1106, 2653, 962, 2798, 531, 846, 1657, 1276, 533, 2009, 714, 2585, 788, 496, 1731, 136, 1699, 1830, 1667, 2066, 1658, 2755, 2756, 767, 1408, 670, 1659, 2558, 961, 2760, 1652, 521, 1641, 2234, 1480, 1399, 2762, 488, 507]. algorithm [2516, 2535, 71, 2193, 1877, 65, 573, 1682]. Algorithmic [951, 678, 2125, 101, 77]. Algorithms [1084, 1484, 1700, 1483, 1625, 1485, 1015, 2472, 1710, 1488, 897, 1026, 445, 1474, 367, 1266, 2544, 825, 2701, 667, 1970, 2556, 1471, 1479, 662, 1661, 602, 877, 1490, 675, 1891, 1686, 1095, 2152, 2583, 1965, 1974, 1886, 1468, 1176, 1458, 1647, 978, 1157, 1467, 1887, 1020, 2498, 1648, 1694, 1627, 2582, 2340, 1770, 764, 1246, 1695, 1629, 2658, 736, 1649, 1714, 2733, 1987, 1697, 1772, 1037, 843, 1630, 844, 1773, 2310, 988, 1988, 1676, 989, 1706, 1894, 1888, 1827, 883, 740, 2051, 936, 2018, 1636, 1637, 1638, 485, 1454, 1786, 679, 1156, 1690, 2272, 1651, 718]. Algorithms [1788, 986, 1663, 1789, 1267, 2631, 800, 429, 451, 2511, 66, 205, 1884, 177, 2187, 2531, 160]. Algorithms-Based [2511]. Alias [1833]. Alignment [916, 1276, 1282]. All-Optical [2577]. Allele [1640]. Allele-Wise [1640]. Allen [906]. Allocating [656]. Allocation [2242, 1059, 1458, 2582, 683, 1050, 2243, 2234]. Allowing [1901]. almost [571]. Alpha [1096, 58]. Alpha-Conversion [1096]. Alphabetic [769]. ALPS [2623]. ALPS-HI [2623]. Alternating [779, 2138, 552]. alternating-direction [552]. Alternative [432, 1594, 1332, 2444]. ALWAN [2077]. Always [1116]. Amalgamation [708]. ambiguity [2800]. AMI [594]. Among [2155, 470, 2589, 1491]. Analog [455, 527, 320]. Analogical [317, 2378]. Analogy [1115, 317, 1521, 158]. Analysed [2388]. Analysers [1820]. Analyses [379]. Analysis [1484, 789, 1809, 1164, 1814, 2453, 438, 912, 773, 1272, 377, 2633, 2035, 2648, 1817, 1818, 862, 910, 1819, 434, 2706, 1285, 588, 1094, 1821, 2682, 1974, 1448, 490, 2657, 704, 338, 2625, 26, 1738, 510, 1512, 44, 1257, 1835, 683, 2665, 2114, 809, 1630, 2666, 1631, 1943, 1810, 810, 599, 408, 1827, 797, 1021, 911, 819, 397, 1120, 1638, 232, 1260, 498, 1454, 1786, 2506, 1640, 2792, 814, 81, 720, 745, 2167, 1833, 1470, 722, 1584, 907, 2578, 2607, 544, 2528, 83, 98, 575, 2525, 2459, 99, 573]. Analytica [1535]. Analytical [1625, 460, 1058]. Analyzer [1834, 1836]. Analyzing [2227, 1179, 1273, 1857, 471]. Anaphora [2160]. Anatomic [1243, 1288]. Anatomical [1283, 1308, 1289]. Anatomy [1289]. AND-EXOR [1773]. And-parallel [1049]. And-Parallelism [1583]. Angiographies [1256]. Angiography [1248]. Animating [1186]. Animation [313, 2761]. Annai [862]. Annealed [489, 430]. Annealing [1669, 1673, 1464, 2020]. Annotated [2623, 1110, 1584]. Anonymity [1839, 1850]. Anonymity-Revoking [1839].

Anonymous [2454, 348]. Answering [1126, 1762]. Ant [1679]. Ant-Q [1679]. Anti [2598, 1683, 517]. Anti-Chains [2598]. Anti-Hebbian [517]. Anti-Universal [1683]. Antidirected [758]. Anytime [730, 625]. API [2233]. appearance [1802, 1798, 1800]. Applicability [2110]. Applicable [392, 215, 399]. Application [1452, 276, 377, 367, 116, 2699, 1437, 2134, 435, 1502, 1297, 1473, 2667, 2547, 837, 1734, 2138, 1372, 1382, 2720, 1694, 2774, 1246, 1513, 2262, 402, 1047, 1994, 593, 1357, 22, 453, 2572, 2771, 1772, 1899, 313, 1943, 2653, 2241, 2770, 2531, 2105, 1314, 1287, 1143, 2612, 1833, 2643, 2797, 575]. Application-Assisted [1047]. Applications [1352, 944, 239, 2542, 912, 1167, 2691, 2633, 2702, 1052, 1949, 1958, 2357, 778, 1938, 673, 2775, 1012, 2081, 1184, 2575, 1770, 794, 1835, 529, 2734, 757, 133, 686, 2063, 811, 1890, 676, 747, 874, 2749, 1889, 2791, 2215, 1601, 2464, 1282, 1054, 2581, 1423, 3, 2776, 2763, 2058, 871, 666, 584, 2534, 198, 1807]. Applied [789, 1754, 2396, 275, 1697, 686, 294, 324, 2193]. Applying [2407, 2255, 1100, 1057, 1357, 2376, 5, 2026, 1396]. Appraisal [1844]. Approach [1966, 1274, 1929, 1568, 873, 1492, 2130, 2155, 2156, 1701, 1324, 1183, 2287, 785, 2005, 1191, 2367, 699, 380, 2706, 1198, 1273, 997, 1500, 2488, 1539, 2707, 1330, 335, 285, 1332, 2140, 1703, 1680, 2609, 2023, 711, 1595, 1365, 1069, 339, 470, 2741, 2003, 1630, 291, 1909, 2747, 2115, 1348, 817, 1243, 1034, 586, 1025, 1546, 2479, 240, 227, 2121, 1070, 485, 2319, 1282, 953, 8, 2117, 2024, 1934, 2175, 442, 746, 1756, 1805, 545, 2424, 2801, 2799, 1801]. Approaches [786, 608, 1993, 1857, 2478, 1046]. Approximability [2548]. Approximate [1491, 659, 2067, 736, 850]. Approximating [1456, 571, 1278]. Approximation [2564, 766, 1466, 383, 1458, 1467, 2549, 1036, 1460, 1459]. Approximations [141, 571]. Approximative [26]. April [266]. Arbitrary [411, 681]. Arbitration [1680]. Arc [690, 758, 1350, 726, 255, 733, 601, 254]. Archimedean [830]. Architectural [466, 1750, 1904, 225, 1728, 1747]. Architecture [2696, 630, 2093, 434, 2218, 1938, 1065, 2235, 2372, 858, 464, 1158, 454, 993, 585, 221, 45, 1023, 2027, 737, 631, 777, 2741, 290, 1720, 2676, 433, 526, 1066, 1736, 2342, 277, 514, 1403]. Architecture-Independent [2235]. Architecture-Invariant [2741]. Architectures [1026, 1195, 2712, 2718, 2737, 2740, 232, 1017, 1028]. Area [1478, 2347, 1212, 2745, 1926]. Areal [817]. ARGOS [1428]. Arguing [2361]. Argumentation [2148]. ARIANE [306]. Arias [2475]. Arithmetic [992, 2308, 824, 1449, 980, 2252, 2040, 1181, 2524, 75]. Arline [2592]. Array [912, 910, 2683, 671, 787, 861, 1020, 453, 2736, 1298, 2699]. Array-Oriented [2683]. Arrays [913, 891, 2738, 2588, 325, 723, 1728]. Arrhythmias [2164]. Art [1891, 321, 2053, 433]. Articulated [2177]. Artificial [1274, 1998, 1783, 1619, 402, 2642, 614, 294, 518, 1547, 461]. Ary [738, 887, 1337, 1646, 2123]. ASCII [336]. ASD [46]. ASF [3]. Asgard [2349]. ASIC [1754]. Aspects [378, 749, 1494, 847, 849, 1145]. Assemblage [1623]. Assemblies [513]. Assembly [2308]. Assertion [612]. Assertional [1324, 2408, 2302]. Assessing [327, 1293]. Assessment [1672, 1246, 2383, 1304, 2436]. Assigned [1941]. Assignment [1693, 763, 781, 2021, 956]. Assignments [280]. Assistance [2648]. Assistant [125]. Assisted [2079, 1047, 1311, 295, 304, 1947]. Associations [433]. Associative [149, 512, 411, 409, 2653, 962, 883, 431, 80]. Associative-Commutative [149, 80]. ASSPEGIQUE0 [40]. Assumption [258, 1138, 600, 2260, 2592]. Assumption-Based [258]. Assumptions [957]. Asteroidal [2559]. Astro [946]. Astro-Gofer [946]. Asymmetric [347]. Asymmetry [1292]. Asymptotic [1626, 417]. Asymptotically [1468, 1032]. Asynchronous [982, 1004, 1744]. Asynchronously [744]. Atlas [1287, 1288]. Atlases [1283]. ATM [2621, 2775, 561, 1744, 1145, 963, 2219, 2764]. Atmospheric [450]. Atomic [1335, 1960, 923, 315]. Atomicity [1912, 88]. Atoms [389]. ATP [163]. ATR [1998]. Atrophy [1293]. Attack [2277]. Attacks [2276, 361, 352, 362, 363, 2431]. ATTEMPT [1737]. ATTEMPT-1 [1737]. Attribute [1586]. audio [82]. Auditory [1300, 1530, 514]. Augmentation [2563]. Augmented [783]. August [2743]. Aunt [2507]. Austrian [2771]. Authenticated [2250, 2442]. Authentication [345, 344, 1967, 1844, 369, 1851, 1846, 2427, 2417]. Author [376]. Authoring [311, 300, 301]. Authorization [1840, 1843]. Autoassociation [497]. Autoassociative [410]. AutoFocus [1451]. Automata [2565, 489, 1677, 48, 1430, 993, 2502, 2413, 1829, 1636, 1431, 2192, 2536]. Automated [1272, 80, 1167, 1309, 1137, 114, 1284, 1824, 1348, 2004, 76, 1314, 194, 176]. Automatic [1274, 1271, 2160, 377, 2648, 1692, 2072, 1415, 69, 2800, 1265, 948, 1517, 905, 449, 629, 92, 1612, 2015, 508, 799, 1421, 2332, 1161, 2167, 816, 2331, 1295, 89]. Automatically [1287]. automaticity [562]. Automating [2312, 2135, 11, 200]. Automation [1926]. Autonomous [1992, 1680, 1361, 631, 291, 2169, 1528, 506, 504]. Autoradiograph [1276]. Available [866, 1964, 926]. Average [683, 1990]. Averaged [441]. Avoid [1992]. Avoidance [1436]. Avoiding [1517]. AVS5 [2682]. Axial [763]. Axiom [17]. Axiomatic [2036]. Axiomatization [16, 2147]. axiomatizations [210]. Axioms [1096].

Back [692, 1139, 154]. Back-up [1139]. Backtrackable [2592]. Backtracking [1501, 738]. Backward [1554, 650, 1830]. Balanced [917, 286]. Balancing [2072, 2727, 1050, 2080, 686, 1705, 860, 2384, 1046, 682]. Banded [968, 2721]. Bandwidth [887, 2711, 896]. Base [332, 990, 2373, 2398, 2382, 1416, 1314, 2389]. Based [2176, 1809, 1962, 2091, 1349, 987, 1497, 1048, 1359, 272, 1956, 276, 2245, 2363, 2156, 2248, 249, 445, 803, 2396, 1550, 2035, 2648, 1259, 317, 1818, 2703, 1270, 1743, 597, 1063, 2242, 696, 2705, 1163, 2367, 2706, 1898, 1590, 2466, 1355, 280, 297, 1407, 1222, 332, 815, 1391, 235, 2218, 1367, 2109, 58, 2370, 31, 2447, 437, 334, 1687, 2775, 335, 1124, 2045, 2371, 2372, 1753, 1382, 2783, 1834, 1280, 1648, 150, 386, 1057, 2603, 258, 1511, 2375, 1268, 114, 2148, 604, 1445, 603, 2410, 948, 2253, 2264, 529, 1757, 943, 288, 1520, 436, 1262]. Based [2055, 2452, 2740, 2003, 738, 2349, 2379, 1983, 2666, 1705, 2380, 2381, 291, 292, 1523, 1173, 225, 1825, 1160, 56, 408, 929, 2383, 47, 797, 2095, 533, 1540, 1950, 342, 1538, 2159, 1278, 1841, 2174, 1120, 1242, 1046, 1602, 281, 369, 2388, 2036, 2163, 2020, 1288, 2630, 2390, 2075, 1765, 1417, 519, 1934, 1660, 147, 2777, 226, 2509, 2686, 2058, 2596, 506, 2374, 633, 924, 2394, 582, 1277, 491, 2511, 1795, 2442, 183, 170, 151, 2437, 1802, 204, 2425, 210, 2198, 2183, 81, 2434, 868, 965, 2337, 884, 358, 1286, 2011, 306, 655, 1738, 2220, 1718, 1721]. Based [2061]. Bases [2669, 837, 118, 2386, 518, 2511]. Basic [992, 1916, 249, 585, 1918, 1518, 1187]. Basis [1767, 1684, 2112, 1906, 1506, 704, 453, 848]. Batch [608, 2234]. Bayesian [415, 495, 470, 2390]. BBDGR [2284]. BC [2206]. BCPR [2285]. BDD [2333]. BDDs [2336]. BDI [630]. Be [875, 1587, 177]. Beam [1903]. Beam-Tracing [1903].

Behaved [2149]. Behavior [1813, 1626, 2633, 2290, 2669, 1680, 510, 1394, 627, 113]. Behaviour [1771, 739]. Behavioural [1231, 1453, 1233, 2436]. Behaviouristic [2369]. Behaviours [2229]. being [1803]. Belief [2166, 1516, 596]. Believing [318]. Bench [1301]. Benchmark [1026, 999]. Benefits [2114]. BEPC [2719]. Best [2354, 1327, 2353, 2392]. Better [1185]. Between [1476, 374, 337, 706, 788, 1237, 477]. Beyond [857, 1594, 1371, 2752, 1807, 2199]. Bias [1259, 1261]. Biconnectivity [2564, 2563]. Bidimensional [799]. Big [2451, 2353, 638]. Billion [1998]. Binary [524, 1368, 681, 769, 2604, 1632, 717, 1766, 2517]. Binding [2071, 2657, 2228, 515]. Binding-Time [2657]. Binocular [439]. Biological [858, 2169, 518]. Biologically [507]. Biology [1000]. Biometrics [1291]. Bioprocess [2362]. Biopsies [1313]. Bipartite [2556, 2563, 757, 752]. Bipartitions [2546]. Biplane [1256]. Bisimilarity [549]. Bisimulation [25, 15, 1231, 66]. bisimulations [81]. BIST [1952]. Bit [997, 2138, 2443]. Bit-True [997]. Bitemporal [2026, 1411]. Bits [355, 2279, 353, 354]. Black [367, 2744, 351]. Black-Box [367, 351]. BLAS [2704]. Blast [438]. Blind [2267, 2603, 2268, 2464, 451]. Blindfolded [359]. Block [913, 1335, 685, 1332, 1165, 2253, 1334, 876, 2580, 1731, 770, 1766, 2607]. Block-Distributed [913]. Block-Matching [1731]. BLOCKER [1081]. Blocking [909, 886]. Blocks [2746, 1663]. Blue [1476]. Blurring [2194]. BMDs [2308]. Board [1949, 1724]. Body [2740]. Boeing [2397]. Boiler [2281, 2305, 2282, 2284, 2285, 2286, 2287, 2288, 2290, 2289, 2291, 2292, 2293, 2294, 2295, 2296, 2297, 2298, 2304, 2299, 2301, 2302, 2303]. Boleean [1428]. Boltzmann [412, 259]. Bookshop [2771]. Boolean [2406, 27, 2265, 63, 1773, 1650, 1328]. Boosting [2195]. Bootstrapping [507]. Border [1569]. Borel [2133]. Borowsky [1986]. Borowsky-Gafni [1986]. Both [1097, 1394]. Bottleneck [2553, 763]. Bottlenecks [1068]. Bottom [2167]. Bottom-Up [2167]. Bound [1488, 762, 1976, 679, 98, 179]. Boundaries [1253]. Boundary [1264, 1272, 1263]. Bounded [1955, 750, 524, 787, 1364, 2043, 38, 1032, 2566, 2601, 2185]. bounded-width [2185]. Boundedness [223]. Bounding [2462]. Bounds [836, 1482, 2582, 898, 742, 2184]. Box [367, 2462, 1675, 351]. BPA [2499]. BPP [2499]. Brain [1271, 1285, 1998, 1297, 1293, 1301, 1294, 1287, 1288, 1267, 1998]. Brains [1282]. Branch [762, 1063, 747, 679]. Branch-and-Bound [679]. Branch-and-Price [747]. Branches [1033]. Branching [2341]. Breakthroughs [2426]. Breast [1261, 323]. Breeder [1711, 1642]. Bridge [2221, 545]. Bridging [1950]. Brief [246, 2333]. Bringing [630]. British [2002]. Brittleness [1614]. Broad [453]. Broadband [2766, 2211, 2046]. Broadcast [374, 1422, 2219]. Broadcasting [1948, 897, 2602, 1440, 1975, 898]. Bronchus [1244]. Brother [2451]. Browsing [448]. Bruijn [898]. Brujn [775]. BSP [1012, 1014, 1056, 1011]. BSP0* [1015]. BSS [2287]. BT [2699]. BTA [2658]. Buckets [1470]. Buffer [1853]. Build [1964]. Building [1115, 203, 2006, 2799, 300, 1369, 1663, 1347, 2125]. Built [1553, 1287]. Built-in [1553]. Bulk [1038, 1013, 2709, 1008, 1010]. Bulk-Synchronous [1008, 1010]. Bundles [2510]. Burst [614]. Bus [1031, 902, 82]. Buses [902]. Business [1937, 2481]. BV [2794]. BV-tree [2794]. BW [2286]. Byzantine [2602].

C [2218, 2079]. C03 [2786]. CA1 [519].

Cache [1813, 921, 1057, 1023, 1951, 1025]. CAD [2337, 667, 1720, 1346, 1801]. CAD/CIM [1346]. CADE [163]. CADE-13 [163]. CADP [106]. Calculate [1578]. Calculating [2638]. Calculation [1695]. Calculational [1610, 2609]. Calculator [1091]. Calculi [1169, 1541, 1574, 128, 393, 204]. Calculus [1443, 1430, 140, 1433, 386, 1434, 1215, 2599, 2500, 2660, 1432, 852, 1531, 2145, 2415, 548, 208, 65, 542]. Calendars [2120]. Calibration [446, 1808]. calibration-free [1808]. Call [1485, 2217]. CAM [1998]. CAM-BRAIN [1998]. CAM2 [993]. Can [2308, 875, 2645, 176, 429]. cancellative [190]. Cancer [1268, 1257]. Candidates [634]. Cannot [2645, 177]. CAP [2079]. Capacitated [1465]. Capacity [410, 493]. CAPpartx [1727]. Capstone [351]. Capturing [2784]. Car [2511]. Card [1743]. Cardiac [2164]. Care [1239]. Caricaturing [632]. Carlo [976, 2689]. carte [1071]. Cartographic [817]. CARVE [488]. Carving [1136]. CASA [1197]. Cascade [792, 244]. Cascade-Correlation [792]. Cascaded [254]. Cascades [2276]. Case [2176, 2281, 1352, 263, 773, 2131, 2363, 1669, 2396, 2400, 1195, 2367, 1220, 1993, 1093, 2096, 1590, 1693, 2369, 2370, 334, 2371, 2372, 2373, 1511, 2375, 2398, 2410, 763, 522, 1228, 683, 2376, 1518, 2771, 2377, 5, 2063, 2379, 2380, 2381, 1106, 627, 2382, 896, 2384, 2481, 2311, 2386, 2388, 2389, 243, 2390, 2391, 2374, 2394, 113, 2200, 203]. Case-Base [2373, 2398, 2389]. Case-Based [2176, 2363, 2396, 2367, 2370, 334, 2371, 2372, 1511, 2375, 2379, 2380, 2381, 2388, 2390, 2374, 2394]. Case-Enhanced [2384]. Cases [2368, 2371, 2549]. Cash [2270, 348, 2586, 2271, 1838]. CAST [229]. CAST.FSM [241]. Cat [469, 1300, 509]. Categorial [1532]. Categorical [1086, 34, 1174, 211]. Categories [2132, 1224, 2144, 1177, 558]. categorization [1863]. Category [1222, 1212]. Category-Based [1222]. Causal [563]. Causality [1445, 929]. Causality-Based [1445]. Cause [2649]. Cause-Effect [2649]. Cautious [869]. Cavity [1272]. Cayley [774, 900]. CBJ [693]. CBL [302]. CBR [2362, 2364, 2397]. cc [741]. CCM [1766]. CCSimP [1746]. CD [2288]. CDMA [1339]. Cell [439, 1435, 1265, 513]. Cells [406, 1540]. Cellular [1677, 993, 1667, 2536]. Censored [2156]. Center [2232, 1465, 2439]. Centered [1912, 1933, 2121, 1185]. Central [756, 1198]. Centred [1914]. Century [217]. Cerebellum [510, 511]. Cerebral [1286, 1256, 1292, 1287]. Certain [2590]. Certified [1847]. CFD [2053]. CFM [2397]. CGAL [1906, 1894]. Chained [2727, 575]. chaining [540]. Chainings [650]. Chains [2598, 505, 478, 84, 2188]. Challenge [1213, 2699, 2713]. Challenges [114, 2172, 2251, 1249, 2031]. Change [1510, 1918, 2359, 805]. Changing [1221, 1665, 1338, 2416, 1652]. Channel [2453, 2457, 1743, 1300, 36, 2264, 2773]. Channels [1948, 1337, 2458, 2451, 423, 1336, 371, 2452, 955, 2463]. Chaotic [1157]. Character [818, 787, 1345, 801, 819]. Characterisation [1294]. Characteristic [1861, 389]. Characteristics [279, 30, 2241, 2240, 1759]. Characterization [34, 2029, 2741]. Characterizing [1131, 1771, 2263, 1859]. Characters [802, 820, 801]. Cheap [1610]. cheater [2420]. Check [1820]. checker [161]. Checking [2318, 2130, 1171, 2787, 1441, 2307, 2406, 830, 2316, 2034, 2617, 2141, 2026, 2323, 2331, 2336, 64, 83, 102, 534, 68, 63, 100, 539, 70, 79, 67, 1947]. Checkpoint [1402]. Checkpointing [1959, 2594]. Checkpoints [2239]. Chemical [2694, 2759]. Chemistry [1768, 620, 2484]. CHESSBOARD [2608]. Chimera [2787]. Chinese [1345, 820, 818]. Chip [527, 1067, 2073, 1951]. Choice [48, 2045, 1341, 627, 305]. Choosing [2509]. Chop [1432]. Chordal [2556, 2557]. Chorus [270, 270]. Chorus/Cool [270]. Chromatic [633]. Church [213]. Cicuits [1692]. CIFI [628]. CIM [1346]. CiME [146]. Cipher [2580, 2607]. Ciphers [2253]. Circle [2536]. Circuit [904, 1029, 2306, 2327, 1145, 1061, 1728, 2637]. Circuit-Switched [904]. Circuits [2307, 331, 320, 1029, 1224, 1697, 2143, 1698, 1752, 1469, 325, 75, 73, 2182]. Circular [1489]. Circulation [450]. Citation [628]. Citizens [2501]. CLA [1721]. Clash [1930]. Class [694, 1191, 1219, 826, 828, 58, 840, 842, 2501, 846, 783, 554, 2205]. Class-Sort [1219]. Classes [626]. Classical [2131, 1525, 230, 2758, 153, 192]. Classification [450, 452, 1415, 1989, 2327, 441, 2381, 2127, 1294, 8, 589, 2153, 1269, 2374]. Classification-Based [2381]. Classifications [414]. Classifier [1344, 1989, 1716]. Classifiers [1504]. Classifying [920]. Classroom [1396]. Client [2357]. Client/Server [2357]. Client/Server-Applications [2357]. Clients [2212]. Climbing [1671, 735]. Clinical [323, 1298]. Clique [1681]. Clock [1974, 2276, 2431]. clock-controlled [2431]. Clocked [775]. Clocks [1969]. Close [1757]. Closed [1061]. closing [2516]. Closure [1138]. CLP [1550, 727, 731, 1583, 1834, 703, 2591, 720, 745]. Cluster [2238, 2059, 1756]. Clustering [1865, 2679, 1034, 419, 1399]. Clusters [2064, 2076, 443, 978, 1320, 963, 2054, 1381, 2092]. CM [2354, 842, 2359]. CM-Fields [842]. CMIST [1736]. CMOS [527, 465, 1698]. CNAPS [525]. Co [937, 946, 2647, 2138, 1720, 1622, 1708, 1765, 1788, 444, 1381, 2203]. Co-evaluation [1720]. Co-Evolution [1622, 1788]. Co-evolving [1708]. Co-inductive [2138]. Co-Learning [2647, 2203]. Co-Occurrence [444]. Co-ordinating [937, 946]. Co-Processor [1765]. Co-working [1381]. Coactivation [473]. Coal [437, 1997]. Coalgebraic [37]. CoCheck [2063]. Cochlear [1300]. Code [2284, 1701, 2071, 535, 382, 1333, 915, 1069, 2075, 2651, 2761, 1875, 1071]. Coded [1648]. Codes [407, 1331, 2278, 1335, 1333, 1332, 1327, 932, 2253, 1334, 2744, 771, 1326, 1325, 770, 855, 1766]. Codesign [1723, 1736]. Codewords [1335]. Coding [985, 1343, 411, 1101, 509, 1990, 408, 2585, 2773, 1637, 1731, 477, 518, 1991]. CoDO [2105]. Coefficient [481, 1740, 2157]. Coercions [2129]. Coevolutionary [1621]. Cofibered [1833]. Cognition [2369]. Cognitive [1506, 2154, 506]. Coherence [1086, 2132, 1031, 130, 1057, 1025]. Coherent [468, 1733]. Collaborative [334, 284, 2686]. Collection [1973]. collections [1806]. Collective [1979, 423]. Collector [1972, 112]. College [2008]. Collicular [476]. Collision [360, 1436, 82]. Collision-Avoidance [1436]. Collision-Free [360]. collisionful [2442, 2441]. Collisions [350]. Collocation [979]. collusion [2459]. Colonoscopy [1296]. Color [1349, 446, 803, 1255, 1464]. Color-Based [1349]. Color-Calibration [446]. Coloring [750, 752, 751]. Colors [1455]. Colossal [1245]. Column [1898]. Column-Based [1898]. Combination [974, 708]. Combinations [2318, 195]. Combinator [1208, 1209]. Combinatorial [1984, 1793, 2096, 2017, 759, 778, 1508, 1517]. Combinatorics [777]. Combinatory [127]. Combined [693, 177]. combiners [2431]. Combines [2027, 2755]. Combining [1535, 1542, 60, 2166, 502, 734, 1225, 607, 1326, 2173, 2319, 2066, 2323, 1831, 100, 77]. commented [1873]. Commercial [2326]. Commitment [360]. Commodity [1320]. Common [2131, 1468, 2549, 850]. Commonsense [615]. Communicate [1198]. Communicating [223, 2667]. Communication [1349, 2465, 2492, 1963, 374, 930, 918, 1970, 1163, 1055, 1424, 889, 967, 2218, 1330, 913, 2455, 1965, 673, 1012, 1041, 2077, 1011, 899, 1036, 1044, 1005, 2473, 1460, 2768, 955, 2214, 348, 2543, 880, 62, 99]. Communication/Computation [967]. Communications [2622, 901, 614, 2773, 2767, 2428]. Commutative [149, 80]. Commutativity [911]. Compact [1670, 851]. Companding [1329]. Comparative [501, 825, 1676, 394, 1789]. Compared [1654]. Comparing [522, 973, 75]. Comparison [1086, 976, 249, 445, 1820, 1339, 1286, 1672, 881, 858, 1696, 2733, 1673, 1674, 1675, 1857, 2243, 2050, 788, 1046, 718, 1111, 1802]. Compatible [1543]. Compensation [1259]. Competence [1933]. Competing [1692]. Competition [2281, 163]. Competitive [1484, 1485, 517, 2578, 175]. Compilation [398, 197, 193]. Compile [1035, 380, 1021]. Compile-Time [1035, 380, 1021]. Compiled [997]. Compiler [677, 918, 382, 1502, 2744, 914, 1107, 2659, 126, 1947]. Compiler-Interpreter-System [1502]. Compilers [905]. Compiling [1590, 1200, 1428, 1826]. complement [2526]. Complementary [759]. Complementing [1572]. Complete [1136, 116, 901, 1007, 2007, 757, 650, 229, 558]. Completeness [2134, 15, 1227, 1228, 1555, 202]. Completion [146, 150, 115, 1544, 1457, 431]. Completion-Based [150]. Complex [2356, 466, 990, 1075, 219, 39, 1363, 2359, 739, 1943, 1905, 227, 1373, 2759]. Complex-Object [1075]. Complexity [2565, 779, 1337, 2535, 608, 729, 2016, 886, 1335, 2794, 1364, 688, 685, 416, 2604, 1475, 749, 372, 850, 1637, 417, 742, 996, 2191, 547, 1865, 2190, 540, 543, 2182]. Complexity-Theoretic [749]. Compliant [2211, 1898, 2057, 2310]. Component [276, 297, 31, 490, 1770, 428, 2662]. Component-Based [276, 297, 31]. Components [1939, 56, 2526, 74]. Componentwise [2563]. Composable [2403]. Composed [829]. Composing [1202]. Composite [1366, 1388]. Composition [1574, 1100]. Compositional [252, 134, 479]. Compositionality [1447]. Compositions [2346]. Compound [1185]. Compression [773, 1341, 986, 2084]. Comprising [1333]. Compromised [1537]. Computability [1112]. Computable [1108]. Computation [937, 1535, 918, 1732, 1147, 1437, 1191, 2178, 889, 877, 967, 1906, 836, 383, 839, 1189, 2728, 1549, 2604, 1486, 935, 1400, 2080, 969, 1905, 2652, 2038, 1639, 680, 403, 1534, 1018, 2090, 2589, 1547, 970, 536, 2521]. Computational [1614, 1557, 1895, 2539, 512, 1012, 646, 711, 842, 2484, 1890, 1894, 1494, 847, 849, 1889, 1637, 2052]. Computations [2689, 1000, 60, 1322, 2583, 268, 923, 1605, 1897, 850, 2759, 196, 2539]. Compute [922]. Computed [1304]. Computer [1614, 239, 2000, 1026, 1168, 2702, 1195, 980, 658, 2483, 321, 2179, 2715, 2447, 2079, 322, 1193, 2477, 2729, 1268, 1188, 216, 1047, 777, 1311, 1187, 288, 1312, 811, 1197, 291, 1727, 295, 1790, 304, 281, 1722, 2759, 324, 231, 226, 244, 2803, 2513]. Computer-Aided [2000, 216, 244]. Computer-Assisted [2079, 1311]. Computer-Based [288, 281, 226]. Computer-Supported [322]. Computers [868, 1780, 330, 2721, 974, 2742, 2243, 2752, 2751, 878, 2755]. Computing [1128, 823, 2693, 355, 825, 2700, 1615, 826, 618, 827, 828, 2554, 1726, 1945, 1144, 2553, 778, 1059, 2714, 857, 837, 838, 1734, 1723, 1482, 1213, 2059, 2733, 841, 1899, 2484, 859, 844, 1983, 988, 1009, 989, 1166, 846, 1317, 2181, 953, 1725, 2052, 2781, 961, 2237, 856, 1185, 1766, 924, 1683]. Concatenated [1333]. Concentration [447]. Concentrators [886]. Concept [607, 2264, 299, 1529, 1028, 589, 2200]. Concepts [623, 1916, 864, 1220, 2483, 2369, 1415, 1184, 607, 609, 2662, 2105, 2219, 217, 2191]. Conceptual [641, 2116, 2106, 648, 649, 646, 644, 642, 2124, 643, 645, 2118, 1414, 215, 2121, 2105, 654, 2103, 636, 651]. concrete [541]. Concurrency [97, 96, 1447, 2212, 1357, 1855, 2595]. Concurrent [1557, 944, 2491, 269, 949, 1567, 1589, 2608, 268, 1223, 939, 115, 2500, 619, 940, 1764, 1527, 2610, 716, 1602, 264, 1607, 97, 112]. Condition [892, 1486]. Conditional [934, 132, 1227, 2277, 61, 2150, 1173, 1177, 1881]. Conditions [20, 1757, 408]. Conducting [453]. Conductors [826]. Conference [2774, 2743, 2803]. Confidence [533]. Confidentiality [1368, 2454, 2036]. Configurable [1746]. Configuration [2350, 2065, 2357, 1921, 2351, 2358, 219, 2352, 2349, 1922, 408, 2384]. Configurations [2339]. Configuring [2343]. Conflict [256]. Confluence [689, 1131, 132, 128]. Confluent [1605]. Confocal [1265]. Conformation [1710]. Congestion [775, 561]. Congruence [16, 851, 854]. Conjecture [830]. conjectures [157]. Conjugate [1026]. conjugation [572]. Conjunction [589]. Conjunctive [2618, 1588]. conjunctively [63]. Connected [1476, 754, 1490, 815, 1467, 412, 755]. Connection [1503, 2642]. Connectionist [462, 460]. Connections [472, 1145]. Connectives [1648]. Connectivity [2785, 1387]. Conquer [1782, 1596, 2631, 2106]. Consensus [1966, 1986, 1964]. Conservative [623]. Conservativity [2139]. Considering [2393]. Consistencies [738]. Consistency [690, 34, 1350, 726, 1542, 255, 1979, 733, 1351, 735, 2316, 601, 715, 254]. consistent [2525]. constancy [1797]. Constant [886, 916, 1740, 1969]. Constant-Degree [916]. Constrained [1128, 1043, 492, 260, 523, 246, 2158, 2660, 2021, 257, 1832, 253, 2764, 2606, 1868, 1871]. Constraint [689, 1809, 2789, 1002, 1962, 691, 1560, 1669, 2787, 1701, 2404, 247, 1550, 727, 2795, 1567, 2786, 728, 762, 2619, 2005, 695, 696, 729, 699, 1815, 1222, 1693, 700, 731, 1561, 248, 2779, 256, 701, 748, 2006, 702, 734, 705, 252, 258, 1835, 708, 2007, 737, 711, 712, 2026, 2003, 250, 738, 2780, 962, 2747, 713, 2796, 715, 716, 1602, 1613, 1831, 743, 1606, 2593, 245, 719, 2792, 2781, 1607, 2668, 722, 254, 253, 2782, 259, 251, 746, 2643, 724, 2197, 1869]. Constraint-Based [1809, 696, 738, 1602]. Constraints [690, 2784, 694, 698, 1279, 1368, 1499, 2790, 1351, 709, 1133, 2412, 1037, 739, 1388, 1828, 1841, 721, 551, 1870, 540]. Construct [2045, 1232]. Constructing [2624]. Construction [1136, 2062, 2016, 832, 564, 2335, 390, 2145, 2342, 2423]. Constructions [840, 2145, 208]. Constructive [18, 2133, 2619, 2134, 766, 495, 1197, 2198, 2147, 1533, 488]. Constructor [119]. Consultations [586]. Consumers [324]. Consumption [2570]. contained [2561]. Containment [1898]. Content [286]. Content-Balanced [286]. Contention [1024, 1055]. Context [1605, 2452, 1235, 2004, 2144, 788, 783, 506, 1868, 562, 1866]. Context-Based [2452, 506]. Context-Free [788, 562, 1866]. Context-Relative [2144]. Context-Sensitive [1605, 783, 1868]. Contexts [1514, 777]. Contiguity [772]. Contiguous [2755, 2234]. Continued [825]. Continuity [597]. Continuous [1625, 1678, 1409, 2625, 1633, 225, 533, 717, 84]. continuous-time [84]. contour [2529]. Contours [809]. Contraction [610, 611]. Contradiction [649]. Control [2305, 2282, 2283, 894, 2284, 1957, 2285, 2217, 1939, 2364, 439, 2286, 1074, 1198, 1945, 1965, 2293, 2294, 2661, 947, 1357, 2297, 389, 2339, 1855, 2642, 2329, 1943, 2605, 2127, 2299, 503, 2348, 2435, 2300, 476, 2301, 2595, 1016, 2302, 2303, 507, 504, 2511, 170]. Controlled [2276, 2625, 1363, 1854, 2431]. Controller [1024, 2292, 22, 2296, 2321]. Controllers [1433, 2616, 1622, 230]. Controlling [475, 1588, 478]. Controls [1841]. Convective [2754]. Convergence [2231]. converse [205]. Conversion [2495, 1096, 2768, 442]. Converters [231]. Convertible [2273]. Converting [192]. Convex [1477, 1901, 501, 2544, 2554, 1032, 2574, 1899, 717, 2537]. Convolution [466]. Convolution-Type [466]. Convolutional [1758]. Cool [270]. Cooling [2020]. COOP [2339]. COOP/ [2339]. Cooperate [1198]. Cooperating [1551]. Cooperation [1184, 1617, 2226]. Cooperative [1929, 921, 1958, 1375, 627, 303, 586, 304, 2595]. Coordinating [2652]. Coordination [1927, 2155, 665, 2230, 941, 2172, 45, 948, 2222]. Coordinator [944]. Coping [1383]. coprocesseur [1071]. Coprocessor [994, 1724]. Coq [2135, 1093, 2138, 1101, 2143]. CORBA [2211, 2213, 2212, 2476, 2220, 7, 2225]. CORBA-Based [2220]. CORDIC [987]. Core [2695, 2486, 507]. Cores [1263]. Cornell [2232]. Corporate [2395]. Corpus [449]. Correcting [1208, 1209]. Correction [786, 1261, 1255]. Corrections [1336]. Correctness [928, 1815, 2330, 939, 1230, 938, 1107]. Correlated [417]. Correlation [2266, 1343, 2276, 2277, 792, 2157, 2431]. Cortex [509, 408, 513, 1530]. Cortical [404, 473, 1279, 1300, 1253, 472, 474]. Cortico [506]. Cortico-Hippocampal [506]. Cortio [469]. Cortio-Tectal [469]. COSPAN [103]. Cost [786, 867, 604, 2193, 2269, 1021, 2385, 1660, 2764, 581, 2393]. Cost-Based [604]. Cost-Effective [2269].

Cost-Efficient [786]. Cost-Sensitive [581, 2193]. count [1871]. counter [2192]. Counterexamples [1115]. Counters [2638]. Counting [822, 1391]. Counting-Based [1391]. Coupled [2703, 2723, 1260]. Coupling [2694, 511, 2629, 2444]. Course [2017, 2010, 322, 300, 302, 2021, 2387, 2052]. Course-Timetabling [2010]. Courses [307]. Courseware [309, 298]. Coverage [1701, 2315, 1961, 1940]. Coverage-Directed [2315]. Covering [1483, 2133, 2168, 1782, 1718]. Covert [2453, 2451, 2452]. CPM [1337]. CPU [1951]. CQL [2783]. Craniofacial [1307]. Crashes [1971]. CRAY [2066, 2064, 1055, 2711]. CRAY-T3D [2066]. CRC [1732]. CREAM [300]. CREAM-Tools [300]. Create [436]. Creating [2368, 621, 622, 433]. Creation [1308, 2511]. Criteria [1769, 2426, 1864]. Criterion [2385, 1865]. critic [155]. critical [1949]. CROQUE [1385]. Cross [1412, 1281]. Cross-Lingual [1412]. Crossbar [1748]. Crossbars [895]. Crossing [745, 2221, 2573]. Crossover [1643, 2727, 1990, 2157, 1709]. Crossovers [1646, 1648]. Crowns [757]. Cryosection [1255]. Cryptanalysis [2255, 364, 2580, 2432]. Cryptographic [2248, 1089, 2278, 2459, 2580, 2260, 2272, 2635]. Cryptography [367, 375, 370, 2251, 347, 2425, 2581, 351]. Cryptosystems [2440, 349]. Crystal [2073]. CS [2096]. CS-2 [2096]. CSCW [2770]. CSF [1253]. CSF/ [1253]. CSP [692, 600, 710, 714, 1850]. CSPs [249, 260, 735, 717, 742]. CT [1274, 1268, 1314, 1241]. CtCoq [52, 173]. CTL [2336]. Cube [776, 966]. cubes [2535]. Cubic [823]. Cumulant [425]. Cumulant-Surrogate [425]. Current [465, 2483, 1318, 1892]. Current-Mode [465]. Curriculum [300]. Curryfied [1229]. Curse [1684]. Curvature [1795]. Curve [2162, 810]. Curves [822, 420, 832, 838, 2679, 840, 2247, 849, 533]. Custom [1734, 1723, 1746]. Custom-Configurable [1746]. Customizable [272]. Cut [1554, 1069, 2666, 2189]. Cuts [2574]. Cutwidth [887]. CW1 [2289]. CW2 [2290]. Cybernetic [1932]. Cycle [779, 1479, 908, 197]. Cycles [1973]. Cyclic [1503, 1041, 1333, 1327, 876, 1972]. cylinders [1792].

D [1726, 1029, 2713, 1164, 2532, 1805, 1026, 445, 1309, 1131, 320, 1029, 2534, 1296, 993, 1246, 791, 1257, 1256, 2521, 807, 657, 1907, 1796, 813, 1287, 2163, 1704, 816, 1791]. D-problems [1704]. D-Schema [320]. D-WFS [1131]. DAC [710]. Dagger [941]. Dags [1038, 1593]. Dangers [2693]. Dark [351]. DASE [237]. Data [1164, 1349, 1955, 1644, 1353, 773, 917, 678, 1315, 1324, 2795, 927, 667, 2107, 435, 1438, 425, 1322, 2669, 1367, 2078, 2682, 2647, 2547, 1343, 2719, 1012, 1409, 2675, 38, 1223, 36, 1157, 1960, 2077, 1020, 2498, 2448, 1320, 2397, 14, 2677, 2671, 1405, 2678, 2609, 935, 261, 1290, 1142, 952, 2220, 1215, 1323, 1362, 1544, 2114, 1855, 2055, 945, 950, 2113, 676, 1386, 874, 1206, 1234, 1825, 1317, 395, 876, 973, 1278, 496, 417, 2684, 2387, 953, 2506, 1722, 2482, 1777, 922, 1953, 654, 1161, 1552, 1389, 519, 518, 1058, 1401, 2686, 956, 544, 2206, 1884]. data [1859, 2205, 1874, 160, 2204, 578]. data-dependent [1884]. Data-flow [36]. Data-Flow-Based [1825]. data-independent [1884]. Data-Mining [2397]. Data-Parallel [1020, 952, 945, 950, 874, 953]. data-types [578]. Database [2670, 2685, 2108, 1352, 2491, 1077, 1350, 726, 1162, 1149, 2786, 818, 698, 2122, 1199, 1196, 648, 1081, 1360, 2790, 2779, 1372, 1371, 1374, 1390, 2679, 2672, 2039, 1395, 1410, 1357, 1365, 1608, 1404, 2098, 1630, 2118, 2681, 845, 1160, 2060, 2480, 2037, 860, 591, 2791, 2796, 2215, 2111, 981, 1422, 2123, 2105, 2024, 1380, 1423, 2110, 2668, 1853, 1399, 1398, 2680, 1403, 1347, 2125, 1379, 2512, 2433]. Databases [2789, 1358, 2042, 2029, 1353, 2041, 1537, 2495, 2040, 1356, 2033, 1392, 649, 748, 2043, 311, 1361, 2023, 2028, 2120, 2674, 1370, 1855, 2026, 2022, 1376, 2628, 1411, 2798, 1377, 1369, 1418, 1346, 2031, 2036, 2792, 2119, 2782]. Dataflow [2650, 912, 2093, 910, 2626]. Datalog [2041, 2025, 2785, 2781]. Datapath [1729]. Datapaths [1739]. Dataset [1249]. Date [2267]. DB [641]. DB-MAT [641]. DBMS [1079, 1075]. DBS3 [1149]. DC [2291]. DCE [2069]. DCS [782]. DCSPs [743]. Dead [915]. Deadlines [2579]. Deadlock [2596]. Dealing [870, 395]. Debugging [263, 2048, 862, 2593]. Decay [494]. December [2803]. Decentralized [1912, 1680, 291, 170]. Decidability [1569, 2142, 1829, 1601]. Decidable [1449, 141, 2499, 554, 187]. Deciding [223, 1089]. Decision [2364, 2648, 2498, 2477, 2165, 1520, 804, 1642, 2393, 2321, 195, 203, 1858, 2190]. Decision-Theoretic [2477]. decision-trees [1858]. Declarative [2041, 950, 2593, 2651, 587]. Declustering [1077, 1150, 1081, 1419]. Decoding [2278, 1502, 1333, 1758, 2464]. Decomposable [1489]. Decompose [2750, 2021]. decomposed [63]. Decomposing [660]. Decomposition [2100, 2692, 2088, 588, 664, 665, 2335, 1420, 1363, 1718, 2739, 2782, 800]. Decompositions [757]. Deconstructionist [281]. Dedicated [1764]. Deduction [1588, 2140, 389, 2629, 194, 76, 176]. Deductive [2491, 2042, 1196, 2669, 2040, 2033, 1392, 2043, 2039, 2023, 1365, 1608, 2026, 591, 2031, 79, 2024, 2102, 101, 77, 206]. deductive-algorithmic [101]. Default [623, 1128, 626, 1126, 618, 1127, 625, 1138, 1519, 1125, 257, 181]. Defeasibility [703]. Defeasible [1510, 2148, 250]. Deficiency [491]. Definability [1579]. Definable [2597, 721]. Defined [1211, 2533]. Definite [2619, 2722]. Definition [7, 1109]. Definitional [124, 191]. definitions [167]. Deforestation [1592]. Deformable [1306]. Deformation [810]. Degree [828, 916, 962, 684, 751, 1419]. Delaunay [1996, 1907]. Delay [2308, 890]. Delays [1041, 1460, 2764]. Deletion [1466]. Deletions [1336]. Delineation [1272]. Delta [1755, 2340, 1760, 767]. DeltaBlue [718]. Deluge [1682]. Demands [2036]. Democratic [1678]. Demodulation [1733]. Demonstration [1539, 51, 45]. Demonstrators [1742]. Demultiplexer [1744]. Dendritic [516]. Dense [2746]. Densities [530]. Density [1635]. Deontic [1220]. Dependability [1958, 1945, 1943, 1946]. Dependable [1952]. Dependence [1164, 908, 1786]. Dependencies [2346, 1383, 1870]. Dependency [1501]. Dependency-Directed [1501]. Dependent [2679, 531, 123, 1884]. Depiction [1248]. Deploying [1936]. Deployment [2769, 2358, 2398]. Depth [886, 1594, 1469, 521]. Depth-First [1594]. Derivation [2325, 1091, 698, 931, 341, 2651]. Derivatives [790]. Derived [1099]. Deriving [460, 1354]. Desaturating [481]. Descent [451]. Described [783]. Description [269, 2230, 760, 2375, 393, 1784, 46, 1795, 182, 181, 184]. Descriptions [1452, 2618, 1791]. Deserve [1914]. Design [2108, 2283, 1754, 2325, 2429, 234, 1768, 933, 228, 1711, 2772, 1741, 2218, 279, 2654, 1119, 1712, 256, 1715, 1333, 858, 1332, 2721, 311, 1374, 1739, 125, 1770, 1932, 2306, 1395, 2730, 1420, 1952, 2327, 1104, 271, 1987, 2298, 1697, 1365, 1724, 749, 1899, 2616, 2098, 2099, 313, 811, 241, 2241, 1666, 295, 1996, 1764, 1889, 2662, 2046, 1759, 1761, 227, 679, 8, 2105, 1690, 2389, 274, 237, 1763, 2612, 231, 325, 315, 1384, 2237, 2637, 871, 1240, 218, 1347, 2125, 244, 2607, 2394, 1379, 2428, 2801, 163]. Designer [236]. Designing [1063, 1685, 1407, 675, 332, 273, 312, 1894]. Designs [1753, 2118, 2329, 1760, 954, 105]. Desirable [2784]. Detecting [649, 470]. Detection [1966, 1479, 1973, 1942, 1426, 1486, 1775, 2596, 113, 2438, 2420]. Detectors [1981]. Determinacy [2666]. Determination [2527]. Determining [2526, 842]. determinism [1239]. Deterministic [2565, 1208, 1974, 2045, 37, 2313, 1565, 843, 1049, 1209, 550, 621, 622]. Develop [1852, 2154]. Developing [239, 2774, 1719, 874, 1622]. Development [2281, 2347, 2698, 1742, 2356, 2357, 437, 39, 1436, 1932, 2398, 1053, 242, 1524, 1666, 1926, 474, 1762, 55, 2075, 2147, 226, 1837, 584, 97, 2799]. Developments [1999, 1423]. Devex [2724]. Devices [406, 1511, 1757]. DFA [1884]. DFTs [1726]. Diagnosable [1963]. Diagnosis [1962, 2091, 328, 2164, 1511, 596, 1961, 323, 1954, 584, 2200]. Diagnostic [1119, 2376]. Diagnostics [436]. Diagonally [968]. diagram [86]. Diagrams [2493, 2613, 129, 1378]. Dialogs [1512]. Diameters [761]. Dichotomic [2166]. dichotomous [1463]. dictionaries [2198]. Dictionary [1412, 2550, 1473, 958]. Didactic [329]. Difference [1477, 2716]. Differences [2361, 1291, 471]. Different [2361, 786, 1216, 1754, 525, 974, 2244, 1637, 1529, 2319, 83, 191]. Differential [2690, 361, 2753, 85]. Difficult [853]. Difficulty [1635]. Diffie [355, 352, 366]. Diffusion [1641]. Digit [441]. Digital [2688, 2465, 2259, 2325, 1839, 1692, 2638, 1158, 1256, 2274, 804, 1, 302, 672, 1752, 325, 2586, 2637, 2526, 2530, 2514]. digitization [2536]. Digraphs [897, 670]. DIIRK [973]. Dijkstra [2653]. Dilation [888]. Dimension [1779, 2571, 441]. Dimension-Exchange [2571].

Dimensional [775, 1948, 48, 2550, 1286, 904, 2783, 931, 763, 499, 1255, 811, 1716, 471, 1298, 1622, 1294, 2311, 476, 1289, 1640, 2516, 1077, 2572]. Dimensionality [1684, 1798]. Dinner [2743]. Diophantine [571]. DiP [1053]. Direct [968, 235, 1246, 2754]. Directed [381, 1590, 1501, 2315, 1647, 1487, 2572, 1718, 61, 757, 1810, 2558, 254, 1881]. direction [552]. Directions [2574, 2031, 2056]. disagreement [1871]. Disambiguation [1532]. DiscAtinf [1169]. DISCIM [1375]. Disciplines [294]. Disclosure [1854]. DISCO [2791, 2796]. Discontinuous [810]. Discourse [597, 2271, 2161]. Discovering [2551, 609]. Discovery [2099, 2506, 200]. Discrete [23, 1873, 2518, 1678, 1134, 734, 2603, 1565, 1042, 220, 852, 1765, 856, 795, 2540, 2535, 2539, 2513, 2527, 2521, 1878, 2515]. Discrete-Event [1042]. Discriminant [496]. Discriminants [826, 835]. Discrimination [1114]. Discussing [933]. Discussion [1893, 293]. Disease [323]. Disjoint [2562]. Disjunction [589, 1533]. Disjunctions [709]. Disjunctive [1130, 1497, 2033, 1392, 1400, 2034, 2660]. Disk [2578]. DisLoP [1497]. Disorders [309]. Display [1293, 1313, 1252]. Disruption [2275]. Distance [2278, 2549, 303, 2445, 304, 755, 2431]. Distance-Hereditary [755]. Distinct [1659]. Distributed [879, 1929, 2769, 272, 1043, 1162, 2217, 1939, 1967, 927, 687, 667, 1970, 928, 662, 1186, 266, 2475, 2587, 1500, 1973, 882, 2218, 1655, 913, 1144, 2357, 665, 1938, 1965, 2712, 857, 866, 1978, 1505, 265, 2224, 2721, 861, 2614, 735, 1451, 1405, 2244, 604, 6, 923, 2223, 1365, 926, 865, 2165, 2737, 2738, 781, 1983, 2628, 938, 966, 2210, 969, 782, 1842, 2605, 293, 2594, 929, 1961, 2348, 1972, 936, 925, 1848, 953, 2229, 743, 2595, 1402, 1969, 277, 147, 2090, 2767, 2686, 2763, 1403, 2596, 87, 2706].

Distributed-Memory [2721, 2244, 938]. Distributed/Parallel [665]. Distribution [1048, 374, 1065, 2081, 685, 2077, 2571, 1142, 371, 2770, 2201]. Distributions [1695, 1632, 1633, 1766, 577]. Distributive [136]. Disturbing [324]. Divide [1782, 1596, 2631]. Divide-and-Conquer [1596, 2631]. dividing [2535]. dividing-cubes [2535]. Divisible [2560]. Division [2310, 71, 72]. Divisor [2131, 850]. DLP [147]. DLW [328]. DNA [536, 1991, 1683]. DNF [2183]. Do [1914, 387, 1718, 2645, 814]. Document [1413, 2366, 2462, 1536, 1185]. Document-Centered [1185]. Documented [2284]. Documents [2359, 643, 1421, 1185]. DOL [950]. Domain [2692, 728, 2088, 664, 665, 1609, 655, 1138, 620, 1108, 715, 1423, 2797, 156, 1883]. Domains [1218, 2622, 258, 1515, 1032, 2376, 1570, 846, 717, 1177, 1833, 2221]. Dominant [968]. Dominating [1467]. Domination [755]. Dominators [1812]. Don’t [1239]. DOOD [1532]. Dopamine [512]. Dorn [2732]. DOSMOS [864]. DOSMOS-Trace [864]. Double [760]. doubly [575]. doubly-chained [575]. Down [820]. DPLGraphs [1080]. Dragon [2508]. Drawing [637, 1455, 814]. Drawings [815, 1904, 522, 816]. Driven [894, 881, 2718, 941, 1436, 1515, 474, 1722, 2529]. DROL [274]. DRS [2325]. DS [1339]. DS/FH [1339]. DS/FH-CDMA [1339]. DSA [1270, 2246, 2247]. DSM [868]. DSM-Based [868]. DSOM [479]. DSP [1742, 997, 244]. DSPACE [2561]. DSS [352, 243, 350]. Dual [1666, 442]. Dual-Agent [1666]. Duality [704]. Dummy [120]. Duration [1433, 1434, 2617]. During [586, 517, 2393, 1883]. Dyadic [770]. Dynamic [1955, 2402, 1048, 1937, 1669, 678, 439, 1063, 1045, 2242, 2550, 1150, 1220, 1815, 1322, 2669, 2112, 652, 1462, 1473, 673, 1978, 1556, 256, 701, 1458, 2614, 2727, 681, 2140, 2244, 1439, 422, 1047, 566, 676, 1705, 424, 596, 1990, 2243, 2570, 1586, 1612, 860, 513, 661, 878, 1691, 2757, 682, 550, 205, 2420, 578]. Dynamical [427]. Dynamically [2368, 1559, 684, 1747, 186]. Dynamicity [945]. Dynamics [2064, 2071, 1179, 459, 423, 2733, 472, 1681, 2759].

E-BSP [1011]. E-Cash [2270, 2271]. E-pattern [568]. E-Unification [1569, 154]. E0* [146]. Eager [1045]. Early [1426, 2780, 242, 418, 1736]. Easier [2021]. East [2011]. Easy [866, 2233]. EBMT [2367]. ECHIDNA [48]. Echo [2467]. Echocardiograms [1272]. ECL0iPS0e [2797]. Economic [1048]. Economic-Based [1048]. Economy [316]. Eddy [2754]. Eden [949, 1597]. Edge [752, 533, 1639, 2562]. Edge-Disjoint [2562]. Edit [2431]. Editing [2549, 645, 2595]. Education [585, 6, 1762, 324, 283]. Educational [310, 312, 282]. EEG [443, 1298]. EER [2109]. EER/GRAL [2109]. Effect [894, 1644, 2649, 280, 1944]. Effective [2269, 2750]. Effectiveness [1545, 852]. Effects [1088, 1655, 1510, 2194, 1638]. Efficiency [2259, 1052, 411, 150, 1207]. Efficient [786, 1968, 919, 2074, 951, 729, 1074, 699, 138, 971, 890, 675, 663, 2192, 1609, 1462, 2271, 357, 1012, 2045, 1019, 2575, 681, 885, 2661, 2335, 388, 1487, 1952, 712, 1370, 1544, 2737, 1037, 1440, 844, 1959, 412, 1666, 876, 925, 2445, 2504, 2585, 783, 2386, 193, 1778, 1728, 1607, 582]. Efficiently [388, 763, 1514]. Effort [1920]. Eiffel [267]. Eiffel/ [267]. Eigenproblems [972, 2760]. Eigensolutions [969, 970]. Eigenvalue [2737]. Elastic [313, 1278, 1282]. elastica [2518]. Elasticity [1247]. Elastography [1247]. Elections [2256]. Electric [1997]. Electrical [1300, 1945]. Electricity [4]. Electro [1993]. Electro-Mechanical [1993]. Electromagnetic [977]. Electron [337, 1695]. Electronic [307, 2256, 2261, 1837, 1847]. Electronics [333, 303, 302]. Element [971, 1667, 2522]. Element-by-element [971]. elementary [179]. Elements [1305, 1094]. Elicit [329]. Eliciting [1934]. Elimination [1170, 120, 164, 975, 915, 1083, 165, 2200, 185]. Elliptic [832, 838, 2247, 854]. Embeddable [1168]. Embedded [4, 2350, 1957, 2477, 1446]. Embedding [775, 1095, 681, 9, 653, 168]. Embeddings [888, 885]. Emergence [31]. Emergent [1928, 1617]. Emphasis [1257]. Emphasizing [533]. Empirical [2071, 701, 702, 2340, 2114, 2243]. Employee [2003]. Employing [1355]. Emulation [1722]. Enables [477]. Enacting [1929]. Enactment [1927, 1913, 1354]. Encapsulation [1151]. Encoding [1755, 652, 1619, 409, 1103, 2334, 598, 2777]. Encodings [2146]. Encryption [374, 2424]. End [2633, 313]. End-to-end [2633]. Ended [1619]. Endoscope [1244]. Endoscopy [1310, 1304]. Energy [244]. Engine [1911, 1738, 1835, 1538]. Engineering [1912, 1536, 1928, 2122, 815, 2345, 338, 298, 6, 1666, 294, 1889, 1907, 2123, 224, 2759, 226, 816, 1396, 218]. Engineers [7]. Engines [2397]. Enhance [1270]. Enhanced [2384]. Enhancement [1254]. Enhancing [1958, 2679, 1854]. Ensure [2658]. Entity [2113, 2117]. Entity-Relationship [2117]. Entropy [1270, 1522, 2383]. Entropy-Based [2383]. Enumerating [2558]. Enumeration [1073, 752]. Envelope [2615]. Environment [2685, 2108, 327, 726, 1937, 328, 40, 2068, 2337, 234, 870, 228, 2356, 297, 2218, 306, 1753, 2723, 39, 2059, 44, 1082, 1053, 1757, 288, 5, 1375, 1935, 1720, 1665, 300, 627, 293, 1175, 295, 296, 2480, 7, 1725, 55, 274, 237, 2644, 1752, 277, 1302, 53, 871, 326, 2064, 97, 110, 109]. Environmental [2165, 226, 2680]. Environments [1912, 2083, 2177, 1615, 1958, 275, 1438, 2357, 1933, 1914, 2345, 866, 285, 1390, 2244, 289, 1890, 1983, 1990, 225, 405, 2024, 1789, 1652, 2767, 2237, 2799]. Envisioning [605]. Ephemeral [1374]. EPIC [148]. Epilepsy [443, 1297, 1298]. Epiphyte [284]. Episodes [2380]. Epistemic [2150]. EPLD [1757]. Equality [2318, 1170, 1821, 2139, 154]. Equation [2086, 979, 977, 853]. Equational [16, 1172, 580, 1499, 1222, 558, 120, 1509, 390, 148, 1239]. Equations [2690, 2710, 1822, 15, 837, 2722, 2067, 1227, 567, 1428, 2753, 1238, 85]. Equitable [750]. Equivalence [2566, 1231, 2664, 568]. Equivalent [2247]. erasing [2202]. ERCW [1005]. Erroneous [1642]. Error [786, 608, 1208, 1942, 1336, 2604, 1758, 1209]. Error-Correcting [1208, 1209]. Errors [784]. Escrow [2469, 2424]. Essentials [1210]. ESSL [2712]. Establishing [1350, 2451]. Establishment [2429]. ESTEREL [1450, 18]. Estimate [1259]. Estimates [2713]. Estimating [414, 1920]. Estimation [1337, 523, 1060, 1994, 1632, 1633, 1731, 1881, 2201]. Estimations [2235, 1940]. eta [122]. eta-Long [122]. Ethernet [1436, 880]. Euclidean [2528, 1808]. EULE2 [1520]. Euler [1973]. Europe [1837]. Evaluate [2718]. Evaluating [1926, 2159, 742, 1017, 282]. Evaluation [1814, 1266, 1218, 381, 1590, 1945, 2715, 1942, 2716, 1941, 1033, 525, 2721, 939, 386, 2615, 2658, 1420, 603, 1258, 1773, 2780, 1705, 301, 302, 2037, 2046, 2385, 232, 398, 399, 1143, 1591, 1417, 1030, 282, 305, 1399, 1148, 1720]. Evaluators [391]. Even [1822]. Event [1367, 734, 1042, 220, 1380, 1765, 1878]. Event/Rule [1380]. Events [1366, 923, 619, 2770]. Every [1654]. EVGC [1453]. Evidential [613, 2438, 797]. Evolution [1614, 1626, 1653, 2126, 1713, 2614, 1616, 2232, 1618, 1619, 1919, 1263, 391, 1620, 1909, 2127, 1622, 1650, 1623, 1778, 1691, 278, 592, 1788]. Evolutionary [1625, 2155, 1701, 1615, 1767, 1768, 2012, 1693, 1686, 236, 1712, 1647, 1782, 1715, 1703, 1695, 1649, 1987, 2158, 1676, 940, 1657, 1681, 2095, 714, 2018, 1636, 1682, 1639, 403, 1707, 1658, 1787, 1704, 2374]. Evolvable [1689]. Evolved [1618]. Evolving [1929, 1670, 1918, 47, 2222, 1777, 1708]. EVP [99]. Exact [492, 2096, 2187]. exactly [2191]. Exam [2013]. Examination [2019, 2002, 1999, 2011]. Example [1176, 1156]. Examples [2551, 1685, 609, 188]. Exception [624]. Exceptionally [692]. Exchange [900, 2571, 2442]. Exchanger [1713]. Exclusion [1527]. Exclusive [1735, 2183]. Exclusive- [1735]. Exclusive-/Inclusive-OR [1735]. exclusive-or [2183]. Exclusively [1939]. Executable [2284, 1598, 2229, 1186]. Execution [1080, 1437, 2615, 932, 952, 1404, 2158, 2053, 973, 1049, 922, 1078]. Executive [1895]. Exercise [2219]. Exercises [308]. Exhaustive [1437, 365]. exist [560]. Existential [28, 647, 715]. Existing [1095, 1369]. EXOR [1773]. Expandable [464]. Expansion [623, 2462, 611, 1811, 2659, 2183]. Expensive [1044]. Experience [4, 2350, 2693, 2068, 2396, 2356, 2011, 1724, 199]. Experiences [1079, 1372, 2351, 734, 1924, 259]. Experiment [1535, 1911, 857, 515, 147, 2437]. Experimental [2108, 1992, 306, 656, 296, 718, 1143, 1752, 1802]. Experimentation [1784]. Experimenting [1176]. Experiments [968, 677, 1218, 2312, 199, 2656, 2051, 343]. Expert [1423, 1408, 222, 584, 172, 174]. Expertext [285]. Explained [472]. Explaining [2371, 2380]. Explanations [27]. Explicit [1122, 1604, 127, 48, 1574, 1603, 128, 1663, 546]. Explicitly [910, 2756, 2651]. Exploiting [1000, 1692, 475, 1176, 710, 90]. Exploration [2672, 1778, 237]. Explorations [2642]. Exploratory [1344]. Explorer [1613]. Explosion [1517]. Exponential [346, 528, 549]. Exponentials [988, 989]. Expressed [955]. Expressibility [220]. Expressing [2785]. Expression [1593, 1777]. Expressions [1088, 397, 783, 1393]. Expressive [2033, 2045, 2590]. Expressiveness [2784, 2044]. Exten [618]. Extended [1441, 836, 1529, 792, 589, 440, 2303, 190, 562, 568]. Extending [1933, 590, 872, 2087]. Extensible [1258]. Extension [2168, 264]. Extensional [128]. Extensionality [2600]. Extensions [1128, 2042, 2622, 618, 155, 388, 1138, 1518, 2014, 2781, 488]. Extensive [1644]. Extent [814]. Extracting [2132, 1904, 2123]. Extraction [1264, 2131, 906, 468, 1596, 669, 1262, 1990, 496, 632, 633, 2534, 2515]. Extremal [2552].

f [1963]. f1 [1963]. f1/f [1963]. F2n [841]. Face [2416, 632, 520, 633]. Faces [457, 633]. Facets [145, 83]. Facial [522, 632]. Facile [1816]. Facilities [1396]. FACT [1720]. Factoring [2279, 843]. Factorization [967, 1475]. Factorizations [1570]. Factory [97]. Fading [1337, 1339]. FAENSY [1524]. Fail [1]. Fail-Stop [1]. Failure [1966, 1981, 619, 582]. Failures [2602]. Fair [1090, 2553, 2271, 93]. Fairness [957, 576]. Fall [596]. FALLDATEN [1511]. Families [143, 1860]. Family [2716, 73]. FAN [2683]. Fast [1264, 824, 368, 1275, 1305, 1767, 662, 1472, 2252, 1748, 2657, 2720, 1740, 2521, 634, 1982, 1244, 2585, 1282, 369, 2760, 1267, 429, 582, 880]. Faster [1180, 1490, 1822, 2432]. FasTraC [170]. Fault [1948, 1701, 1939, 1938, 1942, 1941, 1426, 2490, 1961, 1950, 922, 1954, 584]. Fault-Tolerant [1948, 1961, 922]. Faults [1439]. Faulty [1440, 1975, 1961, 561, 157].

FC [693]. FC2TOOLS [50, 107]. FCCMs [1717]. FDTs [99]. Feasibility [2747, 1639]. Feasible [2494, 1348]. Feature [446, 2618, 669, 422, 1262, 2380, 1775, 471, 342, 496, 498, 444, 2393, 2193, 2194]. Features [2085, 816, 633]. Federated [1916, 1843, 2215, 1381]. Federations [2111]. Feedback [2396, 817, 1950, 1924]. Feedforward [419, 1690]. Feistel [2254]. Fetching [1062]. Few [1256]. FFT [1182, 896]. FH-CDMA [1339]. Fiber [1487]. Fibers [370]. Fibre [1743]. Field [2279, 829, 831, 459, 1261, 1931, 509, 840, 313, 1631, 845, 1340, 476, 1765, 325, 856, 723, 2521, 1728]. Fields [822, 823, 2248, 367, 828, 1199, 834, 468, 835, 837, 346, 840, 848, 851, 854, 521, 1342, 842]. File [921, 1420, 1141, 342, 2341]. Files [1150, 2793]. Fill [2556, 268]. Fill-in [2556]. Filling [2679, 2529]. Filter [2726, 1739, 2277]. Filtering [1755, 583, 1663]. Filters [1330, 1343, 532]. Financial [2220]. Finding [758, 1710, 1806, 2547, 628, 962, 742, 2157, 2225, 2562]. Fine [115, 2339, 1255, 1066, 440]. Fine-Grained [115, 1066]. Fingerprint [812]. Finishing [2717]. Finite [822, 2565, 489, 223, 2062, 1305, 728, 140, 2522, 2716, 346, 2320, 258, 840, 2334, 1340, 1828, 715, 1667, 1954, 1548, 17, 2601, 2187]. Finite-Field [1340]. Finitely [1544]. FIR [1755]. Fired [437]. First [2691, 1079, 2136, 1594, 1570, 2501, 1135, 2146, 135, 723, 2195]. First-Class [2501]. First-Order [135, 2195]. Fish [2386]. Fisher [2201]. Fitness [1654, 2188]. Fitting [1267, 1863, 2514]. Five [1096, 751]. Fixed [886, 2038, 805, 562, 186]. Fixed-Order [2038, 186]. Fizzy [1995]. Fizzy-Fuzzy [1995]. Flage [278]. Flat [1337]. Flat-Fading [1337]. Flattening [2609]. Flattest [1899]. Flexibility [1141]. Flexible [2108, 327, 319, 1938, 673, 1896, 1310, 2173, 1049, 1736, 1302, 570]. FLIP [319]. Floating [2307, 991]. Floating-Point [2307, 991]. Flotter [162]. Flow [894, 523, 1438, 2078, 781, 2329, 1825, 1953, 1704, 521, 36]. Flows [2080, 2754, 1003, 2054]. Flowsheet [2389]. Fluid [1275, 603, 1288]. Fluorescence [498]. fly [91]. Flying [631]. FOCUS [2283]. Fold [1589]. Force [1895]. Fork [2654, 545]. fork-algebraic [545]. Form [1610, 122, 593, 1735, 1061, 191, 839]. Formae [1629]. Formal [2281, 13, 4, 873, 2029, 1452, 33, 2493, 2328, 2337, 1324, 2133, 1186, 2112, 2, 1448, 2230, 2293, 1374, 1426, 931, 2625, 2615, 39, 2306, 6, 2327, 22, 1230, 1365, 1106, 2151, 7, 1651, 278, 1851, 2321, 206, 76]. Formalism [27, 2409, 1582, 1897, 2505]. Formalisms [2319]. Formalization [1444, 2144, 54]. Formalizations [1086]. Formally [1852]. Format [1217]. Formation [984, 624, 471]. Forms [793]. Formula [299]. Formulae [2716]. Formulas [1196, 1456, 2186, 214, 2183]. Forsake [693]. Fortran [861, 2735, 1319, 680]. Forward [758, 650, 540]. found [171]. Foundations [2803, 9, 292, 911]. Founded [2781]. Four [751]. Fourier [975]. Fourier-Motzkin [975]. FPGA [1754, 1741, 1750, 1738, 1724, 1719, 1720, 467, 1760, 1731, 1763, 1749, 1756]. FPGA-Based [1738]. FPGAs [965, 1732, 1726, 1755, 1748, 1739, 1718, 1735, 1733, 1758, 1727, 528, 1761, 1762, 1721, 1730, 1745, 2637]. FPLDs [1746]. Fractal [2466, 2519]. Fractals [688]. Fractions [825]. Fragment [961]. Fragmentation [1162, 1365]. Fragments [1918, 2146]. Framework [2769, 13, 1130, 930, 579, 1115, 1169, 2429, 2474, 1974, 268, 705, 2148, 923, 1719, 2115, 2382, 1, 215, 2662, 1120, 1602, 2611, 2390, 224, 1532, 592, 206]. Free [1984, 2458, 10, 1556, 360, 2559, 529, 1899, 2257, 771, 788, 2762, 562, 1808, 1866]. freely [160]. Freeness [1818]. Frequency [1693, 324]. Friends [1938]. Frontal [2710]. Frontier [2435]. Frontiers [2525]. FS0 [1190]. Full [732, 1314]. Fully [2355, 90]. Function [1653, 1767, 1684, 512, 2555, 1290, 453, 1620, 1825, 848, 797, 1785, 1637, 851, 1109, 854, 1328, 570]. Functional [690, 1575, 1814, 59, 727, 949, 2405, 1092, 41, 1300, 946, 2078, 2667, 1597, 383, 2609, 1323, 1206, 1175, 1236, 1602, 57, 2175]. Functionalities [864]. Functionality [1065, 1507]. Functions [344, 1817, 2266, 2166, 888, 2495, 985, 1569, 357, 346, 1656, 2320, 11, 2253, 2265, 1773, 1784, 1108, 1691, 2442, 2441, 2184, 2423]. Functorial [1580]. Fundamental [2098, 217]. Fundamentals [2171]. Furnace [438]. further [1807]. fusion [564]. Future [2489, 1892, 1540, 2031, 1745, 151]. Fuzzy [912, 1995, 1685, 1993, 1648, 2171, 2180, 1989, 1994, 1987, 821, 1386, 1988, 1990, 230, 485, 222, 746, 651, 1991, 2511]. Fuzzy-Rule [2511]. Fuzzy/Evolutionary [1987].

G [364]. G-DES [364]. GA [2011, 1990]. GA-Based [2011]. Gabor [532]. Gabow [962]. Gafni [1986]. Gaining [150]. Galileo [585]. Galois [346]. Game [2487, 486]. games [70, 67]. Gamma [1981]. Gamma-Accurate [1981]. Gang [2236, 2238, 2237]. Gang-Scheduling [2238]. GANNFL [485]. gap [545, 541]. Garbage [1973, 1972, 112]. GARF [273]. Gas [430, 1772, 1777, 1669]. Gate [1728, 325, 723]. Gaussian [1083, 1641]. GCD [836]. GCWA [1131]. GDK [2293]. Gene [1623]. General [579, 1169, 1126, 831, 2069, 1822, 885, 1011, 372, 483, 1, 1622, 679, 2020, 856]. Generalisation [408]. Generalised [1599]. Generalization [1644, 373, 1687, 1344, 2265, 395, 155]. Generalized [789, 1812, 685, 486, 703, 705, 1603, 1673, 1039, 1493, 2254, 745, 2760, 1792].

Generalized-Reduction [1603]. Generate [1597, 639]. Generated [1504, 612, 1469, 160]. Generates [1989]. Generating [1694, 388, 2246, 184]. Generation [641, 1701, 1147, 1692, 358, 1472, 2072, 382, 2315, 2079, 1628, 1517, 340, 1152, 1774, 2004, 2159, 2111, 1421, 2332, 1161, 229, 1954, 2085, 89, 180, 91, 2523, 193, 1071]. Generational [1789]. Generative [1175]. Generator [785, 1907, 1721, 1749, 2432]. Generators [2277, 1708, 2659]. Generic [1804, 1994, 601, 2768, 2154, 2229, 1416, 2639]. Genes [1632, 1633]. Genetic [1700, 1644, 1710, 480, 2156, 1670, 2697, 2701, 1767, 1684, 662, 1661, 1685, 798, 602, 1711, 2709, 1655, 2010, 1769, 1662, 1702, 1648, 1694, 1656, 1627, 1770, 1617, 1628, 1997, 1629, 1771, 1714, 1664, 1513, 1697, 1772, 2642, 1630, 1773, 1665, 1705, 1988, 1706, 1774, 1775, 2009, 1622, 1637, 485, 1699, 1667, 1786, 1658, 1690, 1668, 1659, 1788, 1663, 1789, 1652, 2188, 1641, 2085, 1709, 2511, 2193]. Genetic-Based [1705]. Genetics [1624]. GENIAL [236]. Geniculocortical [472]. Genome [1786]. Genome-Dependence [1786]. Genomic [2791]. GENREG [1367]. Genus [849]. Geo [1353]. Geo-Data [1353]. Geometric [2552, 1900, 1891, 2553, 1906, 1896, 1886, 1887, 1904, 1892, 1175, 1326, 1888, 1897, 1802]. Geometrical [1550, 2515]. Geometrically [523]. Geometry [1895, 1012, 794, 1890, 1197, 1894, 1889, 2052, 639, 855, 1547, 64, 1793, 174, 2539, 2513, 2567, 166]. Geother [166]. Gesture [456]. Getting [482]. GF [2252, 1327]. Ghost [19]. GI [1241]. Gigaflops [2216]. Gist [2161]. GITE [340]. Given [758, 2574]. Glider [1219]. Global [1780, 997, 2398, 389, 1069, 2226, 739, 1418, 1657, 848, 1070, 2261]. globalized [1723]. Glueware [2357]. GM [2292]. GMRES [2503]. GOAD [1380]. Goal [731, 1924]. Goal-Oriented [1924]. Goals [59]. Gofer [946]. Good [729, 254, 571]. Goodness [1006]. Goppa [1326]. Gossip [902, 899]. Gossiping [774, 904, 903]. GOST [364]. GOTA [1582]. GP [1688]. GRACE [2092]. Graded [26]. Gradient [1026, 1248, 451]. Grain [877]. Grained [115, 2339, 1066]. GRAL [2109]. gram [1872]. Grammar [211, 1480, 1865, 1877]. Grammars [787, 1586, 788, 2592, 1863, 1880, 1869, 1867]. Grammatical [785, 1882, 1512, 799, 1861, 1858, 1866]. Grams [1481]. Grand [1682]. Granular [510]. Granularity [959]. Graph [756, 884, 798, 1973, 637, 2109, 58, 649, 2525, 646, 2720, 642, 794, 2731, 131, 1362, 2653, 2557, 650, 2558, 651, 2510]. Graph-Oriented [1362]. Grapheme [442]. Grapheme-to-Phoneme [442]. Graphic [43, 144]. Graphical [2108, 239, 1444, 2667, 2113, 98, 108]. Graphics [815, 382]. Graphs [1164, 775, 641, 774, 1474, 2556, 754, 750, 888, 893, 648, 1367, 2078, 2553, 2152, 644, 2602, 885, 753, 2559, 757, 1544, 752, 457, 645, 2742, 1464, 647, 1414, 2557, 780, 532, 751, 1584, 1347, 2125, 755, 2321, 566, 759]. GRASP [2748]. Gravitational [529]. Gray [522]. Gray-Level [522]. Greatest [2131, 850]. Greedy [1468, 767]. Grey [1253, 2530]. Grid [1074, 1150, 974, 665]. Grids [775, 662, 2723, 2080, 2055, 2738, 657, 2527]. Gröbner [118]. Ground [196, 1228]. Grounded [339]. Group [1291, 1424, 2273, 846, 929, 2763, 196, 2439, 188]. group-oriented [2439]. Grouping [982]. Groups [2062, 826, 1544, 117, 2440]. Growable [1749]. Growing [443, 669, 499, 529, 1262]. Growth [1677, 2073, 1944]. GSAT [1508]. GTM [432]. Guard [754]. Guess [740]. GUI [1602, 57]. Guidance [1277]. Guide [1498, 2591, 797]. Guided [2367, 1313, 1280, 1310, 1314, 1302].

H [2178]. H-Computation [2178]. h0rm [1985]. Haggis [1202]. Hamiltonian [758, 779, 2572]. Hand [802, 820, 593, 801]. Hand-Printed [802, 820]. Hand-Written [593, 801]. Handling [689, 1701, 1149, 1693, 586]. Handwriting [821]. Handwritten [441]. Hangul [616]. Hard [692, 693, 1438, 1627, 1508, 1050, 353, 354, 1671, 571]. Hardness [355]. Hardware [1087, 1751, 455, 2312, 1093, 2317, 1723, 2171, 1703, 1689, 2059, 1952, 1620, 1764, 227, 2319, 1722, 1111]. Hardware/Software [1723, 1703, 1764, 227]. Hash [344, 346, 2253, 2442, 2441]. Hashing [345, 368, 360, 369, 1470, 1947]. Haskote [1203]. Hausdorff [2445]. HCLP [706]. HDL [2337]. HDL-Based [2337]. Head [439, 1283, 1313, 2508]. Head-Mounted [1313]. Healthcare [2215]. Heap [1207]. Heat [1713, 979]. Hebbian [472, 528, 517]. Height [833]. Heine [2133]. Heine-Borel [2133]. Helical [1268]. Hellman [355, 352, 366]. Her [2470]. Hereditary [755]. Heterogeneity [870, 1361]. Heterogeneous [937, 1560, 33, 2693, 1849, 870, 1052, 2357, 6, 1040, 2480, 227, 1046, 878, 1403]. Heuristic [1498, 2096, 702, 1648, 1673, 1040, 722, 253, 199, 1859]. Heuristic/Stochastic [2096]. Heuristics [1155, 693, 1686, 2711, 701, 764, 2662, 2606]. HI [2623]. HiCoV [2344]. Hidden [2212, 347, 2260, 350]. Hiding [2456, 2467, 2448, 2471, 2461, 1030]. Hierarchial [1290]. Hierarchical [2156, 2613, 1183, 818, 2706, 490, 593, 505, 433, 1727, 959, 479, 2311, 440, 795, 2440]. Hierarchically [2377]. Hierarchies [247, 727, 385, 705, 252, 1582, 1666, 1841, 251]. Hierarchy [1561, 624]. High [987, 984, 2403, 2076, 2700, 1172, 927, 602, 465, 1095, 832, 1022, 857, 1321, 486, 2059, 948, 499, 620, 1104, 905, 2484, 1922, 1940, 471, 2311, 1242, 981, 1110, 2543, 324, 244, 410, 861, 2735, 1319, 680]. High-Dimensional [471]. High-Level [2403, 602, 1104, 1922]. High-Low-Game [486]. High-Performance [984, 927, 857, 1321, 244]. High-Rank [832]. High-School [620]. High-Speed [2076, 465, 2311]. Higher [1814, 48, 140, 1097, 1592, 124, 122, 2265, 1105, 470, 1109, 143, 123]. Higher-Order [140, 124, 122, 143, 123]. Highly [993, 1964, 926, 2642, 2756]. Highly-Parallel [993]. Highways [2280]. Hilbert [828]. Hill [735]. Hill-Climbing [735]. Hills [1671]. Hippocampal [519, 506]. Histologic [1257]. Historical [1367]. History [2446, 403, 2463, 605]. History-Oriented [605]. Hoare [557]. HOL [1086, 1089, 1095, 2324, 1098, 1099, 1102, 9, 1103, 213, 1111, 1112]. Hole [903]. Holes [2574, 2516]. Homeostatic [2669]. Homogeneous [1549, 1166]. Homomorphisms [1019, 564]. Homophonic [2585]. honeycombs [574]. hope [176]. Hopfield [480, 523]. Horizontal [1162, 1420, 1365, 2537]. Horn [1553]. HORNSAT [70]. Hot [1488, 2717, 891]. Hot-Potato [891]. Hough [983, 632, 1807]. Householder [1475]. HPCN [2749]. HPF [918, 933, 2744, 1316]. HPP [1022]. Huge [747]. Hull [2544]. Hulls [1478, 771]. Human [2477, 2262, 1242, 1282, 1923, 461, 633, 1249]. Human-Computer [2477]. Human-Oriented [1923]. Humans [1299]. HW [2294]. Hybrid [1966, 785, 2093, 765, 1339, 454, 37, 1987, 2616, 2151, 740, 1829, 720, 1267, 1805, 2193]. Hyderabad [2803]. Hyper [1113]. Hyperbolic [1017]. Hypercube [1165, 1061]. Hypercubes [888, 889, 882, 681, 885]. Hypergraph [2530]. Hypergraphs [2552]. Hypermedia [2769, 308, 336, 337, 311, 2379, 314]. Hypersphere [917]. Hypertext [1369, 2173, 1421, 315]. Hypertexts [310]. Hypometabolism [1297]. Hypothetical [2623, 2196]. HyTech [2294].

I-WAY [857]. I/O [930, 927, 1142, 2413, 1143, 1239]. IBM [2232, 2687, 2711, 2712, 2715, 1143]. ICAI [620]. ICE [2338]. IDA0* [686]. Idea [1121, 2255, 364]. Ideal [824, 1059, 846, 1059]. Ideas [1518]. identifiable [1860]. Identification [1993, 1884, 812, 2262, 1988, 424, 1785, 2260, 356, 2208]. Identifying [1297, 1068]. Identities [2270]. IDL [2149]. IDRA [1059]. IEEE [2310]. II [2537, 708, 2264, 1228]. II. [1633]. illumination [1798]. Illustrating [1243, 245]. Image [2688, 1349, 1272, 994, 995, 2466, 980, 985, 2713, 1343, 2372, 1157, 1280, 1750, 2180, 993, 1246, 2070, 1284, 1341, 809, 457, 1523, 531, 1166, 1775, 797, 1278, 1294, 1731, 1260, 981, 986, 2084, 1302, 520, 1342, 2509, 1251, 2528, 2525, 1801, 1791, 984]. Image-Based [2509]. Image/Video [980]. Imagery [981]. Images [1274, 1266, 2462, 1275, 1270, 1273, 1297, 1261, 1254, 791, 1268, 1265, 522, 807, 1255, 804, 1298, 2461, 2517, 2530, 1806]. Imaging [1300, 1293]. IMC [2622]. Imitation [1650]. Immune [2266]. Immunity [2580]. Immunoid [1680]. Immunological [1680]. Impact [2097, 59, 1895, 1853]. Impacts [2176]. Imperative [397]. Imperfect [1961, 2174]. Implant [1300]. Implement [2790, 2060]. Implementable [362, 363]. Implementation [1557, 1585, 2692, 1089, 2697, 234, 2252, 985, 998, 659, 2218, 2109, 1095, 1119, 1596, 858, 2721, 2101, 2212, 311, 1374, 14, 2171, 2238, 2788, 2070, 2626, 2314, 1595, 1051, 2740, 1705, 1009, 2151, 657, 940, 467, 964, 863, 925, 1731, 2105, 243, 274, 1765, 1728, 305, 1379, 573, 2067]. Implementations [2313, 2730, 352, 2169, 2758]. Implemented [2078, 1942, 528]. Implementing [2076, 1169, 918, 2795, 1163, 1938, 273, 1887, 2077, 655, 1320, 1224, 1750, 1739, 258, 1205, 972, 1190, 1920, 936, 1760, 1591]. Implicants [2555]. Implication [2592, 587, 2199]. Implications [2623, 1931, 2677]. Implicit [2129, 1654, 2293, 1629, 979, 587]. Importance [2383]. Importing [1102]. Improve [2051, 240, 156, 1872]. Improved [500, 1333, 1649, 2732, 361, 601, 2184, 1982, 2594, 179]. Improvement [233, 425, 1923, 1924, 2586, 2209]. Improvements [488, 1251]. Improving [2116, 1913, 762, 2621, 1594, 1687, 1344, 1507, 1104, 362, 491, 363]. IMPS [182]. Including [794]. Inclusion [1023]. Inclusive-OR [1735]. Incomplete [2141, 548]. Incompressible [2054]. Incorporating [2196, 630, 2724, 1011, 1420, 413]. increased [2417]. Incremental [2544, 1973, 1876, 1687, 1835, 2411, 1827, 54, 1863, 1877]. Independence [694, 1559]. Independent [766, 490, 1583, 2235, 656, 428, 869, 1884]. Index [376, 2795, 1481, 1411, 852]. Indexed [2793]. Indexing [796, 1793, 201, 212]. India [2803]. Indices [2368]. Indicia [2635]. Indirect [2271]. Individualizing [1283]. Individually [2155]. Individually-Motivated [2155]. Indoor [1903]. Induced [473, 897, 913]. Inducing [1869]. Induct [802]. Induction [149, 579, 608, 11, 2197, 804, 519, 1868, 80, 200]. Inductive [1701, 580, 2312, 2135, 1859, 1515, 1878, 2199, 2204, 2196, 167, 2200, 2138, 2209, 155]. Inductively [1211]. Industrial [4, 2691, 2068, 2, 279, 1714, 2053, 5, 1935, 1936, 508, 3]. Industry [2714, 2749, 503]. Inequality [373]. Inexact [798]. Infeasible [1460]. inferable [2205]. Inference [1413, 415, 1088, 785, 1124, 625, 704, 386, 615, 1988, 1538, 799, 1868, 1861, 1876, 1882, 1871, 2209, 1859, 2187, 1877, 2204, 1866]. Inferencing [1995]. Inferring [1880, 1823]. Infinitary [1127]. infinite [62]. Influence [1637]. Informal [2779]. Informatics [306]. Information [1412, 2245, 2280, 468, 2456, 411, 1280, 1512, 1904, 2039, 1284, 1854, 2263, 2676, 784, 2416, 1414, 2173, 2471, 1387, 583, 533, 1540, 2174, 1260, 482, 672, 2461, 2119, 1342, 1384, 1381, 1328, 1881, 554, 1883, 1867]. Information-Clusters [1381]. Information-Theoretic [2263]. Informational [1137]. InforMed [323]. Infrastructure [2347, 2774, 1856]. Infrastucture [2636]. Inheritance [462, 2224, 2124, 1582, 2111, 1347, 2125]. Inherited [391]. Inhomogeneity [1259]. Initial [2089, 1240]. Initialisation [2012]. Initialise [1787]. Initialized [1263]. Injection [1942, 1941]. INKA [180]. Inner [2762]. Innermost [121, 30]. Innovative [307]. Inoculation [1787]. Input [473, 2696, 881, 499, 395]. Input-Output [2696]. Insertion [1311]. Insertions [1336]. Inside-Outside [788]. Inspirations [2169]. Inspired [507]. Instance [700, 1662]. Instance-Oriented [1662]. Instances [692, 1504]. Instantiation [1558, 208]. Instantiations [2320]. Institutions [1216, 1232, 1235]. Instruction [320, 333, 323, 1746, 1070, 1953, 1062]. Instruction-level [1746]. Instruction/Data [1953]. Instructional [279, 298]. Instructions [2308]. Instructor [316]. Instrumentation [307, 867]. Integer [2086, 839, 1181, 768, 747, 117, 2630]. Integral [837, 838, 848, 977]. Integrated [40, 930, 2164, 2306, 1674, 1312, 1377, 304, 1384, 871, 98, 2799]. Integrating [2128, 41, 1188, 1361, 1607]. Integration [2089, 2042, 2337, 2112, 406, 1382, 1512, 2646, 593, 241, 1922, 2111, 2753, 2110, 99]. Integrity [2787, 1368, 2790, 1133, 2026, 1947]. Intel [1165, 1143, 2234]. Intelligence [1274, 594, 294, 1547]. Intelligent [327, 1192, 2477, 585, 287, 1361, 628, 2165, 340, 341, 2170, 300, 293, 314, 2175, 113, 2801]. Intensional [2139]. Intensionality [2600]. intensity [1791]. Intensive [2375, 922, 224]. Intent [1784]. Intention [1502]. Inter [40, 1217, 2528, 2341, 1281, 2221]. Inter-file [2341]. Inter-Operability [40, 1217]. Inter-ORB [2221]. Inter-Patients [1281]. Inter-pixel [2528]. InterACT [42, 1227]. Interaction [1599, 1783, 531, 1003, 113]. Interactions [469, 337, 509, 422]. Interactive [374, 2177, 696, 862, 297, 332, 382, 42, 2672, 1227, 1104, 2771, 1243, 583, 1636, 1548, 325, 326, 584, 1877]. Interchange [1217]. Interconnect [2076, 2075]. Interconnection [1144, 2497, 525, 883, 960]. Interconnections [1029]. Interdependent [919]. Interface [1192, 1202, 872, 336, 2477, 7, 679, 43, 144]. Interfaces [2160, 2667, 1369]. Interfacing [2735, 1194]. Interfailure [1944]. Interference [1057]. Interleave [2631]. Interleavers [1332]. interleaving [1816]. Interlingua [2645]. Interlocking [1943]. Intermediate [382, 390]. InterMUD [2622]. Internal [2137, 158]. International [1790]. INTERNET [334, 2633, 306, 1838, 2251, 303, 283]. Internet-Based [306]. Interoperability [1911, 2223]. Interoperation [2110, 1403]. Interplanetary [2725]. Interplay [506]. Interpretation [1274, 1813, 59, 2230, 1834, 510, 2665, 804, 2036, 2792, 1818]. Interpreted [936]. Interpreter [1502, 387]. Interprocedural [377]. Interprocess [2492]. Intersection [2496]. Intersections [1476]. Interval [1542, 754, 1976, 2547, 753, 2579, 2797, 83, 108]. Intra [1151]. Intra-operation [1151]. Intranasal [1310]. Intravascular [1266]. Intrinsic [526].

Introducing [1870, 2352, 945, 7]. Introducting [1324]. Introduction [2779, 2324, 1317, 1316, 174]. Intron [1624]. intrusion [2438, 2437]. intuitionistic [153, 152, 154]. Intuitive [320]. Invalidation [914]. Invariances [413]. Invariant [1343, 21, 2741, 1325]. Invariants [2617, 2332, 805, 89, 78]. Invention [2044]. Inverse [1711, 659, 2067]. Inverses [1432]. Inversion [976, 2135, 2246]. Inverting [2317]. Investigate [317, 599]. Investigating [1781, 1711, 1066]. Investigation [1647, 2419, 1660]. Investigations [1625, 2005]. Investment [431]. Invisibility [2465]. invisible [90]. invited [2195, 2204, 2182, 2188]. iO2 [2125]. IOGA [1662]. iPSC [1165]. iPSC/860 [1165]. IPSEC [1846]. IPSEN [2799]. IQ [2039]. Irises [632]. Irreducible [1333]. Irregular [2695, 671, 658, 1322, 932, 686, 676, 964, 680, 682, 666]. Irregularly [660]. IS-A [2101]. ISABEL [2770]. Isabelle [1103, 1087, 1190, 213, 168]. Isabelle/HOL [1103, 213]. ISAC [2364]. Isatis [272]. Island [2321]. Iso [2534, 2533, 1920]. Iso-surface [2534]. iso-surfaces [2533]. Isogenies [827, 841]. Isolation [1655]. Isomorphism [1490, 644, 794]. Isomorphisms [1600]. Isosurfaces [1253]. Issues [1916, 317, 933, 2772, 1741, 1095, 310, 1886, 1201, 1619, 807, 1009, 225, 959]. Iterated [897, 1131, 985]. Iteration [16, 558]. Iterations [2724]. Iterative [1157, 2162, 1523, 1034, 2485]. IVTT [1411].

Jacobi [844]. Jacobian [491]. Japanese [801]. Java [1194]. Job [699, 1769, 2231, 724]. Job-Shop [699, 1769, 724]. Jobs [1460, 1457]. Jobshop [762]. Join [1370, 640, 2782, 1148]. Joint [2411, 2434]. JTMS [1501]. Justification [2366, 1513].

Kahn [550]. Kanji [819]. Kennedy [906].

Kernel [2766, 1906, 1448, 274, 2776]. KESO [2482]. Key [2469, 374, 2429, 1741, 2250, 364, 365, 2264, 1517, 1856, 2263, 371, 1857, 2261, 2442, 2424, 2444, 2418]. Key-Schedule [364]. Keying [344, 770]. Keys [355, 575]. kindest [2189]. Kinds [1529]. Kinematic [505]. Kingdom [450]. Kleene [15]. Know [1116]. Knowing [359]. Knowledge [641, 2395, 329, 1407, 588, 332, 2498, 1507, 2375, 2646, 593, 2099, 2104, 1520, 2003, 1522, 1988, 1523, 225, 1524, 804, 1540, 342, 2174, 2506, 1416, 224, 2196, 156, 197, 1872]. Knowledge-Based [1407, 1520, 2003, 225, 1540, 2174]. Kohonen [473, 440]. Kolmogorov [543]. KROSS [1395]. Kummer [847]. Kuperee [2417].

L [2304, 1289]. L-Systems [1289]. l1 [759]. l1-Graphs [759]. Lab [1762]. Labeled [532]. Labeling [690, 1249]. Labelled [1117, 1537]. labelling [189]. Laboratory [307, 304, 1289]. LadMan [2684]. Laguerre [972]. Lake [328]. Lambda [2597, 140, 1215, 2660, 393, 548, 1574, 386, 128, 2567, 1581]. Lambda-Calculi [393, 1574, 128]. Lambda-Calculus [140, 386]. Lambda-Definable [2597]. lambda-geometry [2567]. lambda-Prolog [1581]. Lamda [2599, 2611]. Lanczos [969, 2066, 2756]. Landmark [1291]. Landscapes [1661, 1646]. Language [1814, 2042, 263, 2160, 59, 2106, 942, 1566, 1186, 2683, 1598, 646, 2045, 1200, 2351, 1351, 1214, 941, 935, 261, 2626, 952, 1608, 2030, 749, 5, 945, 950, 1922, 1173, 2115, 295, 2173, 640, 47, 2159, 7, 1529, 2634, 1161, 278, 148, 1396, 2637, 1868, 2800, 565, 92, 2611, 1864, 1872]. Language-Based [47]. Languages [2784, 1088, 2551, 1322, 731, 1462, 1321, 2647, 383, 1559, 939, 606, 2609, 2501, 1577, 940, 936, 783, 17, 721, 1239, 1240, 1860, 1875, 560, 562, 568, 1874, 2204, 1866]. LAPACK [2752]. Larch [1214, 147, 43, 144]. Large [1048, 1024, 2694, 862, 1052, 658, 2682, 38, 2728, 2067, 1649, 448, 1047, 2734, 2737, 2746, 929, 2386, 2684, 2754, 1704, 2760, 970, 1806, 214]. Large-Scale [862, 1047, 2386]. Laser [1265]. Latency [1743, 991, 896, 2075, 1030, 880]. Later [2632]. Lattice [1105, 1841, 653, 2523]. Lattice-Based [1841]. Lattices [834, 855, 210]. Laws [2636, 1014, 1919, 956]. Layer [1687, 2372, 510, 418, 1846, 1996]. layered [2801]. Layering [2608]. Layers [2775]. Layout [1699]. layouts [561]. Laziness [1826]. Lazy [824, 1045, 553]. Lead [805]. Leaf [901]. Leakage [1328]. Lean [2130, 1171]. learn [1777]. Learnability [2183, 2207]. Learned [799]. Learner [582]. Learners [327]. Learning [1868, 802, 489, 328, 1779, 420, 2648, 2177, 317, 455, 2186, 330, 1767, 319, 527, 1685, 280, 425, 156, 1539, 1900, 321, 1686, 235, 512, 1875, 1799, 310, 333, 490, 2647, 334, 1687, 335, 284, 481, 285, 2373, 486, 607, 1783, 1688, 428, 606, 427, 1515, 441, 2202, 2028, 288, 484, 505, 1858, 289, 457, 483, 609, 801, 291, 412, 627, 1621, 303, 1800, 1716, 528, 2479, 304, 1874, 413, 281, 1867, 517, 421, 519, 1016, 440, 520, 282, 254, 431, 800, 2393, 506, 326, 2191, 2206, 2192, 1804, 2200, 2203, 2184, 2185, 1883, 2195, 2198, 1862]. Least [2713, 849, 2762, 2514]. lecture [2195, 2204, 2182, 2188]. Left [1272, 875]. Legal [1421]. Leibinz [336]. Lemma [200, 2331]. Lemmaizing [610]. Lempel [1481]. Length [1043, 902, 2566, 598, 2189, 1870]. Lens [2039]. Lenses [687, 1712]. Lesion [1295]. Lessons [273, 287, 2251]. Level [1962, 2130, 2403, 1063, 2307, 602, 995, 1095, 2765, 612, 993, 947, 948, 448, 522, 1757, 1104, 1773, 1620, 1922, 393, 240, 1953, 237, 1954, 2221, 2433, 1912, 2530, 534, 2657, 1746]. Levels [415, 1967]. Lexical [1863, 500, 599, 598]. Libraries [1163, 2712, 1719, 2751, 672]. Library [1168, 2701, 2088, 58, 2077, 2070, 1051, 1894, 1761, 57, 2797]. Life [2642, 1621]. Life-Time [1621]. Lifting [1176, 1555]. Ligands [1896]. Light [2324, 1838, 2505, 2612]. Light-Weight [1838, 2505, 2612]. lighting [1799]. Lightwave [884]. Like [1331, 1262, 378]. Likelihood [1337, 1881]. Likely [1545]. limit [2208]. Limited [902, 1157]. Limiting [2262]. Limits [2449, 1579, 391, 1232, 2187, 1066, 2038]. Line [1483, 1485, 1476, 773, 897, 818, 2546, 2545, 2271, 990, 753, 2658, 522, 899, 1899, 821, 988, 801, 424, 1676, 989, 616, 528, 806, 517, 814, 420, 1949, 1409, 2184, 2588, 1691, 2578]. Line-Communication [899]. Lineal [669]. Linear [2670, 2784, 992, 968, 823, 223, 2086, 1550, 2088, 1406, 1279, 2706, 971, 550, 985, 1391, 235, 2710, 1330, 1976, 1134, 1335, 1503, 2721, 1165, 2255, 2563, 2067, 2604, 709, 1334, 139, 2617, 2616, 372, 538, 424, 2747, 2750, 1135, 2664, 2580, 2752, 2592, 1459, 2485, 721, 1470, 1766, 1875, 1867, 187]. Linearly [464]. Lines [280, 2524]. Lingual [1412]. Linguistic [1989, 639]. Link [1413]. Linking [422]. Links [1971, 1421, 561]. Liquid [1772]. List [1084, 2587, 1033, 1019, 2575, 960, 2631, 564]. little [2611]. Live [1264]. Live-Wire [1264]. Lived [1982]. Liveness [1090, 1432]. LL [2296]. LM [2297]. LMT [2568]. LMT-skeleton [2568]. Load [1048, 2072, 2081, 2727, 1050, 2080, 313, 686, 1705, 860, 1046, 682]. LoadLeveler [2233]. Local [1542, 697, 1702, 1466, 702, 735, 705, 428, 2334, 424, 2745, 713, 1145, 1950, 1622, 715, 1638, 741, 654, 1652, 1709, 556]. Localised [1783]. Locality [2245, 1063, 1013, 1011, 1682]. Localization [824, 2141, 632]. Location [2455, 2454]. Locomotives [1945]. Logarithm [2603]. Logarithms [988, 989, 852, 856]. Logger [1374]. Logging [1959, 2594]. Logic [2670, 1002, 401, 1122, 1575, 1128, 1130, 1497, 1117, 546, 59, 2404, 1495, 2035, 2795, 1818, 2044, 1743, 2619, 2005, 618, 1815, 1222, 1127, 1561, 1609, 1059, 1137, 1594, 1119, 383, 1559, 10, 1572, 1097, 2623, 2407, 2006, 2171, 2140, 2409, 2032, 947, 1835, 1133, 1105, 2007, 1516, 737, 2034, 1735, 1587, 2665, 28, 1570, 1125, 2022, 619, 242, 647, 56, 395, 1135, 1612, 1827, 1950, 1108, 1538, 1759, 1601, 1120, 230, 257, 485, 1049, 1109, 2036, 1606, 2146, 2593, 719, 1765, 2167, 1747, 544, 205, 557, 170, 1878, 2199, 2197, 108, 92, 152]. logic [154]. Logic-Based [2035, 56, 2036]. Logical [329, 1566, 2030, 1414, 716, 1237, 1969, 2199]. Logics [626, 1441, 1118, 235, 2375, 1519, 2150, 1525, 1851, 172, 554, 209, 203, 211, 538, 204]. Logistic [245]. LOL [1608]. Lola [2637]. Long [2545, 2727, 1627, 2246, 1982, 776, 122]. Long-Chained [2727]. Long-Lived [1982]. Look [692]. Look-Back [692]. Looking [732]. Lookup [1462]. Loops [906, 920]. Loosely [952]. Loss [1349]. Lossless [2782]. LOTOS [1453, 2303]. Low [486, 2075, 324, 2764, 880, 244]. Low-Cost [2764]. Low-Energy [244]. Low-Latency [880]. Lower [1488, 1976, 1482, 898, 179]. LP [2295]. LPG [59]. LR [331]. LRC [331]. LSNs [1355]. LTP [519]. LU [967]. Lukaszewicz [1126]. Lung [1268, 1290]. Lustre [2288, 2283]. LUT [965, 1718]. LUT-Based [965, 1718]. LW [2298]. Lygon [29, 51].

M [1411]. M-IVTT [1411]. MAC [693]. Mach [274]. Machine [802, 2284, 2648, 2367, 1539, 1900, 612, 2626, 2642, 511, 969, 2082, 1762, 869, 1010, 148, 2640, 1683, 552, 171, 186, 2706]. machine-found [171]. Machine-Generated [612]. Machine-Independent [869]. Machines [223, 2494, 921, 1081, 1723, 1044, 2334, 412, 1429, 413, 1954, 1547, 259, 2209, 2802, 2187]. Macro [336, 1749]. Made [2636]. Magnetic [1293, 1247]. Magnetoencephalography [1299]. Magnitude [362, 363]. Mail [1847]. Main [2264, 1410]. Maintaining [1978, 2117]. Maintenance [1358, 258, 2004]. Majesty [2470]. Maker [1810]. Makes [1123]. Making [2648, 1198, 2641, 629, 1348, 1642]. Man [2636, 1250]. Man-Made [2636]. Management [2350, 2395, 276, 995, 2357, 2345, 2069, 2455, 256, 2719, 1921, 1382, 655, 2358, 1931, 2352, 1142, 2220, 1404, 2349, 1922, 1160, 2060, 323, 2684, 1046, 2214, 1389, 2261, 277, 1704, 1853, 1381, 2763, 1379]. Manager [43, 144]. Managers [2069]. Managing [2280, 1514, 2344, 1909, 2094, 2239, 2182]. Manifold [1631, 2526]. Manifolds [1549, 1798]. Manipulating [48]. Manipulation [1084, 2041, 1896]. Manufacture [1889]. Manufacturing [687, 1674, 811, 1890, 2749, 1540]. Many [10, 840, 421, 172]. Many-Take-All [421]. many-valued [172]. Map [789, 432, 435, 529, 474, 471, 476, 498, 444, 440, 506]. Map-Analysis [789]. Mapped [1298]. Mapping [919, 917, 965, 927, 2317, 888, 1075, 920, 463, 1718, 2742, 1288, 1756, 1807]. Maps [530, 446, 1272, 458, 472, 471]. marching [2529]. Marked [2136]. Marker [1277]. Marketplace [1837]. Markets [2226]. Markov [84, 1873, 1871, 1057, 2188, 1342]. MARRELLA [111]. MARS [1942]. MasPar [1181]. Massive [263]. Massively [1084, 944, 1496, 1780, 2696, 658, 466, 2053, 896, 1054, 970]. Master [1669, 1183, 2065, 2620]. Master-Slave [1183, 2065, 2620]. MAT [641]. Match [1151, 398, 1408, 2392, 429, 2427]. Matching [1477, 2472, 1901, 1270, 2556, 2550, 798, 138, 1473, 958, 634, 776, 966, 2570, 1276, 1731, 1281, 1480, 178, 1798, 573]. Matchings [2553]. Material [282]. Materialized [2628]. Mathematical [1196, 1094, 310, 1189, 2677, 314, 2144, 1547]. Mathematics [1102, 2678]. Mating [1622]. Matrices [660, 2711, 262, 2728, 970]. Matrix [879, 976, 660, 839, 2720, 192]. Matrix-Vector [660]. Matroids [768, 767]. Matter [1253]. Matter/White [1253]. Maturity [2350]. MAX [710, 2748]. MAX-CSP [710]. MAX-SAT [2748]. Maxillofacial [1274]. Maximal [640, 751, 254]. Maximum [447, 1337, 1881, 2556, 2554, 766, 931, 736, 1522, 1681]. Mazurkiewicz [538]. MBPC [501]. McCarthy [1121]. MDG [105, 1111]. MDL [1863]. Mealy [1636]. Mean [1631, 1432]. Mean-Value [1432]. Meaningless [1573]. Means [1300, 1713, 1664, 1852, 1665, 325, 2522, 1549, 1961]. Measurable [1925]. Measure [1270, 1635]. Measured [2240]. Measurement [1831, 1924]. Measurements [2633, 307, 1517, 1935]. Measuring [1349, 309, 1545, 470]. Mechanical [1993, 169, 1889]. Mechanically [73, 2331]. Mechanics [319, 322, 326]. Mechanisation [1112]. Mechanism [2107, 863, 2612, 582]. Mechanisms [1916, 673, 1942, 1918, 2087, 1062, 1240, 2427]. Mechanize [56]. Mechanizing [579]. Media [1409, 2768, 583]. Mediators [2661]. Medical [1275, 2372, 1245, 1278, 2200]. MeDLey [2707]. Meet [362, 363]. Meet-in-the-Middle [362, 363]. Meeting [1505]. Meiko [2096]. Membership [1538]. Membrane [516]. Memetic [2013, 2014]. Memoization [1591]. Memory [410, 1968, 2245, 1024, 2074, 2403, 928, 995, 2475, 2587, 882, 2369, 2712, 2721, 2244, 409, 1057, 604, 1023, 1410, 926, 2737, 938, 969, 914, 2594, 929, 2050, 964, 2752, 2751, 869, 907, 2431, 2401, 2402, 2405, 2400, 2399, 2407, 2408, 2411, 2413, 2414, 2415]. Mental [2170, 408]. Merging [1849, 1252, 2576, 2668, 1756]. Merit [797]. Mesh [887, 2575, 2571, 2736, 782, 1907]. Meshes [1948, 903, 2572, 2522]. Meshing [668]. Message [1968, 344, 2074, 1163, 2707, 872, 674, 2049, 941, 2740, 2247, 1959, 2243, 2050, 369, 598, 2189]. Message-Driven [941]. Message-Passing [1968, 872, 2049, 2740, 2050]. Messages [1043]. Meta [2371, 947, 737, 483]. Meta-Cases [2371]. Meta-Level [947]. Metacomputation [384, 400]. Metadata [2673, 2215]. METAFrame [109]. Metalevel [1190, 2681]. Metamodel [58]. Metaobject [2661]. Metastability [1759]. Metasystem [2656, 400]. metatheory [167]. Meteor [614]. Meteor-Burst [614]. Meter [4]. Method [690, 2689, 2703, 1685, 1654, 425, 1170, 1118, 1462, 760, 1165, 885, 974, 1246, 979, 441, 1673, 821, 1244, 969, 1042, 1377, 1326, 959, 2037, 973, 964, 2503, 852, 2341, 2588, 2755, 229, 605, 2054, 2125, 1991, 197, 1146]. Method-Lookup [1462]. Methodological [1348]. Methodology [1754, 1087, 2493, 785, 2317, 2412, 2314, 874, 1072, 226, 1872]. Methods [2281, 976, 4, 1643, 1412, 1077, 1542, 1291, 2337, 2019, 1780, 2088, 1150, 2705, 2096, 637, 664, 1426, 1394, 2306, 2646, 6, 2591, 781, 747, 1698, 7, 1281, 2758, 814, 1659, 2485, 722, 253, 2394, 1859, 214, 76]. Metric [761, 793, 768, 565]. Metrics [1489, 888, 2777]. MFSK [1339]. MGGI [1872]. Micro [1743, 1310, 1838, 1071]. Micro-cash [1838]. micro-code [1071]. Micro-Endoscopy [1310]. Micro-Second [1743]. Microanalysis [648]. Microelectronics [321]. Microkernel [1100]. Microprocessor [1949, 1068]. Microprocessors [2326, 2330, 1953, 1062]. Microscope [337]. Microscopic [1280]. Microscopy [1265, 1260]. Microstructures [1257]. Microsystems [1715]. Microvessels [1257]. MIDAS [1079]. Middle [362, 363, 2011]. Middleware [2214]. Migrating [932]. Migration [1753, 1763]. Millennium [2395]. Million [2368]. Millipede [866]. Millsi [2717]. MIMD [2071]. Min [1069]. Min-Cut [1069]. Mind [1669, 541]. Minding [2249]. Minimal [1488, 2703, 1767, 1226, 1593, 1334, 2567, 1231, 2142, 2203]. Minimizable [1431]. Minimization [1686, 731]. Minimizing [561, 2577]. Minimum [1478, 1043, 2556, 835, 2574, 751, 598, 2569, 2189, 1865, 2187]. Minimum-Area [1478]. Mining [2547, 2675, 2498, 2397, 1376, 2482]. Minor [2724, 484]. Minutiae [812]. Mis [2270]. Mis-representation [2270]. Mismatched [1383]. Missing [609]. Mix [2257]. Mix-Type [2257]. Mixed [1735, 2003]. MIXes [2455]. Mixing [1640, 2188]. Mixins [1568]. Mixture [530]. Mixtures [426]. Mizar [1098]. ML [1580, 1579, 1205, 398, 1210]. ML-Definability [1579]. MLS [1855]. MML [2201]. Mobile [1992, 1816, 275, 1844, 2455, 36, 1390, 1680, 2627, 1387, 2773, 1778, 1397, 2512, 1803, 2428]. Mobility [2768]. Modal [1443, 1117, 1118, 1119, 2412, 1785, 1120, 544, 203, 204]. Mode [465, 1744, 1098, 899, 2665, 1429]. Mode-Machines [1429]. Model [447, 2091, 1840, 1015, 1452, 1912, 480, 1441, 1315, 317, 765, 2307, 1908, 994, 2406, 275, 1843, 320, 1076, 2369, 512, 1367, 2330, 458, 1012, 1004, 36, 2448, 2726, 594, 2582, 338, 509, 932, 1057, 1011, 510, 2731, 1290, 952, 2034, 1362, 2161, 436, 1051, 938, 1386, 1009, 424, 1008, 1842, 2127, 516, 2745, 1414, 2570, 2750, 479, 2652, 1950, 2154, 476, 2323, 2229, 1129, 2145, 1017, 2163, 515, 1018, 996, 1239, 1401, 907, 506, 2374, 2336, 1397, 2338, 2512, 64, 2530, 83, 102, 534, 185, 63, 2444, 1802, 100, 539, 193, 70, 79, 67]. Model-Based [2091, 317, 436, 2163]. Model-Checking [1441, 539]. Modeling [2097, 1677, 2433, 1087, 1604, 785, 695, 2360, 2109, 2715, 1372, 2671, 1885, 1394, 2459, 615, 2104, 1943, 238, 1175, 743, 461]. Modelising [308]. Modelled [2170]. Modelling [1929, 2116, 1913, 335, 655, 1549, 1918, 2114, 1935, 1856, 1944, 314, 2629, 2215, 954, 2121, 232, 1078, 1289, 237, 1003, 324, 514, 2437, 1864, 1872]. Models [2128, 445, 803, 1495, 2694, 2177, 2620, 2107, 1993, 1306, 1925, 2017, 2033, 536, 1921, 502, 2624, 37, 2032, 221, 1308, 620, 1400, 505, 2118, 2113, 2142, 1774, 1243, 1897, 471, 2174, 799, 1946, 281, 1917, 1606, 1552, 1934, 1003, 1267, 315, 1058, 2589, 2643, 1873, 1870, 1800, 541, 184]. Modes [2158]. Modification [473]. Modified [1272, 497, 2732, 969, 498, 718]. Modular [2108, 269, 1101, 2729, 2007, 456, 1824, 2309, 72, 2581, 1238]. Modularity [1568, 131]. Modulation [1755, 2461, 517]. Modulator [1760]. Module [1600, 2403, 68, 683, 242]. Modules [13, 944, 1132, 117, 1210]. Modulo [146, 178]. Mold [1899]. Molecular [2064, 2071, 2733, 2181, 2759]. Molecules [1710, 2727]. Moments [1345]. Mona [1449]. Money [2261]. Monitoring [2048, 864, 2702, 865, 2057]. Monkey [513]. monocular [1792]. Monoidal [2132]. Monoids [2132, 133, 190]. Monomial [347]. Monomorphic [1106]. Monotone [2555, 1463, 2183]. Monotonous [1123]. Monte [976, 2689]. Morphic [785]. morphisms [562, 572]. Morphological [789, 1285, 1507]. Morphology [1290]. Morphometric [1286, 1287]. Moscow [543]. MOSFETs [1695]. Most [355, 1952]. Motion [1492, 459, 1738, 1994, 1069, 2478, 1731, 440]. Motivated [2155]. Motivating [295]. Motor [513, 476, 507]. Motzkin [975]. Mounted [1313]. Movement [1069]. Moving [1992, 1237]. MPEG [2777]. MPEG-2 [2777]. MPI [872, 2049]. MPI-2 [872]. MPP [2730, 2234]. MPSK [1338]. MR [1286, 1261, 1254, 1248, 1294, 1314].

MR-Based [1286]. MRA [986]. MRI [1271, 1259, 1310, 1258, 1269, 1267]. MSparc [1027]. MUCH [1153]. MUDs [2622]. Mulitmedia [2218]. Multi [1929, 1077, 1035, 1615, 1300, 1646, 2256, 2584, 1687, 2657, 674, 1201, 655, 1184, 1246, 2250, 1664, 1047, 290, 1773, 393, 627, 1996, 583, 1298, 1785, 1260, 498, 2755, 2581, 1954, 1269, 2761, 1148, 1709, 2433, 244]. Multi-Agent [1929, 1615, 1184, 290, 627]. Multi-Array [1298]. Multi-Channel [1300]. Multi-Computer [1047]. Multi-dimensional [1077]. Multi-Domain [655]. Multi-DSP [244]. Multi-Fluorescence [498]. Multi-Image [1260]. Multi-join [1148]. Multi-layer [1996]. Multi-Layer-Perceptrons [1687]. Multi-Level [1773, 393, 1954, 2657, 2433]. Multi-Media [583]. Multi-Message [674]. Multi-Modal [1785]. Multi-object [2761]. Multi-Objective [1664]. Multi-Parent [1646]. Multi-Party [2250]. Multi-Phase [1035]. Multi-Recastable [2256]. Multi-Spectral [1269]. Multi-Step [1709]. Multi-tasking [2755]. Multi-threaded [2584]. Multi-variable [2581]. Multi-Volume [1246]. Multiattribute [1081]. Multicast [2765, 2543, 2764, 2606]. Multicasting [674]. Multicomputer [2718, 1404]. Multicontexts [636]. Multidatabase [1076]. Multidimensional [996]. Multifaceted [1414]. Multiflows [768]. Multigrid [1026, 663, 2054]. Multilayer [785, 463, 493, 1016]. Multilevel [2620, 527, 978, 909, 1842, 1853]. Multilevel-Newton [978]. Multimedia [234, 307, 2766, 2211, 2460, 1360, 309, 333, 2775, 312, 2771, 2210, 2773, 2228, 2767, 2776, 2763]. Multimodal [1284]. Multimodality [1314]. Multiobjective [1672, 1987]. Multiobstacles [1992]. Multipacket [891]. Multiparadigm [1200, 225]. Multiparametric [2387]. Multiparty [2275]. Multipermutations [373].

Multiple [345, 1350, 379, 1391, 831, 665, 1617, 947, 1514, 2158, 290, 1377, 2004, 1785, 2588, 1659, 305, 582, 1295, 210]. Multiple-grid [665]. multiple-valued [210]. multipleted [2430]. Multiplication [660, 2311]. Multiplier [1740, 2143, 73]. Multipliers [2309]. Multipoidy [1781]. Multiprocessing [1067]. Multiprocessor [894, 1183, 1737, 777, 1039, 2241]. Multiprocessors [1962, 2245, 1024, 2074, 1031, 2587, 882, 869]. Multiprogrammed [2237]. Multirecombinant [1626]. Multiresolution [2517, 1267]. Multiscale [800]. Multiservice [2046]. Multisets [2501]. Multistage [2497]. Multitemporal [2495]. Multithreaded [2583, 1027, 1068]. Multithreading [1058]. MultiUser [2622]. Multiversion [1346]. Multiway [2321]. MUltlog [172]. Murcursive [95, 74]. Music [1203]. Mutation [1644, 1696, 1650, 1786, 1658]. Mutations [1647, 1649]. Mutual [1280, 1284, 1527, 1881]. Mutual-Information [1280]. Mutually [11, 169]. myocardial [2531]. Myoelectric [1344].

Naive [1696]. Nano [1051]. Nano-Threads [1051]. NAOS [1366]. Narratives [1527]. Narrowing [1611, 124, 61]. NAS [2699]. Natural [2160, 2140, 2173, 640, 2159, 2800, 92]. natural-language [92]. Navigation [2379, 2173]. Navigational [1393]. NCSU [96]. ND [612]. ND-Proofs [612]. Nearest [1504]. Nearest-Neighbor [1504]. Nearly [1488]. Necessary [892]. Need [1593, 2306]. Needed [1611]. Needham [1857]. Needham-Schroeder [1857]. Needle [1313]. Negation [1122, 1130, 2619, 1558]. Negative [1479, 990]. Neglected [2435]. Neighbor [1504]. Neighborhood [1638]. NeMo [2644]. Neocortical [477]. Nerve [406]. Nested [906, 2609, 1370, 2030]. Nesting [1482, 1253]. Net [443, 411, 943, 954, 442, 616]. NETL [263]. Nets [2403, 1599, 130, 2376, 1518, 599, 419, 1623, 2336, 503]. Netters [1623]. Network [447, 2083, 523, 2702, 1743, 450, 597, 765, 527, 1055, 1273, 2772, 465, 235, 236, 437, 458, 2775, 2448, 414, 430, 892, 499, 1036, 615, 505, 483, 781, 467, 526, 1145, 797, 2095, 883, 2046, 1972, 2169, 1046, 1707, 1061, 963, 922, 519, 2577, 1846, 491, 2438, 2437]. Network-Layer [1846]. Networked [1413]. Networks [438, 1485, 1048, 894, 691, 1963, 480, 492, 501, 884, 497, 870, 1767, 1684, 2766, 2621, 1052, 2211, 475, 1472, 904, 998, 900, 902, 2497, 525, 1713, 36, 416, 1694, 422, 1007, 2731, 427, 402, 463, 493, 1515, 2576, 495, 453, 2627, 2053, 449, 2377, 436, 1440, 1975, 2003, 614, 895, 1773, 782, 292, 596, 2254, 898, 2046, 925, 1946, 485, 960, 1690, 421, 2219, 2543, 518, 259, 504, 113, 1880, 550, 561, 1879, 1871]. Neural [447, 438, 492, 803, 501, 523, 497, 455, 1767, 1684, 450, 765, 527, 475, 1273, 425, 998, 235, 236, 437, 458, 500, 459, 416, 1694, 414, 430, 428, 509, 427, 499, 402, 1515, 495, 453, 456, 449, 436, 483, 614, 599, 467, 526, 2095, 2169, 476, 485, 1046, 1690, 518, 442, 491, 1880, 1879]. Neuro [518]. Neuro-Biological [518]. Neurocomputer [525, 464, 463, 1028]. NeuroFuzzy [2194]. neuroidal [2182]. Neuron [1998, 1994]. Neuronal [407, 429]. Neurons [470, 461]. Neurosurgery [1301]. Neurosurgical [1277]. Newsgroups [448]. Newspaper [1673]. Newton [2457, 978]. Next [1779, 2386, 180]. Niche [1781, 1654, 1646, 1657, 2018]. Niching [1654, 1659]. nilpotence [178]. NK [1646]. NK-Landscapes [1646]. NL [641]. NMR [330]. NN [417]. No [2260, 1924, 254]. No-Good [254]. No-Transferable [2260]. Node [1466, 2234]. Node-Deletion [1466]. Nodes [1961]. Noise [415, 2200, 2726, 1634]. Noisy [2264, 371, 2206]. Nomenclature [1361]. Non [1816, 1543, 1186, 1279, 886, 766, 235, 832, 468, 1974, 621, 128, 1228, 622, 529, 1334, 595, 1525, 2573, 717, 1049, 813, 2755, 2418, 229, 986, 2579, 1239, 2234, 2258, 2419, 957, 192, 160, 1808]. Non-Binary [717]. Non-blocking [886]. Non-Classical [1525, 192]. Non-Compatible [1543]. Non-constructive [766]. Non-contiguous [2755, 2234]. Non-crossing [2573]. Non-determinism [1239]. Non-Deterministic [1974, 1049, 621, 622]. Non-Euclidean [1808]. Non-executable [1186]. Non-Extensional [128]. non-freely [160]. Non-ground [1228]. Non-Growing [529]. Non-interleaving [1816]. Non-Linear [1279, 235]. Non-Minimal [1334]. Non-Orthogonal [229]. Non-Overlapping [468]. Non-repudiation [2418, 2258, 2419]. Non-Rigid [813]. Non-Stationary [986]. Non-Symbol [595]. Non-Trivial [832]. Non-uniform [2579]. Nonannotative [953]. Nonclassical [1516]. nonconservative [555]. Nondeterminism [2094]. Nondeterministic [2313, 126]. Nonlinear [501, 509, 2277, 424, 496, 2421]. Nonlinearity [2266]. Nonmonotonic [1122, 1118, 1127, 540]. Nonmonotonicity [1123]. Nonparametric [425]. Nonstationary [1789]. Normal [1495, 1286, 1561, 839, 122, 28, 191, 2521]. Normalisation [127]. Normality [600]. Normalization [2132, 1285, 1391, 1574, 1565, 30]. Normative [2115]. Normed [2499]. Notation [943, 2301]. note [1866]. Novel [1330, 409, 1338, 531, 520]. NP [1671, 560, 571]. NP-Hard [1671, 571]. NTN [1146]. NTN-Method [1146]. Nuclear [2396, 1696]. Nuclei [1265]. Number [1147, 697, 2279, 1199, 829, 831, 834, 990, 837, 747, 845, 1708, 856]. Numbering [2542]. Numbers [825, 764, 842, 2425]. Numeric [1656, 294]. Numerical [1625, 1192, 338, 974, 1990, 225, 2751, 2174]. Nuprl [1102, 9]. Nurse [2630]. NUT [2282].

O [930, 927, 1142, 2413, 1143, 1239]. OB [943]. OBDD [1686, 2334]. OBDD-Representations [2334]. OBDDs [260, 2555, 2185]. Oberon [2640]. Object [2789, 1358, 263, 2097, 1359, 272, 1956, 1353, 1162, 445, 2786, 2701, 2474, 667, 266, 228, 2166, 1355, 2040, 2608, 2370, 1462, 2663, 265, 1075, 1372, 1178, 2224, 454, 1374, 14, 270, 287, 221, 1394, 1395, 2476, 935, 261, 943, 1582, 2118, 1386, 1922, 1842, 2605, 2798, 238, 2299, 1348, 1790, 1236, 2610, 2629, 1346, 240, 2121, 2163, 264, 2644, 2102, 277, 1401, 816, 1399, 871, 1001, 1347, 2125, 1795, 2433, 1799, 1808, 2521, 1800, 2761]. Object-Based [272, 1355]. Object-Orientation [1359, 2629]. Object-Oriented [2789, 1358, 263, 1956, 1353, 2786, 2701, 228, 2166, 2040, 2608, 2663, 1372, 14, 287, 221, 1395, 935, 261, 2118, 1922, 2798, 2299, 1236, 2610, 1346, 264, 2644, 277, 1399, 871, 1001, 2433]. Object-Support [2474]. Objective [1664, 1691]. Objects [879, 269, 2217, 808, 268, 2213, 1902, 1363, 2599, 807, 2141, 1262, 1244, 1388, 1972, 1373, 2228, 2686, 1397, 548, 1806, 1794]. Oblivious [2469, 359, 371]. Observation [16]. Observations [1510, 1513, 2258]. Observers [1446]. Obstacles [1491, 2542, 1492]. Obtaining [1659]. Occlusion [1901]. Occur [1820]. Occur-Check [1820]. Occurrence [47, 444]. Ocean [2745]. OCM [2057]. OCore [935, 261]. Odour [453]. ODP [2227]. Odyssey [2353]. Off [1483, 420, 1949, 2271, 2658]. Off-Line [1483, 2271, 2658, 420]. Off-the-shelf [1949]. Office [1520]. Offline [1814]. offs [374]. Ohta [2586]. Oil [2714]. Okamoto [2586]. OKW [2299]. OLAF [1380]. Olan [276]. Old [843]. Omega [892]. Omega-Omega [892]. OMG [7]. OMIS [2057]. OMT [2128, 2493]. On-Chip [527, 1067, 1951]. On-Line [1483, 1485, 773, 818, 990, 753, 821, 988, 801, 424, 1676, 989, 616, 528, 517, 420, 1949, 1409, 2588, 1691, 2578, 2184]. on-the-fly [91]. One [775, 2259, 1079, 137, 135]. One-Rule [137]. One-Step [135]. One-Time [2259]. Online [385, 399]. Only [1123, 1023]. onto [2736]. Ontology [1406, 603]. OODB [1373]. Opal [41]. Open [1916, 728, 1126, 1214, 1619, 1138, 865, 2226, 2154]. Open-Ended [1619]. Operability [40, 1217]. Operating [1751, 2474, 307, 1075, 1100, 2730, 271]. Operation [2396, 437, 1069, 1151]. Operations [992, 1015, 982, 2252, 913, 2715, 466, 2338, 201]. Operator [1644, 1151, 1696, 640, 1660, 1788, 2631]. Operators [789, 2040, 1702, 1393, 2338, 80, 210]. Opponent [803]. Opportunities [2031]. OPS [884]. OPS-Based [884]. Optical [1485, 691, 523, 1300, 1029, 370, 1007, 1005, 895, 521, 2577]. Optically [1745]. Optima [850, 1785]. Optimal [1073, 1992, 1492, 494, 1147, 2072, 1407, 877, 2546, 882, 1468, 885, 2563, 1245, 1032, 1006, 2298, 1083, 2739, 1975, 784, 210, 2157, 639, 2225, 444, 2508, 2527]. Optimality [906, 487, 1419]. Optimisation [2096, 602, 1714, 1675, 1657, 1635, 1691]. Optimising [994, 1773, 1003]. optimizable [2186]. Optimization [1669, 492, 501, 765, 2705, 967, 1678, 778, 731, 1715, 978, 1713, 2725, 2802, 1385, 1664, 1987, 777, 1044, 1674, 1523, 2798, 1612, 1716, 1698, 1637, 1072, 1699, 1707, 1422, 2758, 2777, 1863, 2188]. Optimizations [2695, 2744]. Optimize [1155]. Optimized [1309, 1504, 1827, 2117, 2146]. Optimizer [2095, 2105]. Optimizers [1672]. Optimizing [873, 2699, 185, 2028, 2334, 1393]. Optimum [769]. Options [2689, 1030]. Optoelectronic [1026]. OR-Parallel [947].

Oracle [1230, 2274]. Oracles [366]. ORB [2221]. Order [1575, 1814, 1986, 138, 2136, 140, 1097, 1592, 124, 12, 122, 2265, 1105, 139, 1570, 470, 1562, 1135, 2479, 2038, 1109, 1110, 2146, 135, 143, 123, 539, 2195, 186]. Order-Sorted [1575, 12, 1562]. Ordered [1898, 2579, 251]. Ordering [701, 1124, 2720, 122]. Ordering-Based [1124]. Orderings [1849]. Orders [772, 590, 809, 362, 1442, 846, 363]. Ordinary [2753, 251]. ordinating [937, 946]. Organic [1710]. Organisation [2352, 2162]. Organization [1300, 1420, 2742, 1524, 1494, 1120, 2147]. Organizational [1933, 1932]. Organizations [1920]. Organizing [446, 1272, 432, 529, 471, 498]. Orientation [1359, 472, 2629, 1238]. Oriented [2789, 1358, 263, 2097, 1956, 1353, 1167, 2404, 2786, 2701, 667, 266, 228, 2166, 2683, 2040, 2608, 1462, 1662, 2663, 265, 1372, 1374, 14, 270, 287, 221, 1395, 935, 261, 1362, 2118, 1386, 1922, 2798, 2299, 1348, 1236, 2610, 1346, 2121, 1923, 1811, 264, 2644, 2102, 277, 1924, 605, 1399, 871, 1001, 2433, 2439, 1801]. Origins [20]. Orm [2339]. Orphan [1973]. Orthodontic [1303]. Orthogonal [1345, 229]. Orthogonality [2503]. Orthography [2470]. Oscillatory [510]. OSI [2448]. Other [2451, 1019, 352, 777]. Out-of-Core [2695]. Outdoors [799]. Outline [1291]. Output [1337, 2544, 2696, 881, 2496, 2265]. Over-Constrained [260, 246, 257, 253]. Overall [2627]. Overhead [1031]. Overlap [2554, 967]. Overlapping [889, 468]. overlay [1808]. Overview [2786, 2474, 2475, 29, 246, 2780]. OWE [526]. OZ [277, 2008, 1613, 2634, 1129]. Ozone [447].

p [2249]. PACA [921]. Package [2053, 2737, 2333]. Packets [774, 38]. Packing [1898, 2236, 2572]. Packings [768].

Page [1418]. Paginating [1673]. Paging [1484, 2584]. Pain [1299]. Pair [767]. Pair-Delta-Matroids [767]. Pairs [2573, 1864]. Paladin [262]. PALLAS [1082]. Palm [993]. Palm-Top [993]. PAN [1836]. Panel [1893]. Paper [1893]. PARA’96 [2743]. Paraconsistent [1140]. Paradigm [2176, 386]. Paradigms [1425]. Paradise [949]. Paragon [2234, 1143]. Parallel [1155, 1002, 2089, 2091, 968, 2689, 937, 263, 868, 1015, 779, 1077, 2062, 1151, 2692, 233, 1474, 1000, 965, 1496, 951, 1149, 1315, 1324, 1079, 1780, 1732, 2695, 2696, 2697, 2698, 942, 1080, 1147, 2701, 2703, 1183, 870, 1038, 927, 660, 687, 2088, 2704, 1182, 2620, 1471, 2242, 552, 2564, 982, 1150, 2705, 862, 910, 921, 994, 2096, 995, 980, 1711, 658, 2587, 1322, 985, 998, 659, 675, 1081, 946, 2711, 778, 665, 2231, 1059, 1473, 958, 2712, 2713, 2714, 2715, 466, 866, 990, 2717, 1075, 978, 688, 1157, 858, 983, 2721, 2722, 2723, 464]. Parallel [656, 1020, 2724, 1371, 2725, 1320, 2728, 2073, 867, 931, 974, 669, 993, 2729, 979, 2070, 2059, 604, 947, 975, 2730, 2733, 1486, 915, 948, 463, 935, 261, 1141, 1153, 1082, 2626, 1142, 1053, 952, 1475, 972, 1404, 1044, 865, 2053, 1083, 2737, 2055, 2738, 2642, 1037, 2741, 2473, 945, 950, 2742, 1042, 962, 2241, 2240, 1008, 874, 2747, 657, 2746, 2243, 1166, 1160, 2060, 2748, 1317, 896, 1152, 526, 2750, 2239, 860, 977, 1021, 973, 936, 2082, 1699, 1667, 2066, 1457, 953, 679, 963, 1054, 1708, 2052, 1028, 878, 2061, 2755, 2756, 1143, 2758, 1010, 2759]. Parallel [670, 1459, 2651, 1003, 986, 2084, 961, 1018, 521, 2631, 2589, 2237, 2054, 970, 1148, 871, 1001, 956, 2762, 666, 924, 668, 1683, 2802, 175, 1049, 947]. parallèle [1071]. Parallelim [1066]. Parallelisation [2709, 2739, 2761]. Parallelising [2734]. Parallelism [263, 1151, 875, 906, 916, 787, 1321, 1583, 1596, 2609, 1323, 1008, 684]. Parallelizable [1681, 2760]. Parallelization [2065, 1019, 905, 2053, 2745, 2050, 2095]. Parallelizing [934, 147]. Parameter [529, 483, 1676]. Parameter-Free [529]. Parameterised [1658]. parameterized [69]. Parameterizing [1178]. Parameters [1147, 379, 2294, 1632, 1633, 2201, 2515]. Parametric [1259, 425, 2023, 1581, 2394, 569]. Parametrised [1719]. Parametrized [1761]. ParaStation [1159]. Parchments [1232]. Parent [1646]. ParFiSys [927]. ParkBench [999]. PARPACK [2737]. PARSE [954]. Parsers [1208, 1209]. Parsimonious [2168]. Parsing [786, 1532]. PARSMI [2724]. Part [449, 2111]. Part-Of-Speech [449]. Partial [1814, 726, 1817, 1218, 1986, 381, 1590, 255, 1490, 2033, 1392, 590, 248, 1588, 464, 939, 386, 2658, 915, 463, 389, 391, 1232, 969, 1442, 539, 398, 399, 1591, 2119, 722, 970, 2562]. Participants [635]. Participatings [638]. Particle [664, 402, 964]. Parting [1899]. Partition [1165, 926, 2574, 65]. Partition-Processing [926]. Partitionable [1958]. Partitioned [2753]. Partitioning [1471, 662, 1973, 2720, 1703, 2244, 2736, 1727, 214, 661, 2755]. Parts [2694, 632]. Party [2250]. Passing [1968, 2074, 2707, 872, 2213, 2049, 2740, 2243, 2050]. Passive [691, 1839]. password [2442]. password-based [2442]. Past [151, 2489, 199]. PATCH [1544]. Patching [157]. Path [1491, 758, 212, 733, 1627, 122, 1393, 670, 551]. Pathologies [1256]. Paths [383, 2573, 2562, 2528, 2515]. Patient [1308]. Patients [1281]. Pattern [2551, 1063, 2072, 1343, 1989, 1689, 634, 784, 408, 883, 2504, 398, 792, 2508, 1480, 568, 2204]. Pattern-Associative [883]. Pattern-Matching [634]. Patterns [473, 450, 140, 1293, 801, 806, 429]. Payload [1742]. Payment [1839, 2269, 1838]. PC [2437]. PCP’s [1459]. PCSP [706]. Peckish [2012]. Peculiarities [809, 2118]. Pedagogical [310, 306, 296]. Pedagogically [289]. Pedicle [1311]. Peeling [1253]. Pentadiagonal [1165]. Pentium [368]. PEP [49]. PEPE [2718]. Perception [459, 521]. Perceptron [494, 463, 1996, 418, 1016]. Perceptrons [785, 1687]. Perceptual [2777]. Performance [2687, 984, 868, 233, 2700, 2702, 1172, 927, 24, 862, 2706, 888, 1407, 1672, 1022, 857, 1321, 2715, 2716, 1687, 2775, 861, 1056, 1405, 2059, 1141, 1718, 905, 2735, 1520, 2484, 1319, 1145, 2046, 680, 2333, 981, 1054, 869, 1417, 2776, 1399, 244, 2221]. Performance-Directed [1718]. Performances [526]. Performing [1313]. Period [2604, 2430]. Periodic [1329, 1330, 2576, 372, 996]. Permission [1220]. Permutation [1645, 1325]. Permutations [1472, 893]. Persistence [1123]. Persistency [2612]. Personal [2768]. Personalized [1413]. Perspective [2176, 1315, 1932, 2172, 1519, 2484, 1810, 2629, 1304, 2388, 1642]. Perspectives [407, 2662, 403]. PET [1297]. Petri [2403, 943, 954, 2336]. Petroleum [1772]. PHANTOMAS [522]. Phase [1035, 733, 1248, 1338, 2018, 770]. Phase-Transition [2018]. phi [95, 74]. Philosophy [585]. Phoneme [442]. Photoelastic [343]. photon [2444]. Phylogeny [2548]. Physical [2122, 475, 1697, 2759]. Physics [317, 319, 402]. Physiological [519]. PICASSO [632]. pictorial [1797]. Picture [2353]. Pictures [1602, 1806]. PIDGETS [1602]. Piece [313]. Pieces [2377]. Piecewise [1902, 1874]. Pins [1277]. Pipe [1714]. Pipe-Route [1714]. Pipeline [1717]. Pipelined [1084, 918, 877, 1729, 983, 991, 1078]. Pipelining [1073, 1072]. Pivoting [2100]. Pixel [993, 2528]. Pixel-Level [993]. Placement [930, 2073, 817]. Placing [1902].

Plan [2151]. Planar [1491, 1980, 793, 2778, 751, 670, 2519]. Plane [1491, 2542, 2567]. Planes [1548, 2524]. Planning [2362, 725, 1002, 1303, 1492, 1309, 2312, 1498, 1500, 1997, 1308, 1188, 1772, 1674, 2380, 2478, 1301, 2024, 245, 1302]. Plans [1080, 630, 2377, 627]. Plant [437, 1997, 1696, 1714, 1001]. Plants [2396]. Plasticity [473, 477]. Plate [1278]. Platform [2211, 77]. Platforms [2071, 1046]. Plausible [1513, 1969]. Play [1718, 627]. Play-Plans [627]. PLDs [2637]. PLS [2088]. Pluggability [1201]. plus [400]. PN [943]. Pocket [487]. POETView [2681]. Point [2307, 2546, 36, 991, 1276, 1278, 1723]. Point-Based [1278]. Point-to-Point [36]. Pointers [1164]. Points [822, 1291, 838, 2727, 840, 2261, 562, 2523]. Policies [1916, 2243, 2434]. Policy [2241]. Pollution [2694, 2642]. poly [560]. Polygon [2574, 2573]. Polygonal [2542]. Polygons [1901, 754, 2554, 2532]. Polyhedra [1901, 761, 2523]. Polyhedral [907]. Polyhedric [1032]. Polyhedron [1899]. Polymers [453]. Polymorphic [1575, 262, 1577, 542]. Polymorphism [1219, 1581]. Polynomial [66, 831, 1823, 562, 2499, 2581, 721, 577, 1861, 177, 2514]. Polynomial-Time [1823, 66]. Polynomially [524]. Polynomials [2090]. polyominoes [2537]. Polytypic [1204]. Polytypically [1578]. Polyvariant [2659]. Population [1655, 1687, 509]. Population-Based [1687]. Populations [1622, 1776]. Portable [2160, 994, 2070, 1082, 2737, 1160, 2060, 1717]. Porting [2071]. Portrait [2728]. Pos [1585]. Pose [2163]. Posets [638]. Position [1893]. Positive [2551, 2647, 2722, 1459, 1859, 2205, 1874, 2204]. Possibilistic [221]. Possible [512, 2590, 2263, 1786]. Possible-is-Certain [2590]. Post [984, 500, 2358, 1258].

Post-Deployment [2358]. Post-Processing [984, 1258]. Post-Processor [500]. Postage [2635]. POSTGRES [2682]. Poststore [1030]. Potato [1488, 891]. potential [151]. pour [1071]. Power [2650, 2396, 129, 2033, 2713, 437, 837, 2590, 1694, 1997, 1696, 22, 324, 1001, 555, 2699]. Power-Plant [1001]. POWER2 [2687]. Powerful [89, 1080]. PPARDB [1160, 2060]. PPARDB/PVM [2060]. Practical [2465, 2693, 378, 1999, 699, 358, 1343, 360, 2582, 2051, 1070, 2341, 1156, 325, 682, 191, 112]. Practicalities [2118]. Practice [18, 1449, 2354, 297, 1931, 1056, 2015, 537]. Practices [2353]. Pragmatics [381]. PRAM [555, 1022, 1004, 1007, 1006]. PRAMs [1005]. Pre [1818]. Pre-interpretation [1818]. Precedence [1043, 1037]. Precise [2564, 1609, 472]. Preconditioner [2065, 659]. Preconditioners [971, 2067]. Predicate [582, 164, 165]. Predicates [2135, 546]. Predicted [509]. Predicting [1273]. Prediction [447, 1813, 233, 452, 985, 1903, 502, 1056, 614, 426, 1054, 2054, 1873]. Predictive [500]. Predictors [1063]. Preemptive [1457]. Preferences [583]. Preferential [1531]. Prefetch [1064, 1030]. Prefetching [1058]. Prefix [16]. prefixing [575]. prefixing-keys [575]. prefixing-tries [575]. Preliminary [626, 730, 1647, 631, 2591, 302, 2221]. Preorder [17]. Preorders [10, 1556]. Preprocessing [971, 663, 1344, 21]. Presence [1971, 1033, 1439, 1338, 1634]. Present [2489, 1540]. Presentation [52, 309, 337, 1185, 173]. presentations [188]. Presented [1544]. Presenting [171]. Preservation [2161]. Preserve [127, 1574, 2207]. Preserving [1221, 1103]. Pressure [1682]. Prevail [2337]. Prevent [2270]. preventive [2437]. Price [747].

Pricing [2724]. Prim [2653]. Prim-Dijkstra [2653]. Primary [339]. Prime [828, 2555, 1570]. Primes [1182]. Primitive [835, 2124, 2600]. Principal [428, 846]. Principle [1100]. Principled [432]. Principles [2691, 529, 1681, 2391]. Printed [802, 820]. Priority [1015, 894, 1463, 142]. Priority-Driven [894]. Privacy [2455, 2269, 2464]. Probabilistic [2565, 839, 416, 565, 1630, 1522, 2478, 591, 1288, 2208, 66]. Probabilities [531, 1526]. Probability [2283]. Probing [1470]. Problem [2401, 2305, 2402, 779, 1167, 1701, 2019, 2405, 2399, 2287, 2288, 2278, 754, 699, 2289, 2368, 766, 916, 2291, 2010, 310, 2548, 1473, 1702, 1466, 2496, 2293, 701, 1041, 1782, 2407, 2408, 931, 1465, 763, 2411, 683, 2295, 2567, 2297, 288, 2378, 2098, 752, 1675, 686, 2381, 2747, 1681, 1527, 1034, 2796, 1538, 2413, 2664, 2300, 137, 2507, 2388, 2414, 2644, 670, 2415, 254, 2394, 724, 554, 568, 2274]. Problem-Oriented [1167]. Problem-Solving [2381]. Problemns [2746]. Problems [2089, 1643, 693, 1645, 765, 1471, 2016, 2096, 2552, 1900, 1769, 2723, 1482, 1627, 1214, 1508, 1989, 736, 629, 2653, 2560, 1464, 2748, 2021, 1635, 715, 257, 743, 1730, 2103, 722, 253, 682, 746, 2643, 2762, 723, 549, 551, 535, 1704]. PROBOT [1312]. Procedure [1293, 1525, 591, 195]. Procedures [632, 203, 169]. proceedings [2803]. Process [1927, 1913, 1971, 1937, 1912, 1939, 1567, 1908, 1928, 1910, 1925, 1933, 1914, 1911, 2069, 1597, 1915, 1921, 1931, 26, 1918, 2500, 2104, 289, 1935, 2738, 241, 1386, 1909, 1936, 1926, 503, 1774, 1923, 1917, 1934, 17, 2196, 2541]. Process-Centered [1912, 1933]. Process-Centred [1914]. Processes [1329, 1929, 25, 1566, 2649, 1589, 946, 1075, 1506, 1224, 1918, 2499, 635, 324, 569, 67]. Processing [2688, 407, 641, 984, 803, 687, 994, 995, 980, 1750, 993, 2070, 2223, 1258, 1036, 593, 1757, 926, 436, 1377, 595, 1166, 640, 1152, 479, 1540, 2481, 1373, 1422, 672, 514, 2800]. Processor [2242, 466, 1941, 500, 891, 1006, 2314, 676, 2240, 595, 2243, 1746, 2332, 2755, 1765, 970]. Processors [1742, 1978, 656, 2053, 2736, 962, 1951, 1040]. Proctoring [1848]. Produced [427]. Product [2762]. Product-Free [2762]. Production [725, 1002, 2156, 1435, 634, 1706, 296, 1143, 198]. Products [2343, 140, 2510]. Professor [323]. Profiling [1207]. ProFun [1598]. Prognoses [2387]. Program [2281, 1809, 2283, 2131, 2698, 1091, 1819, 2608, 1821, 1572, 1823, 385, 867, 931, 1153, 1053, 1587, 1595, 2742, 2656, 1810, 394, 395, 396, 1601, 2792, 2651, 587, 724]. Programmable [1759, 1765, 1728, 1752, 325, 723]. Programmed [1450]. Programming [2282, 1002, 401, 1122, 1575, 1130, 1497, 263, 1644, 2048, 276, 1035, 59, 1088, 1670, 951, 1315, 949, 2795, 1567, 942, 870, 267, 2620, 2005, 266, 699, 1222, 41, 1539, 946, 2709, 2608, 1059, 866, 265, 2047, 336, 2623, 1200, 2006, 29, 51, 1014, 734, 1617, 1628, 1204, 1771, 941, 948, 261, 1133, 1082, 2007, 737, 2617, 943, 1323, 1608, 749, 1051, 242, 2681, 940, 713, 47, 1774, 1775, 1236, 2610, 1021, 716, 1602, 1613, 2634, 1129, 2639, 1606, 2593, 245, 2630, 719, 1459, 1607, 2767, 170, 1878, 2199]. Programming-Based [2630]. Programs [2282, 2670, 1155, 1812, 1164, 873, 233, 377, 1495, 1324, 2695, 1818, 378, 2044, 2619, 862, 910, 1092, 1438, 1815, 2078, 1561, 1609, 1594, 2079, 1320, 2625, 1618, 2032, 947, 1835, 915, 2025, 22, 2034, 1044, 2665, 1605, 865, 28, 390, 1983, 1824, 747, 2094, 957, 896, 2750, 2239, 1612, 1827, 684, 1622, 397, 741, 2300, 1049, 922, 2167, 1833, 651, 545, 2203, 2205, 2197, 186, 576]. Progress [1937, 2458, 1998, 957]. Progressive [687, 668]. Project [4, 1159, 1079, 1998, 1931, 1385, 1318, 2771, 2158, 1936, 2482, 2233]. Projection [481, 1157, 313, 496, 632]. Projections [1344, 1826, 2537]. Projective [1548]. Prolog [1611, 388, 2205, 1595, 1581, 2666, 1824, 1836, 181, 1830, 1584]. Proof [2130, 1171, 2132, 1543, 2312, 2133, 928, 1092, 208, 130, 125, 1188, 134, 2141, 1525, 153, 2323, 1110, 43, 144, 154, 199, 161, 185, 155, 100]. Proof-Checking [2130]. Proof-Manager [43]. Proof-search [154]. Proof-terms [153]. Proofs [1117, 2131, 149, 1115, 2271, 1503, 268, 612, 1824, 1571, 171, 213, 192, 612]. Propagating [649]. Propagation [762, 695, 729, 700, 1903, 705]. Proper [754, 892]. Properties [2784, 626, 1957, 249, 24, 475, 1679, 31, 128, 2025, 2296, 2063, 2118, 2329, 957, 395, 633, 2540, 558, 85, 2527, 1563, 1807]. Property [694, 1023, 1555, 1811, 556]. Property-Oriented [1811]. Proplane [55]. Proposal [2704, 2500, 2163]. Proposed [515]. Propositional [1538, 205, 203]. Prosodic [1512]. Prospect [2714]. Prostate [1308, 1257]. Prostatectomy [1312]. Protect [365]. Protection [2464]. Protein [502, 1897, 1497]. Protocol [1700, 1089, 2622, 1074, 2621, 1163, 2772, 2138, 1201, 38, 1057, 1436, 926, 1857, 87, 231, 2586, 82, 106]. Protocols [2248, 2429, 2230, 2459, 1959, 780, 62, 2427, 2419, 99, 2428]. Proton [330]. Prototype [2663, 1760, 508]. Prototypes [1738]. Prototypical [1520]. Prototyping [308, 1727, 1348, 1736, 1752]. Provable [2580]. Provably [360, 2268, 369]. Provably-Secure [360]. Prove [19, 1521]. Prover [1497, 42, 1227, 2061, 147, 43, 144, 183, 181, 152, 166]. Provers [359, 175]. Provide [875]. Providing [40, 2700]. Proving [2308, 1535, 1957, 1167, 116, 1172, 1090, 1815, 1499, 1500, 121, 150, 359, 2296, 28, 2063, 957, 117, 1548, 1553, 71, 156, 151, 190, 78, 159, 211, 188, 158]. Proxies [1024].

PSAP [725]. PSEEs [1916]. Pseudo [1147, 187]. pseudo-linear [187]. Pseudorandom [1953]. PSOMs [431]. PSPVM [2076]. Psychologically [339]. Psychophysics [2775]. Public [1856, 1857, 2418]. Public-Key [1856, 1857, 2418]. Purchase [1997]. Purely [428]. Purpose [298, 483, 1622]. Pushdown [67]. Pushouts [12]. Putting [1519]. Puzzle [686]. PVM [2089, 2091, 2064, 2048, 2068, 2076, 2083, 2702, 2093, 2078, 2069, 2047, 2079, 2081, 2077, 2049, 2067, 2070, 2063, 2094, 1160, 2060, 2095, 2051, 2066, 2052, 2056, 2061, 2087, 2075, 2084, 2085, 2058, 924, 2092]. PVM-Based [2061]. PVM-GRACE [2092]. PVM-Implementation [2067]. PVM/C [2079]. PVS [1085, 1094, 2408, 100, 2311, 2323, 2302].

Q [703, 2249, 1679]. QDDs [62]. QoS [2767]. QR [1475]. Quadratic [2248, 794, 846, 851]. Quadratic-Time [794]. Quadrature [1755]. Qualification [1531]. Qualitative [2177, 1541, 2148, 603, 1526]. Quality [2116, 1270, 309, 1769, 2114, 2210, 1242, 1907, 2777, 1251]. Quantification [1293, 1295, 2531]. Quantified [1117]. quantifiers [164, 165, 210]. Quantitative [677, 2741, 83, 2436]. Quantization [430]. Quantum [2242, 370, 371, 2444]. Quantum-Based [2242]. Quasi [2538, 823, 2703, 1327, 972, 714]. Quasi-Cyclic [1327]. Quasi-Laguerre [972]. Quasi-Linear [823]. Quasi-Minimal [2703]. Quasi-Shear [2538]. Quasi-Solutions [714]. Quaternary [2253]. Queries [1491, 1404, 2778, 1416, 721, 1148, 2802]. Query [2784, 2106, 1080, 1126, 2107, 2683, 2045, 861, 1351, 1385, 2679, 2028, 2780, 1386, 2185, 2798, 1377, 2037, 1373, 1422, 1078, 721, 1862, 547]. Query-Dependent [2679]. Querying [2793]. Question [305, 1396]. questions [188]. Queue [1015, 1004]. Queue-Read [1004]. Queue-Write [1004]. Queues [1463]. Quickpropagation [483]. Quorums [1980, 1074].

R [1834, 720, 745, 2032, 1807]. R-Stable [2032]. Race [1486, 1725]. Race-Condition [1486]. Radial [1767, 1684, 453, 521]. Radical [1538]. Radio [1903]. Radiodiagnosis [342]. Radiologists [342]. Radiology [804]. Radiometric [1255]. Radiosity [1905, 959, 668]. Radiosurgical [1309]. Radix [987]. Radon [1146]. RAID [1946]. Railroad [745]. Railway [1943]. Ramsey [764]. Random [1331, 2494, 1147, 1472, 1628, 1456, 1631, 1708, 1288, 1010, 1342]. Random-Like [1331]. Randomization [1966, 1461]. Randomized [1484, 2587, 688]. Randomly [1469, 1328]. Range [422, 2733, 1652]. Rank [2278, 832, 2726, 491]. Ranking [2587, 2575, 2559, 960]. Rapid [1727, 1301, 1729, 1071]. Rapid-2 [1071]. Rapidly [1338, 2188]. Rate [1327, 477, 1660]. Rates [1293, 411, 1786]. Ratio [886, 1466]. Rational [822, 1495, 1124, 1181, 708, 840, 1125]. Rationale [2481]. Rationality [630, 2027]. Raw [1743]. Ray [826, 840, 1245, 1897]. Ray-Representation [1897]. Rayleigh [1337, 1339]. RBF [998, 503, 417]. RBF-Nets [503]. RC [331]. RC5 [361]. Reactive [1585, 1450, 35, 267, 2505, 101, 576]. Reactivity [2027]. Read [1004]. Real [2248, 2648, 2494, 1437, 1306, 2211, 2290, 2368, 435, 1438, 1447, 1435, 1444, 1356, 1965, 1448, 2775, 1372, 1252, 1648, 993, 1997, 1445, 1357, 1050, 1824, 1442, 1429, 405, 851, 2595, 274, 1408, 1398, 101, 98, 2192, 108]. Real-Coded [1648]. Real-Quadratic [2248]. Real-Time [2648, 1437, 2211, 2290, 1438, 1447, 1435, 1444, 1356, 1965, 1448, 2775, 1252, 993, 1445, 1357, 1442, 1429, 274, 1398, 2595, 101, 98, 108]. Realisation [1231]. Realistic [1015, 445]. Reality [1308, 1803]. Realizing [2766, 228]. Reals [2133]. Realtime [2680]. Rearrangeable [2731]. Reason [1970, 2479]. Reasoning [2176, 401, 1122, 2363, 1140, 1496, 2396, 580, 1536, 1191, 1406, 1541, 1092, 613, 730, 1554, 1134, 1127, 590, 2370, 1137, 334, 1556, 2371, 2372, 1189, 385, 2477, 1509, 1510, 1511, 2375, 114, 2148, 604, 2039, 603, 1228, 1517, 1772, 1522, 2381, 1526, 2004, 1531, 2388, 2390, 592, 2438, 181]. Reasonings [617]. Reasons [693]. Rebuilding [1250]. Recalibration [508]. Recalling [458]. Recastable [2256]. Receipt [2257]. Receipt-Free [2257]. Receiver [1338]. Receptive [468]. Recognition [802, 820, 445, 1536, 818, 787, 500, 1343, 454, 753, 1689, 1345, 456, 807, 821, 801, 2151, 784, 616, 408, 819, 2445, 799, 783, 632, 792, 2163, 514, 461, 1795, 1804, 1802, 1800]. Recognizing [2524, 813, 2510]. Recombination [1632, 1633, 1658]. Reconfigurable [1743, 1742, 1729, 1748, 2718, 2575, 1739, 1737, 1725, 1747, 1745]. Reconstruction [2537, 1157, 1257, 1256, 532, 1236, 816]. Reconstructions [2713]. Record [2279]. Recorder [1757]. Records [1607]. Recoverable [2594]. Recovering [1792]. Recovery [1355, 1410, 2247]. Recruiting [1623]. Rectangular [313]. Recurrences [2703, 934]. Recurrent [236, 427, 1515, 504, 1880, 1879]. Recursion [2650, 1604, 379, 2107, 2664, 2600, 565, 207]. Recursions [1391]. Recursive [2355, 1569, 2647, 606, 122, 11, 609, 169]. Recursively [2560]. Red [1476, 2744]. Red-Black [2744]. Red-Blue [1476]. Redistribution [876, 878]. Reduce [1024, 1194]. Reduced [1337, 608, 2726, 516]. Reduced-Rank [2726]. Reducing [1031, 2190]. Reduction [1604, 127, 829, 2711, 2715, 416, 2726, 1603, 914, 908, 848, 88, 2193, 539]. Reductions [60]. Redundancy [1342]. Redundant [695]. ReDuX [60, 145]. Reed [1333]. Reference [818, 1908, 2509]. References [2107]. Refined [2282, 836]. Refinement [1091, 1819, 19, 874, 2300, 2415, 88, 65]. Refining [2284]. Reflecting [2209]. Reflection [2139, 271, 206]. Reflective [1173]. Reflexive [684]. Refueling [1696]. Region [1262, 633]. Region-Growing [1262]. Regional [2261]. Regions [1136]. Register [1593, 1072]. Registers [2276]. Registration [1275, 1279, 1293, 1280, 1246, 1278, 1314, 1277]. Regression [415, 435, 2153]. ReGTime [2216]. Regular [2551, 671, 383, 783, 1480, 1868, 1880, 1881, 1876, 1875]. Regularity [2499]. Regularization [495, 418]. Regularized [1269]. Regulation [2269]. Reification [1221]. Reinforcement [301]. Rejection [452]. Related [355, 1471, 759, 2498, 2673, 2377, 1454, 2534]. Relating [1009]. Relation [2049, 299]. Relational [1080, 1661, 2122, 796, 2495, 1371, 1370, 1369, 397, 2117, 2036, 2668]. Relations [1211, 1566, 698, 829, 1821, 2101, 1228, 2414, 2589]. Relationship [2113, 2117, 2001]. Relationships [2100, 2123]. Relative [842, 2144]. Relaxation [258, 259, 795]. Relaxing [1023]. Relevant [2624]. Reliability [414, 1944, 2417]. Reliable [2283, 1971, 2214]. relief [1797]. Relying [2406]. Remarks [2257]. Remembrance [597]. Remote [307]. Remotely [981]. Removal [1961]. Removing [391]. Renaming [1982]. Rendering [1246, 1244, 1304, 1251]. Renovation [2486]. Rent [2216]. Reorganization [1409]. Repartitioning [676]. Replacement [1418]. Replay [863]. Replica [1074]. Replicated [1960, 963, 74]. Replicating [2417]. Replication [1405, 1387]. Report [626, 2350, 2458, 1895, 730, 1998, 631, 2591, 1724, 1936, 1790]. Reporting [1476]. Repositories [1351]. Representability [1125]. Representation [2493, 1080, 818, 2646, 1290, 1187, 2738, 1522, 2113, 962, 1790, 1897, 240, 417, 1530, 851, 2509, 1805, 2517, 2186, 1793, 2270, 2525, 63, 2519]. Representations [1645, 460, 2618, 613, 1124, 1506, 794, 298, 2334, 2117, 2157, 1651, 1799, 1808]. representatives [539]. Represented [620, 117]. Representing [1794, 806, 2341, 2119]. REPRO [2389]. Reproducible [1264]. Reproducing [450]. repudiation [2419, 2418, 2258]. Request [2235, 2221]. Request-Scheduling [2235]. Requirements [2282, 2227, 1928, 1444, 298, 2676, 896, 592]. requiring [549]. Rerouting [782]. Rescheduling [696]. Research [2483, 600, 1318, 5, 1624, 543]. Residual [2703, 1565]. Residuation [1200]. resilient [2423]. RESIS [1072]. Resistant [2468]. Resolution [2160, 59, 610, 1115, 612, 947, 204, 1825, 816, 2596, 196, 209, 2800, 153, 152]. Resolution-based [204]. Resolves [1249]. Resonance [1293, 1247]. Resource [2069, 1059, 2158, 2384, 744, 1078, 98]. resource-bound [98]. Resource-Constrained [2158]. Resources [1060]. Respect [1477]. Responses [532]. Restarted [2756]. Restarting [2502]. Restoration [472]. Restricted [1773, 1785, 560]. Result [1104]. Results [34, 1079, 999, 375, 2152, 1473, 1343, 1146, 1036, 711, 2114, 5, 2780, 1061, 2655, 723, 202]. Retiming [1427]. Retinal [474]. Retransmission [38]. Retrieval [1412, 1413, 2367, 1114, 2376, 1518, 1414, 2173, 2386, 2389]. Retrospective [2022]. Retry [1045]. Return [1947]. Reusable [944]. Reuse [2366, 2370, 2224, 2378, 56, 907, 159].

Reusing [2391]. Reveals [509, 471]. Reverse [1536, 2122, 2123]. Review [2688]. Revised [1910, 2724]. Revision [1501, 1516, 2348, 2159, 1139]. Revision-Based [2159]. Revisitation [394]. Revisited [760, 132, 1919, 1493, 2359, 1564, 1640, 1533, 202]. Revoking [1839]. Rewrite [1218, 132, 114, 1229, 133, 198, 147, 1238]. Rewrite-Based [114]. Rewrite-Rule [147]. Rewriting [1174, 1542, 1183, 1543, 145, 2597, 1599, 120, 141, 1226, 44, 1573, 1228, 131, 134, 2502, 1173, 142, 135, 126]. Ribbon [2090]. Right [217]. Rights [2605]. Rigid [1116, 813, 154]. Rigorous [1909]. Ring [1036, 2577]. Rings [346, 133]. Ripple [820]. Rippling [1509]. RISC [1941, 1025]. Risk [2220, 2275]. RNN [489]. Road [1784]. Roadmap [384]. ROBDD [2335]. ROBDDs [1585]. Robot [1992, 1492, 446, 455, 1688, 508, 506, 504]. Robotic [275]. Robotics [2180, 1778]. Robots [1539, 1680]. Robust [1968, 2460, 357, 430, 1688, 1257, 505, 1276, 1622, 1668]. Robustly [1887]. Robustness [1886, 1985, 1888, 2159]. Role [2363, 1212, 1718, 2503, 1841]. Roles [638, 1529]. Root [356]. Rosser [213]. Rostering [2630, 2001]. Rotation [2538, 987, 408]. Round [2580]. Rounding [987, 2507]. Route [1714, 442]. Routed [691]. Router [883]. Routers [881]. Routes [1768]. Routine [967]. Routines [2706, 1192, 2066]. Routing [894, 691, 888, 893, 890, 1976, 1977, 2456, 1487, 891, 2731, 892, 895, 2570, 574, 2764, 2606]. Royal [1784]. RPC [2401, 2402, 2405, 2400, 2399, 2407, 2408, 2411, 2413, 2414, 2415]. RPC-Memory [2401, 2402, 2405, 2400, 2399, 2407, 2408, 2411, 2413, 2414, 2415]. RS [2687, 2715]. RS/6000 [2687, 2715]. RSA [358, 357, 352]. RSA-Based [358]. RSL [2301]. RTL [105]. Rubber [503]. Rubik [893]. Ruby [168]. Rule [2156, 2648, 608, 490, 481, 2783, 1099, 1990, 137, 1380, 519, 147, 2511, 540, 198, 483]. Rule-Based [2156, 2648, 2783]. Rules [689, 820, 1088, 1354, 1503, 2498, 2373, 1989, 1394, 2028, 615, 2030, 2026, 2003, 650, 1238, 1991, 1379]. Run [862, 380, 2663]. Run-Time [862, 380, 2663]. Running [2092]. Runtime [272, 2240, 963, 274]. RX [342].

S [838, 2301]. S-Integral [838]. SAC [2265]. Saccade [439, 476]. Safe [1772]. SAFER [364]. Safety [1957, 1949, 2296, 112, 85]. Safety-critical [1949]. Salesman [1702]. Salzburg [1146]. samplable [577]. Sampling [1434, 2678, 1940, 2527]. SAR [984]. SAR-Image [984]. SARE [301]. SAT [692, 2748, 740]. Satisfaction [691, 248, 256, 701, 711, 743, 722, 254, 253, 746, 2643]. Satisfiability [1730, 723, 214, 1546]. Saturation [151, 1135, 570]. Saturation-based [151]. Saving [90]. SBC [342]. SBC-RX [342]. Scalability [2765, 959, 2159]. Scalable [862, 465, 1144, 1409, 2490, 1009, 1030]. ScaLAPACK [2706, 967, 2708, 2735]. ScaLAPACK-Based [2706]. Scale [862, 790, 2682, 1047, 2737, 1263, 2386, 2534]. Scale-Space [1263]. Scaling [654]. Scan [1019, 165, 164]. Scanning [337, 1265, 809]. Scans [919]. SCAST [216]. Scattering [977]. Scenarios [1852, 1551]. scene [1803]. Scenes [1360, 799]. Schedule [1769, 364, 1032]. Scheduled [1437, 1559, 1066]. Scheduler [2234]. Schedulers [2231]. Schedules [1073, 2020]. Scheduling [1043, 2245, 1035, 1645, 1038, 762, 1045, 1438, 1815, 2017, 1356, 2011, 2583, 2236, 1594, 1505, 1033, 2235, 1041, 1694, 2238, 1439, 1593, 1047, 1036, 777, 1039, 1404, 2158, 1037, 1674, 2739, 1460, 1042, 2241, 2240, 1706, 1040, 1612, 2021, 1070, 1457, 1049, 744, 996, 2579, 2237, 1398, 2001, 1709, 2578]. Schema [2116, 2122, 2112, 2099, 1387, 2110, 320]. Schemas [2117, 2102]. Schematic [1753, 264]. Schematic-Based [1753]. Scheme [2542, 450, 1074, 2256, 939, 1619, 1050, 1245, 2257, 2247, 657, 2594, 741, 1668, 1871, 2420, 2439, 2421, 938, 264]. Schemes [2650, 374, 355, 2270, 2236, 1977, 360, 1420, 2262, 2269, 2098, 2275, 1376, 1, 2268, 973, 2664, 2260, 1637, 1030, 1881, 2422]. Schizophrenic [1286]. Schlumberger [1924]. Schnorr [373]. Schnorr-Vaudenay [373]. School [2000, 620]. Schroeder [1857]. Schwarz [2692]. SCIDDLE [2064]. SCIDDLE-PVM [2064]. Science [1213, 339, 215, 281, 2803, 2513]. Scientific [377, 2671, 585, 2678, 1317]. Sclerosis [1295]. SCM [2356, 2353]. Scrambling [1329]. Screen [2679]. Screen-Filling [2679]. Screening [1268]. Screws [1311]. Scripting [2611]. SDF [3]. SDM [482]. Search [1413, 697, 2705, 798, 2487, 209, 1693, 1702, 702, 365, 2591, 1675, 1773, 2780, 1666, 1657, 713, 797, 1461, 1110, 1651, 1787, 1668, 1652, 1709, 208, 1882, 185, 197, 154]. Search-Based [2705]. Searches [2588]. Searching [2557, 1634, 1668, 2508]. Second [1743, 138, 139, 295]. Second-Order [138, 139]. Secondary [336, 502]. Secret [355, 2262, 2263, 2420, 2422, 2421]. secret-sharing [2422]. Sections [1257]. Secure [374, 2460, 358, 1938, 360, 2627, 1852, 1855, 2268, 369, 1837, 2433]. Security [1849, 2278, 1844, 1842, 2416, 356, 2426, 1853, 2437, 2417, 2436, 2428, 2434]. SEE [405]. SEE-1 [405]. Seeing [318]. Segment [931]. Segmentation [1271, 1266, 815, 1293, 1252, 1265, 1255, 531, 1775, 1242, 2189]. segmented [1791]. Segments [1476, 2545, 2514]. SEJONG [616]. SEJONG-NET [616]. Select [2392]. selectable [2441]. Selected [999, 1328]. Selection [987, 1015, 1780, 434, 425, 1616, 441, 1631, 1785, 1682, 1776, 1638, 1864, 444, 1642, 2085, 2194]. Selective [83]. Self [446, 1272, 432, 994, 556, 1656, 793, 2316, 529, 1983, 2742, 392, 471, 498, 1658, 1777, 399, 1623, 1947, 504, 189, 2432]. Self-Adapting [504]. Self-Adaptive [1656]. Self-Applicable [392, 399]. Self-Assemblage [1623]. Self-checking [1947]. Self-Consistency [2316]. self-labelling [189]. Self-Optimising [994]. Self-Organization [2742]. Self-Organizing [446, 1272, 432, 529, 471, 498]. self-shrinking [2432]. Self-Similar [793]. Self-similarity [556]. Self-Stabilizing [1983]. SEM [184]. Semantic [2042, 1359, 1174, 610, 1315, 1415, 202, 9, 1361, 1577, 2798, 797, 911, 2645, 46, 196, 573]. Semantics [1575, 18, 2636, 1816, 1536, 1424, 1590, 648, 1222, 1392, 1561, 2497, 1572, 2590, 1434, 2678, 2666, 1810, 1234, 2480, 716, 1601, 142, 1551, 278, 1383, 2781, 1851, 1139, 1240, 2640, 565, 537, 198]. Semantics-Based [1590, 2666]. Semantics-Directed [1810]. Semaphores [1486]. Semi [2160, 1057, 1125, 1243, 137, 210]. Semi-Automatic [2160]. Semi-Interactive [1243]. semi-lattices [210]. Semi-Markov [1057]. Semi-Representability [1125]. Semi-Thue [137]. Semilattices [1828]. Semipositive [2044]. Semiring [249]. Semiring-Based [249]. Sensed [981]. Sensing [453]. Sensitive [2544, 2496, 1605, 1854, 783, 581, 1868, 2193]. Sensitivity [434, 704]. Sensors [436]. Separating [443, 2545]. Separation [451]. Sequence [2041, 1337, 858, 2746, 1873, 1870]. Sequences [1437, 1692, 2604, 372, 457, 1294, 461, 1865, 2430]. Sequencing [1643]. sequent [192, 542]. sequent-style [192]. Sequential [1079, 2143, 2589]. Sequentiality [142]. Serial [1272, 524, 1293, 1257]. Series [1474, 2120]. Server [1845, 845, 305, 2417]. Server-Applications [2357]. Server-Supported [1845]. Servers [1409].

Service [2470, 2218, 2226, 2210, 1961, 2219, 1846, 113]. Services [2769, 655, 2476, 583, 2225, 2777]. Session [2763]. Set [50, 579, 608, 2618, 694, 1986, 766, 1978, 2496, 1782, 1097, 1364, 2043, 1225, 1228, 2338, 107, 208]. Set-Covering [1782]. Set-Relations [1228]. Set-Theoretic [1225]. Sets [1644, 1476, 1471, 2546, 1467, 2598, 1154, 2359, 547, 1861, 556, 2208]. Setting [1013, 1445]. Settings [1788]. Seven [2508]. Seven-Head [2508]. Several [2275, 1304]. Sextic [835]. SGI [2713]. Shape [1291, 1309, 464, 1797, 1716, 819, 806, 1638, 1287, 1267, 1584, 2517, 1793, 1804, 1796]. Shaped [463]. Shapes [1477, 813, 2519]. Shared [2245, 1024, 2074, 1031, 1447, 2475, 882, 1214, 493, 926, 914, 2594, 2346, 929, 2050, 964, 2752, 2751, 744, 869]. Shared-Bus [1031]. Shared-Memory [1024, 2050, 964, 869]. Sharing [1968, 1604, 1060, 1818, 1654, 2360, 1609, 357, 36, 130, 2262, 2264, 1595, 2420, 2422, 2421]. Shattering [2545]. Sheaf [1551]. Shear [2538]. shelf [1949]. Shell [336, 287]. Shellsort [1454]. Shift [2276, 770]. SHMEM [2066]. Shop [699, 1769, 724]. Short [422, 2733, 2646]. Short-Range [2733]. Shortest [1491, 551, 1468, 2573, 670]. Shostak [195]. Should [875, 351]. Shrink [2386]. Shrinking [908, 2432]. Si [1695]. Si-MOSFETs [1695]. sic [313]. SiCoTHEO [175]. SICSDD [2788]. Side [1088, 351]. Siegenthaler [373]. Sieve [2279, 829, 831, 856]. Sigma [1755, 1760]. Sigma-Delta [1755]. Signal [1344, 1771, 1757, 436, 800, 451, 22]. Signalling [2774]. Signals [1337, 1298, 498]. Signature [1420, 2275, 2247, 1, 2268, 2439]. Signatures [2267, 1600, 1845, 2259, 358, 1941, 2603, 2273, 2246, 2274, 1, 513, 2464, 356, 1373, 1795]. Signed [1140]. Significance [2361, 1682]. Significant [355, 1402]. Silicon [406].

SIMFOT [343]. Similar [793, 75]. Similarities [2361, 499]. Similarity [2365, 1270, 1357, 2382, 1135, 556]. Similitude [2383]. Simone [17]. Simple [1140, 677, 2356, 129, 958, 1364, 2043, 311, 952, 609, 2573, 240, 2145, 670, 1641, 1879, 565, 175]. Simplex [2724]. Simplification [741]. Simplified [1982]. Simplifying [475, 913]. Simply [140]. Simulate [473]. Simulated [1669, 331, 1673, 1464, 2020, 315]. Simulating [1971, 343]. Simulation [1303, 1305, 234, 2093, 1986, 1306, 1472, 997, 280, 557, 2717, 734, 1783, 974, 1307, 1757, 2304, 313, 1042, 1706, 225, 964, 1667, 324, 1001, 111, 66]. Simulation-Based [280]. Simulations [2064, 1007, 1006, 2055, 296, 2754]. Simulator [2718, 1258]. Simulators [963, 1765]. simultaneous [571, 154]. Single [1948, 418, 670, 520, 2184]. Single-Source [670]. Sisal [873]. Sites [1772]. Situation [646, 1378]. Size [1052, 886, 877, 467, 1698, 1461, 1638, 1969, 2187]. Sizes [1653]. Skeleton [1296, 2568]. Skeletonization [808, 791, 2531]. Skeletons [951, 678, 761, 1597]. Skew [1149]. Skill [1539]. Skull [1314]. Slack [703]. Slave [1183, 2065, 2620]. Sleepy [1846]. slender [1866]. SLI [992]. Slicing [396, 1830, 1831]. Slovakia [321]. Slow [1738]. Small [1041, 1463, 2208]. Smallest [962]. Smart [2009]. Smith [839]. SML [1176]. Smooth [1902, 2338, 2184]. Snapshot [1163]. Soccer [627]. SOFM [473]. Soft [1337, 216]. Soft-Output [1337]. Software [1929, 1073, 4, 2468, 2350, 1913, 1956, 1937, 2347, 1912, 2217, 2696, 1536, 1195, 2343, 1910, 2252, 2710, 2345, 1942, 335, 1921, 1723, 1931, 1703, 39, 1932, 6, 1919, 313, 242, 1936, 56, 1926, 1764, 1944, 1920, 1072, 7, 227, 1923, 1777, 1831, 343, 2757, 1030, 315, 1185, 2803, 2424, 2799]. Software-Implemented [1942]. Sokoban [629]. Solid [337, 1885, 1897]. Solomon [1333]. Solution [2401, 2405, 2289, 2406, 2354, 971, 658, 2710, 2496, 2723, 2297, 2746, 2750, 977, 714, 2385, 742, 1668]. Solutions [1670, 307, 697, 1460, 215, 714, 639, 1659, 2086]. Solvable [1984, 2062, 763]. Solve [1966, 692, 2019, 714, 853, 2414]. Solver [968, 1550, 728, 1822, 2721, 712]. Solvers [2088, 2096, 1606]. Solving [1560, 1669, 501, 1550, 260, 765, 699, 916, 2010, 310, 1165, 702, 735, 1508, 705, 2067, 288, 2274, 250, 2381, 747, 1681, 2796, 1546, 717, 1831, 742, 743, 1730, 2644, 259, 746, 2643, 723, 2394, 724]. SOM [448]. Some [2540, 2017, 1679, 778, 279, 1343, 1327, 2498, 2632, 1214, 30, 349, 807, 2257, 1890, 2653, 394, 1061, 1060, 202, 990, 2422]. someTimes [2216]. Sort [1219, 590, 187]. Sorted [1575, 1471, 10, 12, 1562]. Sorts [1556]. Sound [289, 650]. Soundness [1555]. Source [1948, 2773, 670]. SP [2715, 1143]. SP2 [2711, 2232, 657]. Space [1136, 729, 790, 2345, 920, 979, 2262, 1263, 2344, 1936, 925, 419, 508, 1207, 1295, 2534, 90]. Space-Time [920]. Spacecraft [1742]. Spaces [1625, 499, 62]. SPAI [2065]. Spanners [1491]. Spanning [756, 962, 2558]. Sparc [1027]. Spario [518]. Spario-Temporal [518]. Sparse [879, 1644, 677, 660, 2704, 971, 658, 659, 2710, 2711, 652, 839, 2720, 974, 2067, 2055, 2738, 2746]. Sparsely [412]. Spass [162]. Spatial [2784, 1136, 1259, 2697, 1406, 1541, 1285, 458, 26, 1395, 2025, 1363, 1772, 1418, 672, 2119, 1003, 2785]. Spatio [519]. Spatio-Temporal [519]. SPC [1018]. Special [2549, 763, 2001]. Specialisation [386, 387, 1587]. Specialization [378, 380, 2657, 385, 395, 396, 397, 581]. Specializations [601]. Specialized [1696]. Speciation [1615]. Specific [1308, 1587]. Specification [2401, 2281, 2305, 2491, 2402, 1956, 2403, 2328, 40, 2285, 2404, 35, 2405, 2399, 2286, 1566, 24, 1179, 1186, 553, 1220, 41, 1127, 2, 1448, 2292, 2293, 2407, 2614, 1374, 2408, 1451, 2409, 1439, 2411, 2295, 2120, 2297, 5, 2501, 289, 1582, 2329, 1577, 1922, 2299, 32, 47, 2413, 1120, 2301, 2505, 2323, 2229, 55, 2414, 2415, 2302, 53, 1240, 98, 557, 211, 100]. Specification-Based [1956]. Specifications [1085, 1955, 2284, 33, 1218, 1212, 1415, 2293, 1598, 42, 1823, 1178, 12, 14, 1225, 37, 1227, 1230, 1453, 1233, 242, 1132, 2103, 1238, 545, 92, 541]. Specificity [453]. Specified [220]. Specifying [2291, 1445, 2298, 1576, 569, 257, 2228, 2767, 2303]. spect [2531, 2713]. Spectral [2728, 661, 1269, 2527]. Spectroscopy [330]. Spectrum [1330]. SPECWARE [39]. Speech [1502, 1507, 449, 514]. Speed [1155, 2076, 465, 2732, 2311, 2543, 478]. Speed-Up [1155, 2732]. Speeding [695]. Speedup [1052]. Speedups [1059]. Sphere [664]. spherical [2201]. Spike [443, 149]. SPIKE-AC [149]. Spike-Clusters [443]. Spin [2090, 2291, 94]. SPIQ [1910]. Spiral [754]. SPIRIT [1522]. Splines [1278]. Split [750]. Splitting [443, 499]. Spontaneous [474]. Spread [1330]. Spreading [497, 597]. Spreadsheet [329]. SPTHEO [2061]. SQL [1079, 2802]. SQL-DBMS [1079]. Square [2338]. Squares [2713, 2762, 2514]. Squeezing [390]. SR [2300]. SRAM [1721]. SRAM-Based [1721]. SRT [71, 72]. Staalmarck [1099]. Stabilizing [1983, 517]. Stable [2032, 1400, 2034]. Stably [1902]. Stack [884]. Stack-Graph [884]. Stage [1685, 514]. Stages [242]. STAN [1834]. Standard [1939, 2049, 1205, 2246, 1210, 2152, 572]. Standardization [1565, 1564, 2426]. Star [888, 15, 2602, 1440, 1975]. STARBASE [2037]. Stardust [870]. Stark [830]. Starting [758]. State [13, 489, 223, 2493, 1891, 321, 1844, 2152, 2293, 1223, 1351, 2053, 2334, 2592, 1954, 1789, 2058, 62, 94, 2187]. State-of-the-Art [2053]. Statecharts [2286, 2489, 2632, 1427]. Statement [2399, 1154]. Statement-Sets [1154]. statements [557]. States [2170, 1635]. Static [2242, 1834, 1040, 578]. Statically [1066]. Station [22]. Stationary [1329, 986]. Statistical [434, 502, 1870, 1858, 2541]. Status [2049, 2172, 1318, 1892, 403, 2056]. Steady [2080, 1789]. Steam [2281, 2305, 2284, 2285, 2286, 2287, 2288, 2290, 2289, 2291, 2292, 2293, 2294, 2295, 2296, 2297, 2298, 2304, 2299, 2301, 2302, 2303]. Steam-Boiler [2305, 2285, 2286, 2288, 2291, 2293, 2295, 2298, 2304, 2299, 2301, 2303]. Steganography [2449, 2466, 2447]. Steiner [2548, 2567, 961, 755]. Stenography [2446]. Step [1653, 2386, 135, 1709, 101]. Steps [1136]. Stereotaxic [1295]. Stereotaxis [1301]. Stick [669]. Stiff [2753]. still [2179]. Stimulation [1300]. Stochastic [1060, 2096, 1672, 1879, 764, 422, 712, 2591, 1698, 1636, 788, 1946, 1606, 746, 1880, 1881]. Stock [2626]. Stop [1]. Stopping [418]. Storage [410, 374, 527, 1458]. store [575]. straight [2524]. Straightforward [950]. Strategies [1604, 1626, 1169, 2012, 1545, 1594, 2654, 1119, 1713, 1616, 1750, 1142, 290, 1707, 1650, 1778, 1651, 209]. Strategy [2108, 233, 1653, 228, 1355, 675, 409, 1363, 1666, 1691]. Stratified [1940, 186]. Stream [1744]. Streaming [1064]. Streams [2414, 2143]. Streets [1493]. Stretching [2449]. Strict [1569, 2583]. Strictly [886]. String [2472, 2041, 371, 776, 966, 2588]. Strings [2588, 2443]. Strip [1898, 2717]. Strong [1122, 127, 30, 142, 1574]. Strongly [1168]. Strongly-Typed [1168]. Structural [2282, 2365, 1000, 2107, 809, 801, 810, 792, 814, 1947, 1870, 1867]. Structure [404, 1164, 1984, 1489, 1993, 1144, 502, 1103, 1515, 617, 1544, 1190, 1244, 1988, 1338, 471, 1325, 421, 1584, 1858, 2422]. Structure-Driven [1515]. Structured [1812, 673, 2723, 311, 493, 2377, 643, 1774, 941]. Structures [678, 2366, 2795, 439, 460, 667, 2649, 1285, 759, 1012, 1020, 1189, 2677, 1565, 1187, 1255, 643, 2127, 1206, 1166, 1417, 805, 636, 575]. Structuring [167, 2217, 1240]. Stubs [2453]. Student [335, 338, 302, 281]. Studies [1352, 702, 1298]. Study [2281, 2108, 1992, 2131, 501, 825, 2400, 1195, 24, 2568, 1993, 1093, 331, 2096, 1590, 1679, 1693, 466, 2775, 701, 656, 1931, 2340, 1405, 2410, 522, 1142, 2771, 2063, 1676, 627, 896, 2311, 2333, 1287, 1789, 315, 2776, 2221, 113, 2200, 203]. Studying [626]. Stuff [217]. StUSPACE [2561]. style [192]. Sub [1743, 1785]. Sub-Optima [1785]. Subdefinite [1552, 2643]. Subexponential [846]. Subgraph [1490, 2569]. Subjective [450, 1616]. Subliminal [2458, 2463]. Subliminal-Free [2458]. Submachine [1013]. Subpopulations [1692]. Subproblems [2021]. Subprograms [2752]. subsequential [1883, 1862]. Subset [751]. Subspace [441, 484]. Substitution [1604, 127, 1603, 201]. Substitutions [1574, 1336, 128, 546]. Substructure [819]. Subsumption [1416]. Subtractive [2310]. Subtree [1461]. Subtrees [2549]. Subtyping [1832, 542]. successes [151]. Sufficient [892]. Suffix [2588]. SUIF [1318]. suites [91]. Sum [931]. Summary [1895]. Sums [346]. Supercompilation [384, 400, 2655]. Supercompiler [2656, 392]. Supercomputers [2231]. Supercover [2532]. Superposition [117, 123]. Superscalar [1065, 2330, 1066, 1068, 1062]. Superstring [1468]. Support [2491, 1353, 1413, 2364, 317, 2474, 927, 862, 1910, 280, 1539, 1915, 1921, 2165, 1520, 940, 1387, 413, 680, 963, 1917, 2434, 226, 1396, 2680, 666, 110]. Supported [1845, 2355, 322]. Supporting [879, 1322, 2069, 2370, 44, 2672, 676, 293, 2768, 2389, 148]. Supports [585]. Suppressed [2336]. Surface [808, 2534, 2521, 2541]. Surfaces [1252, 2540, 2535, 2533]. Surgery [1305, 1306, 1280, 1308, 1307, 1314, 1302]. Surrogate [425, 2747]. Survey [2002, 2327, 1697, 1319, 2481, 1624]. SVMview [868]. SVR4 [2776]. Swinging [1234]. Switch [1744, 2075]. Switched [904]. Switcherland [1144]. Switching [1748, 1061]. Symbiogenesis [1615]. Symbol [460, 844, 595]. Symbolic [64, 1535, 1542, 62, 25, 1536, 982, 1186, 1191, 266, 102, 2315, 1200, 1189, 1549, 1515, 341, 2605, 1526, 804, 1534, 745, 1547, 2336, 63]. symbols [570]. Symideal [118]. Symmetric [1477, 2722, 972]. symmetries [196]. Symmetry [1000, 2045]. Symmetry-Based [2045]. Synapses [477]. Synchronization [379, 675, 1974, 470, 780, 955, 2087]. Synchronized [952]. Synchronous [1450, 1567, 1038, 1013, 2709, 1008, 1010, 1446, 69]. Syndrome [2278]. Synergistic [2096]. Synergy [2608, 218]. Synfire [478]. Synopsis [2400]. Syntactic [821, 784, 2144, 792, 814, 1532]. Syntax [1234, 1881]. syntax-directed [1881]. Synthesis [1084, 2282, 1087, 1452, 2648, 602, 1095, 1753, 2520, 2327, 1104, 1735, 1728, 1747, 1717, 2196, 104, 1796]. Synthesize [2294]. Synthesized [2329]. Synthesizing [1433, 2631]. Synthetic [1768, 1152, 1108]. System [2789, 725, 1002, 1303, 1962, 1128, 1497, 239, 446, 272, 2492, 316, 149, 52, 1140, 2156, 2285, 2364, 1149, 1751, 469, 317, 2786, 48, 2287, 2474, 928, 921, 1993, 618, 1339, 2164, 995, 41, 1435, 1360, 1945, 1502, 333, 437, 2663, 1978, 1505, 1734, 1075, 2081, 2623, 1165, 51, 1280, 1100, 423, 1507, 867, 1405, 1989, 585, 1689, 1268, 2039, 1395, 2730, 1361, 216, 1964, 1050, 1757, 271, 2298, 2627, 456, 865, 2057, 2165, 1520, 2642, 1312, 2349, 2379, 341, 1720, 1244, 1197, 292, 300, 1524, 293, 2014, 1160, 2060, 1666, 2346, 301, 797, 405, 2796]. System [1716, 342, 2215, 304, 2684, 632, 1722, 981, 922, 2634, 2056, 1402, 237, 2020, 2630, 2342, 1389, 1408, 222, 324, 2486, 2302, 2234, 1384, 2761, 514, 2763, 2303, 218, 244, 507, 184, 584, 1277, 1379, 172, 173, 104, 83, 95, 182, 170, 92, 181, 163]. System-Level [1962, 995]. Systematic [233, 1596, 1019, 229, 1934]. Systematically [1232]. Systemic [292]. Systems [2687, 2769, 1557, 968, 119, 2491, 1968, 1174, 1043, 1077, 2129, 1162, 223, 1957, 2086, 2636, 127, 1450, 1550, 2035, 2648, 1060, 35, 1090, 2597, 1839, 927, 1685, 1220, 2406, 1198, 971, 1843, 997, 658, 1407, 985, 1076, 332, 1678, 1554, 310, 1822, 2455, 31, 990, 334, 2779, 284, 132, 2224, 1382, 2722, 655, 1371, 2614, 385, 1374, 219, 1694, 1451, 2625, 37, 141, 2313, 246, 39, 258, 1226, 604, 1445, 44, 708, 427, 603, 352, 620, 1047, 1141, 1229, 1410, 2223, 926, 1404, 1852, 2616, 634, 2473, 950, 781, 2170, 2210, 291, 1576]. Systems [424, 1943, 1842, 2605, 914, 940, 1152, 215, 1764, 1387, 302, 220, 2662, 2222, 860, 936, 1346, 142, 2174, 2154, 227, 232, 2505, 2214, 137, 1054, 2753, 1658, 654, 2595, 243, 1289, 1736, 1383, 1237, 1143, 245, 2757, 1423, 1969, 2759, 277, 17, 720, 961, 2485, 2175, 126, 226, 2391, 1446, 1399, 1398, 251, 1403, 218, 666, 217, 2512, 101, 98, 97, 553, 69, 2438, 74, 93, 1878, 1442, 2194, 540, 192, 198, 81]. Systolic [1029].

T3D [1055, 2711, 2066]. T4 [2492]. Table [2004, 2507]. Tableau [1170, 183]. tableau-based [183]. Tableaux [1113, 205, 1118, 10, 1556, 611, 2090]. Tabled [1594]. Tables [2783]. Tabu [1773, 1882]. Tackling [1684, 2413]. tactics [206]. Tagging [449]. Tailoring [2427]. Take [465, 421]. Talk [2743]. Tamper [2468]. Tap [2264]. TARCA [2164]. Target [1736, 2640]. Task [1035, 942, 1895, 1045, 1437, 875, 1321, 2582, 1050, 781, 973]. Taskers [2069]. tasking [2755]. Tasks [1984, 466, 656, 932, 2376, 1520, 2579, 2394]. Taxonomic [1499]. Taxonomies [652]. TEA [316]. Teaching [310, 306, 339, 294, 314, 342, 1534, 325, 282, 2637]. Team [606]. Teamwork [1500, 2230]. Technical [2011, 1511, 2376]. Technique [1349, 460, 2716, 909, 842]. Techniques [1913, 692, 2396, 2451, 2315, 2788, 2335, 2679, 133, 1427, 1674, 1675, 394, 230, 1304, 2655, 1747, 2110, 1396, 89, 71, 91]. Technological [2644, 2221]. Technologies [1754, 2700, 2486, 282]. Technology [2176, 965, 1908, 332, 321, 270, 2172, 216, 1718, 2416, 1936, 56, 215, 222, 2668, 1756, 217, 2803, 181]. Tectal [469]. Telecommunication [655, 237, 584]. Telematic [239]. Telephone [1762, 113]. Template [1270, 1863]. Tempo [2610]. Temporal [404, 2789, 1353, 2404, 1441, 613, 730, 86, 2495, 2168, 2790, 1134, 2407, 2409, 2120, 408, 1135, 479, 1776, 477, 519, 518, 1401, 209, 538, 92, 204]. Temporally [2377]. Tense [647]. Term [1174, 2703, 1183, 2597, 652, 1599, 141, 1226, 44, 131, 134, 433, 1173, 142, 126]. Terminal [140]. Terminating [132, 957, 576]. Termination [1814, 119, 1817, 2152, 121, 2658, 159, 131, 1229, 2025, 28, 1824, 1562, 160, 137, 189]. Terminological [1531]. Terminology [2471, 589]. Terms [1821, 1364, 2043, 621, 122, 1573, 622, 1110, 153]. TERSE [44]. Tesselation [917]. Test [579, 1692, 2072, 2315, 707, 1517, 1230, 242, 1777, 1954, 91, 2443]. testable [1874]. Testbed [309, 1737]. Testing [1955, 1956, 1949, 286, 1759, 1601, 1953, 1416, 17, 66, 563, 214]. Tests [340, 2190]. Text [815, 1507, 640, 817, 2144]. Text-to-Speech [1507]. Textile [1482]. Texts [2646, 776]. Textual [597, 2039, 98]. Textural [2085]. Texture [444]. Thai [449]. Their [367, 2266, 2667, 2677, 2262, 2464, 2459, 1807]. Thematic [638]. Theorem [1086, 1553, 1497, 2308, 1535, 1167, 116, 1172, 2133, 1499, 1500, 190, 42, 150, 1227, 134, 188, 117, 2061, 1564, 1548, 183, 71, 156, 151, 78, 159, 211, 158, 181, 175, 152, 166]. Theorem-proving [1553]. Theorems [2326, 1521, 88]. Theoretic [1225, 2477, 749, 2263, 784]. Theoretical [1033, 621, 622, 711, 2803]. Theories [1553, 623, 1216, 2318, 149, 1213, 2149, 1125, 719, 2601, 195, 179, 187, 212]. Theory [1443, 2283, 1956, 1136, 18, 1963, 1126, 2168, 2137, 1212, 58, 646, 1097, 1931, 1056, 2139, 2232, 2615, 252, 1138, 216, 1105, 2141, 372, 537, 1233, 2151, 1234, 1494, 395, 847, 1108, 2181, 243, 2146, 135, 2392, 587, 1112, 198]. therapy [2209]. Thermo [2754]. Thermo-Convective [2754]. Thermodynamical [1664, 1665]. Thin [1278]. Thin-Plate [1278]. Think [740]. Thinking [2097, 585]. Thoughts [2632]. Thread [265]. threaded [2584]. Threads [1205, 1051]. Threat [1852]. Three [976, 2019, 1685, 1820, 1998, 1286, 763, 1255, 811, 1716, 1298, 1294, 718, 1289, 1281, 514, 2516]. Three-Dimensional [763, 1255, 1716, 1294, 1289, 2516]. Three-Stage [1685, 514]. Threshold [235, 2603, 1456, 349, 780, 1950]. Throughput [1060]. Thue [837, 137, 853]. Ticket [2256]. Tigger [2474]. Tight [2071, 2235]. Tight-Binding [2071]. tightly [2799]. tile [556]. Time [2089, 823, 1035, 23, 2259, 2648, 1437, 862, 380, 1306, 2211, 2290, 1438, 1447, 1435, 1444, 1356, 1330, 1134, 1965, 1448, 2663, 2775, 2235, 1823, 2657, 920, 1252, 753, 2563, 2615, 993, 979, 794, 1445, 2499, 1357, 1050, 2120, 1006, 1975, 424, 531, 1442, 1621, 1429, 898, 1944, 2004, 1021, 2387, 1457, 274, 1408, 1398, 2303, 84, 66, 101, 549, 551, 98, 2192, 177, 108, 538, 2595, 81, 577]. time-abstracting [81]. Time-Processor [1006]. Time-Variant [424]. Time-Varying [1330]. Timebase [1329]. Timed [25, 1441, 1430, 1431, 553, 81]. Times [1036]. Timestamp [1983, 2596]. Timestamp-Based [2596]. Timetable [2019, 2016]. Timetables [2006]. Timetabling [2000, 2002, 2013, 1999, 2005, 2012, 2010, 2008, 2007, 2003, 2014, 2015, 2009, 2018, 2020, 2001]. Timing [352]. TINA [2218, 2219]. TINA-C [2218]. Tioga [2685]. Tioga-2 [2685]. Tissue [1247]. TkGofer [57]. TLA [2401, 2297]. TLSI [336]. TLT [2289]. TM [1249]. TMN [655]. TMN-Based [655]. Token [2571]. Tolerance [1939, 2490]. Tolerant [1948, 1938, 1961, 922]. Tomography [1304]. Too [299]. Tool [868, 1452, 2698, 864, 696, 1092, 875, 2356, 311, 1056, 1451, 1453, 2676, 1922, 1810, 1727, 238, 342, 1613, 2331, 1240, 244, 111]. Tool-Maker [1810]. toolbox [106]. TOOLBUS [23, 45]. Toolkit [1416]. Tools [2048, 40, 1217, 2337, 1741, 1914, 2626, 1319, 47, 1528, 639, 46, 148, 105, 300]. toolset [98, 108]. Top [2774, 2409, 993]. Topics [294]. Topographic [917]. Topological [530, 435, 2525, 1977, 2785, 2778, 2540, 2527]. Topologically [2533, 2525]. Topologies [1597]. Topology [882, 1549, 1690, 2530, 1794]. tori [574]. Toroidal [775]. Torsion [827, 832]. Torus [904, 2571]. Total [900]. Totally [1505]. Tournament [758, 1631, 1785]. Toyama [134]. TPascal [942]. Trace [2718, 2410, 88, 864]. Trace-Driven [2718]. Traceability [228]. Traced [2450]. Tracer [1380]. traces [535, 538]. Tracing [1903, 20]. Tracking [457, 1676, 1652, 2541]. Tract [1241]. Tractability [707]. Tractable [709, 1828]. Traction [1945]. Trade [374]. Trade-offs [374]. Trader [2226, 2225]. Traditional [784, 1046]. Traffic [894, 2364, 914, 2046, 1704, 170]. Trail [614]. Train [696]. Trainer [236]. Training [420, 307, 279, 1301, 418, 342, 517, 324, 491]. Traitors [2450]. Trajectories [2725]. Trans [1463]. Trans-dichotomous [1463]. Transaction [1076, 1382, 2223, 1398, 2512]. Transaction-Based [1382]. Transactions [1356, 2030, 87]. transcendental [2425]. transducer [1883]. transducers [1862]. Transfer [1744, 2220, 2378, 371, 559]. Transferable [2260]. Transform [983, 1146, 632, 1807]. Transformability [2274]. Transformation [1543, 2078, 1029, 909, 44, 1983, 2656, 394, 2580, 1314, 2167]. Transformational [1104]. Transformations [2670, 2207, 2116, 1359, 1589, 1223, 706, 2334, 810, 1288, 229, 2651, 86]. Transformer [22]. Transforming [1079, 189, 2118, 251]. Transforms [1340]. Transient [1757]. Transistor [1698]. Transition [1090, 2406, 2026, 2018, 93]. Transitions [733, 2656, 477, 400, 90]. Transitive [1529]. Translating [1085, 612]. Translation [1452, 1217, 2367, 2098, 950, 1881, 92]. Translations [2554, 191, 2198]. Transmission [280]. Transmitted [1337]. Transport [2765]. Traveling [1702]. Traversal [1245]. Treatment [1309, 2164, 1577, 2174]. Treatments [1303]. Tree [1492, 2487, 2548, 464, 1628, 463, 2567, 1262, 1037, 2740, 1244, 2570, 804, 2504, 2664, 1417, 2794]. Tree-Based [2740, 1417]. Tree-Like [1262]. Trees [887, 756, 1114, 1490, 901, 2498, 124, 681, 769, 1487, 899, 389, 424, 962, 516, 1461, 2111, 2558, 2601, 755, 202, 201, 1858, 567, 2190, 179, 2562]. Trellis [1335]. Trellises [1334]. Trend [1745]. Trends [2487, 2483, 2481, 2056, 2513]. Trials [2, 2450]. Triangle [1907]. triangles [574]. Triangular [1074]. Triangulations [1996, 2569]. Triangulator [1907]. Trianus [1734]. Tridiagonal [2722, 972]. trie [575]. tries [575]. trigger [1864]. TRIO [2292, 5]. Triple [2559, 961, 364]. Triple-DES [364]. Triple-Free [2559]. Trivial [832]. Troubleshooting [2397]. TruCluster [1158]. True [997]. Truncation [1642]. Trust [351]. trusted [2439]. Trustees [1839]. Truth [1221, 649, 258, 2004]. TSP [2096]. TUDER [341]. Tumours [1294]. Tuning [2687, 868, 2626, 230, 869]. Tunnel [2321]. Tuple [925]. Tupling [1610, 564]. Turbo [1332]. TurboBrain [235]. Turbulence [974]. Tutor [314]. Tutorial [2325, 2708, 2324, 1203]. Tutoring [327, 332, 287, 341, 2170, 292, 300, 2175]. Tutorized [331, 315]. Two [786, 1948, 2130, 1476, 693, 2703, 1063, 2550, 2368, 2554, 2, 525, 2372, 1770, 1835, 2733, 448, 1855, 1857, 1538, 2311, 476, 451, 223, 2192, 75, 2190]. two-counter [2192]. Two-Dimensional [1948, 2550, 2311, 476]. Two-Level [2130, 1063, 448]. Two-Term [2703]. Two-valued [1538]. Type [1600, 1955, 2129, 381, 1590, 2137, 140, 466, 14, 2139, 386, 2264, 1215, 2141, 2055, 2257, 1376, 1233, 1576, 846, 1236, 1108, 2021, 2146]. Type-Directed [381, 1590]. Typed [1168, 140, 2599, 1581]. TYPELAB [53]. Types [1575, 1443, 1819, 2136, 2138, 1223, 1234, 955, 1552, 1832, 123, 160, 542, 578]. Typicality [2392].

UCFG [2566]. Ultimate [2611]. Ultra [1563]. Ultra-properties [1563]. Ultrasound [1266, 1313]. Ultrasound-Guided [1313]. Unassembled [971]. Unbalanced [1011]. Unbounded [2551, 2204]. Uncertain [617, 2024]. Uncertainty [2023, 1526, 586, 1531]. unconditional [559]. Unconditionally [374]. Uncoordinated [1959]. Undecidable [135]. Undeniable [2464, 2439]. Undercut [1899]. Undercut-Free [1899]. Underground [370]. underlying [2427]. Understand [337]. Understandable [2113]. Understanding [377, 1816, 1986, 1502, 2372, 2101, 2180, 2646, 1523, 2115, 2173]. Unfair [1990]. Unfold [1589]. Unfold/Fold [1589]. Unification [116, 1116, 1569, 140, 1364, 2043, 178, 177, 1213, 139, 2141, 2504, 136, 485, 154, 187].

Unified [2472, 1754, 1963, 1466, 909, 2027, 617, 586, 2319]. Unifiers [1114]. Uniform [1038, 1074, 662, 380, 1118, 1187, 1525, 1571, 575, 2579]. Uniformity [1829]. Unifying [407, 1216, 1602, 1425]. Uninterpreted [2320]. Union [2496]. Unions [1860, 2551, 719, 2204]. Unit [2307, 1036]. United [450]. Unity [1101, 2415]. Universal [345, 336, 346, 859, 2645, 2611, 369, 1721, 1683, 2443]. Universally [2257]. Universities [2002]. University [2000, 2013, 2005, 2006, 2007, 2009, 2011]. Unix [2776]. UNLESS [1135]. unlikely [560]. Unraveling [1241]. Unravelings [1563]. Unreliable [1971]. Unsatisfiability [1456]. Unsorted [1562]. Unstructured [663, 2080, 657]. Unsupervised [607, 484, 1269]. Unsymmetric [2485]. Untyped [1833]. Update [2028, 591]. updated [182]. Updates [1537, 2794, 1960]. Updating [1057, 1139, 1277]. upon [529]. UPPAAL [82]. Upper [1482, 742]. Urban [1677]. Use [1329, 1349, 2308, 2689, 2116, 1644, 1309, 2096, 644, 2351, 1394, 710, 405, 936, 283, 1789, 315, 199]. Used [2019, 787, 2580, 1667, 2331]. User [2282, 875, 1741, 1502, 1202, 2667, 336, 2113, 583]. Using [2769, 2089, 802, 820, 2283, 641, 2091, 4, 446, 1585, 1644, 726, 692, 1272, 1604, 1024, 1413, 1669, 2403, 2217, 445, 1939, 2083, 2325, 678, 501, 1550, 579, 260, 2795, 1305, 2613, 1091, 1270, 2227, 2312, 1767, 1684, 2093, 450, 687, 1986, 2307, 2107, 827, 2356, 2706, 1898, 1285, 971, 985, 1499, 2451, 967, 1726, 1755, 1973, 2790, 382, 1330, 91, 1029, 2682, 1119, 1942, 1504, 1556, 2315, 2079, 2081, 2006, 2077, 1280, 2397, 1351, 2008, 2294, 2320, 2408, 974, 2409, 1509, 1617, 2661, 1689, 812, 1771, 258, 2375, 1308, 2679, 6, 1515, 1105, 2007, 1964, 1104, 453]. Using [2591, 1735, 2080, 1724, 449, 1570, 19, 634, 2738, 483, 614, 49, 781, 1460, 1244, 1232, 433, 969, 242, 2240, 596, 1524, 1883, 531, 1905, 2746, 1414, 1166, 2060, 303, 599, 1826, 1996, 503, 47, 1774, 2004, 1278, 1622, 2445, 1346, 1946, 2301, 498, 1707, 1304, 2414, 1606, 1314, 1393, 1143, 598, 1288, 2792, 1003, 986, 2084, 440, 1872, 2302, 2631, 431, 2392, 1446, 1947, 1532, 1139, 506, 723, 1277, 2321, 2221, 1295, 2064, 64, 2442, 2433, 82, 62, 71, 1865, 1882, 78, 271, 1858, 539, 1794, 2157, 2188]. Utility [736, 2388]. Utilization [1078]. Utilize [2274].

Vacillatory [2206]. Validation [1700, 1324, 21, 1308, 2229, 1831, 2272, 1281, 106]. Validity [2318, 600, 554]. Value [2089, 2044, 2213, 1432, 191]. Valued [249, 2618, 172, 1538, 210]. Variability [1944]. Variable [888, 1447, 1081, 10, 1556, 701, 620, 991, 1652, 2596, 1870, 208, 2184, 2190, 2581]. Variable-Dilation [888]. variable-length [1870]. Variables [703, 19, 747, 715]. Variance [1653]. Variant [424]. Variants [1339, 2341]. Variation [462]. variations [1799]. Varieties [822]. Variety [833, 1017]. Various [1769, 809]. Varying [1330]. Vaudenay [373]. VDM [1085, 2488, 2295, 2301]. VDM-SL [1085]. Vector [1483, 2618, 660, 815, 430, 2265, 413, 961, 2521]. Vector-Based [815]. Vectorization [2716, 1156]. Vehicles [631]. Velocity [423]. Ventricle [1287]. Ventricles [1286]. Ventricular [1272]. Verifiable [2257]. Verification [269, 1452, 82, 2328, 2648, 2613, 2312, 113, 24, 2307, 1093, 2317, 20, 75, 2138, 2320, 2408, 2409, 2615, 1439, 1436, 44, 2410, 110, 93, 22, 2314, 49, 2310, 2329, 1106, 1442, 2143, 1107, 2309, 2311, 2319, 1425, 2332, 1111, 2103, 2302, 53, 1446, 2331, 2321, 111, 105, 80, 101, 62, 104, 553, 86, 95, 69, 106, 91, 169, 99, 87, 77, 72, 70]. Verification-Driven [1436]. Verified [1435]. verifiers [560]. Verify [1100]. Verifying [2308, 84, 1450, 71, 2291, 112, 78, 85, 74, 2313, 2303, 83, 73, 569]. Version [2703, 2339, 2344, 2653, 277, 2338, 183, 162].

Versioned [2343]. Versions [2128, 2360, 2790, 1855]. Versus [282, 2565, 2243, 1555, 541, 2600]. Vertex [758, 2559]. Vertical [1420, 2537]. Very [448, 2760]. Vessel [1248]. Vessels [1256]. VH [2302]. VHDL [1452, 2328, 1753, 1757, 2329]. via [824, 1816, 1472, 556, 1335, 1343, 1138, 1345, 2444, 396, 2348, 1623, 504]. Video [980, 1757, 2219, 2777, 1277, 1808]. VIENA [1396]. View [1358, 2472, 445, 1415, 1224, 2452, 2628, 32, 713, 1776, 2121, 2111, 2261, 1934, 1723]. View-Based [445, 1934]. View-Centered [2121]. viewed [556]. Viewing [923, 408]. Viewpoint [1221]. Viewpoints [620]. Views [1350, 607, 2674, 520, 2509, 2433]. Virtual [1948, 1751, 2702, 1296, 1308, 436, 1890, 914, 1825, 1301, 1145, 2082, 1304, 1722, 1289, 1302]. Virtualized [1803, 1244]. Viruses [1614]. VIS [2322, 104]. Viscous [2080]. Visibility [1905]. Visible [1349, 1252, 2262, 1249, 1242, 1250]. Vision [446, 982, 2179, 466, 811, 405, 1792, 1790]. VISLAN [1277]. Visual [469, 2488, 375, 2409, 1246, 509, 44, 2679, 2262, 456, 2501, 238, 408, 2445, 1613, 1803]. Visualisation [1299, 313]. Visualising [1292]. Visualization [2685, 1300, 2483, 2675, 861, 2677, 2679, 1153, 2626, 2673, 1892, 2674, 1247, 2676, 1249, 2681, 1242, 1156, 1288, 2686, 2058, 2680]. Visualized [1253]. Visualizing [1291, 1891, 499, 1245, 1453]. Viterbi [786, 788]. VLSI [667, 671, 1029, 1333, 1697, 241, 1951, 1698, 1699, 244]. VLSI-Layout [1699]. VME [1761]. Voltage [324]. Volume [808, 1246, 2520, 977, 1242, 1304, 1295, 1251]. Volumes [761, 1245]. Volumetric [1306, 1241]. VOMDraw [238]. Voting [2257, 1950]. Vouchers [2605]. voxels [2522]. VRMLise [2622]. vs [2242, 1884, 1723, 2336].

W [1106]. Wait [1984]. Wait-Free [1984]. Waiting [2235]. Waiting-Time [2235]. Waldmeister [1172]. Walking [1180, 475]. Walther [207]. WAM [1107]. Wanted [1116]. Warping [1284]. WASMII [1722]. Water [328]. Watermark [2460]. Watermarking [2462]. Wave [2000]. Wavelength [691, 1487]. Wavelength-Routed [691]. Wavelengths [2577]. Wavelet [1340]. Wavelets [2690, 1341]. Waves [474]. Way [1562, 805, 857]. Weak [1604, 1364, 1432, 1387, 1025]. Weakly [995, 767]. Weaknesses [349, 2580]. Web [1915, 1193, 628, 845]. Weekly [2010]. Weight [494, 527, 1838, 419, 2505, 517, 2612, 2569, 1766]. Weighted [260, 812, 711, 1457, 755]. Weighting [2380]. Weightless [497]. Weights [524, 493, 2393]. Well [2284, 339, 2149, 2781]. Well-Behaved [2149]. Well-Founded [2781]. Well-Grounded [339]. WFS [1131]. Where [830]. Which [585, 814, 1574, 2755]. White [1253]. Whole [1666]. Wide [2347, 2279, 1915, 1193, 845, 1926, 628]. Wide-Area [2347, 1926]. Widgets [1602]. width [2185]. Windows [331, 313]. Windows95 [2082]. Wing [1675]. Wing-Box [1675]. Winner [465]. Winner-Take-All [465]. Wire [1264, 2264, 813]. Wire-Shapes [813]. Wire-Tap [2264]. Wireless [2083, 2627]. Wise [1640]. Within [1502, 2561, 1313, 2225, 2767]. Without [1948, 359, 2246, 1924, 36, 2439, 408, 2418]. Word [2307, 534, 500, 288, 461, 1864]. Word-Level [2307]. WORDMATH [288]. Work [2345, 304, 814]. Workbench [317, 96]. Worker [924]. Worker-Based [924]. Workflow [2126, 1354, 1382, 1926, 2214, 1383, 2103]. Workflows [1840]. working [1381]. Workload [2232, 2741, 2240, 1152]. Workshop [1790]. Workspaces [2355, 1185]. Workstation [1048, 2076, 978, 2238, 2059, 963, 2054, 2092]. Workstationcluster [2084]. Workstations [870, 2053, 925, 922]. World [2279, 2368, 620, 405, 1915, 1193, 628, 845]. World-Wide [628]. Worm [903]. Worm-Hole [903]. Worst [773]. Worst-Case [773]. WPRAM [1009]. Write [1004]. Writers [295]. Writing [387]. Written [593, 801]. WS [2303]. WWW [2348, 305].

X.509 [2621]. XC6000 [1750]. XC6200 [1751, 1740, 1724]. Xilinx [1751, 1724, 1727]. XMOISE [329]. XRay [181]. XVERSA [98].

Year [1998, 2022]. Years [2632]. Yes [299]. Yield [1951].

Z [2286, 1103, 2228]. Z0n [2526]. Zero [2336]. Zero-Suppressed [2336]. Zigzag [771]. Ziv [1481].

References

Pfitzmann:1996:DSS

[1] B. Pfitzmann. Digital signature schemes: General framework and failstop signatures. Lecture Notes in Computer Science, 1100:xvi + 396, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fitzgerald:1996:TIT

[2] J. S. Fitzgerald. Two industrial trials of formal specification. Lecture Notes in Computer Science, 1101:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenBrand:1996:IAA

[3] M. Van den Brand, A. Van Deursen, P. Klint, and S. Klusener. Industrial applications of ASF+SDF. Lecture Notes in Computer Science, 1101:9–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnold:1996:ESE

[4] A. Arnold, D. Begay, and J.-P Radoux. The embedded software of an electricity meter: An experience in using formal methods in an industrial project. Lecture Notes in Computer Science, 1101:19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mandrioli:1996:ARR

[5] D. Mandrioli. Applying research results in the industrial environment: The case of the TRIO specification language. Lecture Notes in Computer Science, 1101:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kraemer:1996:UHF

[6] B. J. Kraemer. Using heterogeneous formal methods in distributed software engineering education. Lecture Notes in Computer Science, 1101:43–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sankar:1996:IFM

[7] S. Sankar. Introducing formal methods to software engineers through OMG’s CORBA environment and interface definition language. Lecture Notes in Computer Science, 1101:52–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:TCA

[8] D. R. Smith. Toward a classification approach to design. Lecture Notes in Computer Science, 1101:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howe:1996:SFE

[9] D. J. Howe. Semantic foundations for embedding HOL in Nuprl. Lecture Notes in Computer Science, 1101:85– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavilanes:1996:FVT

[10] A. Gavilanes, J. Leach, and S. Nieva. Free variable tableaux for a many sorted logic with preorders. Lecture Notes in Computer Science, 1101:102– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1996:AIM

[11] D. Kapur and M. Subramaniam. Automating induction over mutually recursive functions. Lecture Notes in Computer Science, 1101:117–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haxthausen:1996:POS

[12] A. E. Haxthausen and F. Nickl. Pushouts of order-sorted algebraic specifications. Lecture Notes in Computer Science, 1101:132–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ancona:1996:FFM

[13] D. Ancona and E. Zucca. A formal framework for modules with state. Lecture Notes in Computer Science, 1101: 148–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennicker:1996:OOI

[14] R. Hennicker and C. Schmitz. Objectoriented implementation of abstract data type specifications. Lecture Notes in Computer Science, 1101:163–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fokkink:1996:CEK

[15] W. Fokkink. On the completeness of the equations for the Kleene star in bisimulation. Lecture Notes in Computer Science, 1101:180–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aceto:1996:EAO

[16] L. Aceto and A. Ingolfsdottir. An equational axiomatization of observation congruence for prefix iteration. Lecture Notes in Computer Science, 1101: 195–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ulidowski:1996:FAS

[17] I. Ulidowski. Finite axiom systems for testing preorder and De Simone process languages. Lecture Notes in Computer Science, 1101:210–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berry:1996:CSE

[18] G. Berry. Constructive semantics of Esterel: From theory to practice. Lecture Notes in Computer Science, 1101: 225–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcus:1996:UGV

[19] M. Marcus and A. Pnueli. Using ghost variables to prove refinement. Lecture Notes in Computer Science, 1101:226– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraer:1996:TOV

[20] R. Fraer. Tracing the origins of verification conditions. Lecture Notes in Computer Science, 1101:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gribomont:1996:PIV

[21] E. P. Gribomont. Preprocessing for invariant validation. Lecture Notes in Computer Science, 1101:256–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeBorgne:1996:FVS

[22] M. Le Borgne, H. Marchand, E. Rutten, and M. Samaan. Formal verification of SIGNAL programs: Application to a power transformer station controller. Lecture Notes in Computer Science, 1101:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergstra:1996:DTT

[23] J. A. Bergstra and P. Klint. The discrete time TOOLBUS. Lecture Notes in Computer Science, 1101:286–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:SSV

[24] X. J. Chen, F. Corradini, and R. Gorrieri. A study on the specification and verification of performance properties. Lecture Notes in Computer Science, 1101:306–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boreale:1996:SBT

[25] M. Boreale. Symbolic bisimulation for timed processes. Lecture Notes in Computer Science, 1101:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isobe:1996:AAP

[26] Y. Isobe, Y. Sato, and K. Ohmaki. Approximative analysis by process algebra with graded spatial actions. Lecture Notes in Computer Science, 1101: 336–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hehner:1996:BFE

[27] E. C. R. Hehner. Boolean formalism and explanations. Lecture Notes in Computer Science, 1101:351–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchiori:1996:PET

[28] M. Marchiori. Proving existential termination of normal logic programs.

Lecture Notes in Computer Science, 1101:375–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harland:1996:PLO

[29] J. Harland, D. Pym, and M. Winikoff. Programming in Lygon: An overview. Lecture Notes in Computer Science, 1101:391–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

KrishnaRao:1996:SCS

[30] M. R. K. Krishna Rao. Some characteristics of strong innermost normalization. Lecture Notes in Computer Science, 1101:406–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiadeiro:1996:EPC

[31] J. L. Fiadeiro. On the emergence of properties in component-based systems. Lecture Notes in Computer Science, 1101:421–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paech:1996:AVS

[32] B. Paech. Algebraic view specification. Lecture Notes in Computer Science, 1101:444–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernot:1996:THF

[33] G. Bernot, S. Coudert, and P. Le Gall. Towards heterogeneous formal specifications. Lecture Notes in Computer Science, 1101:458–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baier:1996:CCC

[34] C. Baier and M. Majster-Cederbaum. A categorical characterization of consistency results. Lecture Notes in Computer Science, 1101:473–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1996:ASR

[35] M. Broy. Algebraic specification of reactive systems. Lecture Notes in Computer Science, 1101:487–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosu:1996:MMP

[36] R. Grosu and K. Stoelen. A model for mobile point-to-point data-flow networks without channel sharing. Lecture Notes in Computer Science, 1101: 504–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1996:CSM

[37] B. Jacobs. Coalgebraic specifications and models of deterministic hybrid systems. Lecture Notes in Computer Science, 1101:520–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groote:1996:BRP

[38] J. F. Groote and J. Van de Pol. A bounded retransmission protocol for large data packets. Lecture Notes in Computer Science, 1101:536–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juellig:1996:SAE

[39] R. Juellig, Y. Srinivas, and J. Liu. SPECWARE: An advanced environment for the formal development of complex software systems. Lecture Notes in Computer Science, 1101:551– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1996:AIS

[40] M. Bidoit, C. Choppy, and F. Voisin. ASSPEGIQUE0+: An integrated specification environment providing interoperability of tools. Lecture Notes in Computer Science, 1101:555–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Didrich:1996:TIA

[41] K. Didrich, C. Gerke, W. Grieskamp, and C. Maeder. Towards integrating algebraic specification and functional programming: the Opal system. Lecture Notes in Computer Science, 1101:559–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geisler:1996:IIT

[42] R. Geisler, M. Klar, and F. Cornelius. InterACT: An interactive theorem prover for algebraic specifications. Lecture Notes in Computer Science, 1101:563–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voisin:1996:NPMa

[43] F. Voisin. A new proof-manager and graphic interface for the Larch prover. Lecture Notes in Computer Science, 1101:567–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawaguchi:1996:TVE

[44] N. Kawaguchi, T. Sakabe, and Y. Inagaki. TERSE: a visual environment for supporting analysis, verification and transformation of term rewriting systems. Lecture Notes in Computer Science, 1101:571–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klint:1996:TCA

[45] P. Klint and P. Olivier. The TOOLBUS coordination architecture: a demonstration. Lecture Notes in Computer Science, 1101:575–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDeursen:1996:AAS

[46] A. Van Deursen and P. D. Mosses. ASD: The action semantic description tools. Lecture Notes in Computer Science, 1101:579–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poetzsch-Heffter:1996:UOE

[47] A. Poetzsch-Heffter. Using occurrence and evolving algebras for the specification of language-based programming tools. Lecture Notes in Computer Science, 1101:583–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buckland:1996:ESM

[48] R. Buckland and M. Johnson. ECHIDNA: a system for manipulating explicit choice higher dimensional automata.

Lecture Notes in Computer Science, 1101:587–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melzer:1996:VUP

[49] S. Melzer, S. Roemer, and J. Esparza. Verification using PEP. Lecture Notes in Computer Science, 1101: 591–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouali:1996:FSa

[50] A. Bouali, A. Ressouche, V. Roy, and R. De Simone. The FC2TOOLS set. Lecture Notes in Computer Science, 1101:595–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harland:1996:PLS

[51] J. Harland, D. Pym, and M. Winikoff. Programming in Lygon: a system demonstration. Lecture Notes in Computer Science, 1101:599–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertot:1996:CSPa

[52] J. Bertot and Y. Bertot. CtCoq: a system presentation. Lecture Notes in Computer Science, 1101:600–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonHenke:1996:TSV

[53] F. W. Von Henke, M. Luther, H. Pfeiffer, and H. Ruess. The TYPELAB specification and verification environment. Lecture Notes in Computer Science, 1101:604–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steffen:1996:IF

[54] B. Steffen, T. Margaria, A. Classen, and V. Braun. Incremental formalization. Lecture Notes in Computer Science, 1101:608–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Souquieres:1996:PSD

[55] J. Souquieres and N. Levy. Proplane: a specification development environment. Lecture Notes in Computer Science, 1101:612–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parot:1996:LBT

[56] P. Parot. A logic-based technology to mechanize software components reuse. Lecture Notes in Computer Science, 1101:616–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulte:1996:TFG

[57] W. Schulte, T. Schwinn, and T. Vullinghs. TkGofer: a functional GUI library. Lecture Notes in Computer Science, 1101:621–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erdmann:1996:ACL

[58] S. Erdmann and I. Classen. Alpha — A class library for a metamodel based on algebraic graph theory. Lecture Notes in Computer Science, 1101:625–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bert:1996:RGF

[59] D. Bert, R. Echahed, and K. Adi. Resolution of goals with the functional and logic programming language LPG: Impact of abstract interpretation. Lecture Notes in Computer Science, 1101:629–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buendgen:1996:CRC

[60] R. Buendgen and W. Lauterbach. Combining reductions and computations in ReDuX. Lecture Notes in Computer Science, 1101:633–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Limet:1996:CDN

[61] S. Limet and P. Rety. Conditional directed narrowing. Lecture Notes in Computer Science, 1101:637–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boigelot:1996:SVC

[62] B. Boigelot and P. Godefroid. Symbolic verification of communication protocols with infinite state spaces using QDDs. Lecture Notes in Computer Science, 1102:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McMillan:1996:CDB

[63] K. L. McMillan. A conjunctively decomposed Boolean representation for symbolic model checking. Lecture Notes in Computer Science, 1102:13– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avrunin:1996:SMC

[64] G. S. Avrunin. Symbolic model checking using algebraic geometry. Lecture Notes in Computer Science, 1102:26– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pistore:1996:PRA

[65] M. Pistore and D. Sangiorgi. A partition refinement algorithm for the λ(?)- calculus. Lecture Notes in Computer Science, 1102:38–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baier:1996:PTA

[66] C. Baier. Polynomial-time algorithms for testing probabilistic bisimulation and simulation. Lecture Notes in Computer Science, 1102:50–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walukiewicz:1996:PPG

[67] I. Walukiewicz. Pushdown processes: games and model checking. Lecture Notes in Computer Science, 1102:62– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kupferman:1996:MC

[68] O. Kupferman and M. Y. Vardi. Module checking. Lecture Notes in Computer Science, 1102:75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1996:AVP

[69] E. A. Emerson and K. S. Namjoshi. Automatic verification of parameterized synchronous systems. Lecture Notes in Computer Science, 1102:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shukla:1996:HMC

[70] S. K. Shukla, H. B. Hunt, and D. J. Rosenkrantz. HORNSAT, model checking, verification, and games. Lecture Notes in Computer Science, 1102: 99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1996:VSD

[71] E. M. Clarke, S. M. German, and X. Zhao. Verifying the SRT division algorithm using theorem proving techniques. Lecture Notes in Computer Science, 1102:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruess:1996:MVS

[72] H. Ruess, N. Shankar, and M. K. Srivas. Modular verification of SRT division. Lecture Notes in Computer Science, 1102:123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1996:MVF

[73] D. Kapur and M. Subramaniam. Mechanically verifying a family of multiplier circuits. Lecture Notes in Computer Science, 1102:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ip:1996:VSR

[74] C. N. Ip and D. L. Dill. Verifying systems with replicated components in Murcursive phi. Lecture Notes in Computer Science, 1102:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1996:VAC

[75] M. Fujita. Verification of arithmetic circuits by comparing two similar circuits. Lecture Notes in Computer Science, 1102:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rushby:1996:ADF

[76] J. M. Rushby. Automated deduction and formal methods. Lecture Notes in Computer Science, 1102:169–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pnueli:1996:PCD

[77] A. Pnueli and E. Shahar. A platform for combining deductive with algorithmic verification. Lecture Notes in Computer Science, 1102:184–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1996:VIU

[78] S. Graf and H. Saidi. Verifying invariants using theorem proving. Lecture Notes in Computer Science, 1102: 196–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sipma:1996:DMC

[79] H. B. Sipma, T. E. Uribe, and Z. Manna. Deductive model checking. Lecture Notes in Computer Science, 1102:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berregeb:1996:AVI

[80] N. Berregeb, A. Bouhoula, and M. Rusinowitch. Automated verification by induction with associativecommutative operators. Lecture Notes in Computer Science, 1102:220–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tripakis:1996:ATS

[81] S. Tripakis and S. Yovine. Analysis of timed systems based on timeabstracting bisimulations. Lecture Notes in Computer Science, 1102:232– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bengtsson:1996:VAP

[82] J. Bengtsson, W. O. D. Griffioen, K. J. Kristoffersen, and K. G. Larsen. Verification of an audio protocol with bus collision using UPPAAL. Lecture Notes in Computer Science, 1102:244– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campos:1996:SQA

[83] S. Campos and O. Grumberg. Selective quantitative analysis and interval model checking: verifying different facets of a system. Lecture Notes in Computer Science, 1102:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aziz:1996:VCT

[84] A. Aziz, K. Sanwal, V. Singhal, and R. K. Brayton. Verifying continuoustime Markov chains. Lecture Notes in Computer Science, 1102:269–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenstreet:1996:VSP

[85] M. R. Greenstreet. Verifying safety properties of differential equations. Lecture Notes in Computer Science, 1102:277–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAlfaro:1996:TVD

[86] L. De Alfaro and Z. Manna. Temporal verification by diagram transformations. Lecture Notes in Computer Science, 1102:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1996:PVA

[87] S. Park and D. L. Dill. Protocol verification by aggregation of distributed transactions. Lecture Notes in Computer Science, 1102:300–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gribomont:1996:ART

[88] E. P. Gribomont. Atomicity refinement and trace reduction theorems. Lecture Notes in Computer Science, 1102:311– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bensalem:1996:PTA

[89] S. Bensalem, Y. Lakhnech, and H. Saidi. Powerful techniques for the automatic generation of invariants. Lecture Notes in Computer Science, 1102:323–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1996:SSF

[90] H. Miller and S. Katz. Saving space by fully exploiting invisible transitions. Lecture Notes in Computer Science, 1102:336–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1996:UFV

[91] J.-C. Fernandez, C. Jard, T. Jeron, and G. Viho. Using on-the-fly verification techniques for the generation of test suites. Lecture Notes in Computer Science, 1102:348–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nelken:1996:ATN

[92] R. Nelken and N. Francez. Automatic translation of natural-language system specifications into temporal logic. Lecture Notes in Computer Science, 1102: 360–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kupferman:1996:VFT

[93] O. Kupferman and M. Y. Vardi. Verification of fair transition systems. Lecture Notes in Computer Science, 1102: 372–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzmann:1996:SS

[94] G. J. Holzmann and D. Peled. The state of SPIN. Lecture Notes in Computer Science, 1102:385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dill:1996:MPV

[95] D. L. Dill. The Murcursive phi verification system. Lecture Notes in Computer Science, 1102:390–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1996:NCW

[96] R. Cleaveland and S. Sims. The NCSU concurrency workbench. Lecture Notes in Computer Science, 1102: 394–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1996:CFD

[97] R. Cleaveland, P. M. Lewis, S. A. Smolka, and O. Sokolsky. The concurrency factory: a development environment for concurrent systems. Lecture Notes in Computer Science, 1102:398– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1996:XIG

[98] D. Clarke, H. Ben-Abdallah, I. Lee, and H.-L. Xie. XVERSA: an integrated graphical and textual toolset for the specification and analysis of resource-bound real-time systems. Lecture Notes in Computer Science, 1102: 402–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merino:1996:EIF

[99] P. Merino and J. M. Troya. EVP: integration of FDTs for the analysis and verification of communication protocols. Lecture Notes in Computer Science, 1102:406–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Owre:1996:PCS

[100] S. Owre, S. Rajah, J. M. Rushby, and N. Shankar. PVS: combining specification, proof checking, and model checking. Lecture Notes in Computer Science, 1102:411–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjoerner:1996:SDA

[101] N. Bjoerner, A. Browne, E. Chang, and M. Colon. STeP: deductive-algorithmic verification of reactive and real-time systems. Lecture Notes in Computer Science, 1102:415–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1996:SMC

[102] E. M. Clarke, K. L. McMillan, S. Campos, and V. Hartonas-Garmhausen. Symbolic model checking. Lecture Notes in Computer Science, 1102:419– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardin:1996:C

[103] R. H. Hardin, Z. Har’El, and R. P. Kurshan. COSPAN. Lecture Notes in Computer Science, 1102:423–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brayton:1996:VSV

[104] R. K. Brayton, G. D. Hachtel, A. Sangiovanni-Vincentelli, and F. Somenzi. VIS: a system for verification and synthesis. Lecture Notes in Computer Science, 1102:428–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anon:1996:MTV

[105] K. D. Anon, N. Boulerice, E. Cerny, and F. Corella. MDG tools for the verification of RTL designs. Lecture Notes in Computer Science, 1102:433– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1996:CPV

[106] J.-C. Fernandez, H. Garavel, A. Kerbrat, and L. Mounier. CADP: a protocol validation and verification toolbox. Lecture Notes in Computer Science, 1102:437–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouali:1996:FSb

[107] A. Bouali, A. Ressouche, V. Roy, and R. De Simone. The FC2Tools set. Lecture Notes in Computer Science, 1102:441–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moser:1996:RTG

[108] L. E. Moser, P. M. Melliar-Smith, Y. S. Ramakrishna, and G. Kutty. The real-time graphical interval logic toolset. Lecture Notes in Computer Science, 1102:446–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steffen:1996:ME

[109] B. Steffen, T. Margaria, A. Classen, and V. Braun. The METAFrame ’95 environment. Lecture Notes in Computer Science, 1102:450–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koob:1996:VSE

[110] F. Koob, M. Ullmann, and S. Wittmann. Verification support environment. Lecture Notes in Computer Science, 1102: 454–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambroise:1996:MTS

[111] D. Ambroise and B. Rozoy. MAR-RELLA: a tool for simulation and verification. Lecture Notes in Computer Science, 1102:458–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonthier:1996:VSP

[112] G. Gonthier. Verifying the safety of a practical concurrent garbage collector. Lecture Notes in Computer Science, 1102:462–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capellmann:1996:VBA

[113] C. Capellmann, R. Demant, F. Fatahi-Vanani, and R. Galvez-Estrada. Verification by behavior abstraction: a case study of service interaction detection in intelligent telephone networks. Lecture Notes in Computer Science, 1102:466–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1996:RBA

[114] D. Kapur. Rewrite-based automated reasoning: Challenges ahead. Lecture Notes in Computer Science, 1103:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirchner:1996:FGC

[115] C. Kirchner, C. Lynch, and C. Scharff. Fine-grained concurrent completion. Lecture Notes in Computer Science, 1103:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudet:1996:ACU

[116] A. Boudet, E. Contejean, and C. Marche. AC-complete unification and its application to theorem proving. Lecture Notes in Computer Science, 1103:18–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stuber:1996:STP

[117] J. Stuber. Superposition theorem proving for Abelian groups represented as integer modules. Lecture Notes in Computer Science, 1103:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goebel:1996:SGB

[118] M. Goebel. Symideal Gröbner bases. Lecture Notes in Computer Science, 1103:48–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arts:1996:TCS

[119] T. Arts and J. Giesl. Termination of constructor systems. Lecture Notes in Computer Science, 1103:63–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1996:DEE

[120] M. C. F. Ferreira. Dummy elimination in equational rewriting. Lecture Notes in Computer Science, 1103:78– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramlich:1996:PTI

[121] B. Gramlich. On proving termination by innermost termination. Lecture Notes in Computer Science, 1103: 93–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jouannaud:1996:RPO

[122] J.-P. Jouannaud and A. Rubio. A recursive path ordering for higherorder terms in eta-Long — normal form. Lecture Notes in Computer Science, 1103:108–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Virga:1996:HOS

[123] R. Virga. Higher-order superposition for dependent types. Lecture Notes in Computer Science, 1103:123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1996:HON

[124] M. Hanus and C. Prehofer. Higherorder narrowing with definitional trees. Lecture Notes in Computer Science, 1103:138–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huet:1996:DPA

[125] G. Huet. Design of a proof assistant. Lecture Notes in Computer Science, 1103:153–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vittek:1996:CNT

[126] M. Vittek. A compiler for nondeterministic term rewriting systems. Lecture Notes in Computer Science, 1103: 154–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloo:1996:CRS

[127] R. Bloo and K. H. Rose. Combinatory reduction systems with explicit substitution that preserve strong normalisation. Lecture Notes in Computer Science, 1103:169–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesner:1996:CPE

[128] D. Kesner. Confluence properties of extensional and non-extensional lambdacalculi with explicit substitutions. Lecture Notes in Computer Science, 1103: 184–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCosmo:1996:PSD

[129] R. Di Cosmo. On the power of simple diagrams. Lecture Notes in Computer Science, 1103:200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerrini:1996:CSP

[130] S. Guerrini, S. Martini, and A. Masini. Coherence for sharing proof nets. Lecture Notes in Computer Science, 1103: 215–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

KrishnaRao:1996:MTT

[131] M. R. K. Krishna Rao. Modularity of termination in term graph rewriting. Lecture Notes in Computer Science, 1103:230–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramlich:1996:CTC

[132] B. Gramlich and C.-P. Wirth. Confluence of terminating conditional rewrite systems revisited. Lecture Notes in Computer Science, 1103:245–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madlener:1996:ART

[133] K. Madlener. Applications of rewrite techniques in monoids and rings. Lecture Notes in Computer Science, 1103: 260–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lueth:1996:CTR

[134] C. Lueth. Compositional term rewriting: An algebraic proof of Toyama’s theorem. Lecture Notes in Computer Science, 1103:261–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treinen:1996:FOT

[135] R. Treinen. The first-order theory of one-step rewriting is undecidable. Lecture Notes in Computer Science, 1103: 276–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt-Schauss:1996:ADU

[136] M. Schmidt-Schauss. An algorithm for distributive unification. Lecture Notes in Computer Science, 1103:287– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Senizergues:1996:TPO

[137] G. Senizergues. On the termination problem for one-rule semi-Thue systems. Lecture Notes in Computer Science, 1103:302–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curien:1996:ESO

[138] R. Curien, Z. Qian, and H. Shi. Efficient second-order matching. Lecture Notes in Computer Science, 1103:317– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy:1996:LSO

[139] J. Levy. Linear second-order unification. Lecture Notes in Computer Science, 1103:332–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fettig:1996:UHO

[140] R. Fettig and B. Loechner. Unification of higher-order patterns in a simply typed lambda-calculus with finite products and terminal type. Lecture Notes in Computer Science, 1103:347– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquemard:1996:DAT

[141] F. Jacquemard. Decidable approximations of term rewriting systems. Lecture Notes in Computer Science, 1103: 362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakai:1996:SSS

[142] M. Sakai and Y. Toyama. Semantics and strong sequentiality of priority term rewriting systems. Lecture Notes in Computer Science, 1103:377– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanOostrom:1996:HOF

[143] V. Van Oostrom. Higher-order families. Lecture Notes in Computer Science, 1103:392–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voisin:1996:NPMb

[144] F. Voisin. A new proof manager and graphic interface for the Larch prover. Lecture Notes in Computer Science, 1103:408–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buendgen:1996:RNF

[145] R. Buendgen, C. Sinz, and J. Walter. ReDuX 1.5: New facets of rewriting. Lecture Notes in Computer Science, 1103:412–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Contejean:1996:CCM

[146] E. Contejean and C. Marche. CiME: Completion modulo E0*. Lecture Notes in Computer Science, 1103:416–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandevoorde:1996:DLP

[147] M. T. Vandevoorde and D. Kapur. Distributed Larch prover (DLP): An experiment in parallelizing a rewriterule based prover. Lecture Notes in Computer Science, 1103:420–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walters:1996:EEL

[148] H. R. Walters and J. F. T. Kamperman. EPIC: An equational language — abstract machine and supporting tools. Lecture Notes in Computer Science, 1103:424–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berregeb:1996:SAS

[149] N. Berregeb, A. Bouhoula, and M. Rusinowitch. SPIKE-AC: a system for proofs by induction in associativecommutative theories. Lecture Notes in Computer Science, 1103:428–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hillenbrand:1996:GEC

[150] T. Hillenbrand, A. Buch, and R. Fettig. On gaining efficiency in completionbased theorem proving. Lecture Notes in Computer Science, 1103:432–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganzinger:1996:SBT

[151] H. Ganzinger. Saturation-based theorem proving: Past successes and future potential. Lecture Notes in Computer Science, 1104:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tammet:1996:RTP

[152] T. Tammet. A resolution theorem prover for intuitionistic logic. Lecture Notes in Computer Science, 1104:2–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pym:1996:PTC

[153] D. Pym, E. Ritter, and L. Wallen. Proof-terms for classical and intuitionistic resolution. Lecture Notes in Computer Science, 1104:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voronkov:1996:PSI

[154] A. Voronkov. Proof-search in intuitionistic logic with equality, or back to simultaneous rigid E-unification. Lecture Notes in Computer Science, 1104:32– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ireland:1996:EGC

[155] A. Ireland and A. Bundy. Extensions to a generalization critic for inductive proof. Lecture Notes in Computer Science, 1104:47–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denzinger:1996:LDK

[156] J. Denzinger and S. Schulz. Learning domain knowledge to improve theorem proving. Lecture Notes in Computer Science, 1104:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Protzen:1996:PFC

[157] M. Protzen. Patching faulty conjectures. Lecture Notes in Computer Science, 1104:77–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melis:1996:IAT

[158] E. Melis and J. Whittle. Internal analogy in theorem proving. Lecture Notes in Computer Science, 1104:92– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolbe:1996:TTP

[159] T. Kolbe and C. Walther. Termination of theorem proving by reuse. Lecture Notes in Computer Science, 1104: 106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sengler:1996:TAN

[160] C. Sengler. Termination of algorithms over non-freely generated data types. Lecture Notes in Computer Science, 1104:121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giunchiglia:1996:APC

[161] F. Giunchiglia and A. Villafiorita. Absfol: a proof checker with abstraction. Lecture Notes in Computer Science, 1104:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weidenbach:1996:SFV

[162] C. Weidenbach, B. Gaede, and G. Rock. Spass and Flotter version 0.42. Lecture Notes in Computer Science, 1104:141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suttner:1996:DCA

[163] C. Suttner and G. Sutcliffe. The design of the CADE-13 ATP system competition. Lecture Notes in Computer Science, 1104:146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

JuergenOhlbach:1996:SEP

[164] H. Juergen Ohlbach. Scan — elimination of predicate quantifiers. Lecture Notes in Computer Science, 1104:161– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlbach:1996:SEP

[165] H. Juergen Ohlbach. Scan — elimination of predicate quantifiers. Lecture Notes in Computer Science, 1104:161– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:GGT

[166] D. Wang. Geother: a geometry theorem prover. Lecture Notes in Computer Science, 1104:166–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1996:SMI

[167] D. Basin and S. Matthews. Structuring metatheory on inductive definitions. Lecture Notes in Computer Science, 1104:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rasmussen:1996:ERI

[168] O. Rasmussen. An embedding of Ruby in Isabelle. Lecture Notes in Computer Science, 1104:186–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Homeier:1996:MVM

[169] P. V. Homeier and D. F. Martin. Mechanical verification of mutually recursive procedures. Lecture Notes in Computer Science, 1104:201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felici:1996:FDT

[170] G. Felici, G. Rinaldi, and K. Truemper. FasTraC: a decentralized traffic control system based on logic programming. Lecture Notes in Computer Science, 1104:216–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1996:PMF

[171] X. Huang and A. Fiedler. Presenting machine-found proofs. Lecture Notes in Computer Science, 1104:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1996:MTE

[172] M. Baaz, C. G. Fermueller, G. Salzer, and R. Zach. MUltlog 1.0: towards an expert system for many-valued logics. Lecture Notes in Computer Science, 1104:226–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertot:1996:CSPb

[173] J. Bertot and Y. Bertot. CtCoq: a system presentation. Lecture Notes in Computer Science, 1104:231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chou:1996:IGE

[174] S.-C. Chou, X.-S. Gao, and J.-Z. Zhang. An introduction to geometry expert. Lecture Notes in Computer Science, 1104:235–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schumann:1996:SSC

[175] J. Schumann. SiCoTHEO: simple competitive parallel theorem provers. Lecture Notes in Computer Science, 1104: 240–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scott:1996:WCW

[176] D. S. Scott. What can we hope to achieve from automated deduction? Lecture Notes in Computer Science, 1104:245–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermann:1996:UAC

[177] M. Hermann and P. G. Kolaitis. Unification algorithms cannot be combined in polynomial time. Lecture Notes in Computer Science, 1104:246–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guo:1996:UMM

[178] Q. Guo, P. Narendran, and D. A. Wolfram. Unification and matching modulo nilpotence. Lecture Notes in Computer Science, 1104:261–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vorobyov:1996:ILB

[179] S. Vorobyov. An improved lower bound for the elementary theories of trees. Lecture Notes in Computer Science, 1104:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dieter:1996:ING

[180] H. Dieter and C. Sengler. INKA: the next generation. Lecture Notes in Computer Science, 1104:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaub:1996:XPT

[181] T. Schaub, S. Bruening, and P. Nicolas. XRay: a Prolog technology theorem prover for default reasoning: a system description. Lecture Notes in Computer Science, 1104:293–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farmer:1996:IUS

[182] W. M. Farmer, J. D. Guttman, and F. J. Thayer Fabrega. IMPS: an updated system description. Lecture Notes in Computer Science, 1104:298– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1996:TBT

[183] B. Beckert, R. Haehnle, P. Oel, and M. Sulzmann. The tableau-based theorem prover 3T0AP, version 4.0. Lecture Notes in Computer Science, 1104: 303–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1996:SDG

[184] J. Zhang and H. Zhang. System description: generating models by SEM. Lecture Notes in Computer Science, 1104:308–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1996:OPS

[185] J. Harrison. Optimizing proof search in model elimination. Lecture Notes in Computer Science, 1104:313–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sagonas:1996:AMF

[186] K. Sagonas, T. Swift, and D. S. Warren. An abstract machine for fixed-order dynamically stratified programs. Lecture Notes in Computer Science, 1104:328–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weidenbach:1996:UPL

[187] C. Weidenbach. Unification in pseudolinear sort theories is decidable. Lecture Notes in Computer Science, 1104: 343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1996:TPG

[188] U. Martin. Theorem proving with group presentations: examples and questions. Lecture Notes in Computer Science, 1104:358–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Middeldorp:1996:TTS

[189] A. Middeldorp, H. Ohsaki, and H. Zantema. Transforming termination by self-labelling. Lecture Notes in Computer Science, 1104:373–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganzinger:1996:TPC

[190] H. Ganzinger and U. Waldmann. Theorem proving in cancellative Abelian monoids (extended abstract). Lecture Notes in Computer Science, 1104:388– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1996:PVD

[191] U. Egly and T. Rath. On the practical value of different definitional translations to normal form. Lecture Notes in Computer Science, 1104:403–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitt:1996:CNC

[192] S. Schmitt and C. Kreitz. Converting non-classical matrix proofs into sequent-style systems. Lecture Notes in Computer Science, 1104:418–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuetz:1996:EMG

[193] H. Schuetz and T. Geisler. Efficient model generation through compilation. Lecture Notes in Computer Science, 1104:433–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linton:1996:AAD

[194] S. Linton, U. Martin, P. Proehle, and D. Shand. Algebra and automated deduction. Lecture Notes in Computer Science, 1104:448–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cyrluk:1996:SDP

[195] D. Cyrluk, P. Lincoln, and N. Shankar. On Shostak’s decision procedure for combinations of theories. Lecture Notes in Computer Science, 1104:463–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BoydelaTour:1996:GRG

[196] T. Boy de la Tour. Ground resolution with group computations on semantic symmetries. Lecture Notes in Computer Science, 1104:478–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roussel:1996:NMK

[197] O. Roussel and P. Mathieu. A new method for knowledge compilation: the achievement by cycle search. Lecture Notes in Computer Science, 1104:493– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snyder:1996:RSP

[198] W. Snyder and J. G. Schmolze. Rewrite semantics for production rule systems: theory and applications. Lecture Notes in Computer Science, 1104:508–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1996:EHU

[199] M. Fuchs. Experiments in the heuristic use of past proof experience. Lecture Notes in Computer Science, 1104: 523–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1996:LDA

[200] D. Kapur and M. Subramaniam. Lemma discovery in automating induction. Lecture Notes in Computer Science, 1104:538–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1996:AIO

[201] P. Graf and C. Meyer. Advanced indexing operations on substitution trees. Lecture Notes in Computer Science, 1104:553–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fermueller:1996:STR

[202] C. G. Fermueller. Semantic trees revisited: some new completeness results. Lecture Notes in Computer Science, 1104:568–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giunchiglia:1996:BDP

[203] F. Giunchiglia and R. Sebastiani. Building decision procedures for modal logics from propositional decision procedures — the case study of modal K. Lecture Notes in Computer Science, 1104:583–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nonnengart:1996:RBC

[204] A. Nonnengart. Resolution-based calculi for modal and temporal logics. Lecture Notes in Computer Science, 1104: 598–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGiacomo:1996:TAP

[205] G. De Giacomo and F. Massacci. Tableaux and algorithms for propositional dynamic logic with converse. Lecture Notes in Computer Science, 1104:613–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruess:1996:RFT

[206] H. Ruess. Reflection of formal tactics in a deductive reflection framework. Lecture Notes in Computer Science, 1104:628–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McAllester:1996:WR

[207] D. McAllester and K. Arkoudas. Walther recursion. Lecture Notes in Computer Science, 1104:643–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felty:1996:PSS

[208] A. Felty. Proof search with set variable instantiation in the calculus of constructions. Lecture Notes in Computer Science, 1104:658–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dixon:1996:SSR

[209] C. Dixon. Search strategies for resolution in temporal logics. Lecture Notes in Computer Science, 1104:673– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salzer:1996:OAM

[210] G. Salzer. Optimal axiomatizations for multiple-valued operators and quantifiers based on semi-lattices. Lecture Notes in Computer Science, 1104:688– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luz-Filho:1996:GSC

[211] S. F. Luz-Filho. Grammar specification in categorical logics and theorem proving. Lecture Notes in Computer Science, 1104:703–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1996:PIA

[212] P. Graf. Path indexing for ACtheories. Lecture Notes in Computer Science, 1104:718–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1996:MCR

[213] T. Nipkow. More Church–Rosser proofs (in Isabelle/HOL). Lecture Notes in Computer Science, 1104:733–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1996:PMS

[214] Tai Joon Park and A. Van Gelder. Partitioning methods for satisfiability testing on large formulas. Lecture Notes in Computer Science, 1104:748–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pichler:1996:SSS

[215] F. Pichler, H. Schwaertzel, and R. Moreno-Diaz. Systems science and systems technology: From conceptual framework to applicable solutions. Lecture Notes in Computer Science, 1105: 3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klir:1996:SCA

[216] G. J. Klir. Soft computer-aided system theory and technology (SCAST). Lecture Notes in Computer Science, 1105: 13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeigler:1996:FSC

[217] B. P. Zeigler. Fundamental systems concepts: “The Right Stuff” for 21st Century technology. Lecture Notes in Computer Science, 1105:28–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wymore:1996:SDS

[218] A. W. Wymore. SYNERGY: The design of a systems engineering system, I. Lecture Notes in Computer Science, 1105:34–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1996:CCS

[219] G. Hill. The configuration of complex systems. Lecture Notes in Computer Science, 1105:46–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Praehofer:1996:EDE

[220] H. Praehofer and B. P. Zeigler. On the expressibility of discrete event specified systems. Lecture Notes in Computer Science, 1105:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joslyn:1996:OOA

[221] C. Joslyn. An object-oriented architecture for possibilistic models. Lecture Notes in Computer Science, 1105: 80–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuerksen:1996:FES

[222] I. B. Tuerksen. Fuzzy expert system technology. Lecture Notes in Computer Science, 1105:95–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benslimane:1996:DBS

[223] A. Benslimane. Deciding boundedness for systems of two linear communicating finite state machines. Lecture Notes in Computer Science, 1105:108– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomiyama:1996:FKI

[224] T. Tomiyama, Y. Umeda, and T. Kidyama. A framework for knowledge intensive engineering. Lecture Notes in Computer Science, 1105:123– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oeren:1996:MKB

[225] T. I. Oeren and N. Ghasem-Aghaee. Multiparadigm (knowledge-based and numerical) continuous simulation environments: Architectural issues. Lecture Notes in Computer Science, 1105: 151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voss:1996:DMS

[226] M. Voss and G. Schweizer. A development methodology for systems engineering of computer-based systems and its environmental support. Lecture Notes in Computer Science, 1105: 161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaffer:1996:ADC

[227] C. Schaffer. An approach to the design of complex, heterogeneous hardware/ software systems. Lecture Notes in Computer Science, 1105:176–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corriveau:1996:SRT

[228] J.-P. Corriveau and C. Hayashi. A strategy for realizing traceability in an object-oriented design environment. Lecture Notes in Computer Science, 1105:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toledo:1996:TCM

[229] O. B. Toledo, J. C. Q. Losada, R. Moreno-Diaz, and S. C. Sola. Towards a CAST method for a systematic generation of non-orthogonal complete transformations. Lecture Notes in Computer Science, 1105:205–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1996:TFL

[230] M. Santos, S. Dormido, A. P. De Madrid, and F. Morilla. Tuning fuzzy logic controllers by classical techniques. Lecture Notes in Computer Science, 1105:214–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ural:1996:CAD

[231] H. Ural and H. Zeng. Computer aided design of protocol converters. Lecture Notes in Computer Science, 1105:225– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitter:1996:MAE

[232] E. Schmitter. Modelling, analysis and evaluation of systems architectures. Lecture Notes in Computer Science, 1105:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blasco:1996:SSP

[233] R. Blasco. Systematic strategy for performance prediction in improvement of parallel programs. Lecture Notes in Computer Science, 1105:252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bustio:1996:DIM

[234] F. Bustio, P. Corcuera, and E. Mora. Design and implementation of multimedia environment for simulation. Lecture Notes in Computer Science, 1105: 268–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1996:TNN

[235] D. M. Dubois, G. Resconi, and A. Raymondi. TurboBrain: a neural network with direct learning based on linear or non-linear threshold logics. Lecture Notes in Computer Science, 1105:278– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duro:1996:GER

[236] R. J. Duro, J. Santos, and A. Sarmiento. GENIAL: An evolutionary recurrent neural network designer and trainer. Lecture Notes in Computer Science, 1105:295–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanir:1996:DES

[237] O. Tanir, V. K. Agarwal, and P. C. P. Bhatt. DASE: An environment for system level telecommunication design exploration and modelling. Lecture Notes in Computer Science, 1105:302– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oberlik:1996:VTV

[238] G. Oberlik, S. Preishuber, A. M. Tjoa, and R. Wagner. VOMDraw — A tool for visual object modeling. Lecture Notes in Computer Science, 1105:319–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arcieri:1996:CAS

[239] F. Arcieri, M. Fossa, and E. Nardelli. A computer aided system for developing graphical telematic applications. Lecture Notes in Computer Science, 1105: 334–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salviano:1996:SAI

[240] C. F. Salviano, M. A. Nascimento, and M. M. Tanik. A simple approach to improve the abstraction level of object representation. Lecture Notes in Computer Science, 1105:349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller-Wipperfuerth:1996:ICF

[241] T. Mueller-Wipperfuerth. On the integration of CAST.FSM into the VLSI design process. Lecture Notes in Computer Science, 1105:363–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nascimento:1996:ULP

[242] M. A. Nascimento, C. F. Salviano, and M. M. Tanik. Using logic programming to test module specifications in early stages of software development. Lecture Notes in Computer Science, 1105:373–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahara:1996:STS

[243] Y. Takahara and N. Shiba. Systems theory and systems implementation: Case of DSS. Lecture Notes in Computer Science, 1105:388–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeh:1996:CCA

[244] G. K. Yeh, J. B. Burr, K. K. Bagchi, and A. M. Peterson. CAS-CADE: a computer-aided tool for lowenergy high-performance Multi-DSP VLSI system design. Lecture Notes in Computer Science, 1105:409–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thizy:1996:ICP

[245] J.-M. Thizy. Illustrating constraint programming systems in logistic planning. Lecture Notes in Computer Science, 1105:423–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jampel:1996:BOC

[246] M. Jampel. A brief overview of overconstrained systems. Lecture Notes in Computer Science, 1106:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borning:1996:CH

[247] A. Borning, B. Freeman-Benson, and M. Wilson. Constraint hierarchies. Lecture Notes in Computer Science, 1106: 23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freuder:1996:PCS

[248] E. Freuder and R. Wallace. Partial constraint satisfaction. Lecture Notes in Computer Science, 1106:63–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bistarelli:1996:SBC

[249] S. Bistarelli, H. Fargier, U. Montanari, and F. Rossi. Semiring-based CSPs and valued CSPs: Basic properties and comparison. Lecture Notes in Computer Science, 1106:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menezes:1996:DCS

[250] F. Menezes and P. Barahona. Defeasible constraint solving. Lecture Notes in Computer Science, 1106:151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1996:TOC

[251] A. Wolf. Transforming ordered constraint hierarchies into ordinary constraint systems. Lecture Notes in Computer Science, 1106:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jampel:1996:CTC

[252] M. Jampel. A compositional theory of constraint hierarchies. Lecture Notes in Computer Science, 1106:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1996:HMC

[253] R. Wallace and E. Freuder. Heuristic methods for over-constrained constraint satisfaction problems. Lecture Notes in Computer Science, 1106:207– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1996:CDA

[254] R. Wallace. Cascaded directed arc consistency and no-good learning for the maximal constraint satisfaction problem. Lecture Notes in Computer Science, 1106:217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dendris:1996:PAC

[255] N. Dendris, L. Kirousis, Y. Stamatiou, and D. Thilikos. Partial arc consistency. Lecture Notes in Computer Science, 1106:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelle:1996:DCS

[256] E. Gelle and I. Smith. Dynamic constraint satisfaction with conflict management in design. Lecture Notes in Computer Science, 1106:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sattar:1996:SCP

[257] A. Sattar, A. Ghose, and R. Goebel. Specifying over-constrained problems in default logic. Lecture Notes in Computer Science, 1106:253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jussien:1996:ICR

[258] N. Jussien and P. Boizumault. Implementing constraint relaxation over finite domains using assumption-based truth maintenance systems. Lecture Notes in Computer Science, 1106:265– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weissschnur:1996:ESC

[259] R. Weissschnur, J. Hertzberg, and H. W. Guesgen. Experiences in solving constraint relaxation networks with Boltzmann machines. Lecture Notes in Computer Science, 1106:281–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouquet:1996:SCC

[260] F. Bouquet and P. Jegou. Solving over-constrained CSPs using weighted OBDDs. Lecture Notes in Computer Science, 1106:293–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konaka:1996:DPP

[261] H. Konaka, T. Tomokiyo, M. Maeda, and Y. Ishikawa. Data parallel programming in the parallel objectoriented language OCore. Lecture Notes in Computer Science, 1107:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guidec:1996:PMP

[262] F. Guidec and J.-M. Jezequel. Polymorphic matrices in Paladin. Lecture Notes in Computer Science, 1107:18– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baba:1996:PDM

[263] T. Baba, T. Yoshinaga, and T. Furuta. Programming and debugging for massive parallelism: The case for a parallel object-oriented language A-NETL. Lecture Notes in Computer Science, 1107:38–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taura:1996:SCO

[264] K. Taura and A. Yonezawa. Schematic: a concurrent object-oriented extension to Scheme. Lecture Notes in Computer Science, 1107:59–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geib:1996:TOO

[265] J.-M. Geib, C. Gransart, C. Grenot, and P. Merle. (thread and object)- oriented distributed programming. Lecture Notes in Computer Science, 1107: 83–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1996:DOO

[266] K. L. Clark and F. G. McCabe. Distributed and object oriented symbolic programming in April. Lecture Notes in Computer Science, 1107:104–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caromel:1996:RPE

[267] D. Caromel and Y. Roudier. Reactive programming in Eiffel/. Lecture Notes in Computer Science, 1107:125– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galmiche:1996:PCO

[268] D. Galmiche and E. Boudinet. Proofs, concurrent objects, and computations in a FILL framework. Lecture Notes in Computer Science, 1107:148–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bahsoun:1996:MDV

[269] J.-P. Bahsoun, S. Merz, and C. Servieres. Modular description and verification of concurrent objects. Lecture Notes in Computer Science, 1107:168–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquemot:1996:CCC

[270] C. Jacquemot, O. S. Jensen, and S. Carrez. Chorus/Cool: Chorus object oriented technology. Lecture Notes in Computer Science, 1107:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lea:1996:AOS

[271] R. Lea, Y. Yokote, and J.-I Itoh. Adaptive operating system design using reflection. Lecture Notes in Computer Science, 1107:205–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banatre:1996:ICD

[272] M. Banatre, Y. Belhamissi, V. Issarny, and I. Puaut. Isatis: a customizable distributed object-based runtime system. Lecture Notes in Computer Science, 1107:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerraoui:1996:LDI

[273] R. Guerraoui, B. Garbinato, and K. Mazouni. Lessons from designing and implementing GARF. Lecture Notes in Computer Science, 1107:238– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takashio:1996:DID

[274] K. Takashio, H. Shitomi, and M. Tokoro. Design and implementation of DROL runtime environment on real-time Mach kernel. Lecture Notes in Computer Science, 1107:257–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darche:1996:AAM

[275] P. Darche, P.-G. Raverdy, and E. Commelin. ActNet: The Actor model applied to mobile robotic environments. Lecture Notes in Computer Science, 1107:273–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellissard:1996:CBP

[276] L. Bellissard, S. B. Atallah, A. Kerbrat, and M. Riveill. Component-based programming and application management with Olan. Lecture Notes in Computer Science, 1107:290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsukamoto:1996:VMA

[277] M. Tsukamoto, Y. Hamazaki, T. Nishioka, and H. Otokawa. The version management architecture of an object-oriented distributed systems environment: OZ++. Lecture Notes in Computer Science, 1107:310–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tahara:1996:FSA

[278] Y. Tahara, F. Kumeno, A. Ohsuga, and S. Honiden. Formal semantics of agent evolution in language flage. Lecture Notes in Computer Science, 1107: 329–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frasson:1996:SCI

[279] C. Frasson. Some characteristics of instructional design for industrial training. Lecture Notes in Computer Science, 1108:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeJong:1996:SSB

[280] T. De Jong, H. Haertel, J. Swaak, and W. Van Joolingen. Support for simulation-based learning: The effect of assignments in learning about transmission lines. Lecture Notes in Computer Science, 1108:9–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Self:1996:DSM

[281] J. Self. Deconstructionist student models in the computer-based learning of science. Lecture Notes in Computer Science, 1108:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vivet:1996:EET

[282] M. Vivet. Evaluating educational technologies: Evaluation of teaching material versus evaluation of learning. Lecture Notes in Computer Science, 1108: 37–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vanneste:1996:UIE

[283] P. Vanneste, B. De Decker, M. Chrzanowski, and K. Warkentyne. The use of Internet in education. Lecture Notes in Computer Science, 1108:39–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giroux:1996:EAS

[284] S. Giroux, G. Paquette, and J. Girard. Epiphyte advisor systems for collaborative learning. Lecture Notes in Computer Science, 1108:42–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gouarderes:1996:EAL

[285] E. Gouarderes and S. Millet. Expertext approach and learning environments. Lecture Notes in Computer Science, 1108:51–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1996:CBA

[286] S. X. Huang. On content-balanced adaptive testing. Lecture Notes in Computer Science, 1108:60–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jerinic:1996:OOS

[287] L. Jerinic and V. Devedzic. An objectoriented shell for intelligent tutoring lessons. Lecture Notes in Computer Science, 1108:69–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Looi:1996:WCB

[288] C.-K. Looi and B. T. Tan. WORD-MATH: a computer-based environment for learning word problem solving. Lecture Notes in Computer Science, 1108: 78–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marquesuzaa:1996:TPS

[289] C. Marquesuzaa, J. Meyranx, and T. Nodenot. Towards pedagogically sound learning environments: The specification process. Lecture Notes in Computer Science, 1108:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mengelle:1996:MAA

[290] T. Mengelle and C. Frasson. A multiagent architecture for an ITS with multiple strategies. Lecture Notes in Computer Science, 1108:96–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nawarecki:1996:DCL

[291] E. Nawarecki and G. Dobrowolski. Decentralized computer learning systems based on autonomous agent approach. Lecture Notes in Computer Science, 1108:105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

NeiraAlvarez:1996:FAT

[292] A. Neira Alvarez and J. A. Lopez Brugos. Foundations of an adaptive tutoring system based on systemic networks. Lecture Notes in Computer Science, 1108:114–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okamoto:1996:IDS

[293] T. Okamoto and A. Inaba. The intelligent discussion supporting system under the distributed environment. Lecture Notes in Computer Science, 1108:123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel:1996:AAI

[294] A. Patel and Kinshuk. Applied artificial intelligence for teaching numeric topics in engineering disciplines. Lecture Notes in Computer Science, 1108: 132–140, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pemberton:1996:MDC

[295] L. Pemberton, S. Shurville, and T. Hartley. Motivating the design for a computer assisted environment for writers in a second language. Lecture Notes in Computer Science, 1108:141– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pernin:1996:EEP

[296] J. P. Pernin, V. Gueraud, and F. Coudret. An experimental environment for the production of pedagogical simulations. Lecture Notes in Computer Science, 1108:149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeVilliers:1996:CBI

[297] M. R. De Villiers. A component-based interactive practice environment. Lecture Notes in Computer Science, 1108: 158–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kang:1996:RIP

[298] B.-H. Kang and P. Goodyear. Representations of instructional purpose in courseware requirements engineering. Lecture Notes in Computer Science, 1108:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lelouche:1996:FRY

[299] R. Lelouche and J.-F. Morin. The formula: a relation? yes, but A concept too!! Lecture Notes in Computer Science, 1108:176–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nkambou:1996:CTA

[300] R. Nkambou, G. Gauthier, and C. Frasson. CREAM-Tools: An authoring environment for curriculum and course building in an intelligent tutoring system. Lecture Notes in Computer Science, 1108:186–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pescador:1996:ASR

[301] F. Pescador and J. Arriaga. Authoring system for reinforcement and evaluation (SARE). Lecture Notes in Computer Science, 1108:195–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ponta:1996:PSE

[302] D. Ponta, G. Da Bormida, and G. Donzellini. Preliminary student evaluation of a CBL course on digital systems electronics. Lecture Notes in Computer Science, 1108:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parodi:1996:CDL

[303] G. Parodi, D. Ponta, A. M. Scapolla, and M. Taini. Cooperative and distance learning in electronics using Internet. Lecture Notes in Computer Science, 1108:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz:1996:CDL

[304] F. Ruiz, M. Prieto, M. Ortega, and J. Bravo. Cooperative distance learning with an integrated system for computer assisted laboratory work. Lecture Notes in Computer Science, 1108: 220–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Warkentyne:1996:IEW

[305] H. M. K. Warkentyne, I. Smith, and E. Forte. Implementation and evaluation of a WWW multiple choice question server. Lecture Notes in Computer Science, 1108:228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forte:1996:TIA

[306] M. W. Forte, E. McMurray, and F. Simillion. Teaching informatics with ARIANE: An experimental Internet-Based pedagogical environment. Lecture Notes in Computer Science, 1108: 236–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cambiotti:1996:MTR

[307] F. Cambiotti, M. Parvis, and U. Pisani. Multimedia training and remote operating laboratory: Innovative solutions for instrumentation and electronic measurements courses. Lecture Notes in Computer Science, 1108:245– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1996:HEP

[308] J.-P. David, A. Cogne, and A. Dutel. Hypermedia exercises prototyping and modelising. Lecture Notes in Computer Science, 1108:252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dospisil:1996:TMM

[309] J. Dospisil, P. Marshall, E. Kendall, and T. Polgar. Testbed for measuring multimedia presentation quality disorders in courseware. Lecture Notes in Computer Science, 1108:261–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farinetti:1996:HES

[310] L. Farinetti and A. R. Scarafiotti. Hypertexts as educational systems: Pedagogical issues in teaching and learning mathematical problem solving. Lecture Notes in Computer Science, 1108: 270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hendrikx:1996:SHA

[311] K. Hendrikx, J. Elen, and H. Olivie. Structured hypermedia authoring: a simple tool for the design and implementation of structured hypermedia databases. Lecture Notes in Computer Science, 1108:279–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lara:1996:DEM

[312] S. Lara and M. J. Perez-Luque. Designing educational multimedia. Lecture Notes in Computer Science, 1108: 288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinez-Jimenez:1996:DSW

[313] J. M. Martinez-Jimenez, P. Martinez-Jimenez, F. Zafra-Lopez, and F. Casado-Revuelta. Design of a [sic] Windows software for elastic field simulation: Application to visualisation and animation of a rectangular piece in a projection with a load on the end. Lecture Notes in Computer Science, 1108: 298–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quang:1996:HIT

[314] N. Quang, L. Chiang, M. Garay, and D. Garcia. A hypermedia intelligent tutor for mathematical modelling teaching. Lecture Notes in Computer Science, 1108:306–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VillalbaLopez:1996:DSS

[315] R. Villalba Lopez, P. Martinez Jimenez, A. Blanca Pancorbo, and E. Casado-Revuelta. Design of software for the simulated and tutorized study of atomic models through the use of hypertext. Lecture Notes in Computer Science, 1108:313–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belmonte:1996:TAE

[316] M. V. Belmonte, J. Berbel, and R. Conejo. TEA: An agrarian economy instructor system. Lecture Notes in Computer Science, 1108:322–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brna:1996:AMB

[317] P. Brna and D. Duncan. The analogical model-based physics system: a workbench to investigate issues in how to support learning by analogy in physics. Lecture Notes in Computer Science, 1108:331–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Close:1996:SB

[318] A. M. Close and Y. A. Rzhanov. Seeing is believing. Lecture Notes in Computer Science, 1108:340–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1996:FFL

[319] I. Cohen, G. Edvinsson, R. Hsieh, and C. Johannesson. FLIP: Flexible learning in physics and mechanics. Lecture Notes in Computer Science, 1108:349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelRio:1996:SIM

[320] A. Del Rio and A. Rodriguez. 3D-Schema: An intuitive model for analog circuits instruction. Lecture Notes in Computer Science, 1108:356–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donoval:1996:CAL

[321] D. Donoval. Computer aided learning in microelectronics technology in Slovakia (state of the art). Lecture Notes in Computer Science, 1108:365– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haertel:1996:CSC

[322] H. B. Haertel and M. Luedke. A computer-supported course in mechanics. Lecture Notes in Computer Science, 1108:381–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahilly:1996:IPC

[323] T. J. Rahilly, A. Saroyan, J. Greer, and S. P. Lajoie. The InforMed professor: Clinical instruction of breast disease diagnosis and management. Lecture Notes in Computer Science, 1108: 390–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ungureanu:1996:CMS

[324] M. Ungureanu and S. Ioan. Computer modelling and simulation of the high frequency disturbing processes for low voltage consumers applied in power system training and education. Lecture Notes in Computer Science, 1108: 399–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valdes:1996:IPT

[325] M. D. Valdes, M. J. Moure, L. Rodriguez, and A. Del Rio. Interactive practical teaching of digital circuits design by means of field programmable gate arrays. Lecture Notes in Computer Science, 1108:408–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaharia:1996:ILE

[326] S. E. Zaharia and S. Ungureanu. Interactive learning environment in mechanics. Lecture Notes in Computer Science, 1108:415–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baniulis:1996:FIE

[327] K. Baniulis and B. Tamulynas. Flexible intelligent environment for tutoring and assessing learners. Lecture Notes in Computer Science, 1108:424–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bender-Oeberg:1996:DLE

[328] A. Bender-Oeberg and A. Edman. DLW — A learning environment for lake water diagnosis. Lecture Notes in Computer Science, 1108:427–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bersagol:1996:XLS

[329] V. Bersagol, J.-L. Dessalles, F. Kaplan, and J.-C. Marze. XMOISE: a logical spreadsheet to elicit didactic knowledge. Lecture Notes in Computer Science, 1108:430–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cabrol-Bass:1996:LPN

[330] D. Cabrol-Bass, J.-P. Rabine, and M. Rouillard. Learning proton NMR spectroscopy with computers. Lecture Notes in Computer Science, 1108:433– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuevas-Rubino:1996:TSS

[331] M. Cuevas-Rubiño, D. Gonzalez-Caballero, P. Martinez-Jimenez, and E. Casado-Revuelta. Tutorized simulated study of RC, LR and LRC circuits for windows. Lecture Notes in Computer Science, 1108:436–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dikareva:1996:IKB

[332] S. Dikareva and E. Dikarev. Interactive knowledge base for designing new technology based tutoring systems. Lecture Notes in Computer Science, 1108:439–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortet:1996:MSI

[333] P. Fortet, M. J. Bellido, F. Sivianes, and A. V. Medina. Multimedia system for instruction and learning electronics. Lecture Notes in Computer Science, 1108:442–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuji:1996:CLS

[334] T. Fuji, T. Tanigawa, M. Kozeni, and M. Inui. Collaborative learning systems on the INTERNET with casebased reasoning. Lecture Notes in Computer Science, 1108:445–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavrilova:1996:ALS

[335] T. Gavrilova, E. Averbukh, and A. Voinov. An approach to learning software based on student modelling. Lecture Notes in Computer Science, 1108:448–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goppold:1996:LTS

[336] A. Goppold. The Leibinz TLSI: a secondary macro programming interface and universal ASCII user interface shell for hypermedia. Lecture Notes in Computer Science, 1108:451–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamar-Thibault:1996:HPU

[337] S. Hamar-Thibault and R. Hamar. A hypermedia presentation to understand interactions between electron and solid — scanning electron microscope. Lecture Notes in Computer Science, 1108: 454–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huapaya:1996:SMN

[338] C. R. Huapaya and G. M. Arona. A student model in numerical analysis for an actual engineering student. Lecture Notes in Computer Science, 1108:460–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luckin:1996:TPS

[339] R. Luckin. Teaching primary science: a psychologically well-grounded approach. Lecture Notes in Computer Science, 1108:463–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millan:1996:GIG

[340] E. Millan, A. Muñoz-Zea, J. L. Perez de-la Cruz, and F. Triguero-Ruiz. GITE: Intelligent generation of tests. Lecture Notes in Computer Science, 1108:466–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millan:1996:TIT

[341] E. Millan, J. L. Perez de-la Cruz, F. Triguero-Ruiz, and L. Vazquez-Saenz. TUDER: An intelligent tutoring system for symbolic derivation. Lecture Notes in Computer Science, 1108: 469–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RizoAldeguer:1996:SRK

[342] R. Rizo Aldeguer, C. Muñoz, and O. Colomina Pardo. SBC-RX: Knowledge based system for radiodiagnosis and training of radiologists. the teaching file tool. A new feature. Lecture Notes in Computer Science, 1108:472– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soares:1996:SSS

[343] W. A. Soares and A. H. P. De Andrade. SIMFOT: a software for simulating photoelastic experiments. Lecture Notes in Computer Science, 1108:

476–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellare:1996:KHF

[344] Mihir Bellare, Ran Canetti, and Hugo Krawczyk. Keying hash functions for message authentication. Lecture Notes in Computer Science, 1109:1– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090001.
pdf.

Atici:1996:UHM

[345] M. Atici and Douglas R. Stinson. Universal hashing and multiple authentication. Lecture Notes in Computer Science, 1109:16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090016.
pdf.

Helleseth:1996:UHF

[346] Tor Helleseth and Thomas Johansson. Universal hash functions from exponential sums over finite fields and Galois rings. Lecture Notes in Computer Science, 1109:31–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090031.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090031.
                                        pdf.

Patarin:1996:ACH

[347] Jacques Patarin. Asymmetric cryptography with a hidden monomial. Lecture Notes in Computer Science, 1109: 45–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090045.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090045.
                                        pdf.

Simon:1996:ACA

[348] Daniel R. Simon. Anonymous communication and anonymous cash. Lecture Notes in Computer Science, 1109:61– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090061.
                                        pdf.

Langford:1996:WST

[349] Susan K. Langford. Weaknesses in some threshold cryptosystems. Lecture Notes in Computer Science, 1109:74– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090074.
                                        pdf.

Vaudenay:1996:HCD

[350] Serge Vaudenay. Hidden collisions on DSS. Lecture Notes in Computer Science, 1109:83–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090083.
pdf.

Young:1996:DSB

[351] Adam Young and Moti Yung. The dark side of black-box cryptography, or: Should we trust Capstone? Lecture Notes in Computer Science, 1109: 89–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090089.
pdf.

Kocher:1996:TAI

[352] Paul C. Kocher. Timing attacks on implementations of Diffie–Hellman, RSA, DSS, and other systems. Lecture Notes in Computer Science, 1109:104– 113, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090104.
pdf.

Naeslund:1996:ABH

[353] M. Naeslund. All bits in ax + b mod p are hard. Lecture Notes in Computer Science, 1109:114–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naslund:1996:ABH

[354] Mats Näslund. All bits in ax + b mod p are hard. Lecture Notes in Computer Science, 1109:114–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090114.
                                        pdf.

Boneh:1996:HCM

[355] Dan Boneh and Ramarathnam Venkatesan. Hardness of computing the most significant bits of secret keys in Diffie– Hellman and related schemes. Lecture Notes in Computer Science, 1109:129– 142, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090129.
                                        pdf;   http://theory.stanford.edu/
                                        ~dabo/papers/dhmsb.ps.gz.

Schnorr:1996:SRI

[356] Claus P. Schnorr. Security of 2^{t}-root identification and signatures. Lecture Notes in Computer Science, 1109:143– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090143.
pdf.

Gennaro:1996:RES

[357] Rosario Gennaro, Stanislaw Jarecki, Hugo Krawczyk, and Tal Rabin. Robust and efficient sharing of RSA functions. Lecture Notes in Computer Science, 1109:157–172, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090157.
pdf.

Cramer:1996:NGS

[358] Ronald Cramer and Ivan Bjerre Damgård. New generation of secure and practical RSA-Based signatures. Lecture Notes in Computer Science, 1109:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1109/11090173.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1109/11090173.
pdf.

Jakobsson:1996:PKO

[359] Markus Jakobsson and Moti Yung. Proving without knowing: On oblivious, agnostic and blindfolded provers. Lecture Notes in Computer Science, 1109:186–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1109/11090186.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090186.
                                        pdf.

Halevi:1996:PPS

[360] Shai Halevi and Silvio Micali. Practical and provably-secure commitment schemes from collision-free hashing. Lecture Notes in Computer Science, 1109:201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1109/11090201.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090201.
                                        pdf.

Knudsen:1996:IDA

[361] Lars R. Knudsen and Willi Meier. Improved differential attacks on RC5. Lecture Notes in Computer Science, 1109: 216–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090216.
                                        pdf.

Oorschot:1996:IIM

[362] Paul C. van Oorschot and Michael J. Wiener. Improving implementable meet-in-the-middle attacks by orders of magnitude. Lecture Notes in Computer Science, 1109:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1109/11090229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090229.
pdf.

vanOorschot:1996:IIM

[363] Paul C. van Oorschot and Michael J. Wiener. Improving implementable meet-in-the-middle attacks by orders of magnitude. Lecture Notes in Computer Science, 1109:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090229.
pdf.

Kelsey:1996:KSC

[364] John Kelsey, Bruce Schneier, and David Wagner. Key-schedule cryptanalysis of IDEA, G-DES, GOST, SAFER, and Triple-DES. Lecture Notes in Computer Science, 1109:237– 251, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090237.
pdf;   http://www.counterpane.com/
key_schedule.html.

Kilian:1996:HPA

[365] Joe Kilian and Phillip Rogaway. How to protect DES against exhaustive key search. Lecture Notes in Computer Science, 1109:252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090252.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090252.
                                        pdf.

Maurer:1996:DHO

[366] Ueli M. Maurer and Stefan Wolf. Diffie–Hellman oracles. Lecture Notes in Computer Science, 1109:268–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090268.
                                        pdf.

Boneh:1996:ABB

[367] Dan Boneh and Richard J. Lipton. Algorithms for black-box fields and their application to cryptography. Lecture Notes in Computer Science, 1109:283– 297, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090283.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090283.
                                        pdf;   http://theory.stanford.edu/
                                        ~dabo/papers/bbf.ps.gz.

Bosselaers:1996:FHP

[368] Antoon Bosselaers, René Govaerts, and Joos Vandewalle. Fast hashing on the Pentium. Lecture Notes in Computer Science, 1109:298–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090298.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090298.
pdf.

Shoup:1996:FPS

[369] Victor Shoup. On fast and provably secure message authentication based on universal hashing. Lecture Notes in Computer Science, 1109:313– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090313.
pdf.

Hughes:1996:QCU

[370] R. J. Hughes, G. G. Luther, G. L. Morgan, C. G. Peterson, and C. Simmons. Quantum cryptography over underground optical fibers. Lecture Notes in Computer Science, 1109:329– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090329.
pdf.

Mayers:1996:QKD

[371] Dominic Mayers. Quantum key distribution and string oblivious transfer in noisy channels. Lecture Notes in Computer Science, 1109:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090343.
                                        pdf.

Massey:1996:LCP

[372] James L. Massey and Shirlei Serconek. Linear complexity of periodic sequences: a general theory. Lecture Notes in Computer Science, 1109:358– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1109/11090358.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1109/11090358.
                                        pdf.

Camion:1996:GSI

[373] Paul Camion and Anne Canteaut. Generalization of Siegenthaler inequality and Schnorr-Vaudenay multipermutations. Lecture Notes in Computer Science, 1109:372–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1109/11090372.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1109/11090372.pdf.

Blundo:1996:TOB

[374] Carlo Blundo, Luiz A. Frota Mattos, and Douglas R. Stinson. Tradeoffs between communication and storage in unconditionally secure schemes for broadcast encryption and interactive key distribution. Lecture Notes in Computer Science, 1109:387–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090387.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090387.
pdf.

Droste:1996:NRV

[375] Stefan Droste. New results on visual cryptography. Lecture Notes in Computer Science, 1109:401–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1109/11090401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1109/11090401.
pdf.

Anonymous:1996:AI

[376] Anonymous. Author index. Lecture Notes in Computer Science, 1109:417– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1109/1109indx.pdf.

Blazy:1996:AIA

[377] S. Blazy and P. Facon. An automatic interprocedural analysis for the understanding of scientific application programs. Lecture Notes in Computer Science, 1110:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulyonkov:1996:PAS

[378] M. A. Bulyonkov and D. V. Kochetov. Practical aspects of specialization of Algol-like programs. Lecture Notes in Computer Science, 1110:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chin:1996:SAM

[379] W.-N. Chin, S.-C. Khoo, and P. Thiemann. Synchronization analyses for multiple recursion parameters. Lecture Notes in Computer Science, 1110:33– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Consel:1996:UAC

[380] C. Consel, L. Hornof, F. Noel, and J. Noye. A uniform approach for compile-time and run-time specialization. Lecture Notes in Computer Science, 1110:54–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danvy:1996:PTD

[381] O. Danvy. Pragmatics of type-directed partial evaluation. Lecture Notes in Computer Science, 1110:73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Draves:1996:CGI

[382] S. Draves. Compiler generation for interactive graphics using intermediate code. Lecture Notes in Computer Science, 1110:95–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallagher:1996:RAC

[383] J. P. Gallagher and L. Lafave. Regular approximation of computation paths in logic and functional languages. Lecture Notes in Computer Science, 1110:115– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glueck:1996:RMS

[384] R. Glueck and M. H. Soerensen. A roadmap to metacomputation by supercompilation. Lecture Notes in Computer Science, 1110:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hatcliff:1996:RAH

[385] J. Hatcliff and R. Glueck. Reasoning about hierarchies of online program specialization systems. Lecture Notes in Computer Science, 1110:161– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hughes:1996:TSL

[386] J. Hughes. Type specialisation for the lambda-calculus; or, A new paradigm for partial evaluation based on type inference. Lecture Notes in Computer Science, 1110:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1996:WDW

[387] N. D. Jones. What not to do when writing an interpreter for specialisation. Lecture Notes in Computer Science, 1110:216–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joergensen:1996:EGE

[388] J. Joergensen and M. Leuschel. Efficiently generating efficient generating extensions in Prolog. Lecture Notes in Computer Science, 1110:238–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leuschel:1996:GCP

[389] M. Leuschel and B. Martens. Global control for partial deduction through characteristic atoms and global trees. Lecture Notes in Computer Science, 1110:263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miniussi:1996:SIC

[390] A. Miniussi and D. J. Sherman. Squeezing intermediate construction in equational programs. Lecture Notes in Computer Science, 1110:284–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mogensen:1996:EPE

[391] T. A. Mogensen. Evolution of partial evaluators: Removing inherited limits. Lecture Notes in Computer Science, 1110:303–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nemytykh:1996:SAS

[392] A. P. Nemytykh, V. A. Pinchuk, and V. F. Turchin. A self-applicable supercompiler. Lecture Notes in Computer Science, 1110:322–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1996:MLL

[393] F. Nielson and H. R. Nielson. Multilevel lambda-calculi: An algebraic description. Lecture Notes in Computer Science, 1110:338–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pettorossi:1996:CRS

[394] A. Pettorossi and M. Proietti. A comparative revisitation of some program transformation techniques. Lecture Notes in Computer Science, 1110: 355–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pettorossi:1996:TLP

[395] A. Pettorossi and M. Proietti. A theory of logic program specialization and generalization for dealing with input data properties. Lecture Notes in Computer Science, 1110:386–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reps:1996:PSP

[396] T. Reps and T. Turnidge. Program specialization via program slicing. Lecture Notes in Computer Science, 1110:409–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakharov:1996:SIP

[397] A. Sakharov. Specialization of imperative programs through analysis of relational expressions. Lecture Notes in Computer Science, 1110:430–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sestoft:1996:MPM

[398] P. Sestoft. ML pattern match compilation and partial evaluation. Lecture Notes in Computer Science, 1110:446– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sperber:1996:SAO

[399] M. Sperber. Self-applicable online partial evaluation. Lecture Notes in Computer Science, 1110:465–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turchin:1996:MMT

[400] V. F. Turchin. Metacomputation: Metasystem transitions plus supercompilation. Lecture Notes in Computer Science, 1110:481–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferes:1996:RLP

[401] J. J. Alferes and L. Moniz Pereira. Reasoning with logic programming. Lecture Notes in Computer Science, 1111:xx + 326, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolanoski:1996:AAN

[402] H. Kolanoski. Application of artificial neural networks in particle physics. Lecture Notes in Computer Science, 1112:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwefel:1996:ECH

[403] H. P. Schwefel. Evolutionary computation — history, status, and perspectives. Lecture Notes in Computer Science, 1112:15–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abeles:1996:TSC

[404] M. Abeles and Y. Prut. Temporal structure of cortical activity. Lecture Notes in Computer Science, 1112:16– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramacher:1996:SVS

[405] U. Ramacher. SEE-1 — A vision system for use in real world environments. Lecture Notes in Computer Science, 1112:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fromherz:1996:TIN

[406] P. Fromherz. Towards integration of nerve cells and silicon devices. Lecture Notes in Computer Science, 1112: 18–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1996:UPN

[407] C. H. Anderson. Unifying perspectives on neuronal codes and processing. Lecture Notes in Computer Science, 1112:19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perrett:1996:VRB

[408] D. I. Perrett. Visual recognition based on coding in temporal cortex: Analysis of pattern configuration and generalisation across viewing conditions without mental rotation. Lecture Notes in Computer Science, 1112:20–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ji:1996:NES

[409] H.-B. Ji, K.-S. Leung, and Y. Leung. A novel encoding strategy for associative memory. Lecture Notes in Computer Science, 1112:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adeodato:1996:AMH

[410] P. J. L. Adeodato and J. G. Taylor. Autoassociative memory with high storage capacity. Lecture Notes in Computer Science, 1112:29–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1996:IEA

[411] B. Graham and D. Willshaw. Information efficiency of the associative net at arbitrary coding rates. Lecture Notes in Computer Science, 1112:35–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nijman:1996:ELS

[412] M. J. Nijman and H. J. Kappen. Efficient learning in sparsely connected Boltzmann machines. Lecture Notes in Computer Science, 1112:41–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoelkopf:1996:IIS

[413] B. Schoelkopf, C. Burges, and V. Vapnik. Incorporating invariances in support vector learning machines. Lecture Notes in Computer Science, 1112:47– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoekstra:1996:ERN

[414] A. Hoekstra, S. A. Tholen, and R. P. W. Duin. Estimating the reliability of neural network classifications. Lecture Notes in Computer Science, 1112:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bishop:1996:BIN

[415] C. M. Bishop and C. S. Qazaz. Bayesian inference of noise levels in regression. Lecture Notes in Computer Science, 1112:59–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haemaelaeinen:1996:CRP

[416] A. Haemaelaeinen and L. Holmstroem. Complexity reduction in probabilistic neural networks. Lecture Notes in Computer Science, 1112:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sardo:1996:ACR

[417] L. Sardo and J. Kittler. Asymptotic complexity of an RBF NN for correlated data representation. Lecture Notes in Computer Science, 1112:71– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raudys:1996:RES

[418] S. Raudys and T. Cibas. Regularization by early stopping in single layer perceptron training. Lecture Notes in Computer Science, 1112:77–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rueger:1996:CWS

[419] S. M. Rueger and A. Ossen. Clustering in weight space of feedforward nets. Lecture Notes in Computer Science, 1112:83–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boes:1996:LCL

[420] S. Boes. Learning curves of on-line and off-line training. Lecture Notes in Computer Science, 1112:89–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tax:1996:LSM

[421] D. M. J. Tax and H. J. Kappen. Learning structure with many-take-all networks. Lecture Notes in Computer Science, 1112:95–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappen:1996:DFL

[422] H. J. Kappen and P. Varona. Dynamic feature linking in stochastic networks with short range interactions. Lecture Notes in Computer Science, 1112:101– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hillenbrand:1996:CDS

[423] U. Hillenbrand and J. L. Van Hemmen. Collective dynamics of a system of adaptive velocity channels. Lecture Notes in Computer Science, 1112:107– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nelles:1996:LLM

[424] O. Nelles. Local linear model trees for on-line identification of time-variant nonlinear dynamic systems. Lecture Notes in Computer Science, 1112:115– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deco:1996:NDS

[425] G. Deco and B. Schuermann. Nonparametric data selection for improvement of parametric neural learning: a cumulant-surrogate method. Lecture Notes in Computer Science, 1112:121– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pawelzik:1996:PM

[426] K. Pawelzik, K.-R. Mueller, and J. Kohlmorgen. Prediction of mixtures. Lecture Notes in Computer Science, 1112:127–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kimura:1996:LDS

[427] M. Kimura and R. Nakano. Learning dynamical systems produced by recurrent neural networks. Lecture Notes in Computer Science, 1112:133–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hyvaerinen:1996:PLN

[428] A. Hyvaerinen. Purely local neural principal component and independent component learning. Lecture Notes in Computer Science, 1112:139–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuertz:1996:HFC

[429] R. P. Wuertz, W. Konen, and K.-O Behrmann. How fast can neuronal algorithms match patterns. Lecture Notes in Computer Science, 1112:145– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1996:ANG

[430] T. Hoffmann and J. M. Buhmann. An annealed “neural gas” network for robust vector quantization. Lecture Notes in Computer Science, 1112:151– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walter:1996:ACI

[431] J. Walter and H. Ritter. Associative completion and investment learning using PSOMs. Lecture Notes in Computer Science, 1112:157–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bishop:1996:GPA

[432] C. M. Bishop, M. Svensen, and C. K. I. Williams. GTM: a principled alternative to the self-organizing map. Lecture Notes in Computer Science, 1112: 165–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munoz:1996:CTA

[433] A. Muñoz. Creating term associations using a hierarchical ART architecture. Lecture Notes in Computer Science, 1112:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czernichow:1996:AST

[434] T. Czernichow. Architecture selection through statistical sensitivity analysis. Lecture Notes in Computer Science, 1112:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daigremont:1996:RTM

[435] P. Daigremont, H. De Lassus, F. Badran, and S. Thiria. Regression by topological map: Application on real data. Lecture Notes in Computer Science, 1112:185–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marko:1996:SPN

[436] K. A. Marko, J. V. James, T. M. Feldkamp, and G. V. Puskorius. Signal processing by neural networks to create “virtual” sensors and model-based diagnostics. Lecture Notes in Computer Science, 1112:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frenken:1996:DAS

[437] R. M. L. Frenken, C. M. Rozendaal, H. E. Dijk, and P. C. Knoester. Development of an advisory system based on a neural network for the operation of a coal fired power plant. Lecture Notes in Computer Science, 1112:197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angstenberger:1996:BFA

[438] J. Angstenberger. Blast furnace analysis with neural networks. Lecture Notes in Computer Science, 1112:203– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruske:1996:ASC

[439] J. Bruske, M. Hansen, L. Riehn, and G. Soremet. Adaptive saccade control of a binocular head with dynamic cell structures. Lecture Notes in Computer Science, 1112:215–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Versino:1996:LFM

[440] C. Versino and L. M. Gambardella. Learning fine motion by using the hierarchical extended Kohonen map. Lecture Notes in Computer Science, 1112: 221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laaksonen:1996:SDS

[441] J. Laaksonen and E. Oja. Subspace dimension selection and averaged learning subspace method in handwritten digit classification. Lecture Notes in Computer Science, 1112:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolters:1996:DRN

[442] M. Wolters. A dual route neural net approach to grapheme-to-phoneme conversion. Lecture Notes in Computer Science, 1112:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duempelmann:1996:SES

[443] M. Duempelmann and C. E. Elger. Separating EEG spike-clusters in epilepsy by a growing and splitting net. Lecture Notes in Computer Science, 1112:239–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valkealahti:1996:OTF

[444] K. Valkealahti and E. Oja. Optimal texture feature selection for the co-occurrence map. Lecture Notes in Computer Science, 1112:245–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blanz:1996:CVB

[445] V. Blanz, B. Schoelkopf, H. Buelthoff, and C. Burges. Comparison of viewbased object recognition algorithms using realistic 3D models. Lecture Notes in Computer Science, 1112:251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Austermeier:1996:CCR

[446] H. Austermeier, G. Hartmann, and R. Hilker. Color-calibration of robot vision system using self-organizing feature maps. Lecture Notes in Computer Science, 1112:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Acuna:1996:NNM

[447] G. Acuña, H. Jorquera, and R. Perez. Neural network model for maximum ozone concentration prediction. Lecture Notes in Computer Science, 1112: 263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohonen:1996:VLT

[448] T. Kohonen, S. Kaski, K. Lagus, and T. Honkela. Very large twolevel SOM for the browsing of newsgroups. Lecture Notes in Computer Science, 1112:269–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ma:1996:APS

[449] Q. Ma, H. Isahara, and H. Ozaku. Automatic part-of-speech tagging of Thai corpus using neural networks. Lecture Notes in Computer Science, 1112:275– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cawley:1996:RSC

[450] G. C. Cawley and S. R. Dorling. Reproducing a subjective classification scheme for atmospheric circulation patterns over the united kingdom using a neural network. Lecture Notes in Computer Science, 1112:281–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1996:TGD

[451] H. H. Yang and S. Amari. Two gradient descent algorithms for blind signal separation. Lecture Notes in Computer Science, 1112:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLassus:1996:CRP

[452] H. De Lassus, P. Daigremont, F. Badran, and S. Thiria. Classification rejection by prediction. Lecture Notes in Computer Science, 1112:293–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:ARB

[453] D.-H. Lee, J. S. Payne, H.-G. Byun, and K. C. Persaud. Application of radial basis function neural networks to odour sensing using a broad specificity array of conducting polymers. Lecture Notes in Computer Science, 1112:299– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heidemann:1996:HOR

[454] G. Heidemann, F. Kummert, H. Ritter, and G. Sagerer. A hybrid object recognition architecture. Lecture Notes in Computer Science, 1112:305–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buehlmeier:1996:RLA

[455] A. Buehlmeier, G. Manteuffel, M. Rossmann, and K. Goser. Robot learning in analog neural hardware. Lecture Notes in Computer Science, 1112: 311–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Littmann:1996:VGR

[456] E. Littmann, A. Drees, and H. Ritter. Visual gesture recognition by a modular neural system. Lecture Notes in Computer Science, 1112:317–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1996:TLG

[457] T. Maurer and C. Von Der Malsburg. Tracking and learning graphs on image sequences of faces. Lecture Notes in Computer Science, 1112:323–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fukushima:1996:NNM

[458] K. Fukushima, Y. Yamaguchi, and M. Okada. Neural network model recalling spatial maps. Lecture Notes in Computer Science, 1112:329–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giese:1996:NFD

[459] M. A. Giese, G. Schoener, and H. S. Hock. Neural field dynamics for motion perception. Lecture Notes in Computer Science, 1112:335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Callan:1996:ATD

[460] R. E. Callan and D. Palmer-Brown. Analytical technique for deriving connectionist representations of symbol structures. Lecture Notes in Computer Science, 1112:341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wittenburg:1996:MHW

[461] P. Wittenburg, D. Van Kuijk, and T. Dijkstra. Modeling human word recognition with sequences of artificial neurons. Lecture Notes in Computer Science, 1112:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boden:1996:CVI

[462] M. Boden. A connectionist variation on inheritance. Lecture Notes in Computer Science, 1112:353–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolinummi:1996:MMP

[463] P. Kolinummi, T. Haemaelaeinen, H. Klapuri, and K. Kaski. Mapping of multilayer perceptron networks to partial tree shaped parallel neurocomputer. Lecture Notes in Computer Science, 1112:359–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haemaelaeinen:1996:LEP

[464] T. Haemaelaeinen, P. Kolinummi, and K. Kaski. Linearly expandable partial tree shape architecture for parallel neurocomputer. Lecture Notes in Computer Science, 1112:365–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demosthenous:1996:HSS

[465] A. Demosthenous, J. Taylor, and S. Smedley. A high-speed scalable CMOS current-mode winner-take-all network. Lecture Notes in Computer Science, 1112:371–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franz:1996:ASM

[466] M. Franz and R. Schueffny. An architectural study of a massively parallel processor for convolution-type operations in complex vision tasks. Lecture Notes in Computer Science, 1112: 377–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perez-Uribe:1996:FIA

[467] A. Perez-Uribe and E. Sanchez. FPGA implementation of an adaptable-size neural network. Lecture Notes in Computer Science, 1112:383–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Floreano:1996:ECI

[468] D. Floreano. Extraction of coherent information from non-overlapping receptive fields. Lecture Notes in Computer Science, 1112:389–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brecht:1996:CTI

[469] M. Brecht and A. K. Engel. Cortiotectal interactions in the cat visual system. Lecture Notes in Computer Science, 1112:395–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martignon:1996:DMH

[470] L. Martignon, K. Laskey, A. Schwarz, and E. Vaadia. Detecting and measuring higher order synchronization among neurons: A Bayesian approach. Lecture Notes in Computer Science, 1112:401–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riesenhuber:1996:AFS

[471] M. Riesenhuber, H.-U. Bauer, and T. Geisel. Analyzing the formation of structure in high-dimensional selforganizing maps reveals differences to feature map models. Lecture Notes in Computer Science, 1112:409–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pawelzik:1996:PRC

[472] K. Pawelzik, H.-U. Bauer, F. Wolf, and T. Geisel. Precise restoration of cortical orientation maps explained by Hebbian dynamics of geniculocortical connections. Lecture Notes in Computer Science, 1112:415–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andres:1996:MKS

[473] M. Andres, O. Schlueter, F. Spengler, and H. R. Dinse. Modification of Kohonen’s SOFM to simulate cortical plasticity induced by coactivation input patterns. Lecture Notes in Computer Science, 1112:421–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piepenbrock:1996:CMD

[474] C. Piepenbrock, H. Ritter, and K. Obermayer. Cortical map development driven by spontaneous retinal activity waves. Lecture Notes in Computer Science, 1112:427–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cruse:1996:SNN

[475] H. Cruse, C. Bartling, J. Dean, and T. Kindermann. Simplifying neural networks for controlling walking by exploiting physical properties. Lecture Notes in Computer Science, 1112:433– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schierwagen:1996:SCT

[476] A. Schierwagen and H. Werner. Saccade control through the collicular motor map: Two-dimensional neural field model. Lecture Notes in Computer Science, 1112:439–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsodyks:1996:PNS

[477] M. V. Tsodyks and H. Markram. Plasticity of neocortical synapses enables transitions between rate and temporal coding. Lecture Notes in Computer Science, 1112:445–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wennekers:1996:CSS

[478] T. Wennekers and G. Palm. Controlling the speed of synfire chains. Lecture Notes in Computer Science, 1112: 451–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Privitera:1996:TCP

[479] C. M. Privitera and L. Shastri. Temporal compositional processing by a DSOM hierarchical model. Lecture Notes in Computer Science, 1112:457– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertoni:1996:GMH

[480] A. Bertoni, P. Campadelli, M. Carpentieri, and G. Grossi. A genetic model and the Hopfield networks. Lecture Notes in Computer Science, 1112:463– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorodnichy:1996:DCP

[481] D. O. Gorodnichy. Desaturating coefficient for projection learning rule. Lecture Notes in Computer Science, 1112: 469–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sjoedin:1996:GMI

[482] G. Sjoedin. Getting more information out of SDM. Lecture Notes in Computer Science, 1112:477–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCormack:1996:UGP

[483] C. McCormack. Using a general purpose meta neural network to adapt a parameter of the Quickpropagation Learning Rule. Lecture Notes in Computer Science, 1112:483–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luo:1996:ULM

[484] F.-L. Luo and R. Unbehauen. Unsupervised learning of the minor subspace. Lecture Notes in Computer Science, 1112:489–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1996:UGA

[485] M. Schmidt. A unification of genetic algorithms, neural networks and fuzzy logic: The GANNFL approach. Lecture Notes in Computer Science, 1112: 495–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasenjaeger:1996:ALG

[486] M. Hasenjaeger and H. Ritter. Active learning of the generalized highlow-game. Lecture Notes in Computer Science, 1112:501–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muselli:1996:OPA

[487] M. Muselli. Optimality of pocket algorithm. Lecture Notes in Computer Science, 1112:507–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Young:1996:IEC

[488] S. Young and T. Downs. Improvements and extensions to the constructive algorithm CARVE. Lecture Notes in Computer Science, 1112:513–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arai:1996:ARL

[489] K. Arai and R. Nakano. Annealed RNN learning of finite state automata. Lecture Notes in Computer Science, 1112: 519–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freisleben:1996:HLR

[490] B. Freisleben and C. Hagen. A hierarchical learning rule for independent component analysis. Lecture Notes in Computer Science, 1112:525–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1996:INN

[491] G. Zhou and J. Si. Improving neural network training based on Jacobian rank deficiency. Lecture Notes in Computer Science, 1112:531–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biro:1996:NNE

[492] J. Biro, E. Halasz, T. Tron, and M. Boda. Neural networks for exact constrained optimization. Lecture Notes in Computer Science, 1112:537– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroener:1996:CSM

[493] S. Kroener and R. Moratz. Capacity of structured multilayer networks with shared weights. Lecture Notes in Computer Science, 1112:543–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boes:1996:OWD

[494] S. Boes. Optimal weight decay in a perceptron. Lecture Notes in Computer Science, 1112:551–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwok:1996:BRC

[495] T.-Y. Kwok and D.-Y. Yeung. Bayesian regularization in constructive neural networks. Lecture Notes in Computer Science, 1112:557–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SantaCruz:1996:NDA

[496] C. Santa Cruz and J. Dorronsoro. A nonlinear discriminant algorithm for data projection and feature extraction. Lecture Notes in Computer Science, 1112:563–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Browne:1996:MSA

[497] C. Browne, J. De L. Pereira Castro, and A. Storkey. A modified spreading algorithm for autoassociation in weightless neural networks. Lecture Notes in Computer Science, 1112:569– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuenemann:1996:AMF

[498] S. Schuenemann, B. Michaelis, and W. Schubert. Analysis of multifluorescence signals using a modified self-organizing feature map. Lecture Notes in Computer Science, 1112:575– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koehle:1996:VSH

[499] M. Koehle and D. Merkl. Visualizing similarities in high dimensional input spaces with a growing and splitting neural network. Lecture Notes in Computer Science, 1112:581–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia-Salicetti:1996:NLP

[500] S. Garcia-Salicetti. A neural lexical post-processor for improved neural predictive word recognition. Lecture Notes in Computer Science, 1112:587– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botto:1996:SNM

[501] M. A. Botto, H. A. B. Te Braake, and J. Sa Da Costa. Solving nonlinear MBPC through convex optimization: A comparative study using neural networks. Lecture Notes in Computer Science, 1112:593–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guermeur:1996:CSM

[502] Y. Guermeur and P. Gallinari. Combining statistical models for protein secondary structure prediction. Lecture Notes in Computer Science, 1112:599– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pietruschka:1996:URN

[503] U. Pietruschka and R. Brause. Using RBF-Nets in rubber industry process control. Lecture Notes in Computer Science, 1112:605–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziemke:1996:TAR

[504] T. Ziemke. Towards autonomous robot control via self-adapting recurrent networks. Lecture Notes in Computer Science, 1112:611–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mael:1996:HNL

[505] E. Mael. A hierarchical network for learning robust models of kinematic chains. Lecture Notes in Computer Science, 1112:617–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yasuhara:1996:CBC

[506] K. Yasuhara and H. A. Mallot. Context-based cognitive map learning for an autonomous robot using a model of cortico-hippocampal interplay. Lecture Notes in Computer Science, 1112: 623–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zadel:1996:ABC

[507] S. Zadel. An algorithm for bootstrapping the core of a biologically inspired motor control system. Lecture Notes in Computer Science, 1112:629–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RuizDeAngulo:1996:ARS

[508] V. Ruiz De Angulo and C. Torras. Automatic recalibration of a space robot: An industrial prototype. Lecture Notes in Computer Science, 1112:635–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancke:1996:PCC

[509] D. Jancke, A. C. Akhavan, W. Erlhagen, and M. A. Giese. Population coding in cat visual cortex reveals nonlinear interactions as predicted by a neural field model. Lecture Notes in Computer Science, 1112:641–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalia:1996:AIO

[510] L. N. Kalia. An analysis and interpretation of the oscillatory behavior of a model of the granular layer of the cerebellum. Lecture Notes in Computer Science, 1112:655–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mechsner:1996:CCM

[511] F. Mechsner and G. Palm. The cerebellum as a “coupling machine”. Lecture Notes in Computer Science, 1112: 661–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durstewitz:1996:PFD

[512] D. Durstewitz and O. Guentuerkuen. The possible function of dopamine in associative learning: a computational model. Lecture Notes in Computer Science, 1112:667–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riehle:1996:SDC

[513] A. Riehle, S. Gruen, A. Aertsen, and J. Requin. Signatures of dynamic cell assemblies in monkey motor cortex. Lecture Notes in Computer Science, 1112:673–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wesarg:1996:MSP

[514] T. Wesarg, B. Brueckner, and C. Schauer. Modelling speech processing and recognition in the auditory system with a three-stage architecture. Lecture Notes in Computer Science, 1112:679–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Triesch:1996:BPE

[515] J. Triesch and C. Von Der Malsburg. Binding — A proposed experiment and a model. Lecture Notes in Computer Science, 1112:685–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohme:1996:RMD

[516] M. Ohme and A. Schierwagen. A reduced model for dendritic trees with active membrane. Lecture Notes in Computer Science, 1112:691–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tavitian:1996:SCL

[517] S. Tavitian, T. Fomin, and A. Loerincz. Stabilizing competitive learning during on-line training with an Anti-Hebbian weight modulation. Lecture Notes in Computer Science, 1112:697–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaucher:1996:NBB

[518] G. Vaucher. Neuro-biological bases for spario-temporal data coding in artificial neural networks. Lecture Notes in Computer Science, 1112:703–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsukada:1996:STL

[519] M. Tsukada, T. Aihara, H. Saito, and H. Karo. A spatio-temporal learning rule based on the physiological data of LTP induction in the hippocampal CA1 network. Lecture Notes in Computer Science, 1112:709–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vetter:1996:LNV

[520] T. Vetter. Learning novel views to a single face image. Lecture Notes in Computer Science, 1112:715–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogelgesang:1996:PAD

[521] J. Vogelgesang, A. Cozzi, and F. Woergoetter. A parallel algorithm depth perception from radial optical flow fields. Lecture Notes in Computer Science, 1112:721–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konen:1996:CFL

[522] W. Konen. Comparing facial line drawings with gray-level images: A case study on PHANTOMAS. Lecture Notes in Computer Science, 1112:727– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branca:1996:GCO

[523] A. Branca, G. Convertino, E. Stella, and A. Distante. Geometrically constrained optical flow estimation by an Hopfield neural network. Lecture Notes in Computer Science, 1112:735–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cotofana:1996:SBA

[524] S. Cotofana and S. Vassiliadis. Serial binary addition with polynomially bounded weights. Lecture Notes in Computer Science, 1112:741–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Granado:1996:ETD

[525] B. Granado and P. Garda. Evaluation of the two different interconnection networks of the CNAPS neurocomputer. Lecture Notes in Computer Science, 1112:747–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pican:1996:IPP

[526] N. Pican. Intrinsic and parallel performances of the OWE neural network architecture. Lecture Notes in Computer Science, 1112:755–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conti:1996:ACN

[527] M. Conti, G. Guaitini, and C. Turchetti. An analog CMOS neural network with on-chip learning and multilevel weight storage. Lecture Notes in Computer Science, 1112:761–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossmann:1996:EHL

[528] M. Rossmann, T. Jost, K. Goser, and A. Buehlmeier. Exponential Hebbian on-line learning implemented in FP-GAs. Lecture Notes in Computer Science, 1112:767–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1996:PFN

[529] J. S. Lange and H. Freiesleben. A parameter-free non-growing selforganizing map based upon gravitational principles: Algorithm and applications. Lecture Notes in Computer Science, 1112:827–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anouar:1996:TMM

[530] F. Anouar, F. Badran, and S. Thiria. Topological maps for mixture densities. Lecture Notes in Computer Science, 1112:833–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Opara:1996:NAI

[531] R. Opara and F. Woergoetter. A novel algorithm for image segmentation using time dependent interaction probabilities. Lecture Notes in Computer Science, 1112:839–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poetzsch:1996:RGL

[532] M. Poetzsch, T. Maurer, L. Wiskott, and C. Von Der Malsburg. Reconstruction from graphs labeled with responses of Gabor filters. Lecture Notes in Computer Science, 1112:845–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rehse:1996:EIC

[533] O. Rehse, M. Poetzsch, and C. Von Der Malsburg. Edge information: a confidence based algorithm emphasizing continuous curves. Lecture Notes in Computer Science, 1112:851–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1996:WLM

[534] E. M. Clarke and X. Zhao. Word level model checking. Lecture Notes in Computer Science, 1113:1, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekert:1996:CPT

[535] V. Diekert and A. Muscholl. Code problems on traces. Lecture Notes in Computer Science, 1113:2–17, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibbons:1996:MDC

[536] A. Gibbons, M. Amos, and D. Hodgson. Models of DNA computation. Lecture Notes in Computer Science, 1113: 18–36, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mosses:1996:TPA

[537] P. D. Mosses. Theory and practice of action semantics. Lecture Notes in Computer Science, 1113:37–61, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukund:1996:LTT

[538] M. Mukund and P. S. Thiagarajan. Linear time temporal logics over Mazurkiewicz traces. Lecture Notes in Computer Science, 1113:62–92, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peled:1996:POR

[539] D. A. Peled. Partial order reduction: model-checking using representatives. Lecture Notes in Computer Science, 1113:93–112, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Remmel:1996:NRS

[540] J. B. Remmel. Nonmonotonic rule systems: forward chaining, constraints, and complexity. Lecture Notes in Computer Science, 1113:113, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sannella:1996:MGA

[541] D. Sannella and A. Tarlecki. Mind the gap! abstract versus concrete models of specifications. Lecture Notes in Computer Science, 1113:114–134, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiuryn:1996:SCS

[542] J. Tiuryn. A sequent calculus for subtyping polymorphic types. Lecture Notes in Computer Science, 1113:135– 155, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uspensky:1996:KCR

[543] V. A. Uspensky. Kolmogorov complexity: recent research in Moscow. Lecture Notes in Computer Science, 1113:156– 166, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balbiani:1996:MLD

[544] P. Balbiani. A modal logic for data analysis. Lecture Notes in Computer Science, 1113:167–179, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baum:1996:SPF

[545] G. A. Baum, M. F. Frias, A. M. Haeberer, and P. E. Martinez Lopez. From specifications to programs: a fork-algebraic approach to bridge the gap. Lecture Notes in Computer Science, 1113:180–191, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bednarczyk:1996:LPE

[546] M. A. Bednarczyk. Logic of predicates with explicit substitutions. Lecture Notes in Computer Science, 1113:192– 205, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beigel:1996:QCS

[547] R. Beigel, W. Gasarch, M. Kummer, and G. Martin. On the query complexity of sets. Lecture Notes in Computer Science, 1113:206–217, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bono:1996:LCI

[548] V. Bono, M. Bugliesi, and L. Liquori. A lambda calculus of incomplete objects. Lecture Notes in Computer Science, 1113:218–229, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boreale:1996:BPR

[549] M. Boreale and L. Trevisan. Bisimilarity problems requiring exponential time. Lecture Notes in Computer Science, 1113:230–241, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBruin:1996:LDK

[550] A. De Bruin and S. H. Nienhuys-Cheng. Linear dynamic Kahn networks are deterministic. Lecture Notes in Computer Science, 1113:242–254, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1996:SPP

[551] X. Cai, T. Kloks, and C. K. Wong. Shortest path problems with time constraints. Lecture Notes in Computer Science, 1113:255–266, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chlebus:1996:PAD

[552] B. S. Chlebus, A. Czumaj, L. Gasieniec, and M. Kowaluk. Parallel alternating-direction access machine. Lecture Notes in Computer Science, 1113:267–478, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1996:SVT

[553] F. Corradini and M. Pistore. Specification and verification of timed lazy systems. Lecture Notes in Computer Science, 1113:279–290, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demri:1996:CIL

[554] S. Demri. A class of information logics with a decidable validity problem. Lecture Notes in Computer Science, 1113:291–302, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dessmark:1996:PNP

[555] A. Dessmark and A. Lingas. On the power of nonconservative PRAM. Lecture Notes in Computer Science, 1113: 303–311, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1996:SSV

[556] B. Durand. Self-similarity viewed as a local property via tile sets. Lecture Notes in Computer Science, 1113:312– 323, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelhardt:1996:SSS

[557] K. Engelhardt and W.-P. De Roever. Simulation of specification statements in Hoare logic. Lecture Notes in Computer Science, 1113:324–335, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esik:1996:EPI

[558] Z. Esik and A. Labella. Equational properties of iteration in algebraically complete categories. Lecture Notes in Computer Science, 1113:336–347, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernau:1996:UT

[559] H. Fernau. On unconditional transfer. Lecture Notes in Computer Science, 1113:348–359, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fotakis:1996:PPR

[560] D. Fotakis and P. Spirakis. (poly(log log n ), poly (log log n ))-restricted verifiers are unlikely to exist for languages in NP. Lecture Notes in Computer Science, 1113:360–371, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasieniec:1996:MCL

[561] L. Gasieniec, E. Kranakis, D. Krizanc, and A. Pelc. Minimizing congestion of layouts for ATM networks with faulty links. Lecture Notes in Computer Science, 1113:372–381, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glaister:1996:PAC

[562] I. Glaister and J. Shallit. Polynomial automaticity, context-free languages, and fixed points of morphisms (extended abstract). Lecture Notes in Computer Science, 1113:382–393, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goltz:1996:CT

[563] U. Goltz and H. Wehrheim. Causal testing. Lecture Notes in Computer Science, 1113:394–406, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1996:CLH

[564] Z. Hu, H. Iwasaki, and M. Takeichi. Construction of list homomorphisms by tupling and fusion. Lecture Notes in Computer Science, 1113:407–418, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwiatkowska:1996:PMS

[565] M. Kwiatkowska and G. Norman. Probabilistic metric semantics for a simple language with recursion. Lecture Notes in Computer Science, 1113:419– 430, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maggiolo-Schettini:1996:DG

[566] A. Maggiolo-Schettini and J. Winkowski. Dynamic graphs. Lecture Notes in Computer Science, 1113:431–442, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mantaci:1996:ET

[567] S. Mantaci and A. Restivo. Equations on trees. Lecture Notes in Computer Science, 1113:443–456, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlebusch:1996:EPP

[568] E. Ohlebusch and E. Ukkonen. On the equivalence problem for E-pattern languages (extended abstract). Lecture Notes in Computer Science, 1113:457– 468, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pawlowski:1996:SVP

[569] W. Pawlowski, P. Paczkowski, and S. Sokolowski. Specifying and verifying parametric processes. Lecture Notes in Computer Science, 1113:469–481, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1996:SFF

[570] R. Pliuskevicius. On saturation with flexible function symbols. Lecture Notes in Computer Science, 1113:482– 493, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roessner:1996:AGS

[571] C. Roessner and J.-P. Seifert. Approximating good simultaneous Diophantine approximations is almost NPhard. Lecture Notes in Computer Science, 1113:494–505, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seebold:1996:CSM

[572] P. Seebold. On the conjugation of standard morphisms. Lecture Notes in Computer Science, 1113:506–516, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1996:SMA

[573] H. Shi. A semantic matching algorithm: analysis and implementation. Lecture Notes in Computer Science, 1113:517–528, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1996:RTT

[574] J. F. Sibeyn. Routing on triangles, tori and honeycombs. Lecture Notes in Computer Science, 1113:529–541, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelaTorre:1996:UAT

[575] P. De la Torre and D. T. Kao. A uniform analysis of trie structures that store prefixing-keys with application to doubly-chained prefixing-tries. Lecture Notes in Computer Science, 1113:542– 553, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wabenhorst:1996:FTR

[576] A. Wabenhorst. On fairness in terminating and reactive programs. Lecture Notes in Computer Science, 1113:554– 565, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamakami:1996:PTS

[577] T. Yamakami. Polynomial time samplable distributions. Lecture Notes in Computer Science, 1113:566–578, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zucca:1996:SDA

[578] E. Zucca. From static to dynamic abstract data-types. Lecture Notes in Computer Science, 1113:579–590, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouhoula:1996:GFM

[579] A. Bouhoula. A general framework for mechanizing induction using test set. Lecture Notes in Computer Science, 1114:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulmer:1996:IER

[580] M. Bulmer. Inductive equational reasoning. Lecture Notes in Computer Science, 1114:13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Webb:1996:CSS

[581] G. I. Webb. Cost-sensitive specialization. Lecture Notes in Computer Science, 1114:23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1996:EMP

[582] X. Zhang and M. Numao. Efficient multiple predicate learner based on fast failure mechanism. Lecture Notes in Computer Science, 1114:35–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raskutti:1996:AUP

[583] B. Raskutti and A. Beitz. Acquiring user preferences for information filtering in interactive multi-media services. Lecture Notes in Computer Science, 1114:47–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:1996:DIF

[584] M. Zhao and C. Leckie. The development of an interactive fault diagnosis expert system for telecommunication applications. Lecture Notes in Computer Science, 1114:59–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishino:1996:IES

[585] K. Ishino, K. Sugai, and R. Mizoguchi. An intelligent education system which supports scientific thinking: Galileo — philosophy and basic architecture. Lecture Notes in Computer Science, 1114:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raskutti:1996:UAH

[586] B. Raskutti and I. Zukerman. A unified approach to handling uncertainty during cooperative consultations. Lecture Notes in Computer Science, 1114: 85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuwongse:1996:DPT

[587] V. Wuwongse and E. Nantajeewarawat. Declarative program theory with implicit implication. Lecture Notes in Computer Science, 1114:97–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debenham:1996:KDA

[588] J. Debenham. Knowledge decomposition: An analysis. Lecture Notes in Computer Science, 1114:109–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stumme:1996:CCT

[589] G. Stumme. The concept classification of a terminology extended by conjunction and disjunction. Lecture Notes in Computer Science, 1114:121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fall:1996:EPO

[590] A. Fall. Extending partial orders for sort reasoning. Lecture Notes in Computer Science, 1114:132–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Restificar:1996:UPP

[591] A. C. Restificar. An update procedure for A probabilistic deductive database. Lecture Notes in Computer Science, 1114:144–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zowghi:1996:FRA

[592] D. Zowghi, A. K. Ghose, and P. Peppas. A framework for reasoning about requirements evolution. Lecture Notes in Computer Science, 1114:157–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1996:AHK

[593] F.-L. Lai, J.-H. Lim, L. Ding, and A.- H. Tan. An application of hierarchical knowledge integration in hand-written form processing. Lecture Notes in Computer Science, 1114:169–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoang:1996:AMI

[594] D. B. Hoang and M. R. James. AMI: a model of intelligence. Lecture Notes in Computer Science, 1114:181–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohsuga:1996:SPN

[595] S. Ohsuga. Symbol processing by nonsymbol processor. Lecture Notes in Computer Science, 1114:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicholson:1996:FDU

[596] A. E. Nicholson. Fall diagnosis using dynamic belief networks. Lecture Notes in Computer Science, 1114:206– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1996:RDB

[597] S. W. K. Chan and J. Franklin. Remembrance of discourse based on textual continuity: A spreading activation network. Lecture Notes in Computer Science, 1114:218–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1996:LAU

[598] I. Thomas, I. Zukerman, J. Oliver, and B. Raskutti. Lexical access using minimum message length encoding. Lecture Notes in Computer Science, 1114:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel:1996:UNN

[599] M. Patel. Using neural nets to investigate lexical analysis. Lecture Notes in Computer Science, 1114:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwan:1996:VNA

[600] A. C. M. Kwan. Validity of normality assumption in CSP research. Lecture Notes in Computer Science, 1114: 253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1996:IGA

[601] B. Liu. An improved generic arc consistency algorithm and its specializations. Lecture Notes in Computer Science, 1114:264–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daalder:1996:HLS

[602] J. Daalder, P. W. Eklund, and K. Ohmori. High-level synthesis optimisation with genetic algorithms. Lecture Notes in Computer Science, 1114: 276–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitamura:1996:QRB

[603] Y. Kitamura, M. Ikeda, and R. Mizoguchi. A qualitative reasoning based on an ontology of fluid systems and its evaluation. Lecture Notes in Computer Science, 1114:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kato:1996:PCB

[604] S. Kato, H. Seki, and H. Itoh. Parallel cost-based abductive reasoning for distributed memory systems. Lecture Notes in Computer Science, 1114:300– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Washio:1996:HOE

[605] T. Washio and H. Motoda. A historyoriented envisioning method. Lecture Notes in Computer Science, 1114:312– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1996:TLR

[606] S. Jain and A. Sharma. Team learning of recursive languages. Lecture Notes in Computer Science, 1114:324–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1996:CVC

[607] T. Ho. Combining views on concepts in unsupervised concept learning. Lecture Notes in Computer Science, 1114:336–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cameron-Jones:1996:CBA

[608] M. Cameron-Jones. The complexity of batch approaches to reduced error rule set induction. Lecture Notes in Computer Science, 1114:348–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mofizur:1996:LSR

[609] Chowdhury Rahman Mofizur and M. Numao. Learning simple recursive concepts by discovering missing examples. Lecture Notes in Computer Science, 1114:360–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonacina:1996:SRL

[610] M. P. Bonacina and J. Hsiang. On semantic resolution with lemmaizing and contraction. Lecture Notes in Computer Science, 1114:372–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girle:1996:TEC

[611] R. A. Girle. Tableaux for expansion and contraction. Lecture Notes in Computer Science, 1114:387–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1996:TMG

[612] X. Huang. Translating machinegenerated resolution proofs into ND-Proofs at the assertion level. Lecture Notes in Computer Science, 1114:399– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1996:ETR

[613] B. Dai, D. A. Bell, and J. G. Hughes. Evidential temporal representations and reasoning. Lecture Notes in Computer Science, 1114:411–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melville:1996:UAN

[614] S. Melville, G. Sutcliffe, and D. Fraser. Using artificial neural networks for meteor-burst communications trail prediction. Lecture Notes in Computer Science, 1114:423–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Low:1996:MCR

[615] Boon Toh Low. Modeling commonsense rules in an inference network. Lecture Notes in Computer Science, 1114:435–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1996:ASN

[616] H. Park, K. Lee, H. Byun, and Y. Lee. Adaptive SEJONG-NET for on-line Hangul recognition. Lecture Notes in Computer Science, 1114:447–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luo:1996:UAS

[617] X. Luo and C. Zhang. A unified algebraic structure for uncertain reasonings. Lecture Notes in Computer Science, 1114:459–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courtney:1996:ESC

[618] A. P. Courtney, G. Antoniou, and N. Y. Foo. Exten: a system for computing default logic extensions. Lecture Notes in Computer Science, 1114:471–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morley:1996:LCE

[619] D. Morley and L. Sonenberg. A logic for concurrent events and action failure. Lecture Notes in Computer Science, 1114:483–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konishi:1996:DWM

[620] T. Konishi and Y. Itoh. Domain world models represented from variable viewpoints for ICAI systems of high-school chemistry. Lecture Notes in Computer Science, 1114:495–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

HungKwok:1996:CTT

[621] R. B. Hung Kwok. Creating theoretical terms for non-deterministic actions. Lecture Notes in Computer Science, 1114:510–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwok:1996:CTT

[622] R. B. Hung Kwok. Creating theoretical terms for non-deterministic actions. Lecture Notes in Computer Science, 1114:510–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1996:CEC

[623] G. Antoniou, C. K. MacNish, and N. Y. Foo. Conservative expansion concepts for default theories. Lecture Notes in Computer Science, 1114:522–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohara:1996:FEH

[624] K. Ohara, N. Babaguchi, and T. Kitahashi. On formation of exception hierarchy. Lecture Notes in Computer Science, 1114:534–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghose:1996:ADI

[625] A. K. Ghose and R. Goebel. Anytime default inference. Lecture Notes in Computer Science, 1114:546–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1996:SPC

[626] G. Antoniou, T. O’Neill, and J. Thurbon. Studying properties of classes of default logics — preliminary report. Lecture Notes in Computer Science, 1114:558–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noda:1996:LCB

[627] I. Noda, H. Matsubara, and K. Hiraki. Learning cooperative behavior in multi-agent environment — A case study of choice of play-plans in soccer. Lecture Notes in Computer Science, 1114:570–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loke:1996:CIA

[628] S. W. Loke, A. Davison, and L. Sterling. CIFI: An intelligent agent for citation finding on the World-Wide Web. Lecture Notes in Computer Science, 1114:580–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murase:1996:AMS

[629] Y. Murase, H. Matsubara, and Y. Hiraga. Automatic making of Sokoban problems. Lecture Notes in Computer Science, 1114:592–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cavedon:1996:BAR

[630] L. Cavedon and A. Rao. Bringing about rationality: Incorporating plans into a BDI agent architecture. Lecture Notes in Computer Science, 1114:601– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:AAF

[631] L.-F. Lee and A. Kean. An architecture for autonomous flying vehicles: a preliminary report. Lecture Notes in Computer Science, 1114:613–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Segawa:1996:FRT

[632] Y. Segawa, H. Sakai, T. Endoh, and K. Murakami. Face recognition through Hough transform for irises extraction and projection procedures for parts localization — for facial caricaturing system PICASSO. Lecture Notes in Computer Science, 1114:625– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoo:1996:EFR

[633] T.-W. Yoo and I.-S. Oh. Extraction of face region and features based on chromatic properties of human faces. Lecture Notes in Computer Science, 1114: 637–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsushita:1996:FPM

[634] M. Matsushita, M. Umano, I. Hatono, and H. Tamura. A fast patternmatching algorithm using matching candidates for production systems. Lecture Notes in Computer Science, 1114:646–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sowa:1996:PP

[635] J. F. Sowa. Processes and participants. Lecture Notes in Computer Science, 1115:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wille:1996:CSM

[636] R. Wille. Conceptual structures of multicontexts. Lecture Notes in Computer Science, 1115:23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1996:GDM

[637] P. Eades. Graph drawing methods. Lecture Notes in Computer Science, 1115:40–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehmann:1996:BPP

[638] F. Lehmann. Big posets of participatings and thematic roles. Lecture Notes in Computer Science, 1115:50– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stilman:1996:LGT

[639] B. Stilman. Linguistic geometry tools generate optimal solutions. Lecture Notes in Computer Science, 1115:75– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petermann:1996:NLT

[640] H. Petermann. Natural language text processing and the maximal join operator. Lecture Notes in Computer Science, 1115:100–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angelova:1996:DMK

[641] G. Angelova and K. Bontcheva. DB-MAT: Knowledge acquisition, processing and NL generation using conceptual graphs. Lecture Notes in Computer Science, 1115:115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huibers:1996:CGA

[642] T. Huibers, I. Ounis, and J.-P. Chevallet. Conceptual graph aboutness. Lecture Notes in Computer Science, 1115: 130–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1996:CSS

[643] P. Martin and L. Alpay. Conceptual structures and structured documents. Lecture Notes in Computer Science, 1115:145–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guinaldo:1996:CGI

[644] O. Guinaldo. Conceptual graphs isomorphism: Algorithm and use. Lecture Notes in Computer Science, 1115: 160–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moeller:1996:ECG

[645] J.-U. Moeller and D. Wiese. Editing conceptual graphs. Lecture Notes in Computer Science, 1115:175–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghosh:1996:CST

[646] B. Ghosh and V. Wuwongse. Computational situation theory in the conceptual graph language. Lecture Notes in Computer Science, 1115:188–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oehrstroem:1996:EGT

[647] P. Oehrstroem. Existential graphs and tense logic. Lecture Notes in Computer Science, 1115:203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delugach:1996:MAD

[648] H. S. Delugach and T. H. Hinke. Microanalysis: Acquiring database semantics in conceptual graphs. Lecture Notes in Computer Science, 1115:218–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esch:1996:PTD

[649] J. Esch and R. Levinson. Propagating truth and detecting contradiction in conceptual graph databases. Lecture Notes in Computer Science, 1115: 229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salvat:1996:SCF

[650] E. Salvat and M.-L. Mugnier. Sound and complete forward and backward chainings of graph rules. Lecture Notes in Computer Science, 1115:248–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuwongse:1996:TFC

[651] V. Wuwongse and Cao Hoang Tru. Towards fuzzy conceptual graph programs. Lecture Notes in Computer Science, 1115:263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fall:1996:STE

[652] A. Fall. Sparse term encoding for dynamic taxonomies. Lecture Notes in Computer Science, 1115:277–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarbo:1996:LE

[653] J. J. Sarbo. Lattice embedding. Lecture Notes in Computer Science, 1115: 293–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stumme:1996:LSC

[654] G. Stumme. Local scaling in conceptual data systems. Lecture Notes in Computer Science, 1115:308–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1996:MTS

[655] J. Hall. Management of telecommunication systems and services: Modelling and implementing TMN-Based multidomain management. Lecture Notes in Computer Science, 1116:xxi + 229, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagerup:1996:AIT

[656] T. Hagerup. Allocating independent tasks to parallel processors: An experimental study. Lecture Notes in Computer Science, 1117:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliker:1996:PIA

[657] L. Oliker, R. Biswas, and R. C. Strawn. Parallel implementation of an adaptive scheme for 3D unstructured grids on the SP2. Lecture Notes in Computer Science, 1117:35–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dearholt:1996:SLS

[658] W. Dearholt, S. Castillo, and G. Hennigan. Solution of large, sparse, irregular systems on a massively parallel computer. Lecture Notes in Computer Science, 1117:49–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deshpande:1996:PIS

[659] V. Deshpande, M. J. Grote, P. Messmer, and W. Sawyer. Parallel implementation of a sparse approximate inverse preconditioner. Lecture Notes in Computer Science, 1117:63–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Catalyuerek:1996:DIS

[660] U. V. Catalyuerek and C. Aykanat. Decomposing irregularly sparse matrices for parallel matrix-vector multiplication. Lecture Notes in Computer Science, 1117:75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1996:DSP

[661] H. D. Simon and A. Sohn. Dynamic spectral partitioning. Lecture Notes in Computer Science, 1117:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christou:1996:FDG

[662] I. T. Christou and R. R. Meyer. Fast distributed genetic algorithms for partitioning uniform grids. Lecture Notes in Computer Science, 1117:89– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorward:1996:TEU

[663] S. E. Dorward, L. R. Matheson, and R. E. Tarjan. Towards efficient unstructured multigrid preprocessing. Lecture Notes in Computer Science, 1117:105– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egeciouglu:1996:DDP

[664] O. Egeciouglu and A. Srinivasan. Domain decomposition for particle methods on the sphere. Lecture Notes in Computer Science, 1117:119–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Everaars:1996:CDP

[665] C. T. H. Everaars and F. Arbab. Coordination of distributed/parallel multiple-grid domain decomposition. Lecture Notes in Computer Science, 1117:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yelick:1996:SSI

[666] K. Yelick. Systems support for irregular parallel applications. Lecture Notes in Computer Science, 1117:145–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandy:1996:DOO

[667] J. A. Chandy, S. Parkes, and P. Banerjee. Distributed object oriented data structures and algorithms for VLSI CAD. Lecture Notes in Computer Science, 1117:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1996:PPR

[668] Y. Yu, O. H. Ibarra, and T. Yang. Parallel progressive radiosity with adaptive meshing. Lecture Notes in Computer Science, 1117:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunt:1996:LFE

[669] G. C. Hunt and R. C. Nelson. Lineal feature extraction by parallel stick growing. Lecture Notes in Computer Science, 1117:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traeff:1996:SPA

[670] J. L. Traeff and C. D. Zaroliagis. A simple parallel algorithm for the singlesource shortest path problem on planar digraphs. Lecture Notes in Computer Science, 1117:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeDinechin:1996:RVA

[671] F. De Dinechin, D. K. Wilde, S. Rajopadhye, and R. Andonov. A regular VLSI array for an irregular algorithm. Lecture Notes in Computer Science, 1117:195–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:DLS

[672] T. R. Smith. Digital libraries and spatial information processing. Lecture Notes in Computer Science, 1117:201– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fink:1996:FCM

[673] S. J. Fink, S. B. Baden, and S. R. Kohn. Flexible communication mechanisms for dynamic structured applications. Lecture Notes in Computer Science, 1117:203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:1996:MMM

[674] T. F. Gonzalez. Multi-message multicasting. Lecture Notes in Computer Science, 1117:217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diderich:1996:SSD

[675] C. G. Diderich and M. Gengler. Synchronization as a strategy for designing efficient parallel algorithms. Lecture Notes in Computer Science, 1117: 229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreira:1996:SDD

[676] J. E. Moreira, K. Eswar, R. B. Konuru, and V. K. Naik. Supporting dynamic data and processor repartitioning for irregular applications. Lecture Notes in Computer Science, 1117:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bik:1996:SQE

[677] A. J. C. Bik and H. A. G. Wijshoff. Simple quantitative experiments with a sparse compiler. Lecture Notes in Computer Science, 1117:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botorog:1996:UAS

[678] G. H. Botorog and H. Kuchen. Using algorithmic skeletons with dynamic data structures. Lecture Notes in Computer Science, 1117:263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shinano:1996:IDG

[679] Y. Shinano, M. Higaki, and R. Hirabayashi. An interface design for general parallel branch-and-bound algorithms. Lecture Notes in Computer Science, 1117:277– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1996:SIC

[680] R. Schreiber. Support for irregular computation in High Performance Fortran. Lecture Notes in Computer Science, 1117:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heun:1996:EDE

[681] V. Heun and E. W. Mayr. Efficient dynamic embedding of arbitrary binary trees into hypercubes. Lecture Notes in Computer Science, 1117:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watts:1996:PDL

[682] J. Watts, M. Rieffel, and S. Taylor. Practical dynamic load balancing for irregular problems. Lecture Notes in Computer Science, 1117:299–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamari:1996:MAP

[683] M. Lamari and W. F. De la Vega. The module allocation problem: An average case analysis. Lecture Notes in Computer Science, 1117:307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reimer:1996:DAD

[684] N. H. Reimer, S. U. Haenssgen, and W. F. Tichy. Dynamically adapting the degree of parallelism with reflexive programs. Lecture Notes in Computer Science, 1117:313–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grigni:1996:CGB

[685] M. Grigni and F. Manne. On the complexity of the generalized block distribution. Lecture Notes in Computer Science, 1117:319–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melab:1996:ALB

[686] N. Melab, N. Devesa, M. P. Lecouffe, and B. Toursel. Adaptive load balancing of irregular applications IDA0* applied to the 15-Puzzle problem. Lecture Notes in Computer Science, 1117: 327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cela:1996:MPA

[687] J. M. Cela, J. C. Duersteler, and J. Labarta. Manufacturing progressive addition lenses using distributed parallel processing. Lecture Notes in Computer Science, 1117:339–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenlaw:1996:PCR

[688] R. Greenlaw and J. Machta. The parallel complexity of randomized fractals. Lecture Notes in Computer Science, 1117:351–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdennadher:1996:CCH

[689] S. Abdennadher, T. Fruehwirth, and H. Meuss. On confluence of constraint handling rules. Lecture Notes in Computer Science, 1118:1–15, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Affane:1996:LAC

[690] M. S. Affane and H. Bennaceur. A labeling arc consistency method for functional constraints. Lecture Notes in Computer Science, 1118:16–30, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banerjee:1996:CSO

[691] D. Banerjee and J. Frank. Constraint satisfaction in optical routing for passive wavelength-routed networks. Lecture Notes in Computer Science, 1118:31–45, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bayardo:1996:UCL

[692] R. J. Bayardo and R. Schrag. Using CSP look-back techniques to solve exceptionally hard SAT instances. Lecture Notes in Computer Science, 1118: 46–60, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bessiere:1996:MCH

[693] C. Bessiere and J.-C. Regin. MAC and combined heuristics: Two reasons to forsake FC (and CBJ?) on hard problems. Lecture Notes in Computer Science, 1118:61–75, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charatonik:1996:IPC

[694] W. Charatonik and A. Podelski. The independence property of a class of set constraints. Lecture Notes in Computer Science, 1118:76–90, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1996:SCP

[695] B. M. W. Cheng, J. H. M. Lee, and J. C. K. Wu. Speeding up constraint propagation by redundant modeling. Lecture Notes in Computer Science, 1118:91–103, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiu:1996:CBI

[696] C. K. Chiu, C. M. Chou, J. H. M. Lee, and H. F. Leung. A constraintbased interactive train rescheduling tool. Lecture Notes in Computer Science, 1118:104–118, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1996:LSN

[697] D. A. Clark, J. Frank, I. P. Gent, and E. MacIntyre. Local search and the number of solutions. Lecture Notes in Computer Science, 1118:119–133, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1996:DCD

[698] D. Cohen, M. Gyssens, and P. Jeavons. Derivation of constraints and database relations. Lecture Notes in Computer Science, 1118:134–148, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colombani:1996:CPE

[699] Y. Colombani. Constraint programming: An efficient and practical approach to solving the job-shop problem. Lecture Notes in Computer Science, 1118:149–163, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ElSakkout:1996:IAC

[700] H. El Sakkout, M. G. Wallace, and E. B. Richard. An instance of adaptive constraint propagation. Lecture Notes in Computer Science, 1118:164–178, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gent:1996:ESD

[701] I. P. Gent, E. MacIntyre, P. Prosser, and B. M. Smith. An empirical study of dynamic variable ordering heuristics for the constraint satisfaction problem. Lecture Notes in Computer Science, 1118:179–193, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hao:1996:ESH

[702] J.-K. Hao and R. Dorne. Empirical studies of heuristic local search for constraint solving. Lecture Notes in Computer Science, 1118:194–208, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzbaur:1996:DCQ

[703] C. Holzbaur, F. Menezes, and P. Barahona. Defeasibility in CLP(Q) through generalized slack variables. Lecture Notes in Computer Science, 1118:209– 223, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooker:1996:IDB

[704] J. N. Hooker. Inference duality as a basis for sensitivity analysis. Lecture Notes in Computer Science, 1118:224– 236, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hosobe:1996:GLP

[705] H. Hosobe, S. Matsuoka, and A. Yonezawa. Generalized local propagation: a framework for solving constraint hierarchies. Lecture Notes in Computer Science, 1118:237–251, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jampel:1996:TBH

[706] M. Jampel, J.-M. Jacquet, D. Gilbert, and S. Hunt. Transformations between HCLP and PCSP. Lecture Notes in Computer Science, 1118:252–266, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeavons:1996:TT

[707] P. Jeavons, D. Cohen, and M. Gyssens. A test for tractability. Lecture Notes in Computer Science, 1118:267–281, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kepser:1996:CCS

[708] S. Kepser and K. U. Schulz. Combination of constraint systems II: Rational amalgamation. Lecture Notes in Computer Science, 1118:282–296, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koubarakis:1996:TDL

[709] M. Koubarakis. Tractable disjunctions of linear constraints. Lecture Notes in Computer Science, 1118:297–307, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larrosa:1996:EUD

[710] J. Larrosa and P. Meseguer. Exploiting the use of DAC in MAX-CSP. Lecture Notes in Computer Science, 1118:308–322, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1996:NAW

[711] Hoong Chuin Lau. A new approach for weighted constraint satisfaction: Theoretical and computational results. Lecture Notes in Computer Science, 1118:323–337, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:TME

[712] J. H. M. Lee, H.-F. Leung, and H.-W. Won. Towards a more efficient stochastic constraint solver. Lecture Notes in Computer Science, 1118:338–352, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pesant:1996:VLS

[713] G. Pesant and M. Gendreau. A view of local search in constraint programming. Lecture Notes in Computer Science, 1118:353–366, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rojas:1996:QSS

[714] M. C. R. Rojas. From quasi-solutions to solution: An evolutionary algorithm to solve CSP. Lecture Notes in Computer Science, 1118:367–381, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossi:1996:EVL

[715] F. Rossi. Existential variables and local consistency in finite domain constraint problems. Lecture Notes in Computer Science, 1118:382–396, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruet:1996:LSC

[716] P. Ruet. Logical semantics of concurrent constraint programming. Lecture Notes in Computer Science, 1118:397– 409, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sam-Haroud:1996:SNB

[717] D. Sam-Haroud and B. V. Faltings. Solving non-binary convex CSPs in continuous domains. Lecture Notes in Computer Science, 1118:410–424, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1996:ECT

[718] T. Suzuki, N. Kakinuma, and T. Tokuda. An experimental comparison of three modified DeltaBlue algorithms. Lecture Notes in Computer Science, 1118:425–435, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tinelli:1996:CLP

[719] C. Tinelli and M. Harandi. Constraint logic programming over unions of constraint theories. Lecture Notes in Computer Science, 1118:436–450, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Urbina:1996:AHS

[720] L. Urbina. Analysis of hybrid systems in CLP(R). Lecture Notes in Computer Science, 1118:451–467, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandeurzen:1996:QLL

[721] L. Vandeurzen, M. Gyssens, and D. Van Gucht. On query languages for linear queries definable with polynomial constraints. Lecture Notes in Computer Science, 1118:468–481, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1996:AHM

[722] R. J. Wallace. Analysis of heuristic methods for partial constraint satisfaction problems. Lecture Notes in Computer Science, 1118:482–496, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokoo:1996:SSP

[723] M. Yokoo, T. Suyama, and H. Sawada. Solving satisfiability problems using field programmable gate arrays: First results. Lecture Notes in Computer Science, 1118:497–509, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1996:CPS

[724] J. Zhou. A constraint program for solving the job-shop problem. Lecture Notes in Computer Science, 1118:510– 524, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1996:PPS

[725] P. Albers and J. Bellone. PSAP — A planning system for aircraft production. Lecture Notes in Computer Science, 1118:525–526, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Battle:1996:UPA

[726] S. A. Battle. Using partial arc consistency in a database environment. Lecture Notes in Computer Science, 1118:527–528, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouzoubaa:1996:FCH

[727] M. Bouzoubaa. Functional constraint hierarchies in CLP. Lecture Notes in Computer Science, 1118:529–530, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1996:TOF

[728] M. Carlsson, B. Carlson, and G. Ottosson. Towards an open finite domain constraint solver. Lecture Notes in Computer Science, 1118:531–532, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chmeiss:1996:ECP

[729] A. Chmeiss and P. Jegou. Efficient constraint propagation with good space complexity. Lecture Notes in Computer Science, 1118:533–534, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dalal:1996:ATR

[730] M. Dalal and Y. Feng. Anytime temporal reasoning: Preliminary report. Lecture Notes in Computer Science, 1118:535–536, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fages:1996:CMG

[731] F. Fages. From constraint minimization to goal optimization in CLP languages. Lecture Notes in Computer Science, 1118:537–538, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frost:1996:LFL

[732] D. Frost and R. Dechter. Looking at full looking ahead. Lecture Notes in Computer Science, 1118:539–540, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grant:1996:APC

[733] S. A. Grant and B. M. Smith. The arc and path consistency phase transitions. Lecture Notes in Computer Science, 1118:541–542, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hellinck:1996:ECC

[734] W. Hellinck. Experiences with combining constraint programming and discrete event simulation. Lecture Notes in Computer Science, 1118:543–544, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirayama:1996:HCL

[735] K. Hirayama. Hill-climbing with local consistency for solving distributed CSPs. Lecture Notes in Computer Science, 1118:545–546, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juengen:1996:AAM

[736] F. J. Juengen and W. Kowalczyk. Approximate algorithms for maximum utility problems. Lecture Notes in Computer Science, 1118:547–548, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamma:1996:MCL

[737] E. Lamma, P. Mello, and M. Milano. A meta constraint logic programming architecture. Lecture Notes in Computer Science, 1118:549–550, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merel:1996:ACC

[738] P.-P. Merel, Z. Habbas, F. Herrmann, and D. Singer. N-ary consistencies and constraint-based backtracking. Lecture Notes in Computer Science, 1118:551–552, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

NDong:1996:GBC

[739] S. N’Dong and M. Van Caneghem. Global behaviour for complex constraints. Lecture Notes in Computer Science, 1118:553–554, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rish:1996:GTH

[740] I. Rish and R. Dechter. To guess or to think? hybrid algorithms for SAT. Lecture Notes in Computer Science, 1118:555–556, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaechter:1996:LSS

[741] V. Schaechter. A local simplification scheme for cc programs. Lecture Notes in Computer Science, 1118:557–558, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solotorevsky:1996:EUB

[742] G. Solotorevsky. From evaluating upper bounds of the complexity of solving CSPs to finding all the solution of CSPs. Lecture Notes in Computer Science, 1118:559–560, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solotorevsky:1996:MSD

[743] G. Solotorevsky, E. Gudes, and A. Meisels. Modeling and solving distributed constraint satisfaction problems (DCSPs). Lecture Notes in Computer Science, 1118:561–562, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:SAS

[744] D. R. Smith and S. J. Westfold. Scheduling an asynchronously shared resource. Lecture Notes in Computer Science, 1118:563–564, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Urbina:1996:GRC

[745] L. Urbina. The generalized railroad crossing: Its symbolic analysis in CLP(R). Lecture Notes in Computer Science, 1118:565–567, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1996:SAS

[746] J. H. Y. Wong, K.-F. Ng, and H.-F. Leung. A stochastic approach to solving fuzzy constraint satisfaction problems. Lecture Notes in Computer Science, 1118:568–569, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nemhauser:1996:BPS

[747] G. L. Nemhauser. Branch-and-price for solving integer programs with a huge number of variables: Methods and applications. Lecture Notes in Computer Science, 1118:570, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldin:1996:CD

[748] D. Q. Goldin. Constraint databases. Lecture Notes in Computer Science, 1118:571, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mairson:1996:CTA

[749] H. G. Mairson. Complexity-theoretic aspects of programming language design. Lecture Notes in Computer Science, 1118:572, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:EBC

[750] B.-L. Chen, M.-T. Ko, and K.-W. Lih. Equitable and m-bounded coloring of split graphs. Lecture Notes in Computer Science, 1120:1–5, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rolland:1996:FCS

[751] P. Rolland. Four coloring for a subset of maximal planar graphs with minimum degree five. Lecture Notes in Computer Science, 1120:6–17, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsui:1996:EAE

[752] Y. Matsui and T. Matsui. Enumeration algorithm for the edge coloring problem on bipartite graphs. Lecture Notes in Computer Science, 1120:18–26, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1996:LRI

[753] W.-L. Hsu. On-line recognition of interval graphs in O(m + n log n) time. Lecture Notes in Computer Science, 1120: 27–38, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:CPI

[754] C. Chen and C.-C. Chang. Connected proper interval graphs and the guard problem in spiral polygons. Lecture Notes in Computer Science, 1120:39– 47, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeh:1996:WCD

[755] H.-G. Yeh and G. J. Chang. Weighted connected domination and Steiner trees in distance-hereditary graphs. Lecture Notes in Computer Science, 1120:48– 52, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bezrukov:1996:CST

[756] S. Bezrukov, F. Kaderali, and W. Poguntke. On central spanning trees of a graph. Lecture Notes in Computer Science, 1120:53–57, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1996:CBD

[757] C. Lin, J.-J. Lin, and T.-W. Shyu. Complete bipartite decompositions of crowns, with applications to complete directed graphs. Lecture Notes in Computer Science, 1120:58–66, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bampis:1996:FAH

[758] E. Bampis, P. Hell, Y. Manoussakis, and M. Rosenfeld. Finding an antidirected Hamiltonian path starting with a forward arc from a given vertex of a tournament. Lecture Notes in Computer Science, 1120:67–73, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deza:1996:CLG

[759] M. Deza and T. Huang. Complementary l1-Graphs and related combinatorial structures. Lecture Notes in Computer Science, 1120:74–90, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fukuda:1996:DDM

[760] K. Fukuda and A. Prodon. Double description method revisited. Lecture Notes in Computer Science, 1120:91– 111, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deza:1996:SDV

[761] A. Deza, M. Deza, and K. Fukuda. On skeletons, diameters and volumes of metric polyhedra. Lecture Notes in Computer Science, 1120:112–128, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caseau:1996:IBB

[762] Y. Caseau and F. Laburthe. Improving branch and bound for jobshop scheduling with constraint propagation. Lecture Notes in Computer Science, 1120:129–149, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klinz:1996:NES

[763] B. Klinz and G. J. Woeginger. A new efficiently solvable special case of the three-dimensional axial bottleneck assignment problem. Lecture Notes in Computer Science, 1120:150–162, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaam:1996:RNS

[764] J. Jaam. Ramsey numbers by stochastic algorithms with new heuristics. Lecture Notes in Computer Science, 1120: 163–181, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chedid:1996:HNN

[765] F. B. Chedid. On the hybrid neural network model for solving optimization problems. Lecture Notes in Computer Science, 1120:182–193, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demange:1996:CNC

[766] M. Demange and V. Paschos. Constructive — non-constructive approximation and maximum independent set problem. Lecture Notes in Computer Science, 1120:194–207, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takabatake:1996:WGA

[767] T. Takabatake. Weakly greedy algorithm and pair-delta-matroids. Lecture Notes in Computer Science, 1120:208– 217, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcus:1996:IMM

[768] K. Marcus and A. Sebo. On integer multiflows and metric packings in matroids. Lecture Notes in Computer Science, 1120:218–233, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1996:OAB

[769] T. C. Hu and J. D. Morgenthaler. Optimum alphabetic binary trees. Lecture Notes in Computer Science, 1120:234– 243, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sole:1996:BCD

[770] P. Sole and J.-P. Tillich. Block codes for dyadic phase shift keying. Lecture Notes in Computer Science, 1120:244– 262, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patrou:1996:ZCF

[771] B. Patrou. Zigzag codes and Z-free hulls. Lecture Notes in Computer Science, 1120:263–274, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouchitte:1996:CO

[772] V. Bouchitte, A. Hilali, R. Jegou, and J.-X. Rampon. Contiguity orders. Lecture Notes in Computer Science, 1120:275–287, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bekesi:1996:WCA

[773] J. Bekesi, G. Galambos, U. Pferschy, and G. J. Woeginger. Worstcase analysis for on-line data compression. Lecture Notes in Computer Science, 1120:288–300, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bermond:1996:GCG

[774] J.-C. Bermond, T. Kodate, and S. Perennes. Gossiping in Cayley graphs by packets. Lecture Notes in Computer Science, 1120:301–315, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreae:1996:EDT

[775] T. Andreae, M. Noelle, C. Rempel, and G. Schreiber. On embedding 2-dimensional toroidal grids into de Brujn graphs with clocked congestion one. Lecture Notes in Computer Science, 1120:316–327, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossouni:1996:CSM

[776] F. Mossouni and C. Lavault. N-cube string matching algorithm with long texts. Lecture Notes in Computer Science, 1120:328–340, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lennerstad:1996:CMS

[777] H. Lennerstad and L. Lundberg. Combinatorics for multiprocessor scheduling optimization and other contexts in computer architecture. Lecture Notes in Computer Science, 1120:341–347, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Euler:1996:SAC

[778] R. Euler and L. Lemarchand. Some applications of combinatorial optimization in parallel computing. Lecture Notes in Computer Science, 1120:348– 366, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bampis:1996:PCA

[779] E. Bampis, Y. Manoussakis, and I. Milis. On the parallel complexity of the alternating Hamiltonian cycle problem. Lecture Notes in Computer Science, 1120:367–377, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petreschi:1996:TGS

[780] R. Petreschi and A. Sterbini. Threshold graphs and synchronization protocols. Lecture Notes in Computer Science, 1120:378–395, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milis:1996:TAD

[781] I. Milis. Task assignment in distributed systems using network flow methods. Lecture Notes in Computer Science, 1120:396–405, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nace:1996:DRD

[782] D. Nace and J. Carlier. Distributed rerouting in DCS mesh networks. Lecture Notes in Computer Science, 1120: 406–415, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanfeliu:1996:ERC

[783] A. Sanfeliu and R. Alquezar. Efficient recognition of a class of contextsensitive languages described by augmented regular expressions. Lecture Notes in Computer Science, 1121:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oommen:1996:OIT

[784] B. J. Oommen and R. L. Kashyap. Optimal and information theoretic syntactic pattern recognition for traditional errors. Lecture Notes in Computer Science, 1121:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castro:1996:MGG

[785] M. J. Castro and F. Casacuberta. The morphic generator grammatical inference methodology and multilayer perceptrons: a hybrid approach to acoustic modeling. Lecture Notes in Computer Science, 1121:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amengual:1996:TDA

[786] J. C. Amengual and E. Vidal. Two different approaches for cost-efficient Viterbi parsing with error correction. Lecture Notes in Computer Science, 1121:30–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernau:1996:BPA

[787] H. Fernau and R. Freund. Bounded parallelism in array grammars used for character recognition. Lecture Notes in Computer Science, 1121:40–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1996:CBI

[788] J.-A. Sanchez, J.-M. Benedi, and F. Casacuberta. Comparison between the inside-outside algorithm and the Viterbi algorithm for stochastic context-free grammars. Lecture Notes in Computer Science, 1121:50– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agam:1996:GMO

[789] G. Agam and I. Dinstein. Generalized morphological operators applied to map-analysis. Lecture Notes in Computer Science, 1121:60–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cong:1996:DSS

[790] G. Cong and S. D. Ma. Derivatives in scale space. Lecture Notes in Computer Science, 1121:70–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1996:SI

[791] P. P. Jonker and O. Vermeij. On skeletonization in 4D images. Lecture Notes in Computer Science, 1121:79– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sperduti:1996:ECC

[792] A. Sperduti, D. Majidi, and A. Starita. Extended cascade-correlation for syntactic and structural pattern recognition. Lecture Notes in Computer Science, 1121:90–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imiya:1996:MPS

[793] A. Imiya, Y. Fujiwara, and T. Kawashima. A metric of planar self-similar forms. Lecture Notes in Computer Science, 1121:100–??, 1996. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1996:IGG

[794] X. Y. Jiang and H. Bunke. Including geometry in graph representations: a quadratic-time graph isomorphism algorithm and its applications. Lecture Notes in Computer Science, 1121:110–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1996:HDR

[795] R. Wilson and E. R. Hancock. Hierarchical discrete relaxation. Lecture Notes in Computer Science, 1121:120– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costa:1996:RI

[796] M. S. Costa and L. G. Shapiro. Relational indexing. Lecture Notes in Computer Science, 1121:130–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quint:1996:EMF

[797] F. Quint and M. Sties. An evidential merit function to guide search in a semantic network based image analysis system. Lecture Notes in Computer Science, 1121:140–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cross:1996:IGM

[798] A. D. J. Cross and E. R. Hancock. Inexact graph matching with genetic search. Lecture Notes in Computer Science, 1121:150–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanfeliu:1996:ARB

[799] A. Sanfeliu and M. Sainz. Automatic recognition of bidimensional models learned by grammatical inference in outdoors scenes. Lecture Notes in Computer Science, 1121:160–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:SDM

[800] J. W. Wang, C. H. Chen, and J. C. Luo. Signal decomposition by multiscale learning algorithms. Lecture Notes in Computer Science, 1121:170–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakagawa:1996:SLC

[801] M. Nakagawa and L. V. Tu. Structural learning of character patterns for on-line recognition of hand-written Japanese characters. Lecture Notes in Computer Science, 1121:180–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amin:1996:RHPa

[802] A. Amin, A. Rajithan, and P. Compton. Recognition of hand-printed characters using induct machine learning. Lecture Notes in Computer Science, 1121:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bollmann:1996:OCP

[803] M. Bollmann and B. Mertsching. Opponent color processing based on neural models. Lecture Notes in Computer Science, 1121:198–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perner:1996:KAS

[804] P. Perner, T. B. Belikova, and N. I. Yashunskaya. Knowledge acquisition by symbolic decision tree induction for interpretation of digital images in radiology. Lecture Notes in Computer Science, 1121:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGool:1996:IFS

[805] L. Van Gool. Invariants and fixed structures lead the way to change. Lecture Notes in Computer Science, 1121:220– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SannitidiBaja:1996:RSL

[806] G. Sanniti di Baja. Representing shape by line patterns. Lecture Notes in Computer Science, 1121:230–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1996:ROI

[807] B. Liu and P. Wang. Recognition of 3D objects from 2D images — some issues. Lecture Notes in Computer Science, 1121:240–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgefors:1996:SSV

[808] G. Borgefors, I. Nystroem, and G. Sanniti di Baja. Surface skeletonization of volume objects. Lecture Notes in Computer Science, 1121:251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsello:1996:PSA

[809] V. Matsello and M. Schlesinger. Peculiarities of structural analysis of image contours under various orders of scanning. Lecture Notes in Computer Science, 1121:260–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishida:1996:SAC

[810] H. Nishida. A structural analysis of curve deformation by discontinuous transformations. Lecture Notes in Computer Science, 1121:269–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milanova:1996:TDC

[811] M. Milanova, L. Nikolov, and S. Fotev. Three dimensional computer vision for computer aided design and manufacturing applications. Lecture Notes in Computer Science, 1121:279–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johannesen:1996:UWM

[812] F. R. Johannesen, S. Raaschou, O. V. Larsen, and P. Juergensen. Using weighted minutiae for fingerprint identification. Lecture Notes in Computer Science, 1121:289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sossa:1996:RDR

[813] J. H. A. Sossa and J. L. S. Diaz de Leon. Recognizing 2-D rigid and non-rigid wire-shapes. Lecture Notes in Computer Science, 1121:300–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tombre:1996:SSM

[814] K. Tombre. Structural and syntactic methods in line drawing analysis: To which extent do they work? Lecture Notes in Computer Science, 1121:310– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dori:1996:VBS

[815] D. Dori and L. Wenyin. Vectorbased segmentation of text connected to graphics in engineering drawings. Lecture Notes in Computer Science, 1121:322–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weiss:1996:ARO

[816] M. Weiss and D. Dori. Automatic resolution of object features from engineering drawings for 3D reconstruction. Lecture Notes in Computer Science, 1121:332–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinto:1996:FAC

[817] I. Pinto and H. Freeman. The feedback approach to cartographic areal text placement. Lecture Notes in Computer Science, 1121:341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:HRR

[818] J.-W. Chen and S.-Y. Lee. A hierarchical representation for the reference database of on-line Chinese character recognition. Lecture Notes in Computer Science, 1121:351–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rocha:1996:SSA

[819] J. Rocha and H. Fujisawa. Substructure shape analysis for Kanji character recognition. Lecture Notes in Computer Science, 1121:361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amin:1996:RHPb

[820] A. Amin, M. Bamford, A. Hoffmann, and A. Mahidadia. Recognition of hand-printed Chinese characters using ripple down rules. Lecture Notes in Computer Science, 1121:371–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malaviya:1996:FSM

[821] A. Malaviya and R. Klette. A fuzzy syntactic method for on-line handwriting recognition. Lecture Notes in Computer Science, 1121:381–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adleman:1996:CRP

[822] L. M. Adleman and M.-D. Huang. Counting rational points on curves and Abelian varieties over finite fields. Lecture Notes in Computer Science, 1122: 1–16, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belabas:1996:CCF

[823] K. Belabas. Computing cubic fields in quasi-linear time. Lecture Notes in Computer Science, 1122:17–26, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernstein:1996:FIA

[824] D. J. Bernstein. Fast ideal arithmetic via lazy localization. Lecture Notes in Computer Science, 1122:27–34, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brent:1996:CSA

[825] R. P. Brent, A. J. Van der Poorten, and H. J. Te Riele. A comparative study of algorithms for computing continued fractions of algebraic numbers. Lecture Notes in Computer Science, 1122:35–48, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1996:CRC

[826] H. Cohen, F. Diaz y. Diaz, and M. Olivier. Computing ray class groups, conductors and discriminants. Lecture Notes in Computer Science, 1122:49–48, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couveignes:1996:CIU

[827] J.-M. Couveignes. Computing ℓisogenies using the p-torsion. Lecture Notes in Computer Science, 1122:59– 66, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daberkow:1996:CHC

[828] M. Daberkow and M. E. Pohst. On computing Hilbert class fields of prime degree. Lecture Notes in Computer Science, 1122:67–74, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denny:1996:RCR

[829] T. F. Denny and V. Mueller. On the reduction of composed relations from the number field sieve. Lecture Notes in Computer Science, 1122:75–90, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dummit:1996:CAS

[830] D. S. Dummit and D. R. Hayes. Checking the p-adic stark conjecture where p is Archimedean. Lecture Notes in Computer Science, 1122:91–98, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elkenbracht-Huizing:1996:MPG

[831] M. Elkenbracht-Huizing. A multiple polynomial general number field sieve. Lecture Notes in Computer Science, 1122:99–114, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fermigier:1996:CHR

[832] S. Fermigier. Construction of high-rank elliptic curves over Q and Q(t) with non-trivial 2-Torsion. Lecture Notes in Computer Science, 1122:115–120, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fhlathuin:1996:HAV

[833] B. Fhlathuin. The height on an Abelian variety. Lecture Notes in Computer Science, 1122:121–132, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fieker:1996:LNF

[834] C. Fieker and M. E. Pohst. O lattices over number fields. Lecture Notes in Computer Science, 1122:133–140, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ford:1996:MDP

[835] D. Ford. Minimum discriminants of primitive sextic fields. Lecture Notes in Computer Science, 1122:141–144, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ford:1996:NAR

[836] D. Ford and G. Havas. A new algorithm and refined bounds for extended GCD computation. Lecture Notes in Computer Science, 1122:145–150, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaal:1996:ATE

[837] I. Gaal. Application of Thue equations to computing power integral bases in algebraic number fields. Lecture Notes in Computer Science, 1122:151–156, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gebel:1996:CIP

[838] J. Gebel, A. Pethoe, and H. G. Zimmer. Computing S-integral points on elliptic curves. Lecture Notes in Computer Science, 1122:157–172, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giesbrecht:1996:PCS

[839] M. Giesbrecht. Probabilistic computation of the Smith Normal Form of a sparse integer matrix. Lecture Notes in Computer Science, 1122:173–186, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lauter:1996:RCF

[840] K. Lauter. Ray class field constructions of curves over finite fields with many rational points. Lecture Notes in Computer Science, 1122:187–196, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lercier:1996:CIF

[841] R. Lercier. Computing isogenies in F2n. Lecture Notes in Computer Science, 1122:197–212, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Louboutin:1996:CTD

[842] S. Louboutin. A computational technique for determining relative class numbers of CM-Fields. Lecture Notes in Computer Science, 1122:213–216, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKee:1996:OND

[843] J. McKee and R. Pinch. Old and new deterministic factoring algorithms. Lecture Notes in Computer Science, 1122:217–224, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1996:EAC

[844] S. M. Meyer and J. P. Sorenson. Efficient algorithms for computing the Jacobi symbol. Lecture Notes in Computer Science, 1122:225–240, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niklasch:1996:NFD

[845] G. Niklasch. The number field database on the World Wide Web server. Lecture Notes in Computer Science, 1122:241– 242, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulus:1996:AST

[846] S. Paulus. An algorithm of subexponential type computing the class group of quadratic orders over principal ideal domains. Lecture Notes in Computer Science, 1122:243–258, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pohst:1996:CAK

[847] M. E. Pohst. Computational aspects of Kummer theory. Lecture Notes in Computer Science, 1122:259–272, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pohst:1996:IBR

[848] M. E. Pohst and M. Schoernig. On integral basis reduction in global function fields. Lecture Notes in Computer Science, 1122:273–282, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poonen:1996:CAC

[849] B. Poonen. Computational aspects of curves of genus at least 2. Lecture Notes in Computer Science, 1122:283–306, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roessner:1996:CAO

[850] C. Roessner and J.-P. Seifert. The complexity of approximate optima for greatest common divisor computations. Lecture Notes in Computer Science, 1122:307–322, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheidler:1996:CRR

[851] R. Scheidler. Compact representation in real quadratic congruence function fields. Lecture Notes in Computer Science, 1122:323–336, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schirokauer:1996:DLE

[852] O. Schirokauer, D. Weber, and T. Denny. Discrete logarithms: The effectiveness of the index calculus method. Lecture Notes in Computer Science, 1122:337–362, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smart:1996:HDI

[853] N. Smart. How difficult is it to solve a Thue equation? Lecture Notes in Computer Science, 1122:363–374, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stein:1996:ECF

[854] A. Stein. Elliptic congruence function fields. Lecture Notes in Computer Science, 1122:375–384, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsfasman:1996:AGL

[855] M. Tsfasman. Algebraic geometry lattices and codes. Lecture Notes in Computer Science, 1122:385–390, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1996:CDL

[856] D. Weber. Computing discrete logarithms with the general number field sieve. Lecture Notes in Computer Science, 1122:391–403, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foster:1996:HPD

[857] I. Foster. High-performance distributed computing: The I-WAY experiment and beyond. Lecture Notes in Computer Science, 1123:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerdoux-Jamet:1996:DIP

[858] P. Guerdoux-Jamet, D. Lavenier, C. Wagner, and P. Quinton. Design and implementation of a parallel architecture for biological sequence comparison. Lecture Notes in Computer Science, 1123:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McColl:1996:UC

[859] W. F. McColl. Universal computing. Lecture Notes in Computer Science, 1123:25–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahm:1996:DLB

[860] E. Rahm. Dynamic load balancing in parallel database systems. Lecture Notes in Computer Science, 1123:37– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hackstadt:1996:DAQ

[861] S. T. Hackstadt and A. D. Malony. Distributed array query and visualization for High Performance Fortran. Lecture Notes in Computer Science, 1123:55– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clemencon:1996:ASR

[862] C. Clemencon, A. Endo, J. Fritscher, and A. Mueller. Annai scalable runtime support for interactive debugging and performance analysis of large-scale parallel programs. Lecture Notes in Computer Science, 1123:64–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ronsse:1996:IRM

[863] M. A. Ronsse and L. J. Levrouw. On the implementation of a replay mechanism. Lecture Notes in Computer Science, 1123:70–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brunie:1996:CFD

[864] L. Brunie and O. Reymann. Concepts and functionalities of the DOSMOS-Trace monitoring tool. Lecture Notes in Computer Science, 1123:74–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludwig:1996:OMS

[865] T. Ludwig, M. Oberhuber, and R. Wismueller. An open monitoring system for parallel and distributed programs. Lecture Notes in Computer Science, 1123:78–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Friedman:1996:MEP

[866] R. Friedman, M. Goldin, A. Itzkovitz, and A. Schuster. Millipede: Easy parallel programming in available distributed environments. Lecture Notes in Computer Science, 1123:84–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hollingsworth:1996:ACS

[867] J. K. Hollingsworth and B. P. Miller. An adaptive cost system for parallel program instrumentation. Lecture Notes in Computer Science, 1123:88– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badouel:1996:SPT

[868] D. Badouel, T. Priol, and L. Renambot. SVMview: a performance tuning tool for DSM-Based parallel computers. Lecture Notes in Computer Science, 1123:98–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talbot:1996:CMI

[869] S. A. M. Talbot, A. J. Bennett, and P. H. J. Kelly. Cautious, machineindependent performance tuning for shared-memory multiprocessors. Lecture Notes in Computer Science, 1123: 106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cabillic:1996:DHS

[870] G. Cabillic and I. Puaut. Dealing with heterogeneity in Stardust: An environment for parallel programming on networks of heterogeneous workstations. Lecture Notes in Computer Science, 1123:114–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1996:IED

[871] K. Wolf and O. Kraemer-Fuhrmann. An integrated environment to design parallel object-oriented applications. Lecture Notes in Computer Science, 1123:120–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:1996:MEM

[872] A. Geist, W. Gropp, S. Huss-Lederman, and A. Lumsdaine. MPI-2: Extending the message-passing interface. Lecture Notes in Computer Science, 1123:128–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attali:1996:OSP

[873] I. Attali, D. Caromel, R. Guider, and A. L. Wendelborn. Optimizing Sisal programs: a formal approach. Lecture Notes in Computer Science, 1123:136– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nyland:1996:RMD

[874] L. Nyland, J. Prins, A. Goldberg, and P. Mills. A refinement methodology for developing data-parallel applications. Lecture Notes in Computer Science, 1123:145–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crivelli:1996:TPW

[875] S. A. Crivelli and E. R. Jessup. Task parallelism: What a tool can provide and what should be left to the user. Lecture Notes in Computer Science, 1123:151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prylli:1996:EBC

[876] L. Prylli and B. Tourancheau. Efficient block cyclic data redistribution. Lecture Notes in Computer Science, 1123:155–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desprez:1996:OGS

[877] F. Desprez, P. Ramet, and J. Roman. Optimal grain size computation for pipelined algorithms. Lecture Notes in Computer Science, 1123:165–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sueur:1996:DRH

[878] D. Sueur and J.-L. Dekeyser. Dynamic redistribution on heterogeneous parallel computers. Lecture Notes in Computer Science, 1123:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Addison:1996:SDS

[879] C. Addison, T. Oliver, and A. Sunderland. Supporting distributed sparse matrix objects. Lecture Notes in Computer Science, 1123:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welsh:1996:LLC

[880] M. Welsh, A. Basu, and T. Von Eicken. Low-latency communication over Fast Ethernet. Lecture Notes in Computer Science, 1123:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fulgham:1996:CIO

[881] M. L. Fulgham and L. Snyder. A comparison of input and output driven routers. Lecture Notes in Computer Science, 1123:195–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duato:1996:OTD

[882] J. Duato and M. P. Malumbres. Optimal topology for distributed shared memory multiprocessors: Hypercubes again? Lecture Notes in Computer Science, 1123:205–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rice:1996:PAR

[883] D. G. Rice, J. G. Delgado-Frias, and D. H. Summerville. A patternassociative router for interconnection network adaptive algorithms. Lecture Notes in Computer Science, 1123:213– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourdin:1996:SGO

[884] H. Bourdin, A. Ferreira, and K. Marcus. On stack-graph OPS-Based lightwave networks. Lecture Notes in Computer Science, 1123:218–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heun:1996:GME

[885] V. Heun and E. W. Mayr. A general method for efficient embeddings of graphs into optimal hypercubes. Lecture Notes in Computer Science, 1123: 222–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1996:SCS

[886] H. K. Dai. The size complexity of strictly non-blocking fixed ratio concentrators with constant depth. Lecture Notes in Computer Science, 1123: 234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barth:1996:BCM

[887] D. Barth. Bandwidth and cutwidth of the mesh of d-ary trees. Lecture Notes in Computer Science, 1123:243– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAzevedo:1996:VDE

[888] M. M. De Azevedo, N. Bagherzadeh, and S. Latifi. Variable-dilation embeddings of hypercubes into star graphs: Performance metrics, mapping functions, and routing. Lecture Notes in Computer Science, 1123:247–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCerio:1996:OCC

[889] L. D. De Cerio, M. Valero-Garcia, and A. Gonzalez. Overlapping communication and computation in hypercubes. Lecture Notes in Computer Science, 1123:253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiIanni:1996:EDR

[890] M. Di Ianni. Efficient delay routing. Lecture Notes in Computer Science, 1123:258–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaklamanis:1996:MHP

[891] C. Kaklamanis and D. Krizanc. Multipacket hot-potato routing on processor arrays. Lecture Notes in Computer Science, 1123:270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:NSC

[892] M.-K. Kim, H. Yoon, and S.-R. Maeng. A necessary and sufficient condition for proper routing in Omega-Omega network. Lecture Notes in Computer Science, 1123:278–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delorme:1996:RRP

[893] C. Delorme and P. Panaite. Rubik routing permutations on graphs. Lecture Notes in Computer Science, 1123:283– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balakrishnan:1996:EFC

[894] S. Balakrishnan and F. Oezguener. The effect of flow control and routing adaptivity on priority-driven traffic in multiprocessor networks. Lecture Notes in Computer Science, 1123:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MeyeraufderHeide:1996:RNO

[895] F. Meyer auf der Heide, K. Schroeder, and F. Schwarze. Routing on networks of optical crossbars. Lecture Notes in Computer Science, 1123:299–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petrini:1996:LBR

[896] F. Petrini and M. Vanneschi. Latency and bandwidth requirements of massively parallel programs: FFT as a case study. Lecture Notes in Computer Science, 1123:307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bermond:1996:IBA

[897] J.-C. Bermond, X. Muñoz, and A. Marchetti-Spaccamela. Induced broadcasting algorithms in iterated line digraphs. Lecture Notes in Computer Science, 1123:313–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perennes:1996:LBB

[898] S. Perennes. Lower bounds on broadcasting time of de Bruijn networks. Lecture Notes in Computer Science, 1123: 325–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laforest:1996:GTU

[899] C. Laforest. Gossip in trees under linecommunication mode. Lecture Notes in Computer Science, 1123:333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimakopoulos:1996:TEC

[900] V. V. Dimakopoulos and N. J. Dimopoulos. Total exchange in Cayley networks. Lecture Notes in Computer Science, 1123:341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimakopoulos:1996:LCC

[901] V. V. Dimakopoulos and N. J. Dimopoulos. Leaf communications in complete trees. Lecture Notes in Computer Science, 1123:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1996:GAB

[902] S. Fujita, C. Laforest, and S. Perennes. A gossip algorithm for bus networks with buses of limited length. Lecture Notes in Computer Science, 1123:353– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juurlink:1996:WHG

[903] B. H. H. Juurlink, P. S. Rao, and J. F. Sibeyn. Worm-hole gossiping on meshes. Lecture Notes in Computer Science, 1123:361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delmas:1996:CSG

[904] O. Delmas and S. Perennes. Circuitswitched gossiping in 3-dimensional torus networks. Lecture Notes in Computer Science, 1123:370–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lengauer:1996:APH

[905] C. Lengauer. Automatic parallelization and high performance compilers. Lecture Notes in Computer Science, 1123:377–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darte:1996:OAK

[906] A. Darte and F. Vivien. On the optimality of Allen and Kennedy’s algorithm for parallelism extraction in nested loops. Lecture Notes in Computer Science, 1123:379–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilde:1996:MRA

[907] D. Wilde and S. Rajopadhye. Memory reuse analysis in the polyhedral model. Lecture Notes in Computer Science, 1123:389–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okuda:1996:CSD

[908] K. Okuda. Cycle shrinking by dependence reduction. Lecture Notes in Computer Science, 1123:398–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jimenez:1996:UTT

[909] M. Jimenez, J. M. Llaberia, A. Fernandez, and E. Morancho. A unified transformation technique for multilevel blocking. Lecture Notes in Computer Science, 1123:402–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collard:1996:ADA

[910] J.-F. Collard and M. Griebl. Array dataflow analysis for explicitly parallel programs. Lecture Notes in Computer Science, 1123:406–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rinard:1996:SFC

[911] M. C. Rinard and P. C. Diniz. Semantic foundations of commutativity analysis. Lecture Notes in Computer Science, 1123:414–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthou:1996:AFA

[912] D. Barthou, J.-F. Collard, and P. Feautrier. Applications of fuzzy array dataflow analysis. Lecture Notes in Computer Science, 1123:424–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberhart:1996:SCI

[913] A. Eberhart and J. Li. Simplifying communication induced by operations on block-distributed arrays. Lecture Notes in Computer Science, 1123:428– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OBoyle:1996:CRI

[914] M. F. P. O’Boyle, R. W. Ford, and A. P. Nisbet. Compiler reduction of invalidation traffic in virtual shared memory systems. Lecture Notes in Computer Science, 1123:432–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knoop:1996:PDC

[915] J. Knoop. Partial dead code elimination for parallel programs. Lecture Notes in Computer Science, 1123:441– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diderich:1996:SCD

[916] C. G. Diderich and M. Gengler. Solving the constant-degree parallelism alignment problem. Lecture Notes in Computer Science, 1123:451–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besch:1996:TDM

[917] M. Besch and H. W. Pohl. Topographic data mapping by balanced hypersphere tesselation. Lecture Notes in Computer Science, 1123:455–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandes:1996:IPC

[918] T. Brandes and F. Desprez. Implementing pipelined computation and communication in an HPF compiler. Lecture Notes in Computer Science, 1123:459–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barreteau:1996:EMI

[919] M. Barreteau and P. Feautrier. Efficient mapping of interdependent scans. Lecture Notes in Computer Science, 1123:463–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griebl:1996:CLS

[920] M. Griebl and C. Langauer. Classifying loops for space-time mapping. Lecture Notes in Computer Science, 1123: 467–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cortes:1996:PCF

[921] T. Cortes, S. Girona, and J. Labarta. PACA: a cooperative file system cache for parallel machines. Lecture Notes in Computer Science, 1123:477–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:SFT

[922] J. A. Smith and S. K. Shrivastava. A system for fault-tolerant execution of data and compute intensive programs over a network of workstations. Lecture Notes in Computer Science, 1123:487–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kshemkalyani:1996:FVA

[923] A. D. Kshemkalyani. A framework for viewing atomic events in distributed computations. Lecture Notes in Computer Science, 1123:496–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoon:1996:WBP

[924] D.-K. Yoon and J.-L Gaudiot. Workerbased parallel computing on PVM. Lecture Notes in Computer Science, 1123: 506–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rowstron:1996:EDT

[925] A. Rowstron and A. Wood. An efficient distributed tuple space implementation for networks of workstations. Lecture Notes in Computer Science, 1123:510–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1996:HAP

[926] J.-W. Lin and S.-Y. Kuo. A highly available partition-processing protocol for distributed shared memory systems. Lecture Notes in Computer Science, 1123:514–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carretero:1996:DMP

[927] J. Carretero, F. Perez, P. De Miguel, and F. Garcia. I/O data mapping in ParFiSys: Support for highperformance I/O in parallel and distributed systems. Lecture Notes in Computer Science, 1123:522–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cholvi-Juan:1996:CPD

[928] V. Cholvi-Juan and J. M. Bernabeu-Auban. Correctness proof for a distributed memory system. Lecture Notes in Computer Science, 1123:526–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piquer:1996:DSM

[929] J. M. Piquer. Distributed shared memory based on group large causality. Lecture Notes in Computer Science, 1123:532–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bordawekar:1996:FIC

[930] R. Bordawekar, A. Choudhary, and J. Ramanujam. A framework for integrated communication and I/O placement. Lecture Notes in Computer Science, 1123:541–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1996:FDP

[931] Z. Hu, H. Iwasaki, and M. Takeichi. Formal derivation of parallel program for 2-dimensional maximum segment sum problem. Lecture Notes in Computer Science, 1123:553–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jegou:1996:MTE

[932] Y. Jegou. The migrating tasks: An execution model for irregular codes. Lecture Notes in Computer Science, 1123: 563–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coelho:1996:DHD

[933] F. Coelho. Discussing HPF design issues. Lecture Notes in Computer Science, 1123:571–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chin:1996:PCR

[934] W.-N. Chin, J. Darlington, and Y. Guo. Parallelizing conditional recurrences. Lecture Notes in Computer Science, 1123:579–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konaka:1996:ADP

[935] H. Konaka, Y. Itoh, T. Tomokiyo, and M. Maeda. Adaptive data parallel computation in the parallel object-oriented language OCore. Lecture Notes in Computer Science, 1123:587–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodriguez:1996:UIL

[936] N. Rodriguez, C. Ururahy, R. Ierusalimschy, and R. Cerqueira. The use of interpreted languages for implementing parallel algorithms on distributed systems. Lecture Notes in Computer Science, 1123:597–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Au:1996:COH

[937] P. Au, J. Darlington, M. Ghanem, and Y.-K Guo. Co-ordinating heterogeneous parallel computation. Lecture Notes in Computer Science, 1123:601– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreau:1996:CDM

[938] L. Moreau. Correctness of a distributed-memory model for Scheme. Lecture Notes in Computer Science, 1123:615–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hosoya:1996:PES

[939] H. Hosoya, N. Kobayashi, and A. Yonezawa. Partial evaluation scheme for concurrent languages and its correctness. Lecture Notes in Computer Science, 1123:625–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pandey:1996:SIE

[940] R. Pandey and J. C. Browne. Support for implementation of evolutionary concurrent systems in concurrent programming languages. Lecture Notes in Computer Science, 1123:633–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kale:1996:SDC

[941] L. V. Kale and M. A. Bhandarkar. Structured Dagger: a coordination language for message-driven programming. Lecture Notes in Computer Science, 1123:646–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruell:1996:TLT

[942] A. Bruell and H. Kuchen. TPascal — A language for task parallel programming. Lecture Notes in Computer Science, 1123:654–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lilius:1996:OPO

[943] J. Lilius. OB(PN)2: An object based Petri net programming notation. Lecture Notes in Computer Science, 1123: 660–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbab:1996:RCM

[944] F. Arbab, C. L. Blom, F. J. Burger, and C. T. H. Everaars. Reusable coordinator modules for massively concurrent applications. Lecture Notes in Computer Science, 1123:664–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michel:1996:IDD

[945] O. Michel. Introducing dynamicity in the data-parallel language 81/2. Lecture Notes in Computer Science, 1123: 678–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Douglas:1996:AGP

[946] A. Douglas, N. Roejemo, C. Runciman, and A. Wood. Astro-Gofer: Parallel functional programming with coordinating processes. Lecture Notes in Computer Science, 1123:686–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kefalas:1996:MPR

[947] P. Kefalas and I. Vlahavas. Multiple OR-Parallel resolution: Meta-level control of parallel logic programs. Lecture Notes in Computer Science, 1123:694– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knopp:1996:HLP

[948] J. Knopp. High level parallel programming based on automatic coordination. Lecture Notes in Computer Science, 1123:704–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breitinger:1996:EPF

[949] S. Breitinger, R. Loogen, Y. Ortega-Mallen, and R. Peña-Mari. Eden — the paradise of functional concurrent programming. Lecture Notes in Computer Science, 1123:710–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michel:1996:STD

[950] O. Michel. A straightforward translation of DOL systems in the declarative data-parallel language 81/2. Lecture Notes in Computer Science, 1123: 714–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botorog:1996:EPP

[951] G. H. Botorog and H. Kuchen. Efficient parallel programming with algorithmic skeletons. Lecture Notes in Computer Science, 1123:718–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeGuyadec:1996:LSE

[952] Y. Le Guyadec, E. Melin, B. Raffin, and X. Rebeuf. A loosely synchronized execution model for a simple dataparallel language. Lecture Notes in Computer Science, 1123:732–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shafarenko:1996:NAD

[953] A. Shafarenko. A nonannotative approach to distributed data-parallel computing. Lecture Notes in Computer Science, 1123:742–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Russo:1996:PNM

[954] S. Russo, C. Savy, I. Jelly, and P. Collingwood. Petri net modelling of PARSE designs. Lecture Notes in Computer Science, 1123:752–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puntigam:1996:SET

[955] F. Puntigam. Synchronization expressed in types of communication channels. Lecture Notes in Computer Science, 1123:762–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wray:1996:LDP

[956] J. P. Wray. Laws of data parallel assignment. Lecture Notes in Computer Science, 1123:770–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pena:1996:PPP

[957] R. Peña and L. A. Galán. Proving progress properties of non terminating programs under fairness assumptions. Lecture Notes in Computer Science, 1123:775–778, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferragina:1996:SPD

[958] P. Ferragina. A simple parallel dictionary matching algorithm. Lecture Notes in Computer Science, 1123:781– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Podehl:1996:SGI

[959] A. Podehl, T. Rauber, and G. Ruenger. Scalability and granularity issues of the hierarchical radiosity method. Lecture Notes in Computer Science, 1123:789– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1996:LRI

[960] J. F. Sibeyn. List ranking on interconnection networks. Lecture Notes in Computer Science, 1123:799–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Urland:1996:PAC

[961] E. Urland. Parallel algorithm for computing the fragment vector in Steiner triple systems. Lecture Notes in Computer Science, 1123:809–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nepomniaschaya:1996:RGA

[962] A. S. Nepomniaschaya. Representation of the Gabow algorithm for finding smallest spanning trees with a degree constraint on associative parallel processors. Lecture Notes in Computer Science, 1123:813–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shum:1996:RSR

[963] Kam Hong Shum and S.-Y. R. Li. Runtime support for replicated parallel simulators of an ATM network on workstation clusters. Lecture Notes in Computer Science, 1123:818–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1996:SMI

[964] T. Rauber, G. Ruenger, and C. Scholtes. Shared-memory implementation of an irregular particle simulation method. Lecture Notes in Computer Science, 1123:822–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boppana:1996:PAT

[965] V. Boppana, P. Saxena, P. Banerjee, and W. K. Fuchs. A parallel algorithm for the technology mapping of LUT-Based FPGAs. Lecture Notes in Computer Science, 1123:828–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moussouni:1996:DSM

[966] F. Moussouni and C. Lavault. Distributed string matching algorithm on the N-cube. Lecture Notes in Computer Science, 1123:832–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desprez:1996:OSL

[967] F. Desprez, S. Domas, and B. Tourancheau. Optimization of the ScaLAPACK LU factorization routine using communication/computation overlap. Lecture Notes in Computer Science, 1124:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbenz:1996:EPD

[968] P. Arbenz. On experiments with a parallel direct solver for diagonally dominant banded linear systems. Lecture Notes in Computer Science, 1124:11– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murphy:1996:CPE

[969] K. Murphy, M. Clint, M. Szularz, and J. Weston. The computation of partial eigensolutions on a distributed memory machine using a modified Lanczos method. Lecture Notes in Computer Science, 1124:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weston:1996:PCP

[970] J. Weston, M. Szularz, M. Clint, and K. Murphy. The parallel computation of partial eigensolutions of large matrices on a massively parallel processor. Lecture Notes in Computer Science, 1124:26–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dayde:1996:PSU

[971] M. J. Daydé, J.-Y. L’Excellent, and N. I. M. Gould. Preprocessing of sparse unassembled linear systems for efficient solution using element-by-element preconditioners. Lecture Notes in Computer Science, 1124:34–43, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:IPQ

[972] T. Y. Li and X. Zou. Implementing the parallel quasi-Laguerre’s algorithm for symmetric tridiagonal eigenproblems. Lecture Notes in Computer Science, 1124:44–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1996:CTD

[973] T. Rauber and G. Ruenger. Comparing task and data parallel execution schemes for the DIIRK method. Lecture Notes in Computer Science, 1124: 52–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huber:1996:NTS

[974] W. Huber. Numerical turbulence simulation on different parallel computers using the sparse grid combination method. Lecture Notes in Computer Science, 1124:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kessler:1996:PFM

[975] C. W. Kessler. Parallel Fourier-Motzkin elimination. Lecture Notes in Computer Science, 1124:66–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexandrov:1996:CTM

[976] V. N. Alexandrov and S. Lakka. Comparison of three Monte Carlo methods for matrix inversion. Lecture Notes in Computer Science, 1124:72–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahola:1996:PSV

[977] J. Rahola. Parallel solution of the volume integral equation of electromagnetic scattering. Lecture Notes in Computer Science, 1124:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graeb:1996:OPM

[978] R. Graeb, M. Guenther, U. Wever, and Q. Zheng. Optimization of parallel multilevel-Newton algorithms on workstation clusters. Lecture Notes in Computer Science, 1124:91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jezequel:1996:TSP

[979] F. Jezequel. A time and space parallel algorithm for the heat equation: The implicit collocation method. Lecture Notes in Computer Science, 1124: 97–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1996:PIV

[980] L. Davis, J.-M. Delosme, and F. Catthoor. Parallel image/video processing and computer arithmetic. Lecture Notes in Computer Science, 1124:103–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shock:1996:HPI

[981] C. T. Shock, C. Chang, L. Davis, and S. Goward. A high performance image database system for remotely sensed imagery. Lecture Notes in Computer Science, 1124:109–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chung:1996:APA

[982] Y. Chung and V. K. Prasanna. An asynchronous parallel algorithm for symbolic grouping operations in vision. Lecture Notes in Computer Science, 1124:123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guil:1996:PPH

[983] N. Guil and E. L. Zapata. A parallel pipelined Hough transform. Lecture Notes in Computer Science, 1124:131– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appiani:1996:HPS

[984] E. Appiani, M. Corvi, G. Garibotto, and C. Coelho. Highperformance SAR-Image formation and post-processing. Lecture Notes in Computer Science, 1124:139–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DellaVecchia:1996:PII

[985] G. Della Vecchia, R. Distasi, M. Nappi, and D. Vitulano. A parallel implementation of image coding using linear prediction and iterated functions systems. Lecture Notes in Computer Science, 1124:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uhl:1996:PAU

[986] A. Uhl. Parallel algorithms for using non-stationary MRA in image compression. Lecture Notes in Computer Science, 1124:151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antelo:1996:HRC

[987] E. Antelo, J. D. Bruguera, T. Lang, and J. Villalba. High radix CORDIC rotation based on selection by rounding. Lecture Notes in Computer Science, 1124:155–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MunkNielsen:1996:LAC

[988] A. Munk Nielsen and J.-M. Muller. Online algorithms for computing exponentials and logarithms. Lecture Notes in Computer Science, 1124:165–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1996:LAC

[989] A. Munk Nielsen and J.-M. Muller. Online algorithms for computing exponentials and logarithms. Lecture Notes in Computer Science, 1124:165–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frougny:1996:PLA

[990] C. Frougny. Parallel and on-line addition in negative base and some complex number systems. Lecture Notes in Computer Science, 1124:175–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oberman:1996:VLP

[991] S. F. Oberman and M. J. Flynn. A variable latency pipelined floating-point adder. Lecture Notes in Computer Science, 1124:183–192, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anuta:1996:BLA

[992] M. A. Anuta, D. W. Lozier, N. Schabanel, and P. R. Turner. Basic linear algebra operations in SLI arithmetic. Lecture Notes in Computer Science, 1124:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ikenaga:1996:CHP

[993] T. Ikenaga and T. Ogura. CAM2: a highly-parallel 2-D cellular automata architecture for real-time and palm-top pixel-level image processing. Lecture Notes in Computer Science, 1124:203– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crookes:1996:SOC

[994] D. Crookes, T. J. Brown, Y. Dong, and G. McAleese. A self-optimising coprocessor model for portable parallel image processing. Lecture Notes in Computer Science, 1124:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danckaert:1996:SLM

[995] K. Danckaert, F. Catthoor, and H. De Man. System-level memory management for weakly parallel image processing. Lecture Notes in Computer Science, 1124:217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verhaegh:1996:MPS

[996] W. F. J. Verhaegh, P. E. R. Lippens, E. H. L. Aarts, and J. L. Van Meerbergen. Multidimensional periodic scheduling: Model and complexity. Lecture Notes in Computer Science, 1124:226–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCoster:1996:GAC

[997] L. De Coster, M. Engels, R. Lauwereins, and J. A. Peperstraete. Global approach for compiled bit-true simulation of DSP systems. Lecture Notes in Computer Science, 1124:236–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demian:1996:PIR

[998] V. Demian, F. Desprez, H. Paugam-Moisy, and M. Pourzandi. Parallel implementation of RBF neural networks. Lecture Notes in Computer Science, 1124:243–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:SRP

[999] J. J. Dongarra, T. Hey, and E. Strohmaier. Selected results from the ParkBench benchmark. Lecture Notes in Computer Science, 1124:251–254, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BoierMartin:1996:ESP

[1000] I. M. Boier Martin and D. C. Marinescu. Exploiting symmetry in parallel computations for structural biology. Lecture Notes in Computer Science, 1124:255–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1996:OOP

[1001] K. Wolf, A. Mano, S. Prata dos Santos, and J.-M. Letteron. An object-oriented and parallel simulation of a powerplant. Lecture Notes in Computer Science, 1124:259–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1996:PSA

[1002] P. Albers and J. Bellone. A planning system for aircraft production with parallel constraint logic programming. Lecture Notes in Computer Science, 1124:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turton:1996:MOF

[1003] I. Turton and S. Openshaw. Modelling and optimising flows using parallel spatial interaction models. Lecture Notes in Computer Science, 1124: 270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibbons:1996:QRQ

[1004] P. B. Gibbons, Y. Matias, and V. Ramachandran. The queue-read queuewrite asynchronous PRAM model. Lecture Notes in Computer Science, 1124: 279–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacKenzie:1996:EPO

[1005] P. D. MacKenzie and V. Ramachandran. ERCW PRAMs and optical communication. Lecture Notes in Computer Science, 1124:293–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leppaenen:1996:GTP

[1006] V. Leppaenen. Goodness of timeprocessor optimal PRAM simulations. Lecture Notes in Computer Science, 1124:303–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kautonen:1996:SPC

[1007] A. Kautonen, V. Leppaenen, and M. Penttonen. Simulations of PRAM on complete optical networks. Lecture Notes in Computer Science, 1124:307– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nibhanupudi:1996:APB

[1008] M. V. Nibhanupudi and B. K. Szymanski. Adaptive parallelism in the bulksynchronous parallel model. Lecture Notes in Computer Science, 1124:311– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nash:1996:IIR

[1009] J. M. Nash, P. M. Dew, J. R. Davy, and M. E. Dyer. Implementation issues relating to the WPRAM model for scalable computing. Lecture Notes in Computer Science, 1124:319–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiskin:1996:BSP

[1010] A. Tiskin. The bulk-synchronous parallel random access machine. Lecture Notes in Computer Science, 1124:327– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juurlink:1996:BMI

[1011] B. H. H. Juurlink and H. A. G. Wijshoff. The E-BSP model: Incorporating general locality and unbalanced communication into the BSP model. Lecture Notes in Computer Science, 1124:339–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerbessiotis:1996:CED

[1012] A. V. Gerbessiotis and C. J. Siniolakis. Communication efficient data structures on the BSP model with applications in computational geometry. Lecture Notes in Computer Science, 1124:348–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLaTorre:1996:SLB

[1013] P. De La Torre and C. P. Kruskal. Submachine locality in the bulk synchronous setting. Lecture Notes in Computer Science, 1124:352–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1996:ALB

[1014] J. He, Q. Miller, and L. Chen. Algebraic laws for BSP programming. Lecture Notes in Computer Science, 1124: 359–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeumker:1996:RPA

[1015] A. Baeumker, W. Dittrich, F. Meyer, and A. Der Heide. Realistic parallel algorithms: Priority queue operations and selection for the BSP0* model. Lecture Notes in Computer Science, 1124:369–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verley:1996:MPL

[1016] G. Verley and J.-P Asselin De Beauville. Multilayer perceptron learning control. Lecture Notes in Computer Science, 1124:377–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoica:1996:EHM

[1017] I. Stoica, F. Sultan, and D. Keyes. Evaluating the hyperbolic model on a variety of architectures. Lecture Notes in Computer Science, 1124:387– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGemund:1996:SMP

[1018] A. J. C. Van Gemund. SPC: a model of parallel computation. Lecture Notes in Computer Science, 1124:397–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorlatch:1996:SEP

[1019] S. Gorlatch. Systematic efficient parallelization of scan and other list homomorphisms. Lecture Notes in Computer Science, 1124:401–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hains:1996:ASD

[1020] G. Hains and J. Mullins. Array structures and data-parallel algorithms. Lecture Notes in Computer Science, 1124: 409–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rangaswami:1996:CTC

[1021] R. Rangaswami. Compile-time cost analysis for parallel programming. Lecture Notes in Computer Science, 1124: 417–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formella:1996:HHP

[1022] A. Formella, J. Keller, and T. Walle.

HPP: a high performance PRAM. Lecture Notes in Computer Science, 1124: 425–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kong:1996:RIP

[1023] J. Kong and G. Lee. Relaxing the inclusion property in cache only memory architecture. Lecture Notes in Computer Science, 1124:435–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennett:1996:UPR

[1024] A. J. Bennett, P. H. J. Kelly, J. G. Refstrup, and S. A. M. Talbot. Using proxies to reduce controller contention in large shared-memory multiprocessors. Lecture Notes in Computer Science, 1124:445–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Risau:1996:RAW

[1025] J. Risau, A. Mikschl, and W. Damm. A RISC approach to weak cache coherence. Lecture Notes in Computer Science, 1124:453–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Betzos:1996:OCA

[1026] G. A. Betzos and P. A. Mitkas. 3D optoelectronic computer architectures for the conjugate gradient and multigrid benchmark algorithms. Lecture Notes in Computer Science, 1124:457– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikschl:1996:MMS

[1027] A. Mikschl and W. Datum. MSparc: a multithreaded Sparc. Lecture Notes in Computer Science, 1124:461–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strey:1996:NCP

[1028] A. Strey and N. Avellana. A new concept for parallel neurocomputer architectures. Lecture Notes in Computer Science, 1124:470–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fey:1996:TDV

[1029] D. Fey. Transformation of a 2-D VLSI systolic adder circuit in 3-D circuits using optical interconnections. Lecture Notes in Computer Science, 1124:478– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tumuluri:1996:SSL

[1030] C. Tumuluri and A. N. Choudhary. Scalable software latency hiding schemes: Evaluation of the poststore and prefetch options. Lecture Notes in Computer Science, 1124:486–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1996:RCO

[1031] S. Cho and G. Lee. Reducing coherence overhead in shared-bus multiprocessors. Lecture Notes in Computer Science, 1124:492–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeGoueslierdArgence:1996:AOA

[1032] P. Le Goueslier d’Argence. An asymptotically optimal affine schedule on bounded convex polyhedric domains. Lecture Notes in Computer Science, 1124:501–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasperoni:1996:LSP

[1033] F. Gasperoni and U. Schwiegelshohn.

List scheduling in the presence of branches — A theoretical evaluation. Lecture Notes in Computer Science, 1124:515–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rapine:1996:IAC

[1034] C. Rapine and D. Trystram. Iterative approach for the clustering problem. Lecture Notes in Computer Science, 1124:527–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benaini:1996:CTT

[1035] A. Benaini and D. Laiymani. Compiletime task scheduling for multi-phase programming. Lecture Notes in Computer Science, 1124:535–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lahlou:1996:SUP

[1036] C. Lahlou. Scheduling with unit processing and communication times on a ring network: Approximation results. Lecture Notes in Computer Science, 1124:539–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1996:EPA

[1037] E. W. Mayr and H. Stadtherr. Efficient parallel algorithms for scheduling with tree precedence constraints. Lecture Notes in Computer Science, 1124: 543–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calinescu:1996:BSP

[1038] R. Calinescu. Bulk synchronous parallel scheduling of uniform dags. Lecture Notes in Computer Science, 1124: 555–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:GMS

[1039] Z. Li and C. C. Kirkham. Generalized multiprocessor scheduling. Lecture Notes in Computer Science, 1124:563– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oh:1996:SSH

[1040] H. Oh and S. Ha. A static scheduling heuristic for heterogeneous processors. Lecture Notes in Computer Science, 1124:573–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giannakos:1996:CSP

[1041] A. Giannakos, J.-C. Koenig, and A. Munier. On the cyclic scheduling problem with small communication delays. Lecture Notes in Computer Science, 1124:578–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naroska:1996:NSM

[1042] E. Naroska and U. Schwiegelshohn.

A new scheduling method for parallel discrete-event simulation. Lecture Notes in Computer Science, 1124:582– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barcaccia:1996:MLS

[1043] P. Barcaccia, M. A. Bonuccelli, and M. Di Ianni. Minimum length scheduling of precedence constrained messages in distributed systems. Lecture Notes in Computer Science, 1124:594–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loewe:1996:OPP

[1044] W. Loewe, J. Eisenbiegler, and W. Zimmermann. Optimization of parallel programs on machines with expensive communication. Lecture Notes in Computer Science, 1124:602–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:ESL

[1045] H.-L. Chen and C.-T. King. Eager scheduling with lazy retry for dynamic task scheduling. Lecture Notes in Computer Science, 1124:611–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnor:1996:LML

[1046] B. Schnor, S. Petri, and H. Langendoerfer. Load management for load balancing on heterogeneous platforms: a comparison of traditional and neural network based approaches. Lecture Notes in Computer Science, 1124:615– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konuru:1996:AAD

[1047] R. B. Konuru, J. E. Moreira, and V. K. Naik. Application-assisted dynamic scheduling on large-scale multicomputer systems. Lecture Notes in Computer Science, 1124:621–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Backschat:1996:EBD

[1048] M. Backschat, A. Pfaffinger, and C. Zenger. Economic-based dynamic load distribution in large workstation networks. Lecture Notes in Computer Science, 1124:631–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1996:FSN

[1049] K. Shen and M. Hermenegildo. Flexible scheduling for non-deterministic, And-parallel execution of logic programs. Lecture Notes in Computer Science, 1124:635–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lanet:1996:LBT

[1050] J.-L. Lanet. A load balancing task allocation scheme in a hard real time system. Lecture Notes in Computer Science, 1124:640–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martorell:1996:LIN

[1051] X. Martorell, J. Labarta, N. Navarro, and E. Ayguade. A library implementation of the nano-threads programming model. Lecture Notes in Computer Science, 1124:644–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colombet:1996:SEL

[1052] L. Colombet and L. Desbat. Speedup and efficiency of large size applications on heterogeneous networks. Lecture Notes in Computer Science, 1124:653– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Labarta:1996:DPP

[1053] J. Labarta, S. Girona, V. Pillet, and T. Cortes. DiP: a parallel program development environment. Lecture Notes in Computer Science, 1124:665–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1996:APP

[1054] J. Simon and J.-M. Wierum. Accurate performance prediction for massively parallel systems and its applications. Lecture Notes in Computer Science, 1124:675–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cornu:1996:CCT

[1055] T. Cornu and M. Pahud. Contention in the Cray T3D communication network. Lecture Notes in Computer Science, 1124:689–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1996:TPT

[1056] J. M. D. Hill, P. I. Crumpton, and D. A. Burgess. Theory, practice, and a tool for BSP performance prediction. Lecture Notes in Computer Science, 1124:697–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joe:1996:ASM

[1057] K. Joe and A. Fukuda. Applying the semi-Markov memory and cache coherence interference model to an updating based cache coherence protocol. Lecture Notes in Computer Science, 1124:706–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vlassov:1996:AMM

[1058] V. Vlassov and L.-E. Thorelli. Analytical models of multithreading with data prefetching. Lecture Notes in Computer Science, 1124:714–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1996:IIR

[1059] M. J. Fernandez, M. Carro, and M. Hermenegildo. IDRA (IDeal resource allocation): Computing ideal speedups in parallel logic programming. Lecture Notes in Computer Science, 1124:724–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brilman:1996:ETS

[1060] M. Brilman and J.-M. Vincent. Estimation of the throughput for some stochastic resources sharing systems. Lecture Notes in Computer Science, 1124:734–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sharma:1996:SCF

[1061] V. Sharma and E. A. Varvarigos. Some closed form results for circuit switching in a hypercube network. Lecture Notes in Computer Science, 1124:738– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1996:IFM

[1062] S. Wallace and N. Bagherzadeh. Instruction fetching mechanisms for superscalar microprocessors. Lecture Notes in Computer Science, 1124:747– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:DDT

[1063] C.-M. Chen and C.-T. King. Designing dynamic two-level branch predictors based on pattern locality. Lecture Notes in Computer Science, 1124:757– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Temam:1996:SP

[1064] O. Temam. Streaming prefetch. Lecture Notes in Computer Science, 1124: 765–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filho:1996:FDS

[1065] E. M. C. Filho, E. S. T. Fernandes, and A. Wolfe. Functionality distribution on a superscalar architecture. Lecture Notes in Computer Science, 1124:773–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Potter:1996:ILF

[1066] R. Potter and G. Steven. Investigating the limits of fine-grained parallelim in a statically scheduled superscalar architecture. Lecture Notes in Computer Science, 1124:779–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goossens:1996:CM

[1067] B. Goossens and Duc Thang Vu. Onchip multiprocessing. Lecture Notes in Computer Science, 1124:789–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sigmund:1996:IBM

[1068] U. Sigmund and T. Ungerer. Identifying bottlenecks in a multithreaded superscalar microprocessor. Lecture Notes in Computer Science, 1124:797– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lo:1996:AOM

[1069] R. Lo, S. Chan, J. Dehnert, and R. Towle. Aggregate operation movement: a min-cut approach to global code motion. Lecture Notes in Computer Science, 1124:801–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1996:GIS

[1070] S. Schmidt. Global instruction scheduling — A practical approach. Lecture Notes in Computer Science, 1124:815– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winckel:1996:GMC

[1071] L. Winckel. Génération de micro-code parallèle pour la carte coprocesseur Rapid-2. Lecture Notes in Computer Science, 1124:819–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1996:RNM

[1072] F. Sanchez and J. Cortadella. RE-SIS: a new methodology for register optimization in software pipelining. Lecture Notes in Computer Science, 1124:824–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altman:1996:OSP

[1073] E. R. Altman and G. R. Gao. Optimal software pipelining through enumeration of schedules. Lecture Notes in Computer Science, 1124:833–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1996:TGP

[1074] C.-H. Cho and J.-T. Wang. Triangular grid protocol: An efficient scheme for replica control with uniform access quorums. Lecture Notes in Computer Science, 1124:843–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gesmann:1996:MPC

[1075] M. Gesmann. Mapping a parallel complex-object DBMS to operating system processes. Lecture Notes in Computer Science, 1124:852–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devirmis:1996:TMM

[1076] T. Devirmis and O. Ulusoy. A transaction model for multidatabase systems. Lecture Notes in Computer Science, 1124:862–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrena:1996:MDD

[1077] M. Barrena, J. Hernandez, J. M. Martinez, and A. Polo. Multidimensional declustering methods for parallel database systems. Lecture Notes in Computer Science, 1124:866– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spiliopoulou:1996:MRU

[1078] M. Spiliopoulou and J. C. Freytag. Modelling resource utilization in pipelined query execution. Lecture Notes in Computer Science, 1124:872– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozas:1996:TSS

[1079] G. Bozas, M. Jaedicke, A. Listl, and B. Mitschang. On transforming a sequential SQL-DBMS into a parallel one: First results and experiences of the MIDAS project. Lecture Notes in Computer Science, 1124:881–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brunie:1996:DPR

[1080] L. Brunie and H. Kosch. DPLGraphs — A powerful representation of parallel relational query execution plans. Lecture Notes in Computer Science, 1124: 887–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dikenelli:1996:BVM

[1081] O. Dikenelli, M. O. Uenalyr, and E. Ozkarahan. BLOCKER: a variable and multiattribute declustering for parallel database machines. Lecture Notes in Computer Science, 1124:892– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krotz-Vogel:1996:PPP

[1082] W. Krotz-Vogel and H.-C Hoppe. The PALLAS portable parallel programming environment. Lecture Notes in Computer Science, 1124:899–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marrakchi:1996:OPA

[1083] M. Marrakchi. On optimal parallel algorithm for Gaussian elimination. Lecture Notes in Computer Science, 1124:907–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdallah:1996:SMP

[1084] A. E. Abdallah. Synthesis of massively pipelined algorithms for list manipulation. Lecture Notes in Computer Science, 1124:911–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agerholm:1996:TSV

[1085] S. Agerholm. Translating specifications in VDM-SL to PVS. Lecture Notes in Computer Science, 1125:1–16, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agerholm:1996:CHA

[1086] S. Agerholm, I. Beylin, and P. Dybjer. A comparison of HOL and ALF formalizations of a categorical coherence theorem. Lecture Notes in Computer Science, 1125:17–32, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1996:MHS

[1087] D. Basin and S. Friedrich. Modeling a hardware synthesis methodology in Isabelle. Lecture Notes in Computer Science, 1125:33–50, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Black:1996:IRP

[1088] P. E. Black and P. J. Windley. Inference rules for programming languages with side effects in expressions. Lecture Notes in Computer Science, 1125:51–60, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brackin:1996:DCP

[1089] S. H. Brackin. Deciding cryptographic protocol adequacy with HOL: The implementation. Lecture Notes in Computer Science, 1125:61–76, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busch:1996:PLF

[1090] H. Busch. Proving liveness of fair transition systems. Lecture Notes in Computer Science, 1125:77–92, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1996:PDU

[1091] M. Butler and T. Laangbacka. Program derivation using the refinement calculator. Lecture Notes in Computer Science, 1125:93–108, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collins:1996:PTR

[1092] G. Collins. A proof tool for reasoning about functional programs. Lecture Notes in Computer Science, 1125:109– 124, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coupet-Grimal:1996:CHV

[1093] S. Coupet-Grimal and L. Jakubiec. Coq and hardware verification: a case study. Lecture Notes in Computer Science, 1125:125–140, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dutertre:1996:EMA

[1094] B. Dutertre. Elements of mathematical analysis in PVS. Lecture Notes in Computer Science, 1125:141–156, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eisenbiegler:1996:IIA

[1095] D. Eisenbiegler, C. Blumenroehr, and R. Kumar. Implementation issues about the embedding of existing high level synthesis algorithms in HOL. Lecture Notes in Computer Science, 1125: 157–172, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gordon:1996:FAA

[1096] A. D. Gordon and T. Melham. Five axioms of alpha-conversion. Lecture Notes in Computer Science, 1125:173–190, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gordon:1996:STH

[1097] M. Gordon. Set theory, higher order logic or both? Lecture Notes in Computer Science, 1125:191–202, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1996:MMH

[1098] J. Harrison. A Mizar mode for HOL. Lecture Notes in Computer Science, 1125:203–220, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1996:SAH

[1099] J. Harrison. Staalmarck’s algorithm as a HOL derived rule. Lecture Notes in Computer Science, 1125:221–234, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heckman:1996:TAC

[1100] M. R. Heckman, G. Zhang, B. R.

Becker, and D. Peticolas. Towards applying the composition principle to verify a microkernel operating system. Lecture Notes in Computer Science, 1125:235–250, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heyd:1996:MCU

[1101] B. Heyd and P. Cregut. A modular coding of unity in Coq. Lecture Notes in Computer Science, 1125:251–266, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howe:1996:IMH

[1102] D. J. Howe. Importing mathematics from HOL into Nuprl. Lecture Notes in Computer Science, 1125:267–282, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolyang:1996:SPE

[1103] H. T. Kolyang, T. Santen, and B. Wolff. A structure preserving encoding of Z in isabelle/HOL. Lecture Notes in Computer Science, 1125:283–298, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsson:1996:IRH

[1104] M. Larsson. Improving the result of high-level synthesis using interactive transformational design. Lecture Notes in Computer Science, 1125:299–314, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laibinis:1996:ULT

[1105] L. Laibinis. Using lattice theory in higher order logic. Lecture Notes in Computer Science, 1125:315–330, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nazareth:1996:FVA

[1106] D. Nazareth and T. Nipkow. Formal verification of algorithm W: The monomorphic case. Lecture Notes in Computer Science, 1125:331–346, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pusch:1996:VCC

[1107] C. Pusch. Verification of compiler correctness for the WAM. Lecture Notes in Computer Science, 1125:347–362, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reus:1996:SDT

[1108] B. Reus. Synthetic domain theory in type theory: Another logic of computable functions. Lecture Notes in Computer Science, 1125:363–380, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slind:1996:FDH

[1109] K. Slind. Function definition in higher order logic. Lecture Notes in Computer Science, 1125:381–398, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smaill:1996:HOA

[1110] A. Smaill and I. Green. High order annotated terms for proof search. Lecture Notes in Computer Science, 1125:399– 414, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tahar:1996:CMH

[1111] S. Tahar and P. Curzon. A comparison of MDG and HOL for hardware verification. Lecture Notes in Computer Science, 1125:415–430, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zammit:1996:MCT

[1112] V. Zammit. A mechanisation of computability theory in HOL. Lecture Notes in Computer Science, 1125:431– 446, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgartner:1996:HT

[1113] P. Baumgartner, U. Furbach, and I. Niemelae. Hyper tableaux. Lecture Notes in Computer Science, 1126:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNivelle:1996:ARU

[1114] H. De Nivelle. An algorithm for retrieval of unifiers from discrimination trees. Lecture Notes in Computer Science, 1126:18–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourely:1996:BPC

[1115] C. Bourely, G. Defourneaux, and N. Peltier. Building proofs or counterexamples by analogy in a resolution framework. Lecture Notes in Computer Science, 1126:34–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degtyarev:1996:WYA

[1116] A. Degtyarev and A. Voronkov. What you always wanted to know about rigid E-unification. Lecture Notes in Computer Science, 1126:50–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artosi:1996:LPQ

[1117] A. Artosi, P. Benassi, G. Governatori, and A. Rotolo. Labelled proofs for quantified modal logic. Lecture Notes in Computer Science, 1126:70– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donini:1996:UTM

[1118] F. M. Donini, F. Massacci, D. Nardi, and R. Rosati. A uniform tableaux method for nonmonotonic modal logics. Lecture Notes in Computer Science, 1126:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froehlich:1996:DID

[1119] P. Froehlich, W. Nejdl, and M. Schroeder. Design and implementation of diagnostic strategies using modal logic. Lecture Notes in Computer Science, 1126: 104–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1996:MAL

[1120] F. Santos and J. Carmo. A modal action logic based framework for organization specification and analysis. Lecture Notes in Computer Science, 1126: 119–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Genesereth:1996:MI

[1121] M. R. Genesereth. McCarthy’s idea.

Lecture Notes in Computer Science, 1126:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferes:1996:SEN

[1122] J. J. Alferes, L. Moniz Pereira, and T. C. Przymusinski. Strong and explicit negation in nonmonotonic reasoning and logic programming. Lecture Notes in Computer Science, 1126: 143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1996:OPM

[1123] J. Engelfriet. Only persistence makes nonmonotonicity monotonous. Lecture Notes in Computer Science, 1126:164– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgatos:1996:OBR

[1124] K. Georgatos. Ordering-based representations of rational inference. Lecture Notes in Computer Science, 1126: 176–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikitiuk:1996:SRD

[1125] A. Mikitiuk. Semi-representability of default theories in rational default logic. Lecture Notes in Computer Science, 1126:192–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciorba:1996:QAA

[1126] V. Ciorba. A query answering algorithm for Lukaszewicz’ general open default theory. Lecture Notes in Computer Science, 1126:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1996:IDL

[1127] J. Engelfriet, V. W. Marek, J. Treur, and M. Truszczynski. Infinitary default logic for specification of nonmonotonic reasoning. Lecture Notes in Computer Science, 1126:224–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1996:SCC

[1128] G. Antoniou, A. P. Courtney, J. Ernst, and M. A. Williams. A system for computing constrained default logic extensions. Lecture Notes in Computer Science, 1126:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolka:1996:OPM

[1129] G. Smolka. The Oz programming model. Lecture Notes in Computer Science, 1126:251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aravindan:1996:AFN

[1130] C. Aravindan. An abductive framework for negation in disjunctive logic programming. Lecture Notes in Computer Science, 1126:252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brass:1996:CDW

[1131] S. Brass and J. Dix. Characterizing D-WFS: Confluence and iterated GCWA. Lecture Notes in Computer Science, 1126:268–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedro:1996:MS

[1132] V. Pedro and L. Monteiro. Modules and specifications. Lecture Notes in Computer Science, 1126:284–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kowalski:1996:LPI

[1133] R. Kowalski. Logic programming with integrity constraints. Lecture Notes in Computer Science, 1126:301–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Enciso:1996:TRL

[1134] M. Enciso, I. P. De Guzman, and C. Rossi. Temporal reasoning over linear discrete time. Lecture Notes in Computer Science, 1126:303–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1996:SSF

[1135] R. Pliuskevicius. Similarity saturation for first order linear temporal logic with UNLESS. Lecture Notes in Computer Science, 1126:320–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennett:1996:CSS

[1136] B. Bennett. Carving up space: Steps towards construction of an absolutely complete theory of spatial regions. Lecture Notes in Computer Science, 1126: 337–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forcheri:1996:ILA

[1137] P. Forcheri, P. Gentilini, and M. T. Molfino. Informational logic for automated reasoning. Lecture Notes in Computer Science, 1126:354–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaminski:1996:EOD

[1138] M. Kaminski, J. Makowsky, and M. Tiomkin. Extensions for open default theory via the domain closure assumption. Lecture Notes in Computer Science, 1126:373–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1996:RUU

[1139] C. Witteveen and W. Van der Hoek. Revision and updating using a back-up semantics. Lecture Notes in Computer Science, 1126:388–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besnard:1996:SSS

[1140] P. Besnard and T. Schaub. A simple signed system for paraconsistent reasoning. Lecture Notes in Computer Science, 1126:404–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kotz:1996:FPP

[1141] D. Kotz and N. Nieuwejaar. Flexibility and performance of parallel file systems. Lecture Notes in Computer Science, 1127:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwong:1996:SDD

[1142] P. Kwong and S. Majumdar. Study of data distribution strategies for parallel I/O management. Lecture Notes in Computer Science, 1127:12–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thakur:1996:EEP

[1143] R. Thakur, W. Gropp, and E. Lusk.

An experimental evaluation of the parallel I/O systems of the IBM SP and Intel Paragon using a production application. Lecture Notes in Computer Science, 1127:24–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberle:1996:SSI

[1144] H. Eberle. Switcherland — A scalable interconnection structure for distributed computing. Lecture Notes in Computer Science, 1127:36–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pucher:1996:PAV

[1145] F. Pucher and H. A. Huy. Performance aspects of virtual circuit connections over a local ATM network. Lecture Notes in Computer Science, 1127:50– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hausenblas:1996:NRS

[1146] E. Hausenblas. New results of the Salzburg NTN-Method for the Radon transform. Lecture Notes in Computer Science, 1127:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brunner:1996:PCO

[1147] D. Brunner and A. Uhl. Parallel computation of optimal parameters for pseudo random number generation. Lecture Notes in Computer Science, 1127:78–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilschut:1996:PEM

[1148] A. N. Wilschut, J. Flokstra, and P. M. G. Apers. Parallel evaluation of multi-join queries. Lecture Notes in Computer Science, 1127:90–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouganim:1996:SHD

[1149] L. Bouganim, D. Florescu, and B. Dageville. Skew handling in the DBS3 parallel database system. Lecture Notes in Computer Science, 1127: 98–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciaccia:1996:DDM

[1150] P. Ciaccia and A. Veronesi. Dynamic declustering methods for parallel grid files. Lecture Notes in Computer Science, 1127:110–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biscondi:1996:EIO

[1151] N. Biscondi, L. Brunie, A. Flory, and H. Kosch. Encapsulation of intra-operation parallelism in a parallel match operator. Lecture Notes in Computer Science, 1127:124–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfneiszl:1996:SWG

[1152] H. Pfneiszl and G. Kotsis. Synthetic workload generation for parallel processing systems. Lecture Notes in Computer Science, 1127:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kranzlmueller:1996:PPV

[1153] D. Kranzlmueller, R. Koppler, S. Grabner, and C. Holzner. Parallel program visualization with MUCH. Lecture Notes in Computer Science, 1127: 148–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knudsen:1996:SS

[1154] K. S. Knudsen. Statement-sets. Lecture Notes in Computer Science, 1127: 161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilar:1996:HOS

[1155] J. Aguilar. Heuristics to optimize the speed-up of parallel programs. Lecture Notes in Computer Science, 1127:174– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spalt:1996:VVA

[1156] A. Spalt, S. Grabner, and J. Volkert. Vectorization of visualization algorithms — A practical example. Lecture Notes in Computer Science, 1127: 184–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gubareny:1996:PCI

[1157] N. Gubareny and A. Katkov. Parallel chaotic iterative algorithms for image reconstruction with limited projection data. Lecture Notes in Computer Science, 1127:196–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hausberger:1996:DTA

[1158] G. Hausberger. Digital’s TruCluster architecture. Lecture Notes in Computer Science, 1127:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1996:PP

[1159] J. M. Blum, T. M. Warschko, and W. F. Tichy. The ParaStation project. Lecture Notes in Computer Science, 1127:215–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1996:PPP

[1160] N. Papakostas, G. Papakonstantinou, and P. Tsanakas. PPARDB / PVM: a portable PVM based parallel database management system. Lecture Notes in Computer Science, 1127:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szabo:1996:NLA

[1161] E. Szabo, I. Forgacs, C. Bates, and I. Jelly. A new language for automatic data generation. Lecture Notes in Computer Science, 1127:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellatreche:1996:HFD

[1162] L. Bellatreche and A. Simonet. Horizontal fragmentation in distributed object database systems. Lecture Notes in Computer Science, 1127:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clematis:1996:ISP

[1163] A. Clematis, V. Gianuzzi, and G. Sacchetto. Implementing snapshot protocol for message based communication libraries. Lecture Notes in Computer Science, 1127:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amme:1996:DGD

[1164] W. Amme and E. Zehendner. A/D graphs — A data structure for data dependence analysis in programs with pointers. Lecture Notes in Computer Science, 1127:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halada:1996:PMS

[1165] L. Halada and M. Lucka. A partition method for solving block pentadiagonal linear system on Intel Hypercube iPSC/860. Lecture Notes in Computer Science, 1127:231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palenichka:1996:PIP

[1166] R. M. Palenichka and A. Y. Lutsyk. Parallel image processing by using homogeneous computing structures. Lecture Notes in Computer Science, 1127:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bibel:1996:POA

[1167] W. Bibel, D. Korn, C. Kreitz, and S. Schmitt. Problem-oriented applications of automated theorem proving. Lecture Notes in Computer Science, 1128:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bronstein:1996:STE

[1168] M. Bronstein. 0I0T — A stronglytyped embeddable computer algebra library. Lecture Notes in Computer Science, 1128:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourely:1996:DGF

[1169] C. Bourely and N. Peltier. DiscAtinf. A general framework for implementing calculi and strategies. Lecture Notes in Computer Science, 1128:34–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degtyarev:1996:EET

[1170] A. Degtyarev and A. Voronkov. Equality elimination for the tableau method. Lecture Notes in Computer Science, 1128:46–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1996:TLP

[1171] G. Barthe and H. Elbers. Towards lean proof checking. Lecture Notes in Computer Science, 1128:61–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buch:1996:WHP

[1172] A. Buch, T. Hillenbrand, and R. Fettig. Waldmeister: High performance equational theorem proving. Lecture Notes in Computer Science, 1128:63– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Numazawa:1996:RLB

[1173] M. Numazawa, M. Kurihara, and A. Ohuchi. A reflective language based on conditional term rewriting. Lecture Notes in Computer Science, 1128:65– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balestreri:1996:TRS

[1174] G. Balestreri. Term rewriting systems: An h-categorical semantic. Lecture Notes in Computer Science, 1128: 67–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paoluzzi:1996:GGM

[1175] A. Paoluzzi. Generative geometric modeling in a functional environment. Lecture Notes in Computer Science, 1128:79–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehrke:1996:ESE

[1176] W. Gehrke and C. Limongelli. Exploiting SML for experimenting with algebraic algorithms: The example of padic lifting. Lecture Notes in Computer Science, 1128:98–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santas:1996:CCD

[1177] P. S. Santas. Conditional categories and domains. Lecture Notes in Computer Science, 1128:112–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1996:POS

[1178] M. Gogolla. Parameterizing object specifications. Lecture Notes in Computer Science, 1128:126–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1996:ADS

[1179] P. Ciancarini and C. Mascolo. Analyzing the dynamics of a Z specification. Lecture Notes in Computer Science, 1128:138–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amrhein:1996:WF

[1180] B. Amrhein, O. Gloor, and W. Kuechlin. Walking faster. Lecture Notes in Computer Science, 1128:150–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jebelean:1996:IRA

[1181] T. Jebelean. Integer and rational arithmetic on MasPar. Lecture Notes in Computer Science, 1128:162–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cesari:1996:PPF

[1182] G. Cesari and R. Maeder. Parallel 3-Primes FFT algorithm. Lecture Notes in Computer Science, 1128:174– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buendgen:1996:MSA

[1183] R. Buendgen, M. Goebel, and W. Kuechlin. A master-slave approach to parallel term rewriting on a hierarchical multiprocessor. Lecture Notes in Computer Science, 1128:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haugeneder:1996:MAC

[1184] H. Haugeneder and D. Steiner. Multiagent cooperation — concepts and applications. Lecture Notes in Computer Science, 1128:195–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weck:1996:DCP

[1185] W. Weck. Document-centered presentation of computing software: Compound documents are better workspaces. Lecture Notes in Computer Science, 1128:198–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1996:ANE

[1186] P. Ciancarini and S. Cimato. Animating a non-executable formal specification with a distributed symbolic language. Lecture Notes in Computer Science, 1128:200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Limongelli:1996:URB

[1187] C. Limongelli, G. Malerba, and M. Temperini. Uniform representation of basic algebraic structures in computer algebra. Lecture Notes in Computer Science, 1128:202–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerber:1996:ICA

[1188] M. Kerber, M. Kohlhase, and V. Sorge. Integrating computer algebra with proof planning. Lecture Notes in Computer Science, 1128:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamann:1996:SSM

[1189] K. Hamann and J. Calmet. Structures for symbolic mathematical reasoning and computation. Lecture Notes in Computer Science, 1128:216–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthews:1996:IFI

[1190] S. Matthews. Implementing FS0 in Isabelle: Adding structure at the metalevel. Lecture Notes in Computer Science, 1128:228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cioni:1996:ACR

[1191] G. Cioni, A. Colagrossi, and M. Temperini. An approach to class reasoning in symbolic computation. Lecture Notes in Computer Science, 1128:240– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dupee:1996:IIN

[1192] B. J. Dupee and J. H. Davenport.

An intelligent interface to numerical routines. Lecture Notes in Computer Science, 1128:252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hearn:1996:CAW

[1193] A. C. Hearn. Computer algebra and the World Wide Web. Lecture Notes in Computer Science, 1128:263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norman:1996:IRJ

[1194] A. Norman and J. Fitch. Interfacing REDUCE to Java. Lecture Notes in Computer Science, 1128:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1996:SAC

[1195] G. Butler. Software architectures for computer algebra: a case study. Lecture Notes in Computer Science, 1128: 277–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dalmas:1996:DDM

[1196] S. Dalmas, M. Gaetano, and C. Huchet.

A deductive database for mathematical formulas. Lecture Notes in Computer Science, 1128:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mthuk:1996:CSC

[1197] M. Mthuk and F. Winkler. CASA — A system for computer aided constructive algebraic geometry. Lecture Notes in Computer Science, 1128:297–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). Check first author name??

Dalmas:1996:MSC

[1198] S. Dalmas and M. Gaetano. Making systems communicate and cooperate: The central control approach. Lecture Notes in Computer Science, 1128:308– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daberkow:1996:DNF

[1199] M. Daberkow and A. Weber. A database for number fields. Lecture Notes in Computer Science, 1128:320– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grivas:1996:CRM

[1200] G. Grivas and A. Palinginis. Compiling residuation for a multiparadigm symbolic programming language. Lecture Notes in Computer Science, 1128:331– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gray:1996:PIM

[1201] S. Gray, N. Kajler, and P. S. Wang. Pluggability issues in the multi protocol. Lecture Notes in Computer Science, 1128:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finne:1996:CUI

[1202] S. Finne and S. P. Jones. Composing the user interface with Haggis. Lecture Notes in Computer Science, 1129:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudak:1996:HMT

[1203] P. Hudak. Haskote music tutorial. Lecture Notes in Computer Science, 1129: 38–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeuring:1996:PP

[1204] J. Jeuring and P. Jansson. Polytypic programming. Lecture Notes in Computer Science, 1129:68–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:ITS

[1205] P. Lee. Implementing threads in standard ML. Lecture Notes in Computer Science, 1129:115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okasaki:1996:FDS

[1206] C. Okasaki. Functional data structures. Lecture Notes in Computer Science, 1129:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Runciman:1996:HPS

[1207] C. Runciman and N. Roejemo. Heap profiling for space efficiency. Lecture Notes in Computer Science, 1129:159– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DoaitseSwierstra:1996:DEC

[1208] S. Doaitse Swierstra and L. Duponcheel. Deterministic, error-correcting combinator parsers. Lecture Notes in Computer Science, 1129:184–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swierstra:1996:DEC

[1209] S. Doaitse Swierstra and L. Duponcheel. Deterministic, error-correcting combinator parsers. Lecture Notes in Computer Science, 1129:184–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tofte:1996:ESM

[1210] M. Tofte. Essentials of standard ML modules. Lecture Notes in Computer Science, 1129:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burstall:1996:IDR

[1211] R. Burstall. Inductively defined relations. Lecture Notes in Computer Science, 1130:14–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1996:RCT

[1212] H. Ehrig, M. Grosse-Rhode, and U. Wolter. On the role of category theory in the area of algebraic specifications. Lecture Notes in Computer Science, 1130:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoare:1996:UTC

[1213] T. Hoare. Unification of theories: a challenge for computing science. Lecture Notes in Computer Science, 1130: 49–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Homing:1996:LSL

[1214] J. J. Homing. The Larch shared language: Some open problems. Lecture Notes in Computer Science, 1130:58– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lescanne:1996:LCA

[1215] P. Lescanne. The lambda calculus as an abstract data type. Lecture Notes in Computer Science, 1130:74–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arrais:1996:UTD

[1216] M. Arrais and J. L. Fiadeiro. Unifying theories in different institutions. Lecture Notes in Computer Science, 1130: 81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1996:IFI

[1217] M. Bidoit, C. Choppy, and F. Voisin.

Interchange format for inter-operability of tools and translation. Lecture Notes in Computer Science, 1130:102–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buendgen:1996:EPE

[1218] R. Buendgen and W. Lauterbach. Experiments with partial evaluation domains for rewrite specifications. Lecture Notes in Computer Science, 1130: 125–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clerici:1996:CSP

[1219] S. Clerici, R. M. Jimenez, and F. Orejas. Class-sort polymorphism in Glider. Lecture Notes in Computer Science, 1130:143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coscia:1996:DCA

[1220] E. Coscia and G. Reggio. Deontic concepts in the algebraic specification of dynamic systems: The permission case. Lecture Notes in Computer Science, 1130:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denker:1996:RCV

[1221] G. Denker. Reification — changing viewpoint but preserving truth. Lecture Notes in Computer Science, 1130: 182–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaconescu:1996:CBE

[1222] R. Diaconescu. A category-based equational logic semantics to constraint programming. Lecture Notes in Computer Science, 1130:200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosse-Rhode:1996:CST

[1223] M. Grosse-Rhode. Concurrent state transformations on abstract data types. Lecture Notes in Computer Science, 1130:222–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hensel:1996:VIP

[1224] U. Hensel and D. Spooner. A view on implementing processes: Categories of circuits. Lecture Notes in Computer Science, 1130:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hintermeier:1996:CAS

[1225] C. Hintermeier, H. Kirchner, and P. D.

Mosses. Combining algebraic and settheoretic specifications. Lecture Notes in Computer Science, 1130:255–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamperman:1996:MTR

[1226] J. Kamperman and P. Walters. Minimal term rewriting systems. Lecture Notes in Computer Science, 1130:274– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klar:1996:IIT

[1227] M. Klar, R. Geisler, and F. Cornelius. InterACT: An interactive theorem and completeness prover for algebraic specifications with conditional equations. Lecture Notes in Computer Science, 1130:291–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriauciukas:1996:RRS

[1228] V. Kriauciukas and M. Walicki. Rewriting and reasoning with set-relations II: The non-ground case completeness. Lecture Notes in Computer Science, 1130:306–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kristoffersen:1996:TCR

[1229] B. Kristoffersen. Termination of curryfied rewrite systems. Lecture Notes in Computer Science, 1130:322–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeGall:1996:FST

[1230] P. Le Gall and A. Arnould. Formal specifications and test: Correctness and oracle. Lecture Notes in Computer Science, 1130:342–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malcolm:1996:BEB

[1231] G. Malcolm. Behavioural equivalence, bisimulation, and minimal realisation. Lecture Notes in Computer Science, 1130:359–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossakowski:1996:ULP

[1232] T. Mossakowski. Using limits of parchments to systematically construct institutions of partial algebras. Lecture Notes in Computer Science, 1130:379– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mylonakis:1996:BST

[1233] N. Mylonakis. Behavioural specifications in type theory. Lecture Notes in Computer Science, 1130:394–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padawitz:1996:SDT

[1234] P. Padawitz. Swinging data types: Syntax, semantics, and theory. Lecture Notes in Computer Science, 1130:409– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pawlowski:1996:CI

[1235] W. Pawlowski. Context institutions.

Lecture Notes in Computer Science, 1130:436–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qian:1996:OOF

[1236] Z. Qian and B. Krieg-Brueckner.

Object-oriented functional programming and type reconstruction. Lecture Notes in Computer Science, 1130:458– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarlecki:1996:MBL

[1237] A. Tarlecki. Moving between logical systems. Lecture Notes in Computer Science, 1130:478–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voisin:1996:MAS

[1238] F. Voisin and M. Bidoit. Modular algebraic specifications and the orientation of equations into rewrite rules. Lecture Notes in Computer Science, 1130:503–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walters:1996:MEL

[1239] P. Walters and J. Kamperman. A model for I/O in equational languages with don’t care non-determinism. Lecture Notes in Computer Science, 1130: 522–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolz:1996:TDS

[1240] D. Wolz. Tool design for structuring mechanisms for algebraic specification languages with initial semantics. Lecture Notes in Computer Science, 1130:536–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:GTU

[1241] G. Wang, M. W. Vannier, E. McFarland, and J. P. Heiken. GI tract unraveling in volumetric CT. Lecture Notes in Computer Science, 1131:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiemann:1996:SVH

[1242] T. Schiemann, J. Nuthmann, U. Tiede, and K. H. Hoehne. Segmentation of the visible human for high quality volume based visualization. Lecture Notes in Computer Science, 1131:13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preim:1996:IAM

[1243] B. Preim, A. Ritter, and T. Strothotte. Illustrating anatomic models — A semi-interactive approach. Lecture Notes in Computer Science, 1131:23– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mori:1996:FRM

[1244] K. Mori, J.-I. Hasegawa, J.-I. Toriwaki, and H. Anno. A fast rendering method using the tree structure of objects in virtualized bronchus endoscope system. Lecture Notes in Computer Science, 1131:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Law:1996:ORT

[1245] A. Law and R. Yagel. An optimal ray traversal scheme for visualizing colossal medical volumes. Lecture Notes in Computer Science, 1131:43–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacq:1996:DMV

[1246] J.-J. Jacq and C. Roux. A direct multivolume rendering method. application to visual assessment of 3-D image registration algorithms. Lecture Notes in Computer Science, 1131:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manduca:1996:VTE

[1247] A. Manduca, R. Muthupillai, P. J.

Rossman, and J. F. Greenleaf. Visualization of tissue elasticity by magnetic resonance elastography. Lecture Notes in Computer Science, 1131:63– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogeveen:1996:AVD

[1248] R. Hoogeveen, C. Bakker, and M. Viergever. Accurate vessel depiction with phase gradient algorithm in MR angiography. Lecture Notes in Computer Science, 1131:69–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mullick:1996:VLV

[1249] R. Mullick and H. T. Nguyen. Visualization and labeling of the visible Human[TM] dataset: Challenges and resolves. Lecture Notes in Computer Science, 1131:75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stewart:1996:RVM

[1250] J. E. Stewart, W. C. Broaddus, and J. H. Johnson. Rebuilding the visible man. Lecture Notes in Computer Science, 1131:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanScheltinga:1996:IQI

[1251] J. Terwisscha van Scheltinga, M. Bosma, J. Smit, and S. Lobregt. Image quality improvements in volume rendering. Lecture Notes in Computer Science, 1131:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grzeszczuk:1996:RTM

[1252] R. P. Grzeszczuk and C. A. Pelizzari. Real-time merging of visible surfaces for display and segmentation. Lecture Notes in Computer Science, 1131:93– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holmes:1996:CPC

[1253] C. J. Holmes, D. MacDonald, J. G. Sled, and A. W. Toga. Cortical peeling: CSF/ grey matter/white matter boundaries visualized by nesting isosurfaces. Lecture Notes in Computer Science, 1131:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guillemaud:1996:EMI

[1254] R. Guillemaud and M. Brady. Enhancement of MR images. Lecture Notes in Computer Science, 1131:107– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marquez:1996:RCC

[1255] J. Marquez and F. Schmitt. Radiometric correction of color cryosection images for three-dimensional segmentation of fine structures. Lecture Notes in Computer Science, 1131:117–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Launay:1996:RCV

[1256] L. Launay, E. Maurincomme, P. Bouchet, and J.-L. Mallet. 3D reconstruction of cerebral vessels and pathologies from a few biplane digital angiographies. Lecture Notes in Computer Science, 1131: 123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kay:1996:RDR

[1257] P. A. Kay, R. A. Robb, D. G. Bostwick, and J. J. Camp. Robust 3-D reconstruction and analysis of microstructures from serial histologic sections, with emphasis on microvessels in prostate cancer. Lecture Notes in Computer Science, 1131:129–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwan:1996:EMS

[1258] R. K.-S. Kwan, A. C. Evans, and G. B. Pike. An extensible MRI simulator for post-processing evaluation. Lecture Notes in Computer Science, 1131:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brechbuehler:1996:CSI

[1259] C. Brechbuehler, G. Gerig, and G. Szekely. Compensation of spatial inhomogeneity in MRI based on a parametric bias estimate. Lecture Notes in Computer Science, 1131:141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuenemann:1996:ACM

[1260] S. Schuenemann, C. Rethfeldt, F. Mueller, and K. Agha-Amiri. Analysis of coupled multi-image information in microscopy. Lecture Notes in Computer Science, 1131:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilles:1996:BFC

[1261] S. Gilles, M. Brady, J. Declerck, and J.-P. Thirion. Bias field correction of breast MR images. Lecture Notes in Computer Science, 1131:153–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Masutani:1996:RGB

[1262] Y. Masutani, K. Masamune, and T. Dohi. Region-growing based feature extraction algorithm for tree-like objects. Lecture Notes in Computer Science, 1131:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McAuliffe:1996:SSB

[1263] M. J. McAuliffe, D. Eberly, D. S.

Fritsch, and E. L. Chaney. Scalespace boundary evolution initialized by cores. Lecture Notes in Computer Science, 1131:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrett:1996:FAR

[1264] W. A. Barrett and E. N. Mortensen.

Fast, accurate, and reproducible livewire boundary extraction. Lecture Notes in Computer Science, 1131:183– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelemen:1996:ASC

[1265] A. Kelemen, G. Szekely, H.-W. Reist, and G. Gerig. Automatic segmentation of cell nuclei from confocal laser scanning microscopy images. Lecture Notes in Computer Science, 1131:193– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouma:1996:ESA

[1266] C. J. Bouma, W. J. Niessen, K. J.

Zuiderveld, and E. J. Gussenhoven. Evaluation of segmentation algorithms for intravascular ultrasound images. Lecture Notes in Computer Science, 1131:203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vemuri:1996:FAF

[1267] B. C. Vemuri, Y. Guo, S. H. Lai, and C. M. Leonard. Fast algorithms for fitting multiresolution hybrid shape models to brain MRI. Lecture Notes in Computer Science, 1131:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanazawa:1996:CAS

[1268] K. Kanazawa, M. Kubo, N. Niki, and H. Satoh. Computer aided screening system for lung cancer based on helical CT images. Lecture Notes in Computer Science, 1131:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandermeulen:1996:URC

[1269] D. Vandermeulen, X. Descombes, P. Suetens, and G. Marchal. Unsupervised regularized classification of multi-spectral MRI. Lecture Notes in Computer Science, 1131:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buzug:1996:UES

[1270] T. M. Buzug, J. Weese, C. Fassnacht, and C. Lorenz. Using an entropy similarity measure to enhance the quality of DSA images with an algorithm based on template matching. Lecture Notes in Computer Science, 1131:235–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkins:1996:ASB

[1271] M. S. Atkins and B. T. Mackiewich. Automatic segmentation of the brain in MRI. Lecture Notes in Computer Science, 1131:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belohlavek:1996:IAU

[1272] M. Belohlavek, A. Manduca, T. Behrenbeck, and J. B. Seward. Image analysis using modified self-organizing maps: Automated delineation of the left ventricular cavity boundary in serial echocardiograms. Lecture Notes in Computer Science, 1131:247–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBraal:1996:API

[1273] L. De Braal, N. Ezquerra, E. Schwartz, and C. D. Cooke. Analyzing and predicting images through a neural network approach. Lecture Notes in Computer Science, 1131:253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alcaniz:1996:AIA

[1274] M. Alcañiz, V. Grau, C. Knoll, and M. C. Juan. An artificial intelligence approach for automatic interpretation of maxillofacial CT images. Lecture Notes in Computer Science, 1131:259– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bro-Nielsen:1996:FFR

[1275] M. Bro-Nielsen and C. Gramkow. Fast fluid registration of medical images. Lecture Notes in Computer Science, 1131:267–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rangarajan:1996:RPM

[1276] A. Rangarajan, E. Mjolsness, S. Pappu, and L. Davachi. A robust point matching algorithm for autoradiograph alignment. Lecture Notes in Computer Science, 1131:277–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:1996:RUU

[1277] J. Zhao, P. T. E. Roberts, A. C. F. Colchester, and K. Holton-Tainter. Registration updating using marker pins in a video based neurosurgical guidance system (VISLAN). Lecture Notes in Computer Science, 1131:287– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rohr:1996:PBE

[1278] K. Rohr, H. S. Stiehl, R. Sprengel, and W. Beil. Point-based elastic registration of medical image data using approximating thin-plate splines. Lecture Notes in Computer Science, 1131:297– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collins:1996:CCN

[1279] D. L. Collins, G. Le Goualher, R. Venugopal, and A. Caramanos. Cortical constraints for non-linear cortical registration. Lecture Notes in Computer Science, 1131:307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hata:1996:IGM

[1280] N. Hata, W. M. Wells, M. Halle, and S. Nakajima. Image guided microscopic surgery system using mutualinformation based registration. Lecture Notes in Computer Science, 1131:317– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thirion:1996:CVT

[1281] J.-P. Thirion, G. Subsol, and D. Dean. Cross validation of three inter-patients matching methods. Lecture Notes in Computer Science, 1131:327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schormann:1996:NAF

[1282] T. Schormann, S. Henn, and K. Zilles.

A new approach to fast elastic alignment with applications to human brains. Lecture Notes in Computer Science, 1131:337–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1996:IAA

[1283] G. E. Christensen, S. C. Joshi, and M. I. Miller. Individualizing anatomical atlases of the head. Lecture Notes in Computer Science, 1131:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:MIA

[1284] B. Kim, J. L. Boes, K. A. Frey, and C. R. Meyer. Mutual information for automated multimodal image warping. Lecture Notes in Computer Science, 1131:349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davatzikos:1996:MAB

[1285] C. Davatzikos, M. Vaillant, S. Resnick, and J. L. Prince. Morphological analysis of brain structures using spatial normalization. Lecture Notes in Computer Science, 1131:355–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dean:1996:TDM

[1286] D. Dean, P. Buckley, F. L. Bookstein, and J. Kamath. Three dimensional MR-Based morphometric comparison of schizophrenic and normal cerebral ventricles. Lecture Notes in Computer Science, 1131:363–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subsol:1996:AAB

[1287] G. Subsol, J.-P. Thirion, and N. Ayache. Application of an automatically built 3D morphometric brain atlas: Study of cerebral ventricle shape. Lecture Notes in Computer Science, 1131: 373–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thompson:1996:VMA

[1288] P. Thompson and A. W. Toga. Visualization and mapping of anatomic abnormalities using a probabilistic brain atlas based on random fluid transformations. Lecture Notes in Computer Science, 1131:383–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talhami:1996:STD

[1289] H. Talhami. L-systems for threedimensional anatomical modelling: Towards a virtual laboratory in anatomy. Lecture Notes in Computer Science, 1131:393–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriete:1996:HDR

[1290] A. Kriete. Hierarchial data representation of lung to model morphology and function. Lecture Notes in Computer Science, 1131:399–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bookstein:1996:VGD

[1291] F. L. Bookstein. Visualizing group differences in outline shape: Methods from biometrics of landmark points. Lecture Notes in Computer Science, 1131:405–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marais:1996:VCA

[1292] P. Marais, R. Guillemaud, M. Sakuma, and A. Zisserman. Visualising cerebral asymmetry. Lecture Notes in Computer Science, 1131:411–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freeborough:1996:APR

[1293] P. A. Freeborough and N. C. Fox.

Assessing patterns and rates of brain atrophy by serial magnetic resonance imaging: a segmentation, registration, display and quantification procedure. Lecture Notes in Computer Science, 1131:419–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossmanith:1996:CCB

[1294] C. Rossmanith, H. Handels, S. J.

Poeppl, and E. Rinast. Characterisation and classification of brain tumours in three-dimensional MR image sequences. Lecture Notes in Computer Science, 1131:429–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zijdenbos:1996:AQM

[1295] A. Zijdenbos, A. C. Evans, F. Riahi, and J. G. Sled. Automatic quantification of multiple sclerosis lesion volume using stereotaxic space. Lecture Notes in Computer Science, 1131:439– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ge:1996:SVC

[1296] Y. Ge, D. R. Stelts, and D. J. Vining. 3D skeleton for virtual colonoscopy. Lecture Notes in Computer Science, 1131:449–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faber:1996:IHP

[1297] T. L. Faber, J. M. Hoffman, T. R. Henry, and J. R. Votaw. Identifying hypometabolism in PET images of the brain: Application to epilepsy. Lecture Notes in Computer Science, 1131: 457–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rocha:1996:MAE

[1298] C. Rocha, J.-L. Dillenseger, and J.-L. Coatrieux. Multi-array EEG signals mapped with three dimensional images for clinical epilepsy studies. Lecture Notes in Computer Science, 1131:467– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bromm:1996:VPM

[1299] B. Bromm and E. Scharein. Visualisation of pain by magnetoencephalography in humans. Lecture Notes in Computer Science, 1131:477–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinse:1996:VCA

[1300] H. R. Dinse, B. Godde, T. Hilger, and G. Reuter. Visualization of cat auditory cortical functional organization after electrical stimulation with a multichannel cochlear implant by means of optical imaging. Lecture Notes in Computer Science, 1131:483–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poston:1996:BBV

[1301] T. Poston, W. L. Nowinski, L. Serra, and B. C. Chua. The brain bench: Virtual stereotaxis for rapid neurosurgery planning and training. Lecture Notes in Computer Science, 1131:491–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanCleynenbreugel:1996:FEI

[1302] J. Van Cleynenbreugel, K. Verstreken, G. Marchal, and P. Suetens. A flexible environment for image guided virtual surgery planning. Lecture Notes in Computer Science, 1131:501–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alcaniz:1996:ASS

[1303] M. Alcañiz, F. Chinesta, C. Monserrat, and V. Grau. An advanced system for the simulation and planning of orthodontic treatments. Lecture Notes in Computer Science, 1131:511–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahidi:1996:ASV

[1304] R. Shahidi, V. Argiro, S. Napel, and L. Gray. Assessment of several virtual endoscopy techniques using computed tomography and perspective volume rendering. Lecture Notes in Computer Science, 1131:521–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bro-Nielsen:1996:SSU

[1305] M. Bro-Nielsen. Surgery simulation using fast finite elements. Lecture Notes in Computer Science, 1131:529– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cotin:1996:RTV

[1306] S. Cotin, H. Delingette, and N. Ayache. Real time volumetric deformable models for surgery simulation. Lecture Notes in Computer Science, 1131:535– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keeve:1996:CSS

[1307] E. Keeve, S. Girod, and B. Girod. Craniofacial surgery simulation. Lecture Notes in Computer Science, 1131: 541–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kay:1996:CVP

[1308] P. A. Kay, R. A. Robb, R. P. Myers, and B. F. King. Creation and validation of patient specific anatomical models for prostate surgery planning using virtual reality. Lecture Notes in Computer Science, 1131:547–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourland:1996:USA

[1309] J. D. Bourland and Q. R. Wu. Use of shape for automated, optimized 3D radiosurgical treatment planning. Lecture Notes in Computer Science, 1131: 553–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1996:MGI

[1310] D. L. G. Hill, P. N. Poynter Smith, C. L. Emery, and P. E. Summers. MRI guided intranasal flexible microendoscopy. Lecture Notes in Computer Science, 1131:561–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:CAI

[1311] Q. H. Li, H. J. Holdener, L. Zamorano, and P. King. Computer-assisted insertion of pedicle screws. Lecture Notes in Computer Science, 1131:571–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mei:1996:PCI

[1312] Q. Mei, S. J. Harris, F. Arambula-Cosio, and M. S. Nathan. PROBOT — A computer integrated prostatectomy system. Lecture Notes in Computer Science, 1131:581–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1996:TPU

[1313] H. Fuchs, A. State, E. D. Pisano, and W. F. Garrett. Towards performing ultrasound-guided needle biopsies from within a head-mounted display. Lecture Notes in Computer Science, 1131:591–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Studholme:1996:AMR

[1314] C. Studholme, J. A. Little, G. P. Penny, and D. L. G. Hill. Automated multimodality registration using the full affine transformation: Application to MR and CT guided skull base surgery. Lecture Notes in Computer Science, 1131:601–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouge:1996:DPP

[1315] L. Bouge. The data parallel programming model: a semantic perspective. Lecture Notes in Computer Science, 1132:4–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1996:IH

[1316] R. S. Schreiber. An introduction to HPF. Lecture Notes in Computer Science, 1132:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petiton:1996:DPS

[1317] S. G. Petiton and N. Emad. A data parallel scientific computing introduction. Lecture Notes in Computer Science, 1132:45–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1996:CSS

[1318] M. S. Lam. Current status of the SUIF research project. Lecture Notes in Computer Science, 1132:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pazat:1996:THP

[1319] J.-L Pazat. Tools for High Performance Fortran: a survey. Lecture Notes in Computer Science, 1132:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hatcher:1996:IDP

[1320] P. J. Hatcher, R. D. Russell, S. Kumaran, and M. J. Quinn. Implementing data parallel programs on commodity clusters. Lecture Notes in Computer Science, 1132:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foster:1996:TPH

[1321] I. Foster. Task parallelism and highperformance languages. Lecture Notes in Computer Science, 1132:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekeyser:1996:SID

[1322] J.-L. Dekeyser and P. Marquet. Supporting irregular and dynamic computations in data parallel languages. Lecture Notes in Computer Science, 1132: 197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lisper:1996:DPF

[1323] B. Lisper. Data parallelism and functional programming. Lecture Notes in Computer Science, 1132:220–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouge:1996:FVD

[1324] L. Bouge, D. Cachera, Y. Le Guyadec, and G. Utard. Formal validation of data parallel programs: Introducting the assertional approach. Lecture Notes in Computer Science, 1132:252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seguin:1996:ASC

[1325] G. E. Seguin. The algebraic structure of codes invariant under a permutation. Lecture Notes in Computer Science, 1133:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedersen:1996:MCA

[1326] J. P. Pedersen and D. Polemi. A method of combining algebraic geometric Goppa codes. Lecture Notes in Computer Science, 1133:19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gulliver:1996:SBR

[1327] T. A. Gulliver and V. K. Bhargava. Some best rate 1/p quasi-cyclic codes over GF(5). Lecture Notes in Computer Science, 1133:28–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Youssef:1996:ILR

[1328] A. M. Youssef and S. E. Tavares. Information leakage of a randomly selected Boolean function. Lecture Notes in Computer Science, 1133:41–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aakvaag:1996:UPT

[1329] N. D. Aakvaag, B. Lacaze, and A. Duverdier. On the use of periodic timebase companding in the scrambling of stationary processes. Lecture Notes in Computer Science, 1133:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duverdier:1996:NAS

[1330] A. Duverdier, B. Lacaze, and N. D.

Aakvaag. A novel approach to spread spectrum communication using linear periodic time-varying filters. Lecture Notes in Computer Science, 1133:64– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Battail:1996:RLC

[1331] G. Battail. On random-like codes. Lecture Notes in Computer Science, 1133: 76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guinand:1996:AAD

[1332] P. Guinand, J. Lodge, and L. Papke.

An alternative approach to the design of interleavers for block “turbo” codes. Lecture Notes in Computer Science, 1133:95–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gravel:1996:IVD

[1333] D. B. Gravel, G. Drolet, and C. N.

Rozon. Improved VLSI design for decoding concatenated codes comprising an irreducible cyclic code and a Reed–Solomon code. Lecture Notes in Computer Science, 1133:104–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:NMT

[1334] R. C.-K. Lee and F. R. Kschischang.

Non-minimal trellises for linear block codes. Lecture Notes in Computer Science, 1133:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esmaeili:1996:TCL

[1335] M. Esmaeili, T. A. Gulliver, and N. P. Secord. Trellis complexity of linear block codes via atomic codewords. Lecture Notes in Computer Science, 1133:130–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juergensen:1996:ECC

[1336] H. Juergensen and S. Konstantinidis. Error corrections for channels with substitutions, insertions, and deletions. Lecture Notes in Computer Science, 1133:149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudreau:1996:RCS

[1337] D. Boudreau and Y. Viens. Reduced complexity soft-output maximum likelihood sequence estimation of 4-ary CPM signals transmitted over Rayleigh flat-fading channels. Lecture Notes in Computer Science, 1133:164–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nassar:1996:NRS

[1338] C. R. Nassar and M. R. Soleymani. A novel receiver structure for MPSK in the presence of rapidly changing phase. Lecture Notes in Computer Science, 1133:184–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAmours:1996:CMV

[1339] C. D’Amours and A. Yongacoglu. Comparison of MFSK variants for a hybrid DS/FH-CDMA system in Rayleigh fading. Lecture Notes in Computer Science, 1133:200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poor:1996:FFW

[1340] H. V. Poor. Finite-field wavelet transforms. Lecture Notes in Computer Science, 1133:225–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mandal:1996:CWI

[1341] M. K. Mandal, S. Panchanathan, and T. Aboulnasr. Choice of wavelets for image compression. Lecture Notes in Computer Science, 1133:239–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volden:1996:IMR

[1342] E. Volden, G. Giraudon, and M. Berthod.

Information in Markov random fields and image redundancy. Lecture Notes in Computer Science, 1133:250–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gauvin:1996:CID

[1343] J. Gauvin, M. Doucet, D. Gingras, and P. Chevrette. Coding of image data via correlation filters for invariant pattern recognition: Some practical results. Lecture Notes in Computer Science, 1133:269–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallant:1996:IMS

[1344] P. J. Gallant, E. L. Morin, and L. E. Peppard. Improving myoelectric signal classifier generalization by preprocessing with exploratory projections. Lecture Notes in Computer Science, 1133: 278–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liao:1996:CCR

[1345] S. X. Liao and M. Pawlak. Chinese character recognition via orthogonal moments. Lecture Notes in Computer Science, 1133:296–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rykowski:1996:UMO

[1346] J. Rykowski and W. Cellary. Using multiversion object-oriented databases in CAD/CIM systems. Lecture Notes in Computer Science, 1134:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yahia:1996:BIG

[1347] A. Yahia, L. Lakhal, and R. Cicchetti. Building inheritance graphs in object database design. Lecture Notes in Computer Science, 1134:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pastor:1996:OOM

[1348] O. Pastor, V. Pelechano, B. Bonet, and I. Ramos. An object oriented methodological approach for making automated prototyping feasible. Lecture Notes in Computer Science, 1134: 29–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Annamalai:1996:CBT

[1349] M. Annamalai, A. Sundaram, and B. Bhargava. A color-based technique for measuring visible loss for use in image data communication. Lecture Notes in Computer Science, 1134:39– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Battle:1996:EAC

[1350] S. Battle. Establishing arc consistency for multiple database views. Lecture Notes in Computer Science, 1134:49– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henrich:1996:UQL

[1351] A. Henrich and D. Daeberitz. Using a query language to state consistency constraints for repositories. Lecture Notes in Computer Science, 1134:59– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appelrath:1996:CSA

[1352] H.-J. Appelrath, H. Behrends, H. Jasper, and O. Zukunft. Case studies on active database applications. Lecture Notes in Computer Science, 1134:69–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1996:TSG

[1353] L. Becker, A. Voigtmann, and K. H.

Hinrichs. Temporal support for geodata in object-oriented databases. Lecture Notes in Computer Science, 1134: 79–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casati:1996:DAR

[1354] F. Casati, S. Ceri, B. Pernici, and G. Pozzi. Deriving active rules for workflow enactment. Lecture Notes in Computer Science, 1134:94–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFerreiraRezende:1996:EOB

[1355] F. De Ferreira Rezende and T. Baier. Employing object-based LSNs in a recovery strategy. Lecture Notes in Computer Science, 1134:116–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dogdu:1996:SAT

[1356] E. Dogdu. Scheduling adaptive transactions in real-time databases. Lecture Notes in Computer Science, 1134:130– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1996:ASC

[1357] K.-Y. Lam, W.-C. Yau, and V. C. S. Lee. Applying similarity in concurrency control for real-time database application. Lecture Notes in Computer Science, 1134:143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alhajj:1996:VMO

[1358] R. Alhajj and F. Polat. View maintenance in object-oriented databases. Lecture Notes in Computer Science, 1134:153–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bakker:1996:OOB

[1359] J. A. Bakker. Object-orientation based on semantic transformations. Lecture Notes in Computer Science, 1134:163– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djeraba:1996:MSD

[1360] C. Djeraba and K. Hadouda. Multimedia scenes in a database system. Lecture Notes in Computer Science, 1134:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitakami:1996:ISI

[1361] H. Kitakami, Y. Mori, and M. Arikawa.

An intelligent system for integrating autonomous nomenclature databases in semantic heterogeneity. Lecture Notes in Computer Science, 1134:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucarella:1996:GOD

[1362] D. Lucarella and A. Zanzi. A graphoriented data model. Lecture Notes in Computer Science, 1134:197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:CDS

[1363] Y.-J. Lee, D.-M. Lee, S.-J. Ryu, and C.-W. Chung. Controlled decomposition strategy for complex spatial objects. Lecture Notes in Computer Science, 1134:207–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greco:1996:CWU

[1364] S. Greco, C. Mateis, and E. Spadafora. The complexity of weak unification of bounded simple set terms. Lecture Notes in Computer Science, 1134:224– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1996:FAH

[1365] S. J. Lim and Y.-K. Ng. A formal approach for horizontal fragmentation in distributed deductive database design. Lecture Notes in Computer Science, 1134:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collet:1996:CEN

[1366] C. Collet and T. Coupaye. Composite events in NAOS. Lecture Notes in Computer Science, 1134:244–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eaglestone:1996:GHD

[1367] B. Eaglestone, R. Holton, and L. Rold. GENREG: a historical data model based on event graphs. Lecture Notes in Computer Science, 1134:254–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delannoy:1996:BIC

[1368] X. C. Delannoy and C. Del Vigna. Binary integrity constraints against confidentiality. Lecture Notes in Computer Science, 1134:264–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulos:1996:BHI

[1369] A. Papadopoulos, M. Vaitis, and D. Christodoulakis. Building hypertext interfaces to existing relational databases. Lecture Notes in Computer Science, 1134:276–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1996:EJN

[1370] H.-C. Liu and C. Chirathamjaree.

An efficient join for nested relational databases. Lecture Notes in Computer Science, 1134:289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hameurlain:1996:PRD

[1371] A. Hameurlain and F. Morvan. Parallel relational database systems: Why, how and beyond. Lecture Notes in Computer Science, 1134:302–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goeers:1996:EOO

[1372] J. Goeers, K.-P. Lisson, and H.-G.

Linde-Goeers. Experiences in objectoriented modeling of a real database application. Lecture Notes in Computer Science, 1134:313–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shin:1996:SCO

[1373] H. Shin and J. Chang. Signatures of complex objects for query processing in OODB. Lecture Notes in Computer Science, 1134:323–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennessey:1996:FSO

[1374] P. Hennessey, M. T. Ibrahim, and A. M. Fedorec. Formal specification, object oriented design, and implementation of an ephemeral logger for database systems. Lecture Notes in Computer Science, 1134:333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marik:1996:CAD

[1375] V. Marik, O. Stepankova, O. Flek, and J. Kout. Cooperative agents in DISCIM environment. Lecture Notes in Computer Science, 1134:356–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miura:1996:MTS

[1376] T. Miura and I. Shioya. Mining type schemes in databases. Lecture Notes in Computer Science, 1134:369–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishizawa:1996:QPM

[1377] I. Nishizawa, A. Takasu, and J. Adachi. A query processing method for integrated access to multiple databases. Lecture Notes in Computer Science, 1134:385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1996:SD

[1378] P. Lang, W. Obermair, and M. Schrefl. Situation diagrams. Lecture Notes in Computer Science, 1134:400–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zimmermann:1996:DIM

[1379] J. Zimmermann, H. Branding, A. P.

Buchmann, and A. Deutsch. Design, implementation and management of rules in an acitve database system. Lecture Notes in Computer Science, 1134:422–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1996:OGA

[1380] I. S. Thomas and A. C. Jones. OLAF:

The GOAD active database event/rule tracer. Lecture Notes in Computer Science, 1134:436–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiedijk:1996:CWM

[1381] M. Wiedijk, H. Afsarmanesh, and L. O. Hertzberger. Co-working and management of federated informationclusters. Lecture Notes in Computer Science, 1134:446–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guenthoer:1996:TBA

[1382] R. Guenthoer and S. Jablonski.

Transaction-based application integration in workflow management systems. Lecture Notes in Computer Science, 1134:456–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tang:1996:CMS

[1383] J. Tang and S.-Y. Hwang. Coping with mismatched semantics of dependencies in workflow systems. Lecture Notes in Computer Science, 1134:467–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:DII

[1384] B. Wang. The design of an integrated information system. Lecture Notes in Computer Science, 1134:479–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heuer:1996:QOC

[1385] A. Heuer and J. Kroeger. Query optimization in the CROQUE project. Lecture Notes in Computer Science, 1134: 489–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Na:1996:PFQ

[1386] S. Na and S. Park. A process of fuzzy query on new fuzzy object oriented data model. Lecture Notes in Computer Science, 1134:500–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pitoura:1996:RSS

[1387] E. Pitoura. A replication schema to support weak connectivity in mobile information systems. Lecture Notes in Computer Science, 1134:510–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puig:1996:CCO

[1388] V. Puig and C. Oussalah. Constraints and composite objects. Lecture Notes in Computer Science, 1134:521–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tininini:1996:SMA

[1389] L. Tininini, A. Bezenchek, and M. Rafanelli. A system for the management of aggregate data. Lecture Notes in Computer Science, 1134:531– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heuer:1996:DAM

[1390] A. Heuer and A. Lubinski. Database access in mobile environments. Lecture Notes in Computer Science, 1134:544– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Du:1996:CBN

[1391] X. Du, Z. Liu, and N. Ishii. Countingbased normalization for multiple linear recursions. Lecture Notes in Computer Science, 1134:554–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1996:PSD

[1392] T. Eiter, N. Leone, and D. Sacca. Partial semantics for disjunctive deductive databases. Lecture Notes in Computer Science, 1134:564–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tang:1996:OPE

[1393] Z.-H. Tang, G. Gardarin, and V. Smahi. Optimizing path expressions using navigational algebraic operators. Lecture Notes in Computer Science, 1134:574–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappel:1996:MOB

[1394] G. Kappel and M. Schrefl. Modeling object behavior: To use methods or rules or both. Lecture Notes in Computer Science, 1134:584–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:OOD

[1395] Y. H. Kim, M. S. Kim, S.-H. Kwon, and C.-S. Jeong. The object-oriented design of KROSS: An object-oriented spatial database system. Lecture Notes in Computer Science, 1134:603–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winiwarter:1996:ALE

[1396] W. Winiwarter, O. Kagawa, and Y. Kambayashi. Applying language engineering techniques to the question support facilities in VIENA classroom. Lecture Notes in Computer Science, 1134:613–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshida:1996:MMO

[1397] T. Yoshida and M. Takizawa. Model of mobile objects. Lecture Notes in Computer Science, 1134:623–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wietrzyk:1996:RTT

[1398] V. I. Wietrzyk and Y. Ramaswamy. Real-time transaction scheduling in database systems. Lecture Notes in Computer Science, 1134:633–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wietrzyk:1996:PEN

[1399] V. I. Wietrzyk. Performance evaluation of new clustering algorithm in object-oriented database systems. Lecture Notes in Computer Science, 1134: 644–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leone:1996:CDS

[1400] N. Leone, P. Rullo, and F. Scarcello.

On the computation of disjunctive stable models. Lecture Notes in Computer Science, 1134:654–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:ATO

[1401] L. Wang, M. Wing, C. Davis, and N. Revell. An algebra for a temporal object data model. Lecture Notes in Computer Science, 1134:667–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1996:SCD

[1402] K. Tanaka, H. Higaki, and M. Takizawa. Significant checkpoint in distributed system. Lecture Notes in Computer Science, 1134:678–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1996:AID

[1403] X. Wu. An architecture for interoperation of distributed heterogeneous database systems. Lecture Notes in Computer Science, 1134:688–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lo:1996:SQP

[1404] Y.-L. Lo, K. A. Hua, and W. Tavanapong. Scheduling queries for parallel execution on multicomputer database management systems. Lecture Notes in Computer Science, 1134: 698–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1996:DRD

[1405] S.-Y. Hwang, K. K. S. Lee, and Y. H. Chin. Data replication in a distributed system: a performance study. Lecture Notes in Computer Science, 1134:708– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coenen:1996:OLS

[1406] F. Coenen, B. Beattie, T. J. M. Bench-Capon, and M. J. R. Shave. An ontology for linear spatial reasoning. Lecture Notes in Computer Science, 1134:718–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debenham:1996:DKB

[1407] J. Debenham and V. Devedzic. Designing knowledge-based systems for optimal performance. Lecture Notes in Computer Science, 1134:728–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolety:1996:MAR

[1408] S. P. Tolety, U. Garimella, P. E. Bandreddi Middile, and S. P. Rana. Match algorithm for a real time expert system. Lecture Notes in Computer Science, 1134:738–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghandeharizadeh:1996:LRD

[1409] S. Ghandeharizadeh and D. Kim. Online reorganization of data in scalable continuous media servers. Lecture Notes in Computer Science, 1134:751– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:1996:RMM

[1410] V. Kumar. Recovery in main memory database systems. Lecture Notes in Computer Science, 1134:769–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nascimento:1996:MII

[1411] M. A. Nascimento, M. H. Dunham, and R. Elmasri. M-IVTT: An index for bitemporal databases. Lecture Notes in Computer Science, 1134:779–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ballesteros:1996:DMC

[1412] L. Ballesteros and B. Croft. Dictionary methods for cross-lingual information retrieval. Lecture Notes in Computer Science, 1134:791–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1996:PSS

[1413] F. C. Berger and P. Van Bommel.

Personalized search support for networked document retrieval using link inference. Lecture Notes in Computer Science, 1134:802–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ounis:1996:UCG

[1414] L. Ounis and J.-P. Chevallet. Using conceptual graphs in a multifaceted logical model for information retrieval. Lecture Notes in Computer Science, 1134:812–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellmer:1996:ACS

[1415] E. Ellmer, C. Huemer, D. Merkl, and G. Pernul. Automatic classification of semantic concepts in view specifications. Lecture Notes in Computer Science, 1134:824–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staudt:1996:GST

[1416] M. Staudt and K. Von Thadden. A generic subsumption testing toolkit for knowledge base queries. Lecture Notes in Computer Science, 1134:834–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tran:1996:PET

[1417] N. Tran, D. P. Le, B. Srinivasan, and B. Sier. Performance evaluation of tree-based structures. Lecture Notes in Computer Science, 1134:845–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulos:1996:GPR

[1418] A. Papadopoulos and Y. Manolopoulos. Global page replacement in spatial databases. Lecture Notes in Computer Science, 1134:855–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheu:1996:ODD

[1419] S. Sheu, K. A. Hua, and Y. Cai. On the optimality of degree of declustering. Lecture Notes in Computer Science, 1134:865–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitagawa:1996:DES

[1420] H. Kitagawa, N. Watanabe, and Y. Ishikawa. Design and evaluation of signature file organization incorporating vertical and horizontal decomposition schemes. Lecture Notes in Computer Science, 1134:875–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schweighofer:1996:AGH

[1421] E. Schweighofer and D. Scheithauer.

The automatic generation of hypertext links in legal documents. Lecture Notes in Computer Science, 1134:889– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Si:1996:QPO

[1422] A. Si and H. V. Leong. Query processing and optimization for broadcast database. Lecture Notes in Computer Science, 1134:899–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tjoa:1996:DDE

[1423] A. M. Tjoa and R. R. Wagner. Developments in the database and expert systems applications domain. Lecture Notes in Computer Science, 1134:915– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cristian:1996:SGC

[1424] F. Cristian. On the semantics of group communication. Lecture Notes in Computer Science, 1135:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shankar:1996:UVP

[1425] N. Shankar. Unifying verification paradigms. Lecture Notes in Computer Science, 1135:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzmann:1996:FME

[1426] G. J. Holzmann. Formal methods for early fault detection. Lecture Notes in Computer Science, 1135:40–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maggiolo-Schettini:1996:RTS

[1427] A. Maggiolo-Schettini and A. Peron. Retiming techniques for statecharts. Lecture Notes in Computer Science, 1135:55–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maraninchi:1996:CAB

[1428] F. Maraninchi and N. Halbwachs. Compiling ARGOS into boleean equations. Lecture Notes in Computer Science, 1135:72–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paynter:1996:RTM

[1429] S. Paynter. Real-time mode-machines. Lecture Notes in Computer Science, 1135:90–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DArgenio:1996:CTA

[1430] P. R. D’Argenio and E. Brinksma. A calculus for timed automata. Lecture Notes in Computer Science, 1135:110– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Springintveld:1996:MTA

[1431] J. Springintveld and F. Vaandrager.

Minimizable timed automata. Lecture Notes in Computer Science, 1135:130– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pandya:1996:WCI

[1432] P. K. Pandya. Weak chop inverses and liveness in mean-value calculus. Lecture Notes in Computer Science, 1135: 148–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraenzle:1996:SCD

[1433] M. Fraenzle. Synthesizing controllers from duration calculus. Lecture Notes in Computer Science, 1135:168–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hung:1996:SSD

[1434] D. V. Hung and P. H. Giang. Sampling semantics of duration calculus. Lecture Notes in Computer Science, 1135:188– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dierks:1996:PCV

[1435] H. Dierks. The production cell: a verified real-time system. Lecture Notes in Computer Science, 1135:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karsisto:1996:VDD

[1436] K. Karsisto and A. Valmari. Verificationdriven development of a collisionavoidance protocol for the Ethernet. Lecture Notes in Computer Science, 1135:228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choquet-Geniet:1996:ECS

[1437] A. Choquet-Geniet, D. Geniet, and F. Cottet. Exhaustive computation of the scheduled task execution sequences of a real-time application. Lecture Notes in Computer Science, 1135:246– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davoli:1996:SDF

[1438] R. Davoli, F. Tamburini, and L.- A. Giachini. Scheduling data flow programs in hard real-time environments. Lecture Notes in Computer Science, 1135:263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janowski:1996:DSP

[1439] T. Janowski and M. Joseph. Dynamic scheduling in the presence of faults: Specification and verification. Lecture Notes in Computer Science, 1135:279– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mei:1996:EBF

[1440] A. Mei, Y. Igarashi, and N. Shimizu. Efficient broadcasting on faulty star networks. Lecture Notes in Computer Science, 1135:298–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1996:MCE

[1441] A. Bouajjani, Y. Lakhnech, and S. Yovine. Model-checking for extended timed temporal logics. Lecture Notes in Computer Science, 1135:306–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pagani:1996:POV

[1442] F. Pagani. Partial orders and verification of real-time systems. Lecture Notes in Computer Science, 1135:327– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amadio:1996:TMT

[1443] R. M. Amadio and M. Dam. Toward a modal theory of types for the λcalculus. Lecture Notes in Computer Science, 1135:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1996:GFR

[1444] C. Dietz. Graphical formalization of real-time requirements. Lecture Notes in Computer Science, 1135:366–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katoen:1996:SRT

[1445] J.-P. Katoen, R. Langerak, D. Latella, and E. Brinksma. On specifying realtime systems in a causality-based setting. Lecture Notes in Computer Science, 1135:385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Westhead:1996:VES

[1446] M. Westhead and S. Nadjm-Tehrani. Verification of embedded systems using synchronous observers. Lecture Notes in Computer Science, 1135:405– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBoer:1996:CRT

[1447] F. S. De Boer, H. Tej, W.-P. De Roever, and M. Van Hulst. Compositionality in real-time shared variable concurrency. Lecture Notes in Computer Science, 1135:420–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fowler:1996:FAR

[1448] S. Fowler and A. Wellings. Formal analysis of a real-time kernel specification. Lecture Notes in Computer Science, 1135:440–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biehl:1996:MDA

[1449] M. Biehl, N. Klarlund, and T. Rauhe. Mona: Decidable arithmetic in practice. Lecture Notes in Computer Science, 1135:459–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouali:1996:VSR

[1450] A. Bouali, J.-P. Marimorat, R. De Simone, and H. Toma. Verifying synchronous reactive systems programmed in ESTEREL. Lecture Notes in Computer Science, 1135:463–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huber:1996:ATD

[1451] F. Huber, B. Schaetz, A. Schmidt, and K. Spies. AutoFocus: a tool for distributed systems specification. Lecture Notes in Computer Science, 1135:467– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bawa:1996:TTV

[1452] R. K. Bawa and E. Encrenaz. A tool for translation of VHDL descriptions into a formal model and its application to formal verification and synthesis. Lecture Notes in Computer Science, 1135:471–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marrero:1996:ETV

[1453] M. Marrero and A. Suarez. EVGC: a tool for visualizing LOTOS behavioural specifications. Lecture Notes in Computer Science, 1135:475–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sedgewick:1996:ASR

[1454] R. Sedgewick. Analysis of shellsort and related algorithms. Lecture Notes in Computer Science, 1136:1–11, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1996:DC

[1455] A. Garg, R. Tamassia, and P. Vocca.

Drawing with colors. Lecture Notes in Computer Science, 1136:12–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirousis:1996:AUT

[1456] L. M. Kirousis, E. Kranakis, and D. Krizanc. Approximating the unsatisfiability threshold of random formulas. Lecture Notes in Computer Science, 1136:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwiegelshohn:1996:PWC

[1457] U. Schwiegelshohn. Preemptive weighted completion time scheduling of parallel jobs. Lecture Notes in Computer Science, 1136:39–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gergov:1996:AAD

[1458] J. Gergov. Approximation algorithms for dynamic storage allocation. Lecture Notes in Computer Science, 1136: 52–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trevisan:1996:PLP

[1459] L. Trevisan. Positive linear programming, parallel approximation and PCP’s. Lecture Notes in Computer Science, 1136:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moehring:1996:SJC

[1460] R. H. Moehring, M. W. Schaeffter, and A. S. Schulz. Scheduling jobs with communication delays: Using infeasible solutions for approximation. Lecture Notes in Computer Science, 1136: 76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roura:1996:RST

[1461] S. Roura and C. Martinez. Randomization of search trees by subtree size. Lecture Notes in Computer Science, 1136:91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferragina:1996:EDM

[1462] P. Ferragina and S. Muthukrishnan. Efficient dynamic method-lookup in object oriented languages. Lecture Notes in Computer Science, 1136:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raman:1996:PQS

[1463] R. Raman. Priority queues: Small, monotone and trans-dichotomous. Lecture Notes in Computer Science, 1136: 121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolte:1996:SAP

[1464] A. Nolte and R. Schrader. Simulated annealing and its problems to color graphs. Lecture Notes in Computer Science, 1136:138–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khuller:1996:CCP

[1465] S. Khuller and Y. J. Sussmann. The capacitated k-center problem. Lecture Notes in Computer Science, 1136:152– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujito:1996:ULR

[1466] T. Fujito. A unified local ratio approximation of node-deletion problem. Lecture Notes in Computer Science, 1136:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guha:1996:AAC

[1467] S. Guha and S. Khuller. Approximation algorithms for connected dominating sets. Lecture Notes in Computer Science, 1136:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frieze:1996:GAS

[1468] A. Frieze and W. Szpankowski. Greedy algorithms for the shortest common superstring that are asymptotically optimal. Lecture Notes in Computer Science, 1136:194–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsukiji:1996:DRG

[1469] T. Tsukiji and F. Xhafa. On the depth of randomly generated circuits. Lecture Notes in Computer Science, 1136: 208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viola:1996:ALP

[1470] A. Viola and P. V. Poblete. The analysis of linear probing hashing with buckets. Lecture Notes in Computer Science, 1136:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:PAP

[1471] D. Z. Chen, W. Chen, K. Wada, and K. Kawaguchi. Parallel algorithms for partitioning sorted sets and related problems. Lecture Notes in Computer Science, 1136:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czumaj:1996:FGR

[1472] A. Czumaj, P. Kanarek, M. Kutylowski, and K. Lorys. Fast generation of random permutations via networks simulation. Lecture Notes in Computer Science, 1136:246–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferragina:1996:PDD

[1473] P. Ferragina and F. Luccio. On the parallel dynamic dictionary matching problem: New results with application. Lecture Notes in Computer Science, 1136:261–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1996:PAS

[1474] H. L. Bodlaender and B. De Fluiter.

Parallel algorithms for series parallel graphs. Lecture Notes in Computer Science, 1136:277–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leoncini:1996:PCH

[1475] M. Leoncini, G. Manzini, and L. Margara. Parallel complexity of Householder QR factorization. Lecture Notes in Computer Science, 1136:290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basch:1996:RRB

[1476] J. Basch, L. Guibas, and G. D. Ramkumar. Reporting red-blue intersections between two sets of connected line segments. Lecture Notes in Computer Science, 1136:302–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:1996:MCS

[1477] H. Alt, U. Fuchs, G. Rote, and G. Weber. Matching convex shapes with respect to the symmetric difference. Lecture Notes in Computer Science, 1136:320–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arkin:1996:MAH

[1478] E. M. Arkin, Y.-J. Chiang, M. Held, and J. S. B. Mitchell. On minimumarea hulls. Lecture Notes in Computer Science, 1136:334–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherkassky:1996:NCD

[1479] B. V. Cherkassky and A. V. Goldberg. Negative cycle detection algorithms. Lecture Notes in Computer Science, 1136:349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1996:NRG

[1480] B. W. Watson. A new regular grammar pattern matching algorithm. Lecture Notes in Computer Science, 1136: 364–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaerkkaeinen:1996:LZI

[1481] J. Kaerkkaeinen and E. Sutinen.

Lempel–Ziv index for q-grams. Lecture Notes in Computer Science, 1136: 378–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heckmann:1996:CUL

[1482] R. Heckmann and T. Lengauer. Computing upper and lower bounds on textile nesting problems. Lecture Notes in Computer Science, 1136:392–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alon:1996:LLA

[1483] N. Alon, J. Csirik, S. Sevast’janov, and A. P. A. Vestjens. On-line and off-line algorithms for vector covering. Lecture Notes in Computer Science, 1136:406–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achlioptas:1996:CAR

[1484] D. Achlioptas, M. Chrobak, and J. Noga. Competitive analysis of randomized paging algorithms. Lecture Notes in Computer Science, 1136:419– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Awerbuch:1996:LCA

[1485] B. Awerbuch, Y. Azar, A. Fiat, and S. Leonardi. On-line competitive algorithms for call admission in optical networks. Lecture Notes in Computer Science, 1136:431–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1996:RCD

[1486] P. N. Klein, H.-I. Lu, and R. H. B. Netzer. Race-condition detection in parallel computation with semaphores. Lecture Notes in Computer Science, 1136: 445–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaklamanis:1996:EWR

[1487] C. Kaklamanis and P. Persiano. Efficient wavelength routing on directed fiber trees. Lecture Notes in Computer Science, 1136:460–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Aroya:1996:LBN

[1488] I. Ben-Aroya, D. D. Chinn, and A. Schuster. A lower bound for nearly minimal adaptive and hot potato algorithms. Lecture Notes in Computer Science, 1136:471–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christopher:1996:SCD

[1489] G. Christopher, M. Farach, and M. A. Trick. The structure of circular decomposable metrics. Lecture Notes in Computer Science, 1136:486–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dessmark:1996:FAS

[1490] A. Dessmark, A. Lingas, and A. Proskurowski. Faster algorithms for subgraph isomorphism of k-connected partial ktrees. Lecture Notes in Computer Science, 1136:501–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arikati:1996:PSA

[1491] S. Arikati, D. Z. Chen, L. P. Chew, and G. Das. Planar spanners and approximate shortest path queries among obstacles in the plane. Lecture Notes in Computer Science, 1136:514–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auletta:1996:NAO

[1492] V. Auletta, D. Parente, and P. Persiano. A new approach to optimal planning of robot motion on a tree with obstacles. Lecture Notes in Computer Science, 1136:529–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Ortiz:1996:GSR

[1493] A. Lopez-Ortiz and S. Schuierer. Generalized streets revisited. Lecture Notes in Computer Science, 1136:546–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadimitriou:1996:CAO

[1494] C. Papadimitriou. Computational aspects of organization theory. Lecture Notes in Computer Science, 1136:559– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bornscheuer:1996:RMN

[1495] S.-E Bornscheuer. Rational models of normal logic programs. Lecture Notes in Computer Science, 1137:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bornscheuer:1996:MPR

[1496] S.-E. Bornscheuer and T. Seiler. Massively parallel reasoning about actions. Lecture Notes in Computer Science, 1137:5–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aravindan:1996:DDL

[1497] C. Aravindan. DisLoP: a disjunctive logic programming system based on PROTEIN theorem prover. Lecture Notes in Computer Science, 1137:19– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Contzen:1996:AHG

[1498] M. Contzen and K. Moeller. Abstraction as a heuristic to guide planning. Lecture Notes in Computer Science, 1137:25–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denzinger:1996:ETP

[1499] J. Denzinger. Equational theorem proving using taxonomic constraints. Lecture Notes in Computer Science, 1137:29–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denzinger:1996:PDT

[1500] J. Denzinger and M. Kronenburg. Planning for distributed theorem proving: The teamwork approach. Lecture Notes in Computer Science, 1137:43– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dung:1996:RDD

[1501] Truong Quoc Dung. A revision of dependency-directed backtracking for JTMS. Lecture Notes in Computer Science, 1137:57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebersberger:1996:CIS

[1502] M. Ebersberger, J. Mueller, and H. Stahl. A compiler-interpreter-system for decoding the user’s intention within a speech understanding application. Lecture Notes in Computer Science, 1137:61–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fronhoefer:1996:CRL

[1503] B. Fronhoefer. Cyclic rules in linear connection proofs. Lecture Notes in Computer Science, 1137:67–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1996:ONN

[1504] M. Fuchs and A. Abecker. Optimized nearest-neighbor classifiers using generated instances. Lecture Notes in Computer Science, 1137:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garrido-Luna:1996:TTD

[1505] L. Garrido-Luna and K. Sycara. Towards a totally distributed meeting scheduling system. Lecture Notes in Computer Science, 1137:85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1996:RBC

[1506] C. Habel. Representations as basis of cognitive processes. Lecture Notes in Computer Science, 1137:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirschfeld:1996:IFT

[1507] D. Hirschfeld and H.-D. Maas. Improving the functionality of a textto-speech system by adding morphological knowledge. Lecture Notes in Computer Science, 1137:103–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoos:1996:SHC

[1508] H. H. Hoos. Solving hard combinatorial problems with GSAT. Lecture Notes in Computer Science, 1137:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutter:1996:URE

[1509] D. Hutter. Using rippling for equational reasoning. Lecture Notes in Computer Science, 1137:121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jablonowski:1996:RAA

[1510] J. Jablonowski, W. Lukaszewicz, and E. Madalinska-Bugaj. Reasoning about action and change: Defeasible observations and actions with abnormal effects. Lecture Notes in Computer Science, 1137:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamp:1996:FCB

[1511] G. Kamp, P. Pirk, and H.-D. Burkhard.

FALLDATEN: Case-based reasoning for the diagnosis of technical devices. Lecture Notes in Computer Science, 1137:149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasper:1996:IPG

[1512] W. Kasper and H.-U. Krieger. Integration of prosodic and grammatical information in the analysis of dialogs. Lecture Notes in Computer Science, 1137:163–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klopotek:1996:AGA

[1513] Mieczyslaw Alojzy Klopotek, S. T.

Wierzchon, and M. Michalewicz. Application of a genetic algorithm for plausible justification of observations. Lecture Notes in Computer Science, 1137:175–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kraetzschmar:1996:MMC

[1514] G. K. Kraetzschmar and J. Schneeberger. Managing multiple contexts efficiently. Lecture Notes in Computer Science, 1137:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuechler:1996:ILS

[1515] A. Kuechler and C. Goller. Inductive learning in symbolic domains using structure-driven recurrent neural networks. Lecture Notes in Computer Science, 1137:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakemeyer:1996:BRN

[1516] G. Lakemeyer and W. Lang. Belief revision in a nonclassical logic. Lecture Notes in Computer Science, 1137:199– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1996:ACE

[1517] H. Lange, R. Moeller, and B. Neumann. Avoiding combinatorial explosion in automatic test generation: Reasoning about measurements is the key. Lecture Notes in Computer Science, 1137:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenz:1996:CRN

[1518] M. Lenz and H.-D Burkhard. Case retrieval nets: Basic ideas and extensions. Lecture Notes in Computer Science, 1137:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linke:1996:PDL

[1519] T. Linke and T. Schaub. Putting default logics in perspective. Lecture Notes in Computer Science, 1137:241– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Margelisch:1996:EPK

[1520] A. Margelisch, B. Novotny, and U. Reimer. EULE2: a prototypical knowledge-based decision support system for the performance of office tasks. Lecture Notes in Computer Science, 1137:253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melis:1996:WPT

[1521] E. Melis. When to prove theorems by analogy? Lecture Notes in Computer Science, 1137:259–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1996:PKR

[1522] C.-H. Meyer and W. Roedder. Probabilistic knowledge representation and reasoning at maximum entropy by SPIRIT. Lecture Notes in Computer Science, 1137:273–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niemann:1996:KBI

[1523] H. Niemann, V. Fischer, D. Paulus, and J. Fischer. Knowledge based image understanding by iterative optimization. Lecture Notes in Computer Science, 1137:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oertel:1996:KOU

[1524] W. Oertel. Knowledge organization using the development system FAENSY. Lecture Notes in Computer Science, 1137:303–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otten:1996:UPP

[1525] J. Otten and C. Kreitz. A uniform proof procedure for classical and non-classical logics. Lecture Notes in Computer Science, 1137:307–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholczyk:1996:QRU

[1526] D. Pacholczyk and G. Hunault. Qualitative reasoning under uncertainty with symbolic probabilities. Lecture Notes in Computer Science, 1137:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radzikowska:1996:NCA

[1527] A. Radzikowska. Narratives with concurrent actions: The mutual exclusion problem. Lecture Notes in Computer Science, 1137:327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Russell:1996:TAA

[1528] S. Russell. Tools for autonomous agents. Lecture Notes in Computer Science, 1137:331–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sattler:1996:CLE

[1529] U. Sattler. A concept language extended with different kinds of transitive roles. Lecture Notes in Computer Science, 1137:333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheich:1996:RAC

[1530] H. Scheich. Representation in auditory cortex. Lecture Notes in Computer Science, 1137:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnattinger:1996:TQC

[1531] K. Schnattinger and U. Hahn. A terminological qualification calculus for preferential reasoning under uncertainty. Lecture Notes in Computer Science, 1137:349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winiwarter:1996:SDU

[1532] W. Winiwarter, O. Kagawa, and Y. Kambayashi. Syntactic disambiguation by using categorial parsing in a DOOD framework. Lecture Notes in Computer Science, 1137:363–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuertz:1996:CDR

[1533] J. Wuertz and T. Mueller. Constructive disjunction revisited. Lecture Notes in Computer Science, 1137:377–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scott:1996:SCT

[1534] D. S. Scott. Symbolic computation and teaching. Lecture Notes in Computer Science, 1138:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauer:1996:AEC

[1535] A. Bauer, E. Clarke, and X. Zhao. Analytica — an experiment in combining theorem proving and symbolic computation. Lecture Notes in Computer Science, 1138:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1996:DRS

[1536] G. Butler, P. Grogoho, R. Shinghal, and I. Tjandra. Document recognition, semantics, and symbolic reasoning in reverse engineering of software. Lecture Notes in Computer Science, 1138:38–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dargam:1996:CUL

[1537] F. C. C. Dargam. Compromised updates in labelled databases. Lecture Notes in Computer Science, 1138:49– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roanes-Lozano:1996:IEP

[1538] E. Roanes-Lozano, L. M. Laita, and E. Roanes-Macias. An inference engine for propositional two-valued logic based on the radical membership problem. Lecture Notes in Computer Science, 1138:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dillmann:1996:PDM

[1539] R. Dillmann and H. Friedrich. Programming by demonstration: a machine learning approach to support skill acquisition for robots. Lecture Notes in Computer Science, 1138:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reinhart:1996:KBI

[1540] G. Reinhart, R. Diesch, and M. R. Koch. Knowledge-based information processing in manufacturing cells — the present and the future. Lecture Notes in Computer Science, 1138:109– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohn:1996:CQS

[1541] A. G. Cohn. Calculi for qualitative spatial reasoning. Lecture Notes in Computer Science, 1138:124–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benhamou:1996:CLC

[1542] F. Benhamou and L. Granvilliers.

Combining local consistency, symbolic rewriting and interval methods. Lecture Notes in Computer Science, 1138: 144–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buendgen:1996:PTN

[1543] R. Buendgen. Proof transformation for non-compatible rewriting. Lecture Notes in Computer Science, 1138:160– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lynch:1996:PGE

[1544] C. Lynch and P. Strogova. PATCH graphs: An efficient data structure for completion of finitely presented groups. Lecture Notes in Computer Science, 1138:176–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dupee:1996:MLE

[1545] B. J. Dupee. Measuring the likely effectiveness of strategies. Lecture Notes in Computer Science, 1138:191–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodosek:1996:NAS

[1546] R. Rodosek. A new approach on solving 3-Satisfiability. Lecture Notes in Computer Science, 1138:197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:GMA

[1547] D. Wang. Geometry machines: From artificial intelligence to symbolic mathematical computation. Lecture Notes in Computer Science, 1138:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ueberberg:1996:ITP

[1548] J. Ueberberg. Interactive theorem proving and finite projective planes. Lecture Notes in Computer Science, 1138:240–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joswig:1996:TMT

[1549] M. Joswig. Towards modelling the topology of homogeneous manifolds by means of symbolic computation. Lecture Notes in Computer Science, 1138: 258–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouhineau:1996:SGC

[1550] D. Bouhineau. Solving geometrical constraint systems using CLP based on linear constraint solver. Lecture Notes in Computer Science, 1138:274–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sofronie:1996:TSS

[1551] V. Sofronie. Towards a sheaf semantics for cooperating agents scenarios. Lecture Notes in Computer Science, 1138:289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Telerman:1996:DTS

[1552] V. Telerman and D. Ushakov. Data types in subdefinite models. Lecture Notes in Computer Science, 1138:305– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrianarivelo:1996:TPH

[1553] N. Andrianarivelo, W. Bousdira, and J.-M. Talbot. On theorem-proving in Horn theories with built-in algebras. Lecture Notes in Computer Science, 1138:320–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eder:1996:BRS

[1554] E. Eder. Backward reasoning in systems with cut. Lecture Notes in Computer Science, 1138:339–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaza:1996:SCV

[1555] J. A. Plaza. Soundness and completeness versus lifting property. Lecture Notes in Computer Science, 1138:354– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavilanes:1996:RPD

[1556] A. Gavilanes, J. Leach, P. J. Martin, and S. Nieva. Reasoning with preorders and dynamic sorts using free variable tableaux. Lecture Notes in Computer Science, 1138:365–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alouini:1996:TCI

[1557] I. Alouini and C. Kirchner. Toward the concurrent implementation of computational systems. Lecture Notes in Computer Science, 1139:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiPierro:1996:NI

[1558] A. Di Pierro and W. Drabent. On negation as instantiation. Lecture Notes in Computer Science, 1139:32–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GarciadelaBanda:1996:IDS

[1559] M. Garcia de la Banda, M. Hermenegildo, and K. Marriott. Independence in dynamically scheduled logic languages. Lecture Notes in Computer Science, 1139:47–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benhamou:1996:HCS

[1560] F. Benhamou. Heterogeneous constraint solving. Lecture Notes in Computer Science, 1139:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fages:1996:HSN

[1561] F. Fages and R. Gori. A hierarchy of semantics for normal constraint logic programs. Lecture Notes in Computer Science, 1139:77–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oelveczky:1996:OST

[1562] P. C. Oelveczky and O. Lysne. Ordersorted termination: The unsorted way. Lecture Notes in Computer Science, 1139:92–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchiori:1996:UUP

[1563] M. Marchiori. Unravelings and ultraproperties. Lecture Notes in Computer Science, 1139:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1996:STR

[1564] T. Suzuki. Standardization theorem revisited. Lecture Notes in Computer Science, 1139:122–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khasidashvili:1996:DNS

[1565] Z. Khasidashvili and J. Glauert. Discrete normalization and standardization in deterministic residual structures. Lecture Notes in Computer Science, 1139:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caires:1996:LLS

[1566] L. Caires. A language for the logical specification of processes and relations. Lecture Notes in Computer Science, 1139:150–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brim:1996:PAS

[1567] L. Brim, D. Gilbert, J.-M. Jacquet, and M. Kretinsky. A process algebra for synchronous concurrent constraint programming. Lecture Notes in Computer Science, 1139:165–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ancona:1996:AAM

[1568] D. Ancona and E. Zucca. An algebraic approach to mixins and modularity. Lecture Notes in Computer Science, 1139:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fassbender:1996:SBD

[1569] H. Fassbender and S. Maneth. A strict border for the decidability of Eunification for recursive functions. Lecture Notes in Computer Science, 1139: 194–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchiori:1996:PFA

[1570] E. Marchiori. Prime factorizations of abstract domains using first order logic. Lecture Notes in Computer Science, 1139:209–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volpe:1996:AUP

[1571] P. Volpe. Abstractions of uniform proofs. Lecture Notes in Computer Science, 1139:224–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giacobazzi:1996:CLP

[1572] R. Giacobazzi and F. Ranzato. Complementing logic program semantics. Lecture Notes in Computer Science, 1139:238–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennaway:1996:MTR

[1573] R. Kennaway, V. Van Oostrom, and F.-J. De Vries. Meaningless terms in rewriting. Lecture Notes in Computer Science, 1139:254–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1996:LCE

[1574] M. C. F. Ferreira, D. Kesner, and L. Puel. lambda-calculi with explicit substitutions and composition which preserve ?-strong normalization. Lecture Notes in Computer Science, 1139: 284–298, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Almendros-Jimenez:1996:ASF

[1575] J. M. Almendros-Jimenez, A. Gavilanes-Franco, and A. Gil-Luezas. Algebraic semantics for functional logic programming with polymorphic order-sorted types. Lecture Notes in Computer Science, 1139:299–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nazareth:1996:STS

[1576] D. Nazareth. Specifying type systems. Lecture Notes in Computer Science, 1139:314–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nazareth:1996:STP

[1577] D. Nazareth. The semantic treatment of polymorphic specification languages. Lecture Notes in Computer Science, 1139:330–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meertens:1996:CP

[1578] L. Meertens. Calculate polytypically! Lecture Notes in Computer Science, 1140:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahrs:1996:LMD

[1579] S. Kahrs. Limits of ML-Definability. Lecture Notes in Computer Science, 1140:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belle:1996:FM

[1580] G. Belle, C. B. Jay, and E. Moggi. Functorial ML. Lecture Notes in Computer Science, 1140:32–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Louvet:1996:PPT

[1581] P. Louvet and O. Ridoux. Parametric polymorphism for typed Prolog and lambda-Prolog. Lecture Notes in Computer Science, 1140:47–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mateos-Lago:1996:GAS

[1582] J. Mateos-Lago and M. Rodriguez-Artalejo. GOTA algebras: a specification formalism for inheritance and object hierarchies. Lecture Notes in Computer Science, 1140:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GarciadelaBanda:1996:TIP

[1583] M. Garcia de la Banda, F. Bueno, and M. Hermenegildo. Towards independent And-parallelism in CLP. Lecture Notes in Computer Science, 1140:77– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weyer:1996:ASS

[1584] G. Weyer and W. Winsborough. Annotated structure shape graphs for abstract analysis of Prolog. Lecture Notes in Computer Science, 1140:92– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bagnara:1996:RIP

[1585] R. Bagnara. A reactive implementation of Pos using ROBDDs. Lecture Notes in Computer Science, 1140:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parigot:1996:DAG

[1586] D. Parigot, G. Roussel, M. Jourdan, and E. Duris. Dynamic attribute grammars. Lecture Notes in Computer Science, 1140:122–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leuschel:1996:LPS

[1587] M. Leuschel and D. De Schreye.

Logic program specialisation: How to be more specific. Lecture Notes in Computer Science, 1140:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glueck:1996:CCP

[1588] R. Glueck, J. Joergensen, B. Martens, and M. H. Soerensen. Controlling conjunctive partial deduction. Lecture Notes in Computer Science, 1140:152– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFrancesco:1996:UFT

[1589] N. De Francesco and A. Santone. Unfold/fold transformations of concurrent processes. Lecture Notes in Computer Science, 1140:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danvy:1996:SBC

[1590] O. Danvy and R. Vestergaard. Semantics-based compiling: a case study in type-directed partial evaluation. Lecture Notes in Computer Science, 1140:182–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thiemann:1996:IMP

[1591] P. Thiemann. Implementing memoization for partial evaluation. Lecture Notes in Computer Science, 1140:198– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamilton:1996:HOD

[1592] G. W. Hamilton. Higher order deforestation. Lecture Notes in Computer Science, 1140:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kessler:1996:SED

[1593] C. W. Kessler. Scheduling expression DAGs for minimal register need. Lecture Notes in Computer Science, 1140: 228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freire:1996:BDF

[1594] J. Freire, T. Swift, and D. S. Warren. Beyond depth-first: Improving tabled logic programs through alternative scheduling strategies. Lecture Notes in Computer Science, 1140:243– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:PSN

[1595] X. Li. Program sharing: a new implementation approach for Prolog. Lecture Notes in Computer Science, 1140: 259–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorlatch:1996:SEI

[1596] S. Gorlatch. Systematic extraction and implementation of divide-and-conquer parallelism. Lecture Notes in Computer Science, 1140:274–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galan:1996:FSG

[1597] L. A. Galán, C. Pareja, and R. Peña.

Functional skeletons generate process topologies in Eden. Lecture Notes in Computer Science, 1140:289–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehrke:1996:PLE

[1598] T. Gehrke and M. Huhn. ProFun — A language for executable specifications. Lecture Notes in Computer Science, 1140:304–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1996:TRG

[1599] M. Fernandez and I. Mackie. From term rewriting to generalised interaction nets. Lecture Notes in Computer Science, 1140:319–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aponte:1996:TIM

[1600] M.-V. Aponte and R. Di Cosmo. Type isomorphisms for module signatures. Lecture Notes in Computer Science, 1140:334–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruggieri:1996:DLP

[1601] S. Ruggieri. Decidability of logic program semantics and applications to testing. Lecture Notes in Computer Science, 1140:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholz:1996:PUP

[1602] E. Scholz. PIDGETS — unifying pictures and widgets in a constraint-based framework for concurrent functional GUI programming. Lecture Notes in Computer Science, 1140:363–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamareddine:1996:GRE

[1603] F. Kamareddine and A. Rios. Generalizedreduction and explicit substitution. Lecture Notes in Computer Science, 1140:378–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benaissa:1996:MSR

[1604] Z.-E.-A. Benaissa, P. Lescanne, and K. Hoegsbro Rose. Modeling sharing and recursion for weak reduction strategies using explicit substitution. Lecture Notes in Computer Science, 1140:393–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucas:1996:CSC

[1605] S. Lucas. Context-sensitive computations in confluent programs. Lecture Notes in Computer Science, 1140:408– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stuckey:1996:MUS

[1606] P. J. Stuckey and V. Tam. Models for using stochastic constraint solvers in constraint logic programming. Lecture Notes in Computer Science, 1140: 423–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanRoy:1996:IER

[1607] P. Van Roy, M. Mehl, and R. Scheidhauer. Integrating efficient records into concurrent constraint programming. Lecture Notes in Computer Science, 1140:438–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1996:LDD

[1608] M. Liu. The LOL deductive database programming language. Lecture Notes in Computer Science, 1140:454–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fecht:1996:EPS

[1609] C. Fecht. An efficient and precise sharing domain for logic programs. Lecture Notes in Computer Science, 1140: 469–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1996:CTC

[1610] Z. Hu, H. Iwasaki, and M. Takeichi. Cheap tupling in calculational form. Lecture Notes in Computer Science, 1140:471–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoy:1996:NNP

[1611] S. Antoy. Needed narrowing in Prolog. Lecture Notes in Computer Science, 1140:473–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puebla:1996:AOD

[1612] G. Puebla and M. Hermenegildo. Automatic optimization of dynamic scheduling in logic programs. Lecture Notes in Computer Science, 1140:475–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulte:1996:OEV

[1613] C. Schulte. Oz Explorer: a visual constraint programming tool. Lecture Notes in Computer Science, 1140:477– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agapow:1996:CBE

[1614] P.-M. Agapow. Computational brittleness and the evolution of computer viruses. Lecture Notes in Computer Science, 1141:2–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bull:1996:ECM

[1615] L. Bull and T. C. Fogarty. Evolutionary computing in multi-agent environments: Speciation and symbiogenesis. Lecture Notes in Computer Science, 1141:12–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herdy:1996:ESS

[1616] M. Herdy. Evolution strategies with subjective selection. Lecture Notes in Computer Science, 1141:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iba:1996:ECM

[1617] H. Iba. Emergent cooperation for multiple agents using genetic programming. Lecture Notes in Computer Science, 1141:32–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacob:1996:EPE

[1618] C. Jacob. Evolution programs evolved.

Lecture Notes in Computer Science, 1141:42–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakobi:1996:ESI

[1619] N. Jakobi. Encoding scheme issues for open-ended artificial evolution. Lecture Notes in Computer Science, 1141:52– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murakawa:1996:HEF

[1620] M. Murakawa, S. Yoshizawa, I. Kajitani, and T. Furuya. Hardware evolution at function level. Lecture Notes in Computer Science, 1141:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paredis:1996:CLT

[1621] J. Paredis. Coevolutionary life-time learning. Lecture Notes in Computer Science, 1141:72–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ronge:1996:GPC

[1622] A. Ronge and M. G. Nordahl. Genetic programs and co-evolution: Developing robust general purpose controllers using local mating in 2-dimensional populations. Lecture Notes in Computer Science, 1141:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spirov:1996:SAG

[1623] A. V. Spirov. Self-assemblage of gene nets in evolution via recruiting of new netters. Lecture Notes in Computer Science, 1141:91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1996:SIR

[1624] A. S. Wu and R. K. Lindsay. A survey of Intron research in genetics. Lecture Notes in Computer Science, 1141:101–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asselmeyer:1996:ANI

[1625] T. Asselmeyer, W. Ebeling, and H. Rose. Analytical and numerical investigations of evolutionary algorithms in continuous spaces. Lecture Notes in Computer Science, 1141:112–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beyer:1996:ABM

[1626] H.-G. Beyer. On the asymptotic behavior of multirecombinant evolution strategies. Lecture Notes in Computer Science, 1141:122–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoehn:1996:LPP

[1627] C. Hoehn and C. Reeves. Are long path problems hard for genetic algorithms. Lecture Notes in Computer Science, 1141:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iba:1996:RTG

[1628] H. Iba. Random tree generation for genetic programming. Lecture Notes in Computer Science, 1141:144–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jelasity:1996:IFG

[1629] M. Jelasity and J. Dombi. Implicit formae in genetic algorithms. Lecture Notes in Computer Science, 1141:154– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menon:1996:PDA

[1630] A. Menon, K. Mehrotra, C. K. Mohan, and S. Ranka. A probabilistic database approach to the analysis of genetic algorithms. Lecture Notes in Computer Science, 1141:164–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molgedey:1996:MFA

[1631] L. Molgedey. Mean field analysis of tournament selection on a random manifold. Lecture Notes in Computer Science, 1141:174–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muehlenbein:1996:RGEa

[1632] H. Muehlenbein and G. Paass. From recombination of genes to the estimation of distributions: I. binary parameters. Lecture Notes in Computer Science, 1141:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muehlenbein:1996:RGEb

[1633] H. Muehlenbein, J. Bendisch, and H.-M Voigt. From recombination of genes to the estimation of distributions: II. continuous parameters. Lecture Notes in Computer Science, 1141:188–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rana:1996:SPN

[1634] S. Rana, L. D. Whitley, and R. Cogswell. Searching in the presence of noise. Lecture Notes in Computer Science, 1141:198–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rose:1996:DSM

[1635] H. Rose, W. Ebeling, and T. Asselmeyer. The density of states — A measure of the difficulty of optimisation problems. Lecture Notes in Computer Science, 1141:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rudolph:1996:IEA

[1636] G. Rudolph. On interactive evolutionary algorithms and stochastic mealy automata. Lecture Notes in Computer Science, 1141:218–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomon:1996:IDC

[1637] R. Salomon. The influence of different coding schemes on the computational complexity of genetic algorithms in function optimization. Lecture Notes in Computer Science, 1141:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarma:1996:AEN

[1638] J. Sarma and K. De Jong. An analysis of the effects of neighborhood size and shape on local selection algorithms. Lecture Notes in Computer Science, 1141:236–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoenauer:1996:ECE

[1639] M. Schoenauer and Z. Michalewicz. Evolutionary computation at the edge of feasibility. Lecture Notes in Computer Science, 1141:245–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thierens:1996:DAA

[1640] D. Thierens. Dimensional analysis of allele-wise mixing revisited. Lecture Notes in Computer Science, 1141:255– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voget:1996:GDS

[1641] S. Voget. Gaussian diffusion in a simple genetic algorithm. Lecture Notes in Computer Science, 1141:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voigt:1996:ETS

[1642] H.-M. Voigt and H. Muehlenbein. Erroneous truncation selection — A breeder’s decision making perspective. Lecture Notes in Computer Science, 1141:276–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asveren:1996:NCM

[1643] T. Asveren and P. Molitor. New crossover methods for sequencing problems. Lecture Notes in Computer Science, 1141:290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banzhaf:1996:EEU

[1644] W. Banzhaf, F. D. Francone, and P. Nordin. The effect of extensive use of the mutation operator on generalization in genetic programming using sparse data sets. Lecture Notes in Computer Science, 1141:300–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bierwirth:1996:PRS

[1645] C. Bierwirth, D. C. Mattfeld, and H. Kopfer. On permutation representations for scheduling problems. Lecture Notes in Computer Science, 1141: 310–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiben:1996:MPN

[1646] A. E. Eiben and C. A. Schippers.

Multi-parent’s niche: n-ary crossovers on NK-Landscapes. Lecture Notes in Computer Science, 1141:319–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghozeil:1996:PID

[1647] A. Ghozeil and D. B. Fogel. A preliminary investigation into directed mutations in evolutionary algorithms. Lecture Notes in Computer Science, 1141: 329–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrera:1996:HCR

[1648] F. Herrera and M. Lozano. Heuristic crossovers for real-coded genetic algorithms based on fuzzy connectives. Lecture Notes in Computer Science, 1141:336–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappler:1996:EAI

[1649] C. Kappler. Are evolutionary algorithms improved by large mutations. Lecture Notes in Computer Science, 1141:346–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebag:1996:MIB

[1650] M. Sebag and M. Schoenauer. Mutation by imitation in Boolean evolution strategies. Lecture Notes in Computer Science, 1141:356–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Surry:1996:FAF

[1651] P. D. Surry and N. J. Radcliffe. Formal algorithms + formal representations = search strategies. Lecture Notes in Computer Science, 1141:366–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vavak:1996:GAV

[1652] F. Vavak, T. C. Fogarty, and K. Jukes. A genetic algorithm with variable range of local search for tracking changing environments. Lecture Notes in Computer Science, 1141:376–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Born:1996:ESA

[1653] J. Born. An evolution strategy with adaptation of the step sizes by a variance function. Lecture Notes in Computer Science, 1141:388–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darwen:1996:ENM

[1654] P. Darwen and X. Yao. Every niching method has its niche: Fitness sharing and implicit sharing compared. Lecture Notes in Computer Science, 1141: 398–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

East:1996:EID

[1655] I. R. East and J. Rowe. Effects of isolation in a distributed population genetic algorithm. Lecture Notes in Computer Science, 1141:408–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinterding:1996:SAG

[1656] R. Hinterding, Z. Michalewicz, and T. C. Peachey. Self-adaptive genetic algorithm for numeric functions. Lecture Notes in Computer Science, 1141: 420–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedroso:1996:NSE

[1657] J. P. Pedroso. Niche search: An evolutionary algorithm for global optimisation. Lecture Notes in Computer Science, 1141:430–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:APE

[1658] J. E. Smith and T. C. Fogarty. Adaptive parameterised evolutionary systems: Self adaptive recombination and mutation in a genetic algorithm. Lecture Notes in Computer Science, 1141: 441–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turner:1996:OMD

[1659] A. Turner, D. Corne, G. Ritchie, and P. Ross. Obtaining multiple distinct solutions with genetic algorithm niching methods. Lecture Notes in Computer Science, 1141:451–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuson:1996:CBO

[1660] A. Tuson and P. Ross. Cost based operator rate adaption: An investigation. Lecture Notes in Computer Science, 1141:461–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collard:1996:GAR

[1661] P. Collard, C. Escazut, and A. Gaspar. Genetic algorithms and relational landscapes. Lecture Notes in Computer Science, 1141:472–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forsyth:1996:IIO

[1662] R. S. Forsyth. IOGA: An instanceoriented genetic algorithm. Lecture Notes in Computer Science, 1141:482– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanKemenade:1996:EFB

[1663] K. C. H. M. Van Kemenade. Explicit filtering of building blocks for genetic algorithms. Lecture Notes in Computer Science, 1141:494–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kita:1996:MOO

[1664] H. Kita, Y. Yabumoto, N. Mori, and Y. Nishikawa. Multi-objective optimization by means of the thermodynamical genetic algorithm. Lecture Notes in Computer Science, 1141:504– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mori:1996:ACE

[1665] N. Mori, H. Kita, and Y. Nishikawa.

Adaption to a changing environment by means of the thermodynamical genetic algorithm. Lecture Notes in Computer Science, 1141:513–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parmee:1996:DDA

[1666] I. C. Parmee. The development of a dual-agent strategy for efficient search across whole system engineering design hierarchies. Lecture Notes in Computer Science, 1141:523–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoneveld:1996:PCG

[1667] A. Schoneveld, J. F. De Ronde, P. M. A. Sloot, and J. A. Kaandorp. A parallel cellular genetic algorithm used in finite element simulation. Lecture Notes in Computer Science, 1141:533–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsutsui:1996:RSS

[1668] S. Tsutsui, A. Ghosh, and Y. Fujimoto.

A robust solution searching scheme in genetic search. Lecture Notes in Computer Science, 1141:543–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernier:1996:SMM

[1669] J. L. Bernier, C. I. Herraiz, J. J. Merelo, and S. Olmeda. Solving master mind using GAs and simulated annealing: a case of dynamic constraint optimization. Lecture Notes in Computer Science, 1141:554–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blickle:1996:ECS

[1670] T. Blickle. Evolving compact solutions in genetic programming. Lecture Notes in Computer Science, 1141:564– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duvivier:1996:CNH

[1671] D. Duvivier, P. Preux, and E.-G. Talbi. Climbing up NP-Hard hills. Lecture Notes in Computer Science, 1141:574– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fonseca:1996:PAC

[1672] C. M. Fonseca and P. J. Fleming. On the performance assessment and comparison of stochastic multiobjective optimizers. Lecture Notes in Computer Science, 1141:584–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lagus:1996:PGN

[1673] K. Lagus, I. Karanta, and J. Ylae-Jaeaeski. Paginating the generalized newspaper — A comparison of simulated annealing and a heuristic method. Lecture Notes in Computer Science, 1141:594–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McIlhagga:1996:COT

[1674] M. McIlhagga, P. Husbands, and R. Ives. A comparison of optimization techniques for integrated manufacturing planning and scheduling. Lecture Notes in Computer Science, 1141:604– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McIlhagga:1996:CST

[1675] M. McIlhagga, P. Husbands, and R. Ives. A comparison of search techniques on a wing-box optimisation problem. Lecture Notes in Computer Science, 1141:614–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neubauer:1996:CSE

[1676] A. Neubauer. A comparative study of evolutionary algorithms for on-line parameter tracking. Lecture Notes in Computer Science, 1141:624–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeck:1996:MUG

[1677] T. Baeck, H. Doernemann, U. Hammel, and P. Frankhauser. Modeling urban growth by cellular automata. Lecture Notes in Computer Science, 1141: 636–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dittes:1996:DOD

[1678] F.-M. Dittes. Democratic optimization for discrete and continuous systems. Lecture Notes in Computer Science, 1141:646–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorigo:1996:SSP

[1679] M. Dorigo and L. M. Gambardella. A study of some properties of Ant-Q. Lecture Notes in Computer Science, 1141:656–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishiguro:1996:IIA

[1680] A. Ishiguro, T. Kondo, Y. Watanabe, and Y. Uchikawa. Immunoid: An immunological approach to decentralized behavior arbitration of autonomous mobile robots. Lecture Notes in Computer Science, 1141:666–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pelillo:1996:PED

[1681] M. Pelillo and I. M. Bomze. Parallelizable evolutionary dynamics principles for solving the maximum clique problem. Lecture Notes in Computer Science, 1141:676–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rudolph:1996:SLS

[1682] G. Rudolph and J. Sprave. Significance of locality and selection pressure in the Grand Deluge Evolutionary Algorithm. Lecture Notes in Computer Science, 1141:686–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zauner:1996:PCD

[1683] K.-P. Zauner and M. Conrad. Parallel computing with DNA: Toward the antiuniversal machine. Lecture Notes in Computer Science, 1141:696–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carse:1996:TCD

[1684] B. Carse and T. C. Fogarty. Tackling the “curse of dimensionality” of radial basis function neural networks using a genetic algorithm. Lecture Notes in Computer Science, 1141:710–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordon:1996:TSM

[1685] O. Cordon, F. Herrera, and M. Lozano.

A three-stage method for designing genetic fuzzy systems by learning from examples. Lecture Notes in Computer Science, 1141:720–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drechsler:1996:LHO

[1686] R. Drechsler, N. Goeckel, and B. Becker. Learning heuristics for OBDD minimization by evolutionary algorithms. Lecture Notes in Computer Science, 1141:730–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galic:1996:IGP

[1687] E. Galic and M. Hoehfeld. Improving the generalization performance of multi-layer-perceptrons with population-based incremental learning. Lecture Notes in Computer Science, 1141:740–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hondo:1996:RGR

[1688] N. Hondo, H. Iba, and Y. Kakazu. Robust GP in robot learning. Lecture Notes in Computer Science, 1141:751– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwata:1996:PRS

[1689] M. Iwata, I. Kajitani, H. Yamada, and H. Iba. A pattern recognition system using evolvable hardware. Lecture Notes in Computer Science, 1141:761– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stepniewski:1996:TDF

[1690] S. W. Stepniewski and A. J. Keane. Topology design of feedforward neural networks by genetic algorithms. Lecture Notes in Computer Science, 1141: 771–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sullivan:1996:ESL

[1691] C. C. W. Sullivan and A. G. Pipe. An evolution strategy for on-line optimisation of dynamic objective functions. Lecture Notes in Computer Science, 1141:781–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corno:1996:ECS

[1692] F. Corno, P. Prinetto, M. Rebaudengo, and M. S. Reorda. Exploiting competing subpopulations for automatic generation of test sequences for digital cicuits. Lecture Notes in Computer Science, 1141:792–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorne:1996:CHE

[1693] R. Dorne and J.-K. Hao. Constraint handling in evolutionary search: a case study of the frequency assignment. Lecture Notes in Computer Science, 1141:801–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hillermeier:1996:AGA

[1694] C. Hillermeier and J. Keppler. An application of genetic algorithms and neural networks to scheduling power generating systems. Lecture Notes in Computer Science, 1141:811–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakumeit:1996:EAC

[1695] J. Jakumeit. Evolutionary algorithms for the calculation of electron distributions in Si-MOSFETs. Lecture Notes in Computer Science, 1141:819–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappler:1996:RNP

[1696] C. Kappler, T. Baeck, J. Heistermann, and A. Van Der Velde. Refueling of a nuclear power plant: Comparison of a naive and a specialized mutation operator. Lecture Notes in Computer Science, 1141:829–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lienig:1996:GAA

[1697] J. Lienig and J. P. Cohoon. Genetic algorithms applied to the physical design of VLSI circuits: a survey. Lecture Notes in Computer Science, 1141:839–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rogenmoser:1996:SMT

[1698] R. Rogenmoser, H. Kaeslin, and T. Blickle. Stochastic methods for transistor size optimization of CMOS VLSI circuits. Lecture Notes in Computer Science, 1141:849–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnecke:1996:APG

[1699] V. Schnecke and O. Vornberger. An adaptive parallel genetic algorithm for VLSI-Layout optimization. Lecture Notes in Computer Science, 1141:859– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alba:1996:GAP

[1700] E. Alba and J. M. Troya. Genetic algorithms for protocol validation. Lecture Notes in Computer Science, 1141:870–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilchev:1996:CHF

[1701] G. Bilchev and I. Parmee. Constraint handling for the fault coverage code generation problem: An inductive evolutionary approach. Lecture Notes in Computer Science, 1141:880–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freisleben:1996:NGL

[1702] B. Freisleben and P. Merz. New genetic local search operators for the traveling salesman problem. Lecture Notes in Computer Science, 1141:890–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1996:EAH

[1703] X. Hu, G. W. Greenwood, and J. G.

D’Ambrosio. An evolutionary approach to hardware/software partitioning. Lecture Notes in Computer Science, 1141:900–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanKemenade:1996:EAT

[1704] K. C. H. M. Van Kemenade, J. M. Van Den Akker, and J. N. Kok. Evolutionary air traffic flow management for large 3D-problems. Lecture Notes in Computer Science, 1141:910–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munetomo:1996:GBD

[1705] M. Munetomo, Y. Takai, and Y. Sato.

Genetic-based dynamic load balancing: Implementation and evaluation. Lecture Notes in Computer Science, 1141: 920–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niemeyer:1996:PSG

[1706] G. Niemeyer and P. Shiroma. Production scheduling with genetic algorithms and simulation. Lecture Notes in Computer Science, 1141:930–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schweitzer:1996:NOU

[1707] F. Schweitzer, W. Ebeling, H. Rose, and O. Weiss. Network optimization using evolutionary strategies. Lecture Notes in Computer Science, 1141:940– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sipper:1996:CEP

[1708] M. Sipper and M. Tomassini. Coevolving parallel random number generators. Lecture Notes in Computer Science, 1141:950–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamada:1996:SGL

[1709] T. Yamada and R. Nakano. Scheduling by genetic local search with multi-step crossover. Lecture Notes in Computer Science, 1141:960–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beiersdoerfer:1996:FCO

[1710] S. Beiersdoerfer, J. Schmitt, M. Sauer, and A. Schulz. Finding the conformation of organic molecules with genetic algorithms. Lecture Notes in Computer Science, 1141:972–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFalco:1996:IPB

[1711] I. De Falco, A. Della Cioppa, R. Del Balio, and E. Tarantino. Investigating a parallel breeder genetic algorithm on the inverse aerodynamic design. Lecture Notes in Computer Science, 1141:982–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujimoto:1996:EDL

[1712] Y. Fujimoto, M. Nishiguchi, K. Nomoto, and K. Takahashi. An evolutionary design for f-lenses. Lecture Notes in Computer Science, 1141:992–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gross:1996:OHE

[1713] B. Gross, U. Hammel, P. Maldaner, and A. Meyer. Optimization of heat exchanger networks by means of evolution strategies. Lecture Notes in Computer Science, 1141:1002–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:IPP

[1714] Dae Gyu Kim, D. Corne, and P. Ross.

Industrial plant pipe-route optimisation with genetic algorithms. Lecture Notes in Computer Science, 1141:1012– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorges-Schleuter:1996:EAD

[1715] M. Gorges-Schleuter, W. Jakob, S. Meinzer, and A. Quinte. An evolutionary algorithm for design optimization of microsystems. Lecture Notes in Computer Science, 1141:1022–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richards:1996:LCS

[1716] R. A. Richards and S. D. Sheppard. A learning classifier system for three-dimensional shape optimization. Lecture Notes in Computer Science, 1141:1032–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weinhardt:1996:PPS

[1717] M. Weinhardt. Portable pipeline synthesis for FCCMs. Lecture Notes in Computer Science, 1142:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Legl:1996:PDT

[1718] C. Legl, K. Eckl, and B. Wurth. Performance-directed technology mapping for LUT-Based FPGAs — what role do decomposition and covering play? Lecture Notes in Computer Science, 1142:14–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luk:1996:FDP

[1719] W. Luk, S. Guo, N. Shirazi, and N. Zhuang. A framework for developing parametrised FPGA libraries. Lecture Notes in Computer Science, 1142: 24–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyazaki:1996:FCE

[1720] T. Miyazaki, A. Tsutsui, K. Ishii, and N. Ohta. FACT: Co-evaluation environment for FPGA architecture and CAD system. Lecture Notes in Computer Science, 1142:34–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stohmann:1996:UCA

[1721] J. Stohmann and E. Barke. An universal CLA adder generator for SRAM-Based FPGAs. Lecture Notes in Computer Science, 1142:44–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shibata:1996:ESW

[1722] Y. Shibata, X.-P. Ling, and H. Amano.

An emulation system of the WASMII: a data driven computer on a virtual hardware. Lecture Notes in Computer Science, 1142:55–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartenstein:1996:CCM

[1723] R. W. Hartenstein, J. Becker, and R. Kress. Custom computing machines vs. hardware/software codesign: From a globalized point of view. Lecture Notes in Computer Science, 1142:65– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludwig:1996:DCB

[1724] S. H.-M. Ludwig. The design of a coprocessor board using Xilinx’s XC6200 FPGA — an experience report. Lecture Notes in Computer Science, 1142: 77–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:RRA

[1725] D. Smith and D. Bhatia. RACE:

Reconfigurable and adaptive computing environment. Lecture Notes in Computer Science, 1142:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dick:1996:CDD

[1726] C. H. Dick. Computing 2-D DFTs using FPGAs. Lecture Notes in Computer Science, 1142:96–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ober:1996:CCA

[1727] U. Ober, H.-J. Herpel, and M. Glesner. CAPpartx: Computer aided prototyping partitioning for Xilinx FPGAs, a hierarchical partitioning tool for rapid prototyping. Lecture Notes in Computer Science, 1142:106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trainor:1996:ASE

[1728] D. W. Trainor and R. F. Woods. Architectural synthesis and efficient circuit implementation for Field Programmable Gate Arrays. Lecture Notes in Computer Science, 1142:116–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebeling:1996:RRP

[1729] C. Ebeling, D. C. Cronquist, and P. Franklin. RaPiD — reconfigurable pipelined datapath. Lecture Notes in Computer Science, 1142:126–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suyama:1996:SSP

[1730] T. Suyama, M. Yokoo, and H. Sawada. Solving satisfiability problems on FP-GAs. Lecture Notes in Computer Science, 1142:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanz:1996:FIB

[1731] C. Sanz, L. De Zulueta, and J. M. Meneses. FPGA implementation of the block-matching algorithm for motion estimation in image coding. Lecture Notes in Computer Science, 1142:146– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braun:1996:PCC

[1732] M. Braun, J. Friedrich, T. Gruen, and J. Lembert. Parallel CRC computation in FPGAs. Lecture Notes in Computer Science, 1142:156–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer-Baese:1996:CDF

[1733] U. Meyer-Baese. Coherent demodulation with FPGAs. Lecture Notes in Computer Science, 1142:166–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehring:1996:TSA

[1734] S. Gehring and S. Ludwig. The Trianus system and its application to custom computing. Lecture Notes in Computer Science, 1142:176–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lester:1996:LSF

[1735] N. Lester and J. Saul. Logic synthesis for FPGAs using A mixed exclusive-/ inclusive-OR form. Lecture Notes in Computer Science, 1142:185–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tammemaee:1996:FCT

[1736] K. Tammemaee, M. O’Nils, and A. Hemani. Flexible codesign target architecture for early prototyping of CMIST systems. Lecture Notes in Computer Science, 1142:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inoue:1996:ARM

[1737] K. Inoue, T. Kisuki, M. Okuno, and E. Shimizu. ATTEMPT-1: a reconfigurable multiprocessor testbed. Lecture Notes in Computer Science, 1142: 200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janzen:1996:SME

[1738] N. Janzen and F. J. Rammig. A slow motion engine for the analysis of FPGA-Based prototypes. Lecture Notes in Computer Science, 1142:210– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hesener:1996:IRD

[1739] A. Hesener. Implementing reconfigurable datapaths in FPGAs for adaptive filter design. Lecture Notes in Computer Science, 1142:220–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kean:1996:FCC

[1740] T. Kean, B. New, and B. Slous. A fast constant coefficient multiplier for the XC6200. Lecture Notes in Computer Science, 1142:230–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ditzinger:1996:KIU

[1741] A. Ditzinger and R. Remme. Key issues for user acceptance of FPGA design tools. Lecture Notes in Computer Science, 1142:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Combridge:1996:RDD

[1742] B. L. Combridge, P. S. Cornfield, and S. Naunton. Reconfigurable DSP demonstrators for the development of spacecraft payload processors. Lecture Notes in Computer Science, 1142:242– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casselman:1996:RLB

[1743] S. Casselman. Reconfigurable logic based fibre channel network card with sub 2 micro-second raw latency. Lecture Notes in Computer Science, 1142: 252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddy:1996:ATM

[1744] J. R. Haddy and D. J. Skellern. An asynchronous transfer mode (ATM) stream demultiplexer and switch. Lecture Notes in Computer Science, 1142: 260–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasilko:1996:ORF

[1745] M. Vasilko and D. Ait-Boudaoud. Optically reconfigurable FPGAs: Is this a future trend? Lecture Notes in Computer Science, 1142:270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salcic:1996:CIL

[1746] Z. Salcic and B. Maunder. CC-SimP — an instruction-level customconfigurable processor for FPLDs. Lecture Notes in Computer Science, 1142: 280–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasilko:1996:AST

[1747] M. Vasilko and D. Ait-Boudaoud. Architectural synthesis techniques for dynamically reconfigurable logic. Lecture Notes in Computer Science, 1142:290– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eggers:1996:FRC

[1748] H. Eggers, P. Lysaght, H. Dick, and G. McGregor. Fast reconfigurable crossbar switching in FPGAs. Lecture Notes in Computer Science, 1142:297– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yasar:1996:GFM

[1749] G. Yasar, J. Devins, Y. Tsyrkina, and G. Stadtlander. Growable FPGA macro generator. Lecture Notes in Computer Science, 1142:307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heron:1996:ASI

[1750] J. P. Heron and R. F. Woods. Architectural strategies for implementing an image processing algorithm on XC6000 FPGA. Lecture Notes in Computer Science, 1142:317–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brebner:1996:VHO

[1751] G. Brebner. A virtual hardware operating system for the Xilinx XC6200. Lecture Notes in Computer Science, 1142: 327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trost:1996:EPE

[1752] A. Trost, R. Kuznar, A. Zemva, and B. Zajc. An experimental programmable environment for prototyping digital circuits. Lecture Notes in Computer Science, 1142:337–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gschwind:1996:MSB

[1753] M. Gschwind and C. Mautner. Migration from schematic-based designs to a VHDL synthesis environment. Lecture Notes in Computer Science, 1142: 346–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balboni:1996:ADF

[1754] A. Balboni and L. Valenti. ASIC design and FPGA design: a unified design methodology applied to different technologies. Lecture Notes in Computer Science, 1142:356–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dick:1996:FFF

[1755] C. H. Dick and F. Harris. FIR filtering with FPGAs using quadrature sigma-delta modulation encoding. Lecture Notes in Computer Science, 1142:361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1996:NFT

[1756] K. Yi and C.-S. Jhon. A new FPGA technology mapping approach by cluster merging. Lecture Notes in Computer Science, 1142:366–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsson:1996:EBT

[1757] L. Larsson. An EPLD based transient recorder for simulation of video signal processing devices in a VHDL environment close to system level conditions. Lecture Notes in Computer Science, 1142:371–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer-Baese:1996:CED

[1758] U. Meyer-Baese. Convolutional error decoding with FPGAs. Lecture Notes in Computer Science, 1142:376– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rogina:1996:MCT

[1759] B. M. Rogina, K. Skala, and B. Vojnovic. Metastability characteristics testing for programmable logic design. Lecture Notes in Computer Science, 1142:381–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rowley:1996:ISD

[1760] K. Rowley and C. Lyden. Implementing sigma delta modulator prototype designs on an FPGA. Lecture Notes in Computer Science, 1142:389–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz:1996:DVP

[1761] J. L. Ruiz, Y. Torroja, and J. L. Garcia. Design of a VME parametrized library for FPGAs. Lecture Notes in Computer Science, 1142:394–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schumacher:1996:DTA

[1762] G. Schumacher, B. Josko, G. Wagner, and M. Radetzki. Development of a telephone answering machine in a lab — FPGAs in education. Lecture Notes in Computer Science, 1142:400–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tchoumatchenko:1996:FDM

[1763] V. Tchoumatchenko, T. Vassileva, R. Ribas, and A. Guyot. FPGA design migration. Lecture Notes in Computer Science, 1142:405–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pillement:1996:CDH

[1764] S. Pillement, L. Torres, M. Robert, and G. Cambon. Concurrent design of hardware/software dedicated systems. Lecture Notes in Computer Science, 1142:410–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Touhafi:1996:IFP

[1765] A. Touhafi, W. Brissinck, and E. F.

Dirkx. The implementation of a field programmable logic based co-processor for the acceleration of discrete event simulators. Lecture Notes in Computer Science, 1142:415–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weinhardt:1996:CWD

[1766] M. Weinhardt. Computing weight distributions of binary linear block codes on a CCM. Lecture Notes in Computer Science, 1142:425–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carse:1996:FEL

[1767] B. Carse and T. C. Fogarty. Fast evolutionary learning of minimal radial basis function neural networks using a genetic algorithm. Lecture Notes in Computer Science, 1143:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cartwright:1996:EDS

[1768] H. M. Cartwright and J. A. Hopkins. Evolutionary design of synthetic routes in chemistry. Lecture Notes in Computer Science, 1143:23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fang:1996:GAJ

[1769] H.-L. Fang, D. Come, and P. Ross. A genetic algorithm for job-shop problems with various schedule quality criteria. Lecture Notes in Computer Science, 1143:39–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Husbands:1996:TAG

[1770] P. Husbands, G. Jenny, M. McIlhagga, and R. Ives. Two applications of genetic algorithms to component design. Lecture Notes in Computer Science, 1143:50–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1996:CSB

[1771] P. Jonsson and J. Barklund. Characterizing signal behaviour using genetic programming. Lecture Notes in Computer Science, 1143:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lunn:1996:SRG

[1772] K. Lunn and C. Johnson. Spatial reasoning with genetic algorithms — an application in planning of safe liquid petroleum gas sites. Lecture Notes in Computer Science, 1143:73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1996:REG

[1773] J. F. Miller and P. Thompson.

Restricted evaluation genetic algorithms with tabu search for optimising Boolean functions as multi-level AND-EXOR networks. Lecture Notes in Computer Science, 1143:85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pohlheim:1996:GSP

[1774] H. Pohlheim and P. Marenbach. Generation of structured process models using genetic programming. Lecture Notes in Computer Science, 1143:102– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poli:1996:GPF

[1775] R. Poli. Genetic programming for feature detection and image segmentation. Lecture Notes in Computer Science, 1143:110–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandqvist:1996:TVS

[1776] S. Sandqvist. A temporal view of selection and populations. Lecture Notes in Computer Science, 1143:126–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:EST

[1777] J. Smith and T. C. Fogarty. Evolving software test data — GA’s learn self expression. Lecture Notes in Computer Science, 1143:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sullivan:1996:EES

[1778] J. C. W. Sullivan and A. G. Pipe. Efficient evolution strategies for exploration in mobile robotics. Lecture Notes in Computer Science, 1143:147– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilchev:1996:LND

[1779] G. Bilchev and I. C. Parmee. Learning the “next” dimension. Lecture Notes in Computer Science, 1143:162–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branke:1996:GSM

[1780] J. Branke, H. C. Andersen, and H. Schmeck. Global selection methods for massively parallel computers. Lecture Notes in Computer Science, 1143:175–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corne:1996:IMN

[1781] D. Corne, E. Collingwood, and P. Ross.

Investigating Multipoidy’s niche. Lecture Notes in Computer Science, 1143: 189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GonzalezHernandez:1996:EDC

[1782] L. F. Gonzalez Hernandez and D. W.

Corne. Evolutionary divide and conquer for the set-covering problem. Lecture Notes in Computer Science, 1143: 198–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1996:SLI

[1783] R. Hoffmann and N. Waring. The simulation of localised interaction and learning in artificial adaptive agents. Lecture Notes in Computer Science, 1143: 209–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quick:1996:RRF

[1784] R. J. Quick, V. J. Rayward-Smith, and G. D. Smith. The Royal Road functions: Description, intent and experimentation. Lecture Notes in Computer Science, 1143:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roy:1996:ART

[1785] R. Roy and I. C. Parmee. Adaptive restricted tournament selection for the identification of multiple sub-optima in a multi-modal function. Lecture Notes in Computer Science, 1143:236– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sendhoff:1996:APG

[1786] B. Sendhoff and M. Kreutz. Analysis of possible genome-dependence of mutation rates in genetic algorithms. Lecture Notes in Computer Science, 1143: 257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Surry:1996:IIE

[1787] P. D. Surry and N. J. Radcliffe. Inoculation to initialise evolutionary search. Lecture Notes in Computer Science, 1143:269–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuson:1996:CEO

[1788] A. Tuson and P. Ross. Co-evolution of operator settings in genetic algorithms. Lecture Notes in Computer Science, 1143:286–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vavak:1996:CSS

[1789] F. Vavak and T. C. Fogarty. A comparative study of steady state and generational genetic algorithms for use in nonstationary environments. Lecture Notes in Computer Science, 1143:297– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ponce:1996:RIW

[1790] J. Ponce, M. Hebert, and A. Zisserman. Report on the 1996 International Workshop on Object Representation in Computer Vision. Lecture Notes in Computer Science, 1144:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zerroug:1996:IID

[1791] M. Zerroug and R. Nevatia. From an intensity image to 3-D segmented descriptions. Lecture Notes in Computer Science, 1144:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sayd:1996:RGC

[1792] P. Sayd, M. Dhome, and J.-M Lavest.

Recovering generalized cylinders by monocular vision. Lecture Notes in Computer Science, 1144:25–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1996:CGS

[1793] S. Carlsson. Combinatorial geometry for shape representation and indexing. Lecture Notes in Computer Science, 1144:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rothwell:1996:ROU

[1794] C. Rothwell, J. Mundy, and B. Hoffman. Representing objects using topology. Lecture Notes in Computer Science, 1144:79–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angelopoulou:1996:CBS

[1795] E. Angelopoulou, J. P. Williams, and L. B. Wolff. Curvature based signatures for object description and recognition. Lecture Notes in Computer Science, 1144:109–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shum:1996:SS

[1796] H.-Y. Shum, M. Hebert, and K. Ikeuchi. On 3D shape synthesis. Lecture Notes in Computer Science, 1144:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koenderink:1996:SCP

[1797] J. J. Koenderink, A. J. Van Doorn, C. Christou, and J. S. Lappin. Shape constancy in pictorial relief. Lecture Notes in Computer Science, 1144:151– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nayar:1996:DIM

[1798] S. K. Nayar and H. Murase. Dimensionality of illumination manifolds in appearance matching. Lecture Notes in Computer Science, 1144:165–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Epstein:1996:LOR

[1799] R. Epstein, A. L. Yuille, and P. N. Belhumeur. Learning object representations from lighting variations. Lecture Notes in Computer Science, 1144:179– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pope:1996:LAM

[1800] A. R. Pope and D. G. Lowe. Learning appearance models for object recognition. Lecture Notes in Computer Science, 1144:201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmid:1996:IOC

[1801] C. Schmid, P. Bobet, B. Lamiroy, and R. Mohr. An image oriented CAD approach. Lecture Notes in Computer Science, 1144:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundy:1996:ECA

[1802] J. Mundy, A. Liu, N. Pillow, and A. Zisserman. An experimental comparison of appearance and geometric model based recognition. Lecture Notes in Computer Science, 1144:247– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanade:1996:VRB

[1803] T. Kanade, P. J. Narayanan, and P. W.

Rander. Virtualized reality: being mobile in a visual scene. Lecture Notes in Computer Science, 1144:273–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Francois:1996:GSL

[1804] A. R. J. François and G. Medioni.

Generic shape learning and recognition. Lecture Notes in Computer Science, 1144:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayoung-Chee:1996:HAR

[1805] N. Ayoung-Chee, G. Dudek, and F. P.

Ferrie. A hybrid approach to 3D representation. Lecture Notes in Computer Science, 1144:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forsyth:1996:FPO

[1806] D. A. Forsyth, J. Malik, M. M. Fleck, and H. Greenspan. Finding pictures of objects in large collections of images. Lecture Notes in Computer Science, 1144:335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wright:1996:BHT

[1807] M. Wright, A. Fitzgibbon, P. J. Giblin, and R. B. Fisher. Beyond the Hough transform: further properties of the R mapping and their applications. Lecture Notes in Computer Science, 1144:361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kutulakos:1996:NEO

[1808] K. N. Kutulakos and J. R. Vallino. Non-Euclidean object representations for calibration-free video overlay. Lecture Notes in Computer Science, 1144: 381–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiken:1996:CBP

[1809] A. Aiken. Constraint-based program analysis. Lecture Notes in Computer Science, 1145:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1996:SDP

[1810] F. Nielson. Semantics-directed program analysis: a tool-maker’s perspective. Lecture Notes in Computer Science, 1145:2–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steffen:1996:POE

[1811] B. Steffen. Property-oriented expansion. Lecture Notes in Computer Science, 1145:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alstrup:1996:GDS

[1812] S. Alstrup, P. W. Lauridsen, and M. Thorup. Generalized dominators for structured programs. Lecture Notes in Computer Science, 1145:42–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:1996:CBP

[1813] M. Alt, C. Ferdinand, F. Martin, and R. Wilhelm. Cache behavior prediction by abstract interpretation. Lecture Notes in Computer Science, 1145: 52–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1996:TAO

[1814] P. H. Andersen and C. K. Holst. Termination analysis for offline partial evaluation of a higher order functional language. Lecture Notes in Computer Science, 1145:67–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBoer:1996:PCC

[1815] F. S. De Boer, M. Gabbrielli, and C. Palamidessi. Proving correctness of constraint logic programs with dynamic scheduling. Lecture Notes in Computer Science, 1145:83–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgia:1996:UMA

[1816] R. Borgia, P. Degano, C. Priami, and L. Leth. Understanding mobile agents via a non-interleaving semantics for Facile. Lecture Notes in Computer Science, 1145:98–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauburger:1996:TAP

[1817] J. Brauburger and J. Giesl. Termination analysis for partial functions. Lecture Notes in Computer Science, 1145: 113–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruynooghe:1996:FSA

[1818] M. Bruynooghe, B. Demoen, D. Boulanger, and M. Denecker. A freeness and sharing analysis of logic programs based on a pre-interpretation. Lecture Notes in Computer Science, 1145:128–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coppo:1996:RTP

[1819] M. Coppo, F. Damiani, and P. Giannini. Refinement types for program analysis. Lecture Notes in Computer Science, 1145:143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crnogorac:1996:CTO

[1820] L. Crnogorac, A. D. Kelly, and H. Søndergaard. A comparison of three occur-check analysers. Lecture Notes in Computer Science, 1145:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emelianov:1996:AER

[1821] P. G. Emelianov. Analysis of the equality relations for the program terms. Lecture Notes in Computer Science, 1145:174–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fecht:1996:EFS

[1822] C. Fecht and H. Seidl. An even faster solver for general systems of equations. Lecture Notes in Computer Science, 1145:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Givan:1996:IPS

[1823] R. Givan. Inferring program specifications in polynomial-time. Lecture Notes in Computer Science, 1145:205– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1996:AMT

[1824] M. Mueller, T. Glass, and K. Stroetmann. Automated modular termination proofs for real Prolog programs. Lecture Notes in Computer Science, 1145:220–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pande:1996:DFB

[1825] H. D. Pande and B. G. Ryder. Dataflow-based virtual function resolution. Lecture Notes in Computer Science, 1145:238–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterson:1996:CLU

[1826] R. Paterson. Compiling laziness using projections. Lecture Notes in Computer Science, 1145:255–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puebla:1996:OAI

[1827] G. Puebla and M. Hermenegildo. Optimized algorithms for incremental analysis of logic programs. Lecture Notes in Computer Science, 1145:270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rehof:1996:TCF

[1828] J. Rehof and T. A. Mogensen. Tractable constraints in finite semilattices. Lecture Notes in Computer Science, 1145:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roux:1996:UDH

[1829] O. Roux and V. Rusu. Uniformity for the decidability of hybrid automata. Lecture Notes in Computer Science, 1145:301–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoenig:1996:BSA

[1830] S. Schoenig and M. Ducasse. A backward slicing algorithm for Prolog. Lecture Notes in Computer Science, 1145:317–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snelting:1996:CSC

[1831] G. Snelting. Combining slicing and constraint solving for validation of measurement software. Lecture Notes in Computer Science, 1145:332–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trifonov:1996:SCT

[1832] V. Trifonov and S. Smith. Subtyping constrained types. Lecture Notes in Computer Science, 1145:349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Venet:1996:ACD

[1833] A. Venet. Abstract cofibered domains:

Application to the alias analysis of untyped programs. Lecture Notes in Computer Science, 1145:366–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Handjieva:1996:SSA

[1834] M. Handjieva. STAN: a static analyzer for CLP(R) based on abstract interpretation. Lecture Notes in Computer Science, 1145:383–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelly:1996:TAI

[1835] A. D. Kelly, K. Marriott, H. Søndergaard, and P. J. Stuckey. Two applications of an incremental analysis engine for (constraint) logic programs. Lecture Notes in Computer Science, 1145:385– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1996:PPA

[1836] M. Mueller, T. Glass, and K. Stroetmann. PAN — the Prolog analyzer. Lecture Notes in Computer Science, 1145:387–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waidner:1996:DSE

[1837] M. Waidner. Development of a secure electronic marketplace for Europe. Lecture Notes in Computer Science, 1146:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mao:1996:LWM

[1838] W. Mao. Light-weight micro-cash payment for the Internet. Lecture Notes in Computer Science, 1146:15–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camenisch:1996:DPS

[1839] J. Camenisch, U. Maurer, and M. Stadler. Digital payment systems with passive anonymity-revoking trustees. Lecture Notes in Computer Science, 1146:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atluri:1996:AMW

[1840] V. Atluri and W.-K Huang. An authorization model for workflows. Lecture Notes in Computer Science, 1146: 44–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandhu:1996:RHC

[1841] R. Sandhu. Role hierarchies and constraints for lattice-based access controls. Lecture Notes in Computer Science, 1146:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicomette:1996:MSM

[1842] V. Nicomette and Y. Deswarte. A multilevel security model for distributed object systems. Lecture Notes in Computer Science, 1146:80–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCapitanidiVimercati:1996:AMF

[1843] S. De Capitani di Vimercati and P. Samarati. An authorization model for federated systems. Lecture Notes in Computer Science, 1146:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farmer:1996:SMA

[1844] W. Farmer, J. Guttman, and V. Swarup.

Security for mobile agents: Authentication and state appraisal. Lecture Notes in Computer Science, 1146:118– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asokan:1996:SSS

[1845] N. Asokan, G. Tsudik, and M. Waidner. Server-supported signatures. Lecture Notes in Computer Science, 1146: 131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1996:SNL

[1846] S. F. Wu. Sleepy network-layer authentication service for IPSEC. Lecture Notes in Computer Science, 1146: 146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1996:CEM

[1847] J. Zhou and D. Gollmann. Certified electronic mail. Lecture Notes in Computer Science, 1146:160–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneier:1996:DP

[1848] B. Schneier, J. Kelsey, and J. Walker. Distributed proctoring. Lecture Notes in Computer Science, 1146:172–182, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://www.counterpane.com/
                                        distributed_proctoring.html.

Bonatti:1996:MHS

[1849] P. Bonatti, M. L. Sapino, and V. S. Subrahmanian. Merging heterogeneous security orderings. Lecture Notes in Computer Science, 1146:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1996:CA

[1850] S. Schneider and A. Sidiropoulos. CSP and anonymity. Lecture Notes in Computer Science, 1146:198–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wedel:1996:FSA

[1851] G. Wedel and V. Kessler. Formal semantics for authentication logics. Lecture Notes in Computer Science, 1146: 219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lotz:1996:TSM

[1852] V. Lotz. Threat scenarios as a means to formally develop secure systems. Lecture Notes in Computer Science, 1146: 242–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Warner:1996:IMS

[1853] A. Warner, Q. Li, T. Keefe, and S. Pal. The impact of multilevel security on database buffer management. Lecture Notes in Computer Science, 1146:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marks:1996:ECD

[1854] D. G. Marks, A. Motro, and S. Jajodia.

Enhancing the controlled disclosure of sensitive information. Lecture Notes in Computer Science, 1146:290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mancini:1996:SCC

[1855] L. V. Mancini and I. Ray. Secure concurrency control in MLS databases with two versions of data. Lecture Notes in Computer Science, 1146:304– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1996:MPK

[1856] U. Maurer. Modelling a publickey infrastructure. Lecture Notes in Computer Science, 1146:325–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meadows:1996:ANS

[1857] C. A. Meadows. Analyzing the Needham-Schroeder public-key protocol: A comparison of two approaches. Lecture Notes in Computer Science, 1146:351–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magerman:1996:LGS

[1858] D. M. Magerman. Learning grammatical structure using statistical decisiontrees. Lecture Notes in Computer Science, 1147:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knuutila:1996:IIP

[1859] T. Knuutila. Inductive inference from positive data: from heuristic to characterizing methods. Lecture Notes in Computer Science, 1147:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apsitis:1996:UIF

[1860] K. Apsitis, R. Freivalds, R. Simanovskis, and J. Smotrovs. Unions of identifiable families of languages. Lecture Notes in Computer Science, 1147:48–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelaHiguera:1996:CSP

[1861] C. De la Higuera. Characteristic sets for polynomial grammatical inference. Lecture Notes in Computer Science, 1147:59–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vilar:1996:QLS

[1862] J. M. Vilar. Query learning of subsequential transducers. Lecture Notes in Computer Science, 1147:72–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brent:1996:LCF

[1863] M. R. Brent and T. A. Cartwright.

Lexical categorization: fitting template grammars by incremental MDL optimization. Lecture Notes in Computer Science, 1147:84–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tillmann:1996:SCW

[1864] C. Tillmann and H. Ney. Selection criteria for word trigger pairs in language modelling. Lecture Notes in Computer Science, 1147:95–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fred:1996:CSU

[1865] A. L. N. Fred. Clustering of sequences using minimum grammar complexity criterion. Lecture Notes in Computer Science, 1147:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takada:1996:NGI

[1866] Y. Takada and T. Y. Nishida. A note on grammatical inference of slender context-free languages. Lecture Notes in Computer Science, 1147:117– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sempere:1996:LLG

[1867] J. M. Sempere and A. Fos. Learning linear grammars from structural information. Lecture Notes in Computer Science, 1147:126–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alquezar:1996:LCS

[1868] R. Alquezar, A. Sanfeliu, and J. Cueva. Learning of context-sensitive language acceptors through regular inference and constrained induction. Lecture Notes in Computer Science, 1147:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samuelsson:1996:ICG

[1869] C. Samuelsson, P. Tapanainen, and A. Voutilainen. Inducing constraint grammars. Lecture Notes in Computer Science, 1147:146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deligne:1996:ISD

[1870] S. Deligne, F. Yvon, and F. Bimbot. Introducing statistical dependencies and structural constraints in variable-length sequence models. Lecture Notes in Computer Science, 1147: 156–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregor:1996:DCS

[1871] J. Gregor and M. G. Thomason. A disagreement count scheme for inference of constrained Markov networks. Lecture Notes in Computer Science, 1147: 168–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vidal:1996:UKI

[1872] E. Vidal and D. Llorens. Using knowledge to improve N-gram language modelling through the MGGI methodology. Lecture Notes in Computer Science, 1147:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blasig:1996:DSP

[1873] R. Blasig. Discrete sequence prediction with commented Markov models. Lecture Notes in Computer Science, 1147:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz:1996:LPT

[1874] J. Ruiz and P. Garcia. Learning kpiecewise testable languages from positive data. Lecture Notes in Computer Science, 1147:203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerald:1996:LCR

[1875] J. D. Emerald, K. G. Subramanian, and D. G. Thomas. Learning code regular and code linear languages. Lecture Notes in Computer Science, 1147: 211–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dupont:1996:IRI

[1876] P. Dupont. Incremental regular inference. Lecture Notes in Computer Science, 1147:222–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parekh:1996:IIA

[1877] R. Parekh and V. Honavar. An incremental interactive algorithm for grammar inference. Lecture Notes in Computer Science, 1147:238–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lorenzo:1996:ILP

[1878] D. Lorenzo. Inductive logic programming for discrete event systems. Lecture Notes in Computer Science, 1147: 250–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golea:1996:SSR

[1879] M. Golea, M. Matsuoka, and Y. Sakakibara. Stochastic simple recurrent neural networks. Lecture Notes in Computer Science, 1147:262–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrasco:1996:ISR

[1880] R. C. Carrasco, M. L. Forcada, and L. Santamaria. Inferring stochastic regular grammars with recurrent neural networks. Lecture Notes in Computer Science, 1147:274–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casacuberta:1996:MMI

[1881] F. Casacuberta. Maximum mutual information and conditional maximum likelihood estimation of stochastic regular syntax-directed translation schemes. Lecture Notes in Computer Science, 1147:282–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordano:1996:GIU

[1882] J. Y. Giordano. Grammatical inference using tabu search. Lecture Notes in Computer Science, 1147:292–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oncina:1996:UDI

[1883] J. Oncina and M. A. Varo. Using domain information during the learning of a subsequential transducer. Lecture Notes in Computer Science, 1147:301– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelaHiguera:1996:IDD

[1884] C. De la Higuera, J. Oncina, and E. Vidal. Identification of DFA: datadependent vs. data-independent algorithms. Lecture Notes in Computer Science, 1147:313–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1996:HSS

[1885] C. M. Hoffmann. How solid is solid modeling? Lecture Notes in Computer Science, 1148:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortune:1996:RIG

[1886] S. Fortune. Robustness issues in geometric algorithms. Lecture Notes in Computer Science, 1148:9–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guibas:1996:IGA

[1887] L. J. Guibas. Implementing geometric algorithms robustly. Lecture Notes in Computer Science, 1148:15–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preparata:1996:RGA

[1888] F. P. Preparata. Robustness in geometric algorithms. Lecture Notes in Computer Science, 1148:23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pratt:1996:ACG

[1889] M. J. Pratt. Applications of computational geometry in mechanical engineering design and manufacture. Lecture Notes in Computer Science, 1148: 25–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1996:SAC

[1890] J. S. B. Mitchell. On some applications of computational geometry in manufacturing and virtual environments. Lecture Notes in Computer Science, 1148:37–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobkin:1996:VGA

[1891] D. Dobkin. Visualizing geometric algorithms — state of the art. Lecture Notes in Computer Science, 1148:41– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:GAV

[1892] D. T. Lee. Geometric algorithm visualization, current status and future. Lecture Notes in Computer Science, 1148: 45–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehlhorn:1996:PPP

[1893] K. Mehlhorn. Position paper for panel discussion. Lecture Notes in Computer Science, 1148:51–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Overmars:1996:DCG

[1894] M. H. Overmars. Designing the computational geometry algorithms library CGAL. Lecture Notes in Computer Science, 1148:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chazelle:1996:CGI

[1895] B. Chazelle. The computational geometry impact task force report: An executive summary. Lecture Notes in Computer Science, 1148:59–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finn:1996:GMF

[1896] P. W. Finn, D. Halperin, L. E. Kavraki, and J.-C Latombe. Geometric manipulation of flexible ligands. Lecture Notes in Computer Science, 1148:67– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prisant:1996:RRF

[1897] M. G. Prisant. Ray-representation formalism for geometric computations on protein solid models. Lecture Notes in Computer Science, 1148:79–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniels:1996:CBS

[1898] K. Daniels and V. J. Milenkovic.

Column-based strip packing using ordered and compliant containment. Lecture Notes in Computer Science, 1148: 91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Majhi:1996:CFU

[1899] J. Majhi, P. Gupta, and R. Janardan. Computing a flattest, undercutfree parting line for a convex polyhedron, with application to mold design. Lecture Notes in Computer Science, 1148:109–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobkin:1996:GPM

[1900] D. Dobkin and D. Gunopulos. Geometric problems in machine learning. Lecture Notes in Computer Science, 1148:121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basri:1996:MCP

[1901] R. Basri and D. Jacobs. Matching convex polygons and polyhedra, allowing for occlusion. Lecture Notes in Computer Science, 1148:133–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hung:1996:SPP

[1902] C.-K. Hung and D. Ierardi. Stably placing piecewise smooth objects. Lecture Notes in Computer Science, 1148:149– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortune:1996:BTA

[1903] S. Fortune. A beam-tracing algorithm for prediction of indoor radio propagation. Lecture Notes in Computer Science, 1148:157–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kernighan:1996:EGI

[1904] B. W. Kernighan and C. J. Van Wyk.

Extracting geometric information from architectural drawings. Lecture Notes in Computer Science, 1148:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orti:1996:UVC

[1905] R. Orti, F. Durand, S. Riviere, and C. Puech. Using the visibility complex for radiosity computation. Lecture Notes in Computer Science, 1148: 177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fabri:1996:CKB

[1906] A. Fabri, G.-J. Giezeman, L. Kettner, and S. Schirra. The CGAL kernel: a basis for geometric computation. Lecture Notes in Computer Science, 1148:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shewchuk:1996:TEQ

[1907] J. R. Shewchuk. Triangle: Engineering a 2D quality mesh generator and Delaunay triangulator. Lecture Notes in Computer Science, 1148:203–222, 1996. CODEN LNCSD9. ISBN 3-540-61785-X (softcover), 3-540-70680-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). URL https:

//link.springer.com/chapter/10.
1007/BFb0014497.

Christie:1996:RMP

[1908] A. M. Christie, A. N. Earl, M. I. Kellner, and W. E. Riddle. A reference model for process technology. Lecture Notes in Computer Science, 1149:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1996:TRA

[1909] M. N. Nguyen and R. Conradi. Towards a rigorous approach for managing process evolution. Lecture Notes in Computer Science, 1149:18–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1996:SRA

[1910] R. Conradi. SPIQ: a revised agenda for software process support. Lecture Notes in Computer Science, 1149:36– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estublier:1996:PEI

[1911] J. Estublier and S. Dami. Process engine interoperability: An experiment. Lecture Notes in Computer Science, 1149:43–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Shaul:1996:LAM

[1912] I. Z. Ben-Shaul and G. T. Heineman. A 3-level atomicity model for decentralized process-centered software engineering environments. Lecture Notes in Computer Science, 1149:61–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avrilionis:1996:ISP

[1913] D. Avrilionis, N. Belkhatir, and P.- Y. Cunin. Improving software process modelling and enactment techniques. Lecture Notes in Computer Science, 1149:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emmerich:1996:DPC

[1914] W. Emmerich and A. Finkelstein. Do process-centred environments deserve process-centred tools? Lecture Notes in Computer Science, 1149:75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenwood:1996:PWP

[1915] R. M. Greenwood and B. Warboys.

Process Web — process support for the World Wide Web. Lecture Notes in Computer Science, 1149:82–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basile:1996:MPF

[1916] C. Basile, S. Calanna, E. Di Nitto, and A. Fuggetta. Mechanisms and policies for federated PSEEs: Basic concepts and open issues. Lecture Notes in Computer Science, 1149:86–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snowdon:1996:AMP

[1917] B. Snowdon. Active models and process support. Lecture Notes in Computer Science, 1149:93–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaba:1996:MPC

[1918] A. B. Kaba and J.-C. Derniame. Modelling processes for change: Basic mechanisms for evolving process fragments. Lecture Notes in Computer Science, 1149:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehman:1996:LSE

[1919] M. M. Lehman. Laws of software evolution revisited. Lecture Notes in Computer Science, 1149:108–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahhal:1996:EEI

[1920] S. Rahhal and N. H. Madhavji. Estimating the effort of implementing ISO 9001 in software organizations. Lecture Notes in Computer Science, 1149:125– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruhn:1996:CMS

[1921] V. Gruhn. Configuration management support for software process models. Lecture Notes in Computer Science, 1149:132–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neumann:1996:HLO

[1922] O. Neumann, S. Sachweh, and W. Schaefer. A high-level objectoriented specification language for configuration management and tool integration. Lecture Notes in Computer Science, 1149:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sherdil:1996:HOI

[1923] K. Sherdil and N. H. Madhavji.

Human-oriented improvement in the software process. Lecture Notes in Computer Science, 1149:145–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLatum:1996:NIF

[1924] F. Van Latum, M. Oivo, B. Hoisl, and G. Ruhe. No improvement without feedback: Experiences from goal-oriented measurement at Schlumberger. Lecture Notes in Computer Science, 1149:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBunje:1996:TMP

[1925] T. De Bunje, G. Engels, L. Groenewegen, and M. Heus. Towards measurable process models. Lecture Notes in Computer Science, 1149:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perry:1996:EWP

[1926] D. E. Perry, A. Porter, L. G. Votta, and M. W. Wade. Evaluating workflow and process automation in wide-area software development. Lecture Notes in Computer Science, 1149:188–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreoli:1996:PEC

[1927] J.-M. Andreoli, J.-L. Meunier, and D. Pagani. Process enactment and coordination. Lecture Notes in Computer Science, 1149:195–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cisse:1996:REP

[1928] A. Cisse and M. S. Pimenta. Requirements engineering process as an emergent process. Lecture Notes in Computer Science, 1149:217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alloui:1996:MAA

[1929] I. Alloui, S. Latrous, and F. Oquendo. A multi-agent approach for modelling, enacting and evolving distributed cooperative software processes. Lecture Notes in Computer Science, 1149:225– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nuseibeh:1996:WAC

[1930] B. Nuseibeh. When agents clash. Lecture Notes in Computer Science, 1149: 237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hesse:1996:TPS

[1931] W. Hesse. Theory and practice of the software process: a field study and its implications for project management. Lecture Notes in Computer Science, 1149:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawalek:1996:ODS

[1932] P. Kawalek and D. G. Wastell. Organizational design for software development: a cybernetic perspective. Lecture Notes in Computer Science, 1149: 257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellmer:1996:EPC

[1933] E. Ellmer. Extending process-centered environments with organizational competence. Lecture Notes in Computer Science, 1149:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turgeon:1996:SVB

[1934] J. Turgeon and N. H. Madhavji. A systematic, view-based approach to eliciting process models. Lecture Notes in Computer Science, 1149:276–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsinger:1996:PMM

[1935] A. A. J. Matsinger. Process modelling and measurements in an industrial environment. Lecture Notes in Computer Science, 1149:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panaroni:1996:RDP

[1936] P. Panaroni. Report on deploying process technology on an industrial space software project. Lecture Notes in Computer Science, 1149:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beims:1996:SPP

[1937] M. A. Beims and L. B. Strader. Software process progress in a dynamic business environment. Lecture Notes in Computer Science, 1149:289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fabre:1996:FFA

[1938] J.-C. Fabre and T. Perennou. Friends — A flexible architecture for implementing fault tolerant and secure distributed applications. Lecture Notes in Computer Science, 1150:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bohne:1996:AFT

[1939] J. Bohne and R. Groenberg. Adaptable fault tolerance for distributed process control using exclusively standard components. Lecture Notes in Computer Science, 1150:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Powell:1996:SSH

[1940] D. Powell, M. Cukier, and J. Arlat. On stratified sampling for high coverage estimations. Lecture Notes in Computer Science, 1150:37–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furtado:1996:FIE

[1941] P. Furtado and H. Madeira. Fault injection evaluation of assigned signatures in a RISC processor. Lecture Notes in Computer Science, 1150:55–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1996:EED

[1942] E. Fuchs. An evaluation of the error detection mechanisms in MARS using software-implemented fault injection. Lecture Notes in Computer Science, 1150:73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nelli:1996:DMA

[1943] M. Nelli, A. Bondavalli, and L. Simoncini. Dependability modeling and analysis of complex control systems: An application to railway interlocking. Lecture Notes in Computer Science, 1150:93–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Popov:1996:EIT

[1944] P. Popov. The effect of interfailure time variability on the software reliability growth modelling. Lecture Notes in Computer Science, 1150:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Draber:1996:DEC

[1945] S. Draber and B. Eschermann. Dependability evaluation of a computing system for traction control of electrical locomotives. Lecture Notes in Computer Science, 1150:129–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santonja:1996:DMR

[1946] V. Santonja, M. Alonso, J. Molero, and J. J. Serrano. Dependability models of RAID using stochastic activity networks. Lecture Notes in Computer Science, 1150:141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wildner:1996:CAS

[1947] U. Wildner. Compiler assisted selfchecking of structural integrity using return address hashing. Lecture Notes in Computer Science, 1150:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avresky:1996:SSF

[1948] D. R. Avresky and C. M. Cunningham. Single source fault-tolerant broadcasting for two-dimensional meshes without virtual channels. Lecture Notes in Computer Science, 1150:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corno:1996:LTS

[1949] F. Corno, P. Prinetto, M. Rebaudengo, and M. Sonza Reorda. Online testing of an off-the-shelf microprocessor board for safety-critical applications. Lecture Notes in Computer Science, 1150:190–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Renovell:1996:LTB

[1950] M. Renovell, P. Huc, and Y. Bertrand. The logic threshold based voting: a model for local feedback bridging fault. Lecture Notes in Computer Science, 1150:205–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikolos:1996:YVP

[1951] D. Nikolos and H. T. Vergos. On the yield of VLSI processors with onchip CPU cache. Lecture Notes in Computer Science, 1150:214–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krasniewski:1996:DDH

[1952] A. Krasniewski. Design of dependable hardware: What BIST is most efficient? Lecture Notes in Computer Science, 1150:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sosnowski:1996:PTM

[1953] J. Sosnowski and A. Kusmierczyk.

Pseudorandom testing of microprocessors at instruction/data flow level. Lecture Notes in Computer Science, 1150: 246–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ubar:1996:MLT

[1954] R. Ubar and M. Brik. Multi-level test generation and fault diagnosis for finite state machines. Lecture Notes in Computer Science, 1150:264–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnould:1996:DTB

[1955] A. Arnould, B. Marre, and P. Le Gall.

Dynamic testing from bounded data type specifications. Lecture Notes in Computer Science, 1150:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbey:1996:TSB

[1956] S. Barbey, D. Buchs, and C. Peraire.

A theory of specification-based testing for object-oriented software. Lecture Notes in Computer Science, 1150:303– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernardeschi:1996:PSP

[1957] C. Bernardeschi, A. Fantechi, S. Gnesi, and G. Mongardi. Proving safety properties for embedded control systems. Lecture Notes in Computer Science, 1150:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosquer:1996:EDC

[1958] F. J. N. Cosquer, P. Antunes, and P. Verissimo. Enhancing dependability of cooperative applications in partitionable environments. Lecture Notes in Computer Science, 1150:335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mostefaoui:1996:EML

[1959] A. Mostefaoui and M. Raynal. Efficient message logging for uncoordinated checkpointing protocols. Lecture Notes in Computer Science, 1150:353– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerraoui:1996:AUR

[1960] R. Guerraoui, R. Oliveira, and A. Schiper. Atomic updates of replicated data. Lecture Notes in Computer Science, 1150:365–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Postma:1996:RAF

[1961] A. Postma, G. Hartman, and T. Krol. Removal of all faulty nodes from a fault-tolerant service by means of distributed diagnosis with imperfect fault coverage. Lecture Notes in Computer Science, 1150:385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altmann:1996:CBS

[1962] J. Altmann, A. Pataricza, T. Bartha, and P. Urban. Constraint based system-level diagnosis of multiprocessors. Lecture Notes in Computer Science, 1150:403–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthet:1996:UTF

[1963] G. G. Berthet and H. J. Nussbaumer.

A unified theory for f1/f2-Diagnosable communication networks. Lecture Notes in Computer Science, 1150:421– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lampson:1996:HBH

[1964] B. W. Lampson. How to build a highly available system using consensus. Lecture Notes in Computer Science, 1151:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1996:DAC

[1965] D. Ferrari. Distributed admission control algorithms for real-time communication. Lecture Notes in Computer Science, 1151:18–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilera:1996:RFD

[1966] M. K. Aguilera and S. Toueg. Randomization and failure detection: a hybrid approach to solve consensus. Lecture Notes in Computer Science, 1151: 29–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borcherding:1996:LAD

[1967] M. Borcherding. Levels of authentication in distributed agreement. Lecture Notes in Computer Science, 1151:40– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attiya:1996:ERS

[1968] H. Attiya. Efficient and robust sharing of memory in message-passing systems. Lecture Notes in Computer Science, 1151:56–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torres-Rojas:1996:PCC

[1969] F. J. Torres-Rojas and M. Ahamad. Plausible clocks: Constant size logical clocks for distributed systems. Lecture Notes in Computer Science, 1151:71– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charpentier:1996:ACR

[1970] M. Charpentier, M. Filali, P. Mauran, and G. Padiou. Abstracting communication to reason about distributed algorithms. Lecture Notes in Computer Science, 1151:89–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basu:1996:SRL

[1971] A. Basu, B. Charron-Bost, and S. Toueg. Simulating reliable links with unreliable links in the presence of process crashes. Lecture Notes in Computer Science, 1151:105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodrigues:1996:CDG

[1972] H. Rodrigues and R. Jones. A cyclic distributed garbage collector for network objects. Lecture Notes in Computer Science, 1151:123–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dickman:1996:IDO

[1973] P. Dickman. Incremental, distributed orphan detection and Actor garbage collection using graph partitioning and Euler cycles. Lecture Notes in Computer Science, 1151:141–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fonseca:1996:FAN

[1974] P. Fonseca and Z. Mammeri. A framework for the analysis of nondeterministic clock synchronization algorithms. Lecture Notes in Computer Science, 1151:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mei:1996:OTB

[1975] A. Mei, F. Bao, Y. Hamada, and Y. Igarashi. Optimal time broadcasting in faulty star networks. Lecture Notes in Computer Science, 1151:175– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eilam:1996:LBL

[1976] T. Eilam, S. Moran, and S. Zaks. A lower bound for linear interval routing. Lecture Notes in Computer Science, 1151:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gambosi:1996:TRS

[1977] G. Gambosi and P. Vocca. Topological routing schemes. Lecture Notes in Computer Science, 1151:206–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1996:MDS

[1978] S. Fujita and M. Yamashita. Maintaining a dynamic set of processors in a distributed system. Lecture Notes in Computer Science, 1151:220–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dwork:1996:CC

[1979] C. Dwork, C.-T. Ho, and R. Strong. Collective consistency. Lecture Notes in Computer Science, 1151:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bazzi:1996:PQ

[1980] R. A. Bazzi. Planar quorums. Lecture Notes in Computer Science, 1151: 251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerraoui:1996:GAF

[1981] R. Guerraoui and A. Schiper. “gammaaccurate” failure detectors. Lecture Notes in Computer Science, 1151:269– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moir:1996:FLL

[1982] M. Moir and J. A. Garay. Fast longlived renaming improved and simplified. Lecture Notes in Computer Science, 1151:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizuno:1996:TBT

[1983] M. Mizuno and H. Kakugawa. A timestamp based transformation of selfstabilizing programs for distributed computing environments. Lecture Notes in Computer Science, 1151:304– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attiya:1996:CSW

[1984] H. Attiya and S. Rajsbaum. The combinatorial structure of wait-free solvable tasks. Lecture Notes in Computer Science, 1151:322–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moran:1996:RH

[1985] S. Moran and L. Rappoport. On the robustness of h0rm. Lecture Notes in Computer Science, 1151:344–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1996:USC

[1986] S. Chaudhuri and P. Reiners. Understanding the set consensus partial order using the Borowsky-Gafni simulation. Lecture Notes in Computer Science, 1151:362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:DHF

[1987] M. A. Lee, H. Esbensen, and L. Lemaitre. The design of hybrid fuzzy/evolutionary multiobjective optimization algorithms. Lecture Notes in Computer Science, 1152:1–20, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakanishi:1996:SIA

[1988] S. Nakanishi, A. Ohtake, R. R. Yager, and S. Ohtani. Structure identification of acquired knowledge in fuzzy inference by genetic algorithms. Lecture Notes in Computer Science, 1152:21– 34, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishibuchi:1996:FCS

[1989] H. Ishibuchi, T. Nakashima, and T. Murata. A fuzzy classifier system that generates linguistic rules for pattern classification problems. Lecture Notes in Computer Science, 1152:35– 54, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nomura:1996:NCU

[1990] T. Nomura and T. Miyoshi. Numerical coding and unfair average crossover in GA for fuzzy rule extraction in dynamic environments. Lecture Notes in Computer Science, 1152:55–72, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshikawa:1996:AFR

[1991] T. Yoshikawa, T. Furuhashi, and Y. Uchikawa. Acquisition of fuzzy rules from DNA coding method. Lecture Notes in Computer Science, 1152:73– 88, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aoki:1996:ESA

[1992] T. Aoki, T. Oka, S. Hayakawa, and T. Suzuki. Experimental study on acquisition of optimal action for autonomous mobile robot to avoid moving multiobstacles. Lecture Notes in Computer Science, 1152:89–103, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CostaBranco:1996:NAS

[1993] P. J. Costa Branco, N. Lori, and J. A.

Dente. New approaches on structure identification of fuzzy models: Case study in an electro-mechanical system. Lecture Notes in Computer Science, 1152:104–143, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kouzani:1996:GFN

[1994] A. Z. Kouzani and A. Bouzerdoum.

A generic fuzzy neuron and its application to motion estimation. Lecture Notes in Computer Science, 1152:144– 171, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buller:1996:FFI

[1995] A. Buller. Fizzy-fuzzy inferencing. Lecture Notes in Computer Science, 1152: 172–187, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perez-Mithana:1996:MLP

[1996] E. Perez-Mithana, P. Ross, and J. Hallam. Multi-layer perceptron design using Delaunay triangulations. Lecture Notes in Computer Science, 1152:188– 197, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inuiguchi:1996:GAP

[1997] M. Inuiguchi, T. Miyake, M. Sakawa, and I. Shiromaru. A genetic algorithm for planning coal purchase of a real electric power plant. Lecture Notes in Computer Science, 1152:198–214, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGaris:1996:CBA

[1998] H. De Garis. “CAM-BRAIN” ATR’s billion neuron artificial brain project — A three year progress report. Lecture Notes in Computer Science, 1152:215– 234, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carter:1996:RDP

[1999] M. W. Carter and G. Laporte. Recent developments in practical examination timetabling. Lecture Notes in Computer Science, 1153:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bardadym:1996:CAS

[2000] V. A. Bardadym. Computer-aided school and university timetabling: The new wave. Lecture Notes in Computer Science, 1153:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wren:1996:STR

[2001] A. Wren. Scheduling, timetabling and rostering — A special relationship? Lecture Notes in Computer Science, 1153:46–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burke:1996:ETB

[2002] E. K. Burke, D. G. Elliman, P. H.

Ford, and R. F. Weare. Examination timetabling in British universities — A survey. Lecture Notes in Computer Science, 1153:76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meisels:1996:ETC

[2003] A. Meisels, E. Gudes, and G. Solotorevsky. Employee timetabling, constraint networks and knowledge-based rules: a mixed approach. Lecture Notes in Computer Science, 1153:93– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ram:1996:ATT

[2004] V. Ram and C. Scogings. Automated time table generation using multiple context reasoning with truth maintenance. Lecture Notes in Computer Science, 1153:106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1996:ICL

[2005] C. Cheng, L. Kang, N. Leung, and G. M. White. Investigations of a constraint logic programming approach to university timetabling. Lecture Notes in Computer Science, 1153:112–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gueret:1996:BUT

[2006] C. Gueret, N. Jussien, P. Boizumault, and C. Prins. Building university timetables using constraint logic programming. Lecture Notes in Computer Science, 1153:130–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lajos:1996:CUM

[2007] G. Lajos. Complete university modular timetabling using constraint logic programming. Lecture Notes in Computer Science, 1153:146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henz:1996:UOC

[2008] M. Henz and J. Wuertz. Using Oz for college timetabling. Lecture Notes in Computer Science, 1153:162–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rich:1996:SGA

[2009] D. C. Rich. A smart genetic algorithm for university timetabling. Lecture Notes in Computer Science, 1153: 181–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erben:1996:GAS

[2010] W. Erben and J. Keppler. A genetic algorithm solving a weekly coursetimetabling problem. Lecture Notes in Computer Science, 1153:198–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erguel:1996:GBE

[2011] A. Erguel. GA-Based examination scheduling experience at Middle East Technical University. Lecture Notes in Computer Science, 1153:212–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corne:1996:PIS

[2012] D. Corne and P. Ross. Peckish initialisation strategies for evolutionary timetabling. Lecture Notes in Computer Science, 1153:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burke:1996:MAU

[2013] E. K. Burke, J. P. Newall, and R. F.

Weare. A memetic algorithm for university exam timetabling. Lecture Notes in Computer Science, 1153:241– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paechter:1996:EMT

[2014] B. Paechter, A. Cumming, M. G.

Norman, and H. Luchian. Extensions to a memetic timetabling system. Lecture Notes in Computer Science, 1153:251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rankin:1996:ATP

[2015] R. C. Rankin. Automatic timetabling in practice. Lecture Notes in Computer Science, 1153:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:1996:CTC

[2016] T. B. Cooper and J. H. Kingston. The complexity of timetable construction problems. Lecture Notes in Computer Science, 1153:283–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeWerra:1996:SCM

[2017] D. De Werra. Some combinatorial models for course scheduling. Lecture Notes in Computer Science, 1153:296– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ross:1996:PTN

[2018] P. Ross, D. Corne, and H. Terashima-Marin. The phase-transition niche for evolutionary algorithms in timetabling. Lecture Notes in Computer Science, 1153:309–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boufflet:1996:TMU

[2019] J. P. Boufflet and S. Negre. Three methods used to solve an examination timetable problem. Lecture Notes in Computer Science, 1153:327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thompson:1996:GCS

[2020] J. Thompson and K. A. Dowsland. General cooling schedules for a simulated annealing based timetabling system. Lecture Notes in Computer Science, 1153:345–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robert:1996:HDC

[2021] V. Robert and A. Hertz. How to decompose constrained course scheduling problems into easier assignment type subproblems. Lecture Notes in Computer Science, 1153:364–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minker:1996:LDY

[2022] J. Minker. Logic and databases: a 20 year retrospective. Lecture Notes in Computer Science, 1154:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakshmanan:1996:PAD

[2023] L. V. S. Lakshmanan and N. Shiri.

A parametric approach to deductive databases with uncertainty. Lecture Notes in Computer Science, 1154:61– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subrahmanian:1996:DDA

[2024] V. S. Subrahmanian and C. Ward. A deductive database approach to planning in uncertain environments. Lecture Notes in Computer Science, 1154: 83–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuijpers:1996:TPS

[2025] B. Kuijpers, J. Paredaens, M. Sinits, and J. Van den Bussche. Termination properties of spatial datalog programs. Lecture Notes in Computer Science, 1154:101–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1996:ATR

[2026] C. Martin and J. Sistac. Applying transition rules to bitemporal deductive databases for integrity constraint checking. Lecture Notes in Computer Science, 1154:117–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kowalski:1996:TUA

[2027] R. Kowalski and F. Sadri. Towards a unified agent architecture that combines rationality with reactivity. Lecture Notes in Computer Science, 1154: 137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laurent:1996:LQR

[2028] D. Laurent and C. Vrain. Learning query rules for optimizing databases with update rules. Lecture Notes in Computer Science, 1154:153–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baral:1996:FCA

[2029] C. Baral and J. Lobo. Formal characterization of active databases. Lecture Notes in Computer Science, 1154: 175–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludaescher:1996:NTL

[2030] B. Ludaescher, W. May, and G. Lausen. Nested transactions in a logical language for active rules. Lecture Notes in Computer Science, 1154:197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siebes:1996:DDC

[2031] A. Siebes, S. Tsur, J. Ullman, and L. Vielle. Deductive databases: Challenges, opportunities and future directions. Lecture Notes in Computer Science, 1154:225–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakobovits:1996:RSM

[2032] H. Jakobovits and D. Vermeir. Rstable models for logic programs. Lecture Notes in Computer Science, 1154: 233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1996:EPP

[2033] T. Eiter, N. Leone, and D. Sacca. The expressive power of partial models in disjunctive deductive databases. Lecture Notes in Computer Science, 1154: 245–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leone:1996:SMC

[2034] N. Leone, P. Rullo, and F. Scarcello.

Stable model checking for disjunctive logic programs. Lecture Notes in Computer Science, 1154:265–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulanger:1996:ALB

[2035] D. Boulanger. Analysis of logic-based systems. Lecture Notes in Computer Science, 1154:281–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spalka:1996:AIC

[2036] A. Spalka and A. B. Cremers. An axiomatic interpretation of confidentiality demands in logic-based relational databases. Lecture Notes in Computer Science, 1154:303–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pulido:1996:DQE

[2037] E. Pulido. Database query evaluation with the STARBASE method. Lecture Notes in Computer Science, 1154:323– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sagonas:1996:LFO

[2038] K. Sagonas, T. Swift, and D. S. Warren. The limits of fixed-order computation. Lecture Notes in Computer Science, 1154:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kero:1996:ISD

[2039] B. Kero and S. Tsur. The IQ system: a deductive database information lens for reasoning about textual information. Lecture Notes in Computer Science, 1154:367–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobbie:1996:AAO

[2040] G. Dobbie and R. Topor. Arithmetic and aggregate operators in deductive object-oriented databases. Lecture Notes in Computer Science, 1154: 387–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonner:1996:SDD

[2041] A. Bonner and G. Mecca. Sequence datalog: Declarative string manipulation in databases. Lecture Notes in Computer Science, 1154:399–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asirelli:1996:LES

[2042] P. Asirelli, C. Renso, and F. Turini.

Language extensions for semantic integration of deductive databases. Lecture Notes in Computer Science, 1154: 415–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greco:1996:UBS

[2043] S. Greco, C. Mateis, and E. Spadafora.

Unification of bounded simple set terms in deductive databases. Lecture Notes in Computer Science, 1154:435–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cabibbo:1996:ESL

[2044] L. Cabibbo. Expressiveness of semipositive logic programs with value invention. Lecture Notes in Computer Science, 1154:457–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gire:1996:MED

[2045] F. Gire and H. K. Hoang. A more expressive deterministic query language with efficient symmetry-based choice construct. Lecture Notes in Computer Science, 1154:475–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roberts:1996:BNT

[2046] J. Roberts, U. Mocci, and J. Virtamo.

Broadband network traffic: Performance evaluation and design of broadband multiservice networks. Lecture Notes in Computer Science, 1155:xxii + 584, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:1996:APP

[2047] G. A. Geist. Advanced programming in PVM. Lecture Notes in Computer Science, 1156:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beguelin:1996:TMD

[2048] A. Beguelin and V. Sunderam. Tools for monitoring, debugging, and programming in PVM. Lecture Notes in Computer Science, 1156:7–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hempel:1996:SMM

[2049] R. Hempel. The status of the MPI message-passing standard and its relation to PVM. Lecture Notes in Computer Science, 1156:14–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pokorny:1996:CMP

[2050] S. Pokorny. A comparison of messagepassing parallelization to sharedmemory parallelization. Lecture Notes in Computer Science, 1156:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roda:1996:PEI

[2051] J. Roda, J. Herrera, J. Gonzalez, and C. Rodriguez. Practical experiments to improve PVM algorithms. Lecture Notes in Computer Science, 1156:30– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soch:1996:PCG

[2052] M. Soch, J. Trdlicka, and P. Tvrdik.

PVM, computational geometry, and parallel computing course. Lecture Notes in Computer Science, 1156:38– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luksch:1996:PSA

[2053] P. Luksch, U. Maier, S. Rathmayer, and M. Weidmann. Parallelization of a state-of-the-art industrial CFD package for execution on networks of workstations and massively parallel processors. Lecture Notes in Computer Science, 1156:45–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wechsler:1996:PMM

[2054] K. Wechsler, M. Breuer, and F. Durst.

A parallel multigrid method for the prediction of incompressible flows on workstation clusters. Lecture Notes in Computer Science, 1156:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:1996:ADT

[2055] M. May and T. Schiekofer. An abstract data type for parallel simulations based on sparse grids. Lecture Notes in Computer Science, 1156:59–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sunderam:1996:PSS

[2056] V. Sunderam. The PVM system: Status, trends, and directions. Lecture Notes in Computer Science, 1156:68– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludwig:1996:OOC

[2057] T. Ludwig, R. Wismueller, and M. Oberhuber. OCM — an OMIS compliant monitoring system. Lecture Notes in Computer Science, 1156:81– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wismueller:1996:SBV

[2058] R. Wismueller. State based visualization of PVM applications. Lecture Notes in Computer Science, 1156:91– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahlert:1996:WCP

[2059] M. Kahlert, U. Wever, and Q. Zheng. Workstation cluster as a parallel hardware environment for high performance computing. Lecture Notes in Computer Science, 1156:100–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1996:UPI

[2060] N. Papakostas, G. Papakonstantinou, and P. Tsanakas. Using PVM to implement PPARDB/PVM, a portable parallel database management system. Lecture Notes in Computer Science, 1156:108–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suttner:1996:SPB

[2061] C. B. Suttner. SPTHEO — A PVM-Based parallel theorem prover. Lecture Notes in Computer Science, 1156:116– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Betten:1996:PCF

[2062] A. Betten. Parallel construction of finite solvable groups. Lecture Notes in Computer Science, 1156:126–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menden:1996:PPP

[2063] J. Menden and G. Stellner. Proving properties of PVM applications — A case study with CoCheck. Lecture Notes in Computer Science, 1156:134– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbenz:1996:MDS

[2064] P. Arbenz, M. Billeter, P. Guentert, and P. Luginbuehl. Molecular dynamics simulations on Cray clusters using the SCIDDLE-PVM environment. Lecture Notes in Computer Science, 1156: 142–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cela:1996:PSP

[2065] J. M. Cela and J. M. Alfonso. Parallelization of the SPAI preconditioner in a master-slave configuration. Lecture Notes in Computer Science, 1156: 150–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuele:1996:PLA

[2066] J. Schuele. Parallel Lanczos algorithm on a CRAY-T3D combining PVM and SHMEM routines. Lecture Notes in Computer Science, 1156:158–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huckle:1996:PIS

[2067] T. Huckle. PVM-Implementation of sparse approximate inverse preconditioners for solving large sparse linear equations. Lecture Notes in Computer Science, 1156:166–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaszczyk:1996:EPI

[2068] A. Blaszczyk and C. Trinitis. Experience with PVM in an industrial environment. Lecture Notes in Computer Science, 1156:174–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagg:1996:TGR

[2069] G. E. Fagg, K. S. London, and J. J. Dongarra. Taskers and general resource managers: PVM supporting DCE process management. Lecture Notes in Computer Science, 1156:180–187, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juhasz:1996:PIP

[2070] Z. Juhasz and D. Crookes. A PVM implementation of a portable parallel image processing library. Lecture Notes in Computer Science, 1156:188–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Celino:1996:PET

[2071] M. Celino. Porting of an empirical tight-binding molecular dynamics code on MIMD platforms. Lecture Notes in Computer Science, 1156:197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahmen:1996:ATP

[2072] H.-C. Dahmen, U. Glaeser, and H. T.

Vierhaus. Automatic test pattern generation with optimal load balancing. Lecture Notes in Computer Science, 1156:205–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirik:1996:PCG

[2073] E. Hirik and I. Exman. Parallel crystal growth for chip placement. Lecture Notes in Computer Science, 1156:213– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernaschi:1996:EMP

[2074] M. Bernaschi. Efficient message passing on shared memory multiprocessors. Lecture Notes in Computer Science, 1156:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Touhafi:1996:DPC

[2075] A. Touhafi, W. Brissinck, and E. F.

Dirkx. Development of PVM code for a low latency switch based interconnect. Lecture Notes in Computer Science, 1156:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1996:PIP

[2076] J. M. Blum, T. M. Warschko, and W. F. Tichy. PSPVM: Implementing PVM on a high-speed interconnect for workstation clusters. Lecture Notes in Computer Science, 1156:235–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haechler:1996:IAC

[2077] G. Haechler and H. Burkhart. Implementing the ALWAN communication and data distribution library using PVM. Lecture Notes in Computer Science, 1156:243–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebner:1996:TFP

[2078] R. Ebner and A. Pfaffinger. Transformation of functional programs into data flow graphs implemented with PVM. Lecture Notes in Computer Science, 1156:251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gennart:1996:CAG

[2079] B. A. Gennart, J. Tarraga Gimenez, and R. D. Hersch. Computer-assisted generation of PVM/C++ programs using CAP. Lecture Notes in Computer Science, 1156:259–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levin:1996:LBC

[2080] E. I. Levin and A. I. Zhmakin. Load balancing for computation of steady viscous flows using adaptive unstructured grids. Lecture Notes in Computer Science, 1156:270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gold:1996:UAL

[2081] C. Gold and T. Schnekenburger. Using the ALDY load distribution system for PVM applications. Lecture Notes in Computer Science, 1156:278–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santana:1996:PVM

[2082] M. S. Santana, R. C. Santana, P. S.

Souza, and S. S. Souza. Parallel virtual machine for Windows95. Lecture Notes in Computer Science, 1156:288– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonnet:1996:UPW

[2083] C. Bonnet. Using PVM in wireless network environments. Lecture Notes in Computer Science, 1156:296–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uhl:1996:PIC

[2084] A. Uhl and J. Haemmerle. Parallel image compression on a workstationcluster using PVM. Lecture Notes in Computer Science, 1156:301–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1996:GSG

[2085] T. Wagner, C. Kueblbeck, and C. Schittko. Genetic selection and generation of textural features with PVM. Lecture Notes in Computer Science, 1156:305–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Betten:1996:SIL

[2086] A. Betten and A. Wassermann. 00,1-Solutions of integer linear equation systems. Lecture Notes in Computer Science, 1156:311–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Theodoropoulos:1996:ESP

[2087] P. Theodoropoulos, G. Manis, P. Tsanakas, and G. Papakonstantinou. Extending synchronization PVM mechanisms. Lecture Notes in Computer Science, 1156:315–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cela:1996:PPL

[2088] J. M. Cela, J. M. Alfonso, and J. Labarta. PLS: a parallel linear solvers library for domain decomposition methods. Lecture Notes in Computer Science, 1156:319–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:1996:PIA

[2089] R. Alt and J. L. Lamotte. Parallel integration across time of initial value problems using PVM. Lecture Notes in Computer Science, 1156:323–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veigneau:1996:DCR

[2090] S. Veigneau. Distributed computation of ribbon tableaux and spin polynomials. Lecture Notes in Computer Science, 1156:327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anglano:1996:PMB

[2091] C. Anglano and L. Portinale. Parallel model-based diagnosis using PVM. Lecture Notes in Computer Science, 1156:331–??, 1996. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuasa:1996:RPG

[2092] F. Yuasa, S. Kawabata, T. Ishikawa, and D. Perret-Gallix. Running PVM-GRACE on workstation clusters. Lecture Notes in Computer Science, 1156: 335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cavenaghi:1996:UPS

[2093] M. A. Cavenaghi, R. Spolon, J. E. M.

Perea-Martins, and S. G. Domingues. Using PVM in the simulation of a hybrid dataflow architecture. Lecture Notes in Computer Science, 1156:343– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oberhuber:1996:MNP

[2094] M. Oberhuber. Managing nondeterminism in PVM programs. Lecture Notes in Computer Science, 1156:347– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ragg:1996:PEN

[2095] T. Ragg. Parallelization of an evolutionary neural network optimizer based on PVM. Lecture Notes in Computer Science, 1156:351–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAcierno:1996:SUP

[2096] A. D’Acierno and S. Palma. On the synergistic use of parallel exact solvers and heuristic/stochastic methods for combinatorial optimisation problems: a study case for the TSP on the Meiko CS-2. Lecture Notes in Computer Science, 1156:355–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachman:1996:IOO

[2097] C. W. Bachman. Impact of object oriented thinking on ER modeling. Lecture Notes in Computer Science, 1157:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowsky:1996:TSF

[2098] J. A. Makowsky and E. V. Ravve. Translation schemes and the fundamental problem of database design. Lecture Notes in Computer Science, 1157:5–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mannila:1996:SDK

[2099] H. Mannila. Schema design and knowledge discovery. Lecture Notes in Computer Science, 1157:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biskup:1996:DRT

[2100] J. Biskup, R. Menzel, T. Polle, and Y. Sagiv. Decomposition of relationships through pivoting. Lecture Notes in Computer Science, 1157:28– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hainaut:1996:UIR

[2101] J.-L. Hainaut, J.-M. Hick, V. Englebert, and J. Henrard. Understanding the implementation of IS-A relations. Lecture Notes in Computer Science, 1157:42–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Theodoratos:1996:DOO

[2102] D. Theodoratos. Deductive object oriented schemas. Lecture Notes in Computer Science, 1157:58–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

TerHofstede:1996:VPC

[2103] A. H. M. Ter Hofstede, M. E. Orlowska, and J. Rajapakse. Verification problems in conceptual workflow specifications. Lecture Notes in Computer Science, 1157:73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maret:1996:PKM

[2104] P. Maret and J.-M. Pinon. Process knowledge modeling. Lecture Notes in Computer Science, 1157:89–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steeg:1996:CDD

[2105] M. Steeg. The conceptual database design optimizer CoDO — concepts, implementation, application. Lecture Notes in Computer Science, 1157:105– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloesch:1996:CCQ

[2106] A. C. Bloesch and T. A. Halpin.

ConQuer: a conceptual query language. Lecture Notes in Computer Science, 1157:121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clauss:1996:USR

[2107] W. Clauss. Using structural recursion as query mechanism for data models with references. Lecture Notes in Computer Science, 1157:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altus:1996:MDS

[2108] M. Altus. A modular design strategy for a flexible graphical database design environment: An experimental study. Lecture Notes in Computer Science, 1157:146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebert:1996:GBM

[2109] J. Ebert, A. Winter, P. Dahrn, and A. Franzke. Graph based modeling and implementation with EER/ GRAL. Lecture Notes in Computer Science, 1157:163–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vermeer:1996:ASI

[2110] M. W. W. Vermeer and P. M. G. Apers. On the applicability of schema integration techniques to database interoperation. Lecture Notes in Computer Science, 1157:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitt:1996:IIT

[2111] I. Schmitt and G. Saake. Integration of inheritance trees as part of view generation for database federations. Lecture Notes in Computer Science, 1157:195–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ekenberg:1996:FBD

[2112] L. Ekenberg and P. Johannesson. A formal basis for dynamic schema integration. Lecture Notes in Computer Science, 1157:211–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moody:1996:GER

[2113] D. Moody. Graphical entity relationship models: Towards a more user understandable representation of data. Lecture Notes in Computer Science, 1157:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maier:1996:BQD

[2114] R. Maier. Benefits and quality of data modelling — results of an empirical analysis. Lecture Notes in Computer Science, 1157:245–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ortner:1996:NLA

[2115] E. Ortner and B. Schienmann. Normative language approach — A framework for understanding. Lecture Notes in Computer Science, 1157:261–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Assenova:1996:IQC

[2116] P. Assenova and P. Johannesson. Improving quality in conceptual modelling by the use of schema transformations. Lecture Notes in Computer Science, 1157:277–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SoaresDaSilva:1996:AMO

[2117] A. Soares Da Silva, A. H. F. Laender, and M. A. Cassanova. An approach to maintaining optimized relational representations of entity-relationship schemas. Lecture Notes in Computer Science, 1157:292–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mok:1996:TCM

[2118] Wai Yin Mok and D. W. Ernbley. Transforming conceptual models to object-oriented database designs: Practicalities, properties, and peculiarities. Lecture Notes in Computer Science, 1157:309–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Topaloglou:1996:RPS

[2119] T. Topaloglou and J. Mylopoulos. Representing partial spatial information in databases. Lecture Notes in Computer Science, 1157:325–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:SCT

[2120] Jae Young Lee, R. Elmasri, and J. Won. Specification of calendars and time series for temporal databases. Lecture Notes in Computer Science, 1157: 341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schewe:1996:VCC

[2121] K.-D. Schewe and B. Schewe. Viewcentered conceptual modelling — an object oriented approach. Lecture Notes in Computer Science, 1157:357– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comyn-Wattiau:1996:RER

[2122] I. Comyn-Wattiau and J. Akoka. Reverse engineering of relational database physical schema. Lecture Notes in Computer Science, 1157:372–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soutou:1996:EAR

[2123] C. Soutou. Extracting N-ary relationships through database reverse engineering. Lecture Notes in Computer Science, 1157:392–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaschek:1996:ICP

[2124] R. Kaschek. Inheritance as a conceptual primitive. Lecture Notes in Computer Science, 1157:406–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yahia:1996:IAM

[2125] A. Yahia, L. Lakhal, R. Cicchetti, and J.-P. Bordal. iO2: An algorithmic method for building inheritance graphs in object database design. Lecture Notes in Computer Science, 1157:422–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casali:1996:WE

[2126] F. Casali, S. Ceri, B. Pernici, and G. Pozzi. Workflow evolution. Lecture Notes in Computer Science, 1157:

438–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norrie:1996:MCS

[2127] M. C. Norrie, A. Steiner, A. Wuergler, and M. Wunderli. A model for classification structures with evolution control. Lecture Notes in Computer Science, 1157:456–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andonoff:1996:IVO

[2128] E. Andonoff, G. Hubert, A. Le Parc, and G. Zurfluh. Integrating versions in the OMT models. Lecture Notes in Computer Science, 1157:472–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1996:ICT

[2129] G. Barthe. Implicit coercions in type systems. Lecture Notes in Computer Science, 1158:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1996:TLA

[2130] G. Barthe, M. Ruys, and H. Barendregt. A two-level approach towards lean proof-checking. Lecture Notes in Computer Science, 1158:16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1996:GCD

[2131] U. Berger and H. Schwichtenberg. The greatest common divisor: a case study for program extraction from classical proofs. Lecture Notes in Computer Science, 1158:36–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beylin:1996:EPC

[2132] I. Beylin and P. Dybjer. Extracting a proof of coherence for monoidal categories from a proof of normalization for monoids. Lecture Notes in Computer Science, 1158:47–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cederquist:1996:CPH

[2133] J. Cederquist and S. Negri. A constructive proof of the Heine-Borel covering theorem for formal reals. Lecture Notes in Computer Science, 1158:62– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coquand:1996:ACC

[2134] T. Coquand and J. Smith. An application of constructive completeness. Lecture Notes in Computer Science, 1158: 76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cornes:1996:AII

[2135] C. Cornes and D. Terrasse. Automating inversion of inductive predicates in Coq. Lecture Notes in Computer Science, 1158:85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curmin:1996:FOM

[2136] P. Curmin. First order marked types.

Lecture Notes in Computer Science, 1158:105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dybjer:1996:ITT

[2137] P. Dybjer. Internal type theory. Lecture Notes in Computer Science, 1158: 120–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimenez:1996:ACI

[2138] E. Gimenez. An application of coinductive types in Coq: Verification of the alternating bit protocol. Lecture Notes in Computer Science, 1158:135– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1996:CER

[2139] M. Hofmann. Conservativity of equality reflection over intensional type theory. Lecture Notes in Computer Science, 1158:153–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honsell:1996:NDA

[2140] F. Honsell and M. Miculan. A natural deduction approach to dynamic logic. Lecture Notes in Computer Science, 1158:165–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magnusson:1996:ACI

[2141] L. Magnusson. An algorithm for checking incomplete proof objects in type theory with localization and unification. Lecture Notes in Computer Science, 1158:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padovani:1996:DAM

[2142] V. Padovani. Decidability of all minimal models. Lecture Notes in Computer Science, 1158:201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulin-Mohring:1996:CSC

[2143] C. Paulin-Mohring. Circuits as streams in Coq: Verification of a sequential multiplier. Lecture Notes in Computer Science, 1158:216–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranta:1996:CRS

[2144] A. Ranta. Context-relative syntactic categories and the formalization of mathematical text. Lecture Notes in Computer Science, 1158:231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stefanova:1996:SMC

[2145] M. Stefanova and H. Geuvers. A simple model construction for the calculus of constructions. Lecture Notes in Computer Science, 1158:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tammet:1996:OEF

[2146] T. Tammet and J. Smith. Optimized encodings of fragments of type theory in first order logic. Lecture Notes in Computer Science, 1158:265–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonPlato:1996:ODC

[2147] J. Von Plato. Organization and development of a constructive axiomatization. Lecture Notes in Computer Science, 1158:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karacapilidis:1996:ABF

[2148] N. I. Karacapilidis, D. Papadias, and T. F. Gordon. An argumentation based framework for defeasible and qualitative reasoning. Lecture Notes in Computer Science, 1159:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martins:1996:WBI

[2149] A. T. C. Martins, M. Pequeno, and T. Pequeno. Well-behaved IDL theories. Lecture Notes in Computer Science, 1159:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monteiro:1996:ECL

[2150] A. M. Monteiro and J. Wainer. Epistemic conditional logics. Lecture Notes in Computer Science, 1159:21– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nogueira:1996:HFT

[2151] J. H. M. Nogueira, A. L. Furtado, and J. Y. P. Alcazar. A hybrid formal theory of plan recognition and its implementation. Lecture Notes in Computer Science, 1159:31–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farreny:1996:ASS

[2152] H. Farreny. Algorithms ρ for ?- standard state graphs: Results about termination and admissibility. Lecture Notes in Computer Science, 1159:41– 50, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torgo:1996:RC

[2153] L. Torgo and J. Gama. Regression by classification. Lecture Notes in Computer Science, 1159:51–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scalabrin:1996:GMC

[2154] E. E. Scalabrin, L. Vandenberghe, H. De Azevedo, and J.-P. A. Barthes. A generic model of cognitive agent to develop open systems. Lecture Notes in Computer Science, 1159:61–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bazzan:1996:CAI

[2155] A. L. C. Bazzan. Coordination among individually-motivated agents: An evolutionary approach. Lecture Notes in Computer Science, 1159:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bharadwaj:1996:AHC

[2156] K. K. Bharadwaj, N. M. Hewahi, and M. A. Brandao. Adaptive hierarchical censored production rulebased system: a genetic algorithm approach. Lecture Notes in Computer Science, 1159:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spiessens:1996:FOR

[2157] P. Spiessens and B. Manderick. Finding optimal representations using the crossover correlation coefficient. Lecture Notes in Computer Science, 1159: 91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1996:EAR

[2158] O. C. Lopez, R. M. Barcia, O. Eyada, and F. O. Gauthier. An evolutionary algorithm for resource-constrained project scheduling and multiple execution modes. Lecture Notes in Computer Science, 1159:101–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robin:1996:ERS

[2159] J. Robin. Evaluating the robustness and scalability of revision-based natural language generation. Lecture Notes in Computer Science, 1159:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barros:1996:SAA

[2160] F. A. Barros. Semi-automatic anaphora resolution in portable natural language interfaces. Lecture Notes in Computer Science, 1159:121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MachadoRino:1996:DMG

[2161] L. H. Machado Rino and D. Scott. A discourse model for gist preservation. Lecture Notes in Computer Science, 1159:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leite:1996:ICO

[2162] J. A. F. Leite and E. R. Hancock. Iterative curve organisation with the EM algorithm. Lecture Notes in Computer Science, 1159:141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tagare:1996:MBP

[2163] H. Tagare and D. McDermott. Modelbased pose proposal for 2-D object recognition. Lecture Notes in Computer Science, 1159:151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaSilva:1996:TIS

[2164] A. J. F. Da Silva and E. Oliveira.

TARCA — an integrated system for diagnosis and treatment of cardiac arrhythmias. Lecture Notes in Computer Science, 1159:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malheiro:1996:IDE

[2165] B. Malheiro and E. Oliveira. Intelligent distributed environmental decision support system. Lecture Notes in Computer Science, 1159:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaSilva:1996:OOA

[2166] W. T. Da Silva and P. A. D. De Rezende. Object-oriented algorithm for combining dichotomic belief functions. Lecture Notes in Computer Science, 1159:181–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasconcelos:1996:ABA

[2167] W. W. Vasconcelos, M. A. T. Aragao, and N. E. Fuchs. Automatic bottom-up analysis and transformation of logic programs. Lecture Notes in Computer Science, 1159:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMeloRezende:1996:TEP

[2168] A. De Melo Rezende and J. Wainer. A temporal extension to the parsimonious covering theory. Lecture Notes in Computer Science, 1159:201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roisenberg:1996:BIN

[2169] M. Roisenberg, J. M. Barreto, and F. M. Azevedo. Biological inspirations in neural network implementations of autonomous agents. Lecture Notes in Computer Science, 1159:211–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moussalle:1996:ITS

[2170] N. M. Moussalle, R. M. Viccari, and M. Correa. Intelligent tutoring systems modelled through the mental states. Lecture Notes in Computer Science, 1159:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirota:1996:FFL

[2171] K. Hirota. Fundamentals of fuzzy logic and its hardware implementation. Lecture Notes in Computer Science, 1159:231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1996:SCC

[2172] M. Klein. Status and challenges for coordination technology: An AI perspective. Lecture Notes in Computer Science, 1159:232–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

PereiraLopes:1996:CNL

[2173] J. G. Pereira Lopes. Combining natural language understanding and information retrieval for flexible hypertext navigation. Lecture Notes in Computer Science, 1159:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandri:1996:NMT

[2174] S. Sandri. Numerical models for the treatment of imperfect information in knowledge-based systems. Lecture Notes in Computer Science, 1159:234– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viccari:1996:ITS

[2175] R. M. Viccari and L. M. Giraffa. Intelligent tutoring systems: Functional approach × agents approach. Lecture Notes in Computer Science, 1159:235– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aamodt:1996:CBR

[2176] A. Aamodt. The case-based reasoning paradigm: Perspective, technology and impacts. Lecture Notes in Computer Science, 1159:236–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bredeweg:1996:QMI

[2177] B. Bredeweg. Qualitative models for interactive articulated learning environments. Lecture Notes in Computer Science, 1159:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaCosta:1996:HC

[2178] N. C. A. Da Costa and F. A. Doria. H-computation. Lecture Notes in Computer Science, 1159:238–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1996:CVS

[2179] R. B. Fisher. Is computer vision still AI? Lecture Notes in Computer Science, 1159:239–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirota:1996:FRF

[2180] K. Hirota. Fuzzy robotics and fuzzy image understanding. Lecture Notes in Computer Science, 1159:240–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rocha:1996:TTM

[2181] A. F. Rocha. Toward a theory of molecular computing. Lecture Notes in Computer Science, 1159:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valiant:1996:MCN

[2182] L. G. Valiant. Managing complexity in neuroidal circuits (invited lecture). Lecture Notes in Computer Science, 1160:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takimoto:1996:LEE

[2183] E. Takimoto, Y. Sakai, and A. Maruoka. Learnability of exclusive-or expansion based on monotone DNF formulas. Lecture Notes in Computer Science, 1160: 12–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Long:1996:IBA

[2184] P. M. Long. Improved bounds about on-line learning of smooth functions of a single variable. Lecture Notes in Computer Science, 1160:26–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1996:QLB

[2185] A. Nakamura. Query learning of bounded-width OBDDs. Lecture Notes in Computer Science, 1160:37–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buening:1996:LRO

[2186] H. K. Buening and T. Lettmann.

Learning a representation for optimizable formulas. Lecture Notes in Computer Science, 1160:51–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliveira:1996:LEA

[2187] A. L. Oliveira and S. Edwards. Limits of exact algorithms for inference of minimum size finite state machines. Lecture Notes in Computer Science, 1160: 59–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitanyi:1996:GFO

[2188] P. Vitanyi. Genetic fitness optimization using rapidly mixing Markov chains (invited lecture). Lecture Notes in Computer Science, 1160:67–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baxter:1996:KCM

[2189] R. A. Baxter and J. J. Oliver. The kindest cut: minimum message length segmentation. Lecture Notes in Computer Science, 1160:83–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pearson:1996:RCD

[2190] R. A. Pearson and E. K. T. Smith. Reducing complexity of decision trees with two variable tests. Lecture Notes in Computer Science, 1160:91– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1996:CEL

[2191] V. Arvind and N. V. Vinodchandran. The complexity of exactly learning algebraic concepts. Lecture Notes in Computer Science, 1160:100–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fahmy:1996:ELR

[2192] A. F. Fahmy and R. S. Roos. Efficient learning of real time two-counter automata. Lecture Notes in Computer Science, 1160:113–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lavrac:1996:CSF

[2193] N. Lavrac, D. Gamberger, and P. Turney. Cost-sensitive feature reduction applied to a hybrid genetic algorithm. Lecture Notes in Computer Science, 1160:127–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piramuthu:1996:EFS

[2194] S. Piramuthu. Effects of feature selection with blurring’ on NeuroFuzzy systems. Lecture Notes in Computer Science, 1160:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quinlan:1996:BFO

[2195] J. R. Quinlan. Boosting first-order learning (invited lecture). Lecture Notes in Computer Science, 1160:143– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baerzdins:1996:IHK

[2196] J. Baerzdins and U. Sarkans. Incorporating hypothetical knowledge into the process of inductive synthesis. Lecture Notes in Computer Science, 1160:156– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1996:ICL

[2197] L. Martin and C. Vrain. Induction of constraint logic programs. Lecture Notes in Computer Science, 1160:169– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sugimoto:1996:CLT

[2198] N. Sugimoto, K. Hirata, and H. Ishizaka.

Constructive learning of translations based on dictionaries. Lecture Notes in Computer Science, 1160:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1996:ILP

[2199] J. Lu and J. Arima. Inductive logic programming beyond logical implication. Lecture Notes in Computer Science, 1160:185–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gamberger:1996:NEI

[2200] D. Gamberger, N. Lavrac, and S. Dzeroski. Noise elimination in inductive concept learning: a case study in medical diagnosis. Lecture Notes in Computer Science, 1160:199–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dowe:1996:MEP

[2201] D. L. Dowe, J. J. Oliver, and C. S. Wallace. MML estimation of the parameters of the spherical Fisher distribution. Lecture Notes in Computer Science, 1160:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1996:LE

[2202] S. Lange, R. Wiehagen, and T. Zeugmann. Learning by erasing. Lecture Notes in Computer Science, 1160:228– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1996:LCL

[2203] S. Jain, E. Kinber, and R. Wiehagen. On learning and co-learning of minimal programs. Lecture Notes in Computer Science, 1160:242–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shinohara:1996:IIU

[2204] T. Shinohara and H. Arimura. Inductive inference of unbounded unions of pattern languages from positive data (invited lecture). Lecture Notes in Computer Science, 1160:256–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

KrishnaRao:1996:CPP

[2205] M. R. K. Krishna Rao. A class of Prolog programs inferable from positive data. Lecture Notes in Computer Science, 1160:272–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1996:VBL

[2206] J. Case, S. Jain, and F. Stephan.

Vacillatory and BC learning on noisy data. Lecture Notes in Computer Science, 1160:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1996:TPL

[2207] A. Ambainis and R. Freivalds. Transformations that preserve learnability. Lecture Notes in Computer Science, 1160:299–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viksna:1996:PLI

[2208] J. Viksna. Probabilistic limit identification up to “small” sets. Lecture Notes in Computer Science, 1160:312– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grieser:1996:RII

[2209] G. Grieser. Reflecting inductive inference machines and its improvement by therapy. Lecture Notes in Computer Science, 1160:325–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mullender:1996:QSD

[2210] S. J. Mullender and P. Sijben. Quality of service in distributed multimedia systems. Lecture Notes in Computer Science, 1161:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coulson:1996:CCR

[2211] G. Coulson and D. G. Waddington. A CORBA compliant real-time multimedia platform for broadband networks. Lecture Notes in Computer Science, 1161:14–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hellemans:1996:IHC

[2212] P. Hellemans, F. Steegmans, H. Vanderstraeten, and H. Zuidweg. Implementation of hidden concurrency in CORBA clients. Lecture Notes in Computer Science, 1161:30–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grasso:1996:POV

[2213] E. Grasso. Passing objects by value in CORBA. Lecture Notes in Computer Science, 1161:43–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuster:1996:CMR

[2214] H. Schuster. Communication middleware for reliable workflow management systems. Lecture Notes in Computer Science, 1161:57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roantree:1996:MMH

[2215] M. Roantree, P. Hickey, A. Crilly, and J. Murphy. Metadata modelling for healthcare applications in a federated database system. Lecture Notes in Computer Science, 1161:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dreier:1996:RRG

[2216] B. Dreier, A. Huber, M. Zahn, and H. Karl. ReGTime — Rent Gigaflops someTimes. Lecture Notes in Computer Science, 1161:84–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blair:1996:SCC

[2217] H. Blair, H. Green, S. J. Caughey, and S. K. Shrivastava. Structuring call control software using distributed objects. Lecture Notes in Computer Science, 1161:94–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durmosch:1996:DIM

[2218] M. K. Durmosch, C. Egelhaaf, K.-D.

Engel, and P. Schoo. Design and implementation of a mulitmedia communication service in a distributed environment based on the TINA-C architecture. Lecture Notes in Computer Science, 1161:108–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

TroncoFudoli:1996:ETC

[2219] T. R. Tronco Fudoli and E. Najm. Exercise of TINA concepts for a video broadcast service over ATM networks. Lecture Notes in Computer Science, 1161:122–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leclerc:1996:CBD

[2220] M. Leclerc, T. Muessener, C. Linnhoff-Popien, and S. Lipperts. CORBA-Based data transfer for financial risk management. Lecture Notes in Computer Science, 1161:136–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zielinski:1996:CTD

[2221] K. Zielinski, A. Uszok, and M. Steinder. Crossing technological domains using the Inter-ORB request level bridge — preliminary performance study. Lecture Notes in Computer Science, 1161: 148–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radestock:1996:CES

[2222] M. Radestock and S. Eisenbach. Coordination in evolving systems. Lecture Notes in Computer Science, 1161: 162–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunkelmann:1996:IDT

[2223] T. Kunkelmann, H. Vogler, and S. Thomas. Interoperability of distributed transaction processing systems. Lecture Notes in Computer Science, 1161:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruender:1996:RID

[2224] H. Gruender and K. Geihs. Reuse and inheritance in distributed object systems. Lecture Notes in Computer Science, 1161:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thissen:1996:FOS

[2225] D. Thissen and C. Linnhoff-Popien.

Finding optimal services within a CORBA trader. Lecture Notes in Computer Science, 1161:200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1996:GTC

[2226] S. Mueller, K. Mueller-Jones, W. Lamersdorf, and T. Tu. Global trader cooperation in open service markets. Lecture Notes in Computer Science, 1161: 214–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caneschi:1996:ARU

[2227] F. Caneschi. Analyzing requirements using ODP. Lecture Notes in Computer Science, 1161:228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sinnott:1996:SMB

[2228] R. O. Sinnott and K. J. Turner. Specifying multimedia binding objects in Z. Lecture Notes in Computer Science, 1161:244–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sidou:1996:GEM

[2229] D. Sidou. A generic and executable model for the specification and validation of distributed behaviours. Lecture Notes in Computer Science, 1161:258– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frick:1996:FDI

[2230] O. Frick. Formal description and interpretation of coordination protocols for teamwork. Lecture Notes in Computer Science, 1161:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feitelson:1996:TCJ

[2231] D. G. Feitelson and L. Rudolph. Towards convergence in job schedulers for parallel supercomputers. Lecture Notes in Computer Science, 1162:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hotovy:1996:WEC

[2232] S. Hotovy. Workload evolution on the Cornell Theory Center IBM SP2. Lecture Notes in Computer Science, 1162: 27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skovira:1996:ELA

[2233] J. Skovira, W. Chan, H. Zhou, and D. Lifka. The EASY — LoadLeveler API project. Lecture Notes in Computer Science, 1162:41–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wan:1996:BSI

[2234] M. Wan, R. Moore, G. Kremenek, and K. Steube. A batch scheduler for the Intel Paragon MPP system with a non-contiguous node allocation algorithm. Lecture Notes in Computer Science, 1162:48–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehring:1996:AIR

[2235] J. Gehring and F. Ramme. Architectureindependent request-scheduling with tight waiting-time estimations. Lecture Notes in Computer Science, 1162:65– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feitelson:1996:PSG

[2236] D. G. Feitelson. Packing schemes for gang scheduling. Lecture Notes in Computer Science, 1162:89–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:GSD

[2237] F. Wang, H. Franke, M. Papaefthymiou, and P. Pattnaik. A gang scheduling design for multiprogrammed parallel computing environments. Lecture Notes in Computer Science, 1162:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hori:1996:IGS

[2238] A. Hori, H. Tezuka, Y. Ishikawa, and N. Soda. Implementation of gangscheduling on workstation cluster. Lecture Notes in Computer Science, 1162: 126–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pruyne:1996:MCP

[2239] J. Pruyne and M. Livny. Managing checkpoints for parallel programs. Lecture Notes in Computer Science, 1162: 140–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1996:URM

[2240] T. D. Nguyen, R. Vaswani, and J. Zahorjan. Using runtime measured workload characteristics in parallel processor scheduling. Lecture Notes in Computer Science, 1162:155–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1996:PAC

[2241] T. D. Nguyen, R. Vaswani, and J. Zahorjan. Parallel application characteristics for multiprocessor scheduling policy design. Lecture Notes in Computer Science, 1162:175–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiang:1996:DVS

[2242] S.-H. Chiang and M. Vernon. Dynamic vs. static quantum-based parallel processor allocation. Lecture Notes in Computer Science, 1162:200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padhye:1996:DVA

[2243] J. D. Padhye and L. W. Dowdy. Dynamic versus adaptive processor allocation policies for message passing parallel computers: An empirical comparison. Lecture Notes in Computer Science, 1162:224–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Islam:1996:DPD

[2244] N. Islam, A. Prodromidis, and M. S.

Squillante. Dynamic partitioning in different distributed-memory environments. Lecture Notes in Computer Science, 1162:244–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellosa:1996:LIB

[2245] F. Bellosa. Locality information based scheduling in shared memory multiprocessors. Lecture Notes in Computer Science, 1162:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenstra:1996:GSD

[2246] A. K. Lenstra. Generating standard DSA signatures without long inversion. Lecture Notes in Computer Science, 1163:57–64, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyaji:1996:MRS

[2247] A. Miyaji. A message recovery signature scheme equivalent to DSA over elliptic curves. Lecture Notes in Computer Science, 1163:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biehl:1996:CPB

[2248] I. Biehl, B. Meyer, and C. Thiel.

Cryptographic protocols based on realquadratic A-fields. Lecture Notes in Computer Science, 1163:15–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1996:MYP

[2249] R. Anderson and S. Vaudenay. Minding your p’s and q’s. Lecture Notes in Computer Science, 1163:26–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Just:1996:AMP

[2250] M. Just and S. Vaudenay. Authenticated multi-party key agreement. Lecture Notes in Computer Science, 1163: 36–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCurley:1996:CIL

[2251] K. S. McCurley. Cryptography and the Internet: Lessons and challenges. Lecture Notes in Computer Science, 1163: 50–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeWin:1996:FSI

[2252] E. De Win, A. Bosselaers, S. Vandenberghe, and P. De Getsem. A fast software implementation for arithmetic operations in GF(20n). Lecture Notes in Computer Science, 1163:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knudsen:1996:HFB

[2253] L. Knudsen and B. Preneel. Hash functions based on block ciphers and quaternary codes. Lecture Notes in Computer Science, 1163:77–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nyberg:1996:GFN

[2254] K. Nyberg. Generalized Feistel networks. Lecture Notes in Computer Science, 1163:91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawkes:1996:ALC

[2255] P. Hawkes and L. O’Connor. On applying linear cryptanalysis to IDEA. Lecture Notes in Computer Science, 1163: 105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fan:1996:MRT

[2256] C.-I. Fan and C.-L. Lei. A multirecastable ticket scheme for electronic elections. Lecture Notes in Computer Science, 1163:116–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michels:1996:SRR

[2257] M. Michels and P. Horster. Some remarks on a receipt-free and universally verifiable mix-type voting scheme. Lecture Notes in Computer Science, 1163: 125–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1996:ONR

[2258] J. Zhou and D. Gollmann. Observations on non-repudiation. Lecture Notes in Computer Science, 1163:133– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bleichenbacher:1996:EOT

[2259] D. Bleichenbacher and U. Maurer. On the efficiency of one-time digital signatures. Lecture Notes in Computer Science, 1163:145–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakurai:1996:HCA

[2260] K. Sakurai. A hidden cryptographic assumption in no-transferable identification schemes. Lecture Notes in Computer Science, 1163:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsujii:1996:EMK

[2261] S. Tsujii. Electronic money and key management from global and regional points of view. Lecture Notes in Computer Science, 1163:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobara:1996:LVS

[2262] K. Kobara and H. Imai. Limiting the visible space visual secret sharing schemes and their application to human identification. Lecture Notes in Computer Science, 1163:185–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1996:TCW

[2263] U. Maurer and S. Wolf. Towards characterizing when information-theoretic secret key agreement is possible. Lecture Notes in Computer Science, 1163: 196–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korjik:1996:KSB

[2264] V. Korjik and D. Kushnir. Key sharing based on the wire-tap channel type II concept with noisy main channel. Lecture Notes in Computer Science, 1163:210–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurosawa:1996:GHO

[2265] K. Kurosawa and T. Satoh. Generalization of higher order SAC to vector output Boolean functions. Lecture Notes in Computer Science, 1163:218– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chee:1996:CIF

[2266] S. Chee, S. Lee, D. Lee, and Soo Hak Sung. On the correlation immune functions and their nonlinearity. Lecture Notes in Computer Science, 1163:232– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abe:1996:HDB

[2267] M. Abe and E. Fujisaki. How to date blind signatures. Lecture Notes in Computer Science, 1163:244–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pointcheval:1996:PSB

[2268] D. Pointcheval and J. Stern. Provably secure blind signature schemes. Lecture Notes in Computer Science, 1163: 252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MRaihi:1996:CEP

[2269] D. M’Raihi. Cost-effective payment schemes with privacy regulation. Lecture Notes in Computer Science, 1163: 266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1996:MRI

[2270] A. Chan, Y. Frankel, P. MacKenzie, and Y. Tsiounis. Mis-representation of identities in E-cash schemes and how to prevent it. Lecture Notes in Computer Science, 1163:276–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frankel:1996:IDP

[2271] Y. Frankel, Y. Tsiounis, and M. Yung. “indirect discourse proofs”: Achieving efficient fair off-line E-cash. Lecture Notes in Computer Science, 1163:286– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stern:1996:VCA

[2272] J. Stern. The validation of cryptographic algorithms. Lecture Notes in Computer Science, 1163:301–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:CGS

[2273] Seung Joo Kim, Sung Jun Park, and Dong Ho Won. Convertible group signatures. Lecture Notes in Computer Science, 1163:311–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mambo:1996:HUT

[2274] M. Mambo, K. Sakurai, and E. Okamoto.

How to utilize the transformability of digital signatures for solving the Oracle Problem. Lecture Notes in Computer Science, 1163:322–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michels:1996:RDS

[2275] M. Michels and P. Horster. On the risk of disruption in several multiparty signature schemes. Lecture Notes in Computer Science, 1163:334–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geiselmann:1996:CAC

[2276] W. Geiselmann and D. Gottmann. Correlation attacks on cascades of clock controlled shift registers. Lecture Notes in Computer Science, 1163:346–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:CCA

[2277] Sangjin Lee, Seongtaek Chee, Sangjoon Park, and Sungmo Park. Conditional correlation attack on nonlinear filter generators. Lecture Notes in Computer Science, 1163:360–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chabaud:1996:CSS

[2278] F. Chabaud and J. Stern. The cryptographic security of the syndrome decoding problem for rank distance codes. Lecture Notes in Computer Science, 1163:368–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cowie:1996:WWN

[2279] J. Cowie, B. Dodson, R. M. Elkenbracht-Huizing, and A. K. Lenstra. A world wide number field sieve factoring record: on to 512 bits. Lecture Notes in Computer Science, 1163:382–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berquist:1996:MIH

[2280] K. Berquist and A. Berquist. Managing information highways. Lecture Notes in Computer Science, 1164:1–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abrial:1996:ASB

[2281] J.-R. Abrial, E. Boerger, and H. Langmaack. ABL: The steam boiler case study: Competition of formal program specification and development methods. Lecture Notes in Computer Science, 1165:1–12, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Addibpour:1996:SSP

[2282] M. Addibpour and E. Tyugu. AT: Structural synthesis of programs from refined user requirements (programming boiler control in NUT). Lecture Notes in Computer Science, 1165:13– 34, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andriessens:1996:AUF

[2283] C. Andriessens and T. Lindner. AL: Using FOCUS, LUSTRE, and probability theory for the design of a reliable control program. Lecture Notes in Computer Science, 1165:35–51, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beierle:1996:BRA

[2284] C. Beierle, E. Boerger, I. Durdanovic, and U. Glaesser. BBDGR: Refining abstract machine specifications of the steam boiler control to well documented executable code. Lecture Notes in Computer Science, 1165:52– 78, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1996:BAS

[2285] M. Bidoit, C. Chevenier, C. Pellen, and J. Ryckbosch. BCPR: An algebraic specification of the steam-boiler control system. Lecture Notes in Computer Science, 1165:79–108, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buessow:1996:BSB

[2286] R. Buessow and M. Weber. BW: a steam-boiler control specification with statecharts and Z. Lecture Notes in Computer Science, 1165:109–128, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1996:BAS

[2287] M. Butler, E. Sekerinski, and K. Sere.

BSS: An action system approach to the steam boiler problem. Lecture Notes in Computer Science, 1165:129–148, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattel:1996:CSB

[2288] T. Cattel and G. Duval. CD: The steam-boiler problem in lustre. Lecture Notes in Computer Science, 1165:149– 164, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuellar:1996:CSB

[2289] J. Cuellar and I. Wildgruber. CW1: The steam boiler problem — A TLT solution. Lecture Notes in Computer Science, 1165:165–183, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuellar:1996:CRT

[2290] J. Cuellar and I. Wildgruber. CW2: The real-time behavior of the steam boiler. Lecture Notes in Computer Science, 1165:184–202, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duval:1996:DSV

[2291] G. Duval and T. Cattel. DC: Specifying and verifying the steam-boiler problem with SPIN. Lecture Notes in Computer Science, 1165:203–217, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gargantini:1996:GTS

[2292] A. Gargantini and A. Morzenti. GM: TRIO specification of a steam boiler controller. Lecture Notes in Computer Science, 1165:218–232, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaudel:1996:GFS

[2293] M.-C. Gaudel, P. Dauchy, and C. Khoury. GDK: a formal specification of the steam-boiler control problem by algebraic specifications with implicit state. Lecture Notes in Computer Science, 1165:233–264, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1996:HUH

[2294] T. A. Henzinger and H. Wong-Toi.

HW: Using HyTech to synthesize control parameters for a steam boiler. Lecture Notes in Computer Science, 1165: 265–282, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ledru:1996:LVS

[2295] Y. Ledru and M.-L. Potet. LP: a VDM specification of the steam-boiler problem. Lecture Notes in Computer Science, 1165:283–317, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leeb:1996:LPS

[2296] G. Leeb and N. Lynch. LL: Proving safety properties of the steam boiler controller. Lecture Notes in Computer Science, 1165:318–338, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lesske:1996:LSB

[2297] F. Lesske and S. Merz. LM: Steam boiler control specification problem: a TLA solution. Lecture Notes in Computer Science, 1165:339–358, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:LSO

[2298] X. Li and J. Wang. LW: Specifying optimal design of a steam-boiler system. Lecture Notes in Computer Science, 1165:359–378, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oelveczky:1996:OOO

[2299] P. C. Oelveczky, P. Kosiuczenko, and M. Wirsing. OKW: An object-oriented algebraic steam-boiler control specification. Lecture Notes in Computer Science, 1165:379–402, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schenke:1996:SRC

[2300] M. Schenke and A. P. Ravn. SR: Refinement from a control problem to programs. Lecture Notes in Computer Science, 1165:403–427, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schinagl:1996:VSS

[2301] C. P. Schinagl. S: VDM specification of the steam-boiler control using RSL notation. Lecture Notes in Computer Science, 1165:428–452, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitt:1996:VAS

[2302] J. Vitt and J. Hooman. VH: Assertional specification and verification using PVS of the steam boiler control system. Lecture Notes in Computer Science, 1165:453–472, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willig:1996:WSV

[2303] A. Willig and I. Schieferdecker. WS: Specifying and verifying the steamboiler control system with time extended LOTOS. Lecture Notes in Computer Science, 1165:473–492, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loetzbeyer:1996:SSB

[2304] A. Loetzbeyer. L: Simulation of a steam-boiler. Lecture Notes in Computer Science, 1165:493–499, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abrial:1996:SBC

[2305] J.-R. Abrial. A steam-boiler control specification problem. Lecture Notes in Computer Science, 1165:500–509, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keutzer:1996:NFM

[2306] K. Keutzer. The need for formal methods for integrated circuit design. Lecture Notes in Computer Science, 1166: 1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:VAC

[2307] Y.-A. Chen, E. Clarke, P.-H. Ho, Y. Hoskote, T. Kam, M. Khaira, J. O. Leary, and X. Zhao. Verification of all circuits in a floating-point unit using word-level model checking. Lecture Notes in Computer Science, 1166:19– 33, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arditi:1996:BCD

[2308] L. Arditi. BMDs can delay the use of theorem proving for verifying arithmetic assembly instructions. Lecture Notes in Computer Science, 1166:34– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravi:1996:MVM

[2309] K. Ravi, A. Pardo, G. D. Hachtel, and F. Somenzi. Modular verification of multipliers. Lecture Notes in Computer Science, 1166:49–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miner:1996:VIC

[2310] P. S. Miner and J. F. Leathrum. Verification of IEEE compliant subtractive division algorithms. Lecture Notes in Computer Science, 1166:64–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruess:1996:HVT

[2311] H. Ruess. Hierarchical verification of two-dimensional high-speed multiplication in PVS: a case study. Lecture Notes in Computer Science, 1166:79– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cantu:1996:EAH

[2312] F. J. Cantu, A. Bundy, A. Smaill, and D. Basin. Experiments in automating hardware verification using inductive proof planning. Lecture Notes in Computer Science, 1166:94–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1996:VNI

[2313] A. Jain, K. Nelson, and R. E. Bryant. Verifying nondeterministic implementations of deterministic systems. Lecture Notes in Computer Science, 1166: 109–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewin:1996:MPI

[2314] D. M. Lewin, D. H. Lorenz, and S. Ur.

A methodology for processor implementation verification. Lecture Notes in Computer Science, 1166:126–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:1996:CDT

[2315] D. Geist, M. Farkas, A. Landver, and Y. Lichtenstein. Coverage-directed test generation using symbolic techniques. Lecture Notes in Computer Science, 1166:143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1996:SCC

[2316] R. B. Jones, C.-J. H. Seger, and D. L. Dill. Self-consistency checking. Lecture Notes in Computer Science, 1166:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cyrluk:1996:IAM

[2317] D. Cyrluk. Inverting the abstraction mapping: a methodology for hardware verification. Lecture Notes in Computer Science, 1166:172–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrett:1996:VCC

[2318] C. Barrett, D. L. Dill, and J. Levitt.

Validity checking for combinations of theories with equality. Lecture Notes in Computer Science, 1166:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1996:UAC

[2319] K. Schneider and T. Kropf. A unified approach for combining different formalisms for hardware verification. Lecture Notes in Computer Science, 1166:202–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hojati:1996:VUU

[2320] R. Hojati, A. Isles, D. Kirkpatrick, and R. K. Brayton. Verification using uninterpreted functions and finite instantiations. Lecture Notes in Computer Science, 1166:218–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1996:FVI

[2321] Z. Zhou, X. Song, S. Tahar, and E. Cerny. Formal verification of the island tunnel controller using multiway decision graphs. Lecture Notes in Computer Science, 1166:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brayton:1996:V

[2322] R. K. Brayton, G. D. Hachtel, A. Sangiovanni-Vincentelli, and F. Somenzi. VIS. Lecture Notes in Computer Science, 1166:248–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shankar:1996:PCS

[2323] N. Shankar. PVS: Combining specification, proof checking, and model checking. Lecture Notes in Computer Science, 1166:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1996:HLT

[2324] J. Harrison. HOL light: a tutorial introduction. Lecture Notes in Computer Science, 1166:265–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1996:TDD

[2325] B. Bose, M. Esen Tuna, and V. Choppella. A tutorial on digital design derivation using DRS. Lecture Notes in Computer Science, 1166:270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brock:1996:ATA

[2326] B. Brock, M. Kaufmann, and J. S.

Moore. ACL2 theorems about commercial microprocessors. Lecture Notes in Computer Science, 1166:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:1996:FSC

[2327] R. Kumar, C. Blumenroehr, D. Eisenbiegler, and D. Schmid. Formal synthesis in circuit design — a classification and survey. Lecture Notes in Computer Science, 1166:294–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bickford:1996:FSV

[2328] M. Bickford and D. Jamsek. Formal specification and verification of VHDL. Lecture Notes in Computer Science, 1166:310–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narasimhan:1996:SCF

[2329] N. Narasimhan and R. Vemuri. Specification of control flow properties for verification of synthesized VHDL designs. Lecture Notes in Computer Science, 1166:327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1996:AMC

[2330] A. C. J. Fox and N. A. Harman. An algebraic model of correctness for superscalar microprocessors. Lecture Notes in Computer Science, 1166:346– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Windley:1996:MCL

[2331] P. J. Windley and J. R. Burch. Mechanically checking a lemma used in an automatic verification tool. Lecture Notes in Computer Science, 1166:362– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Su:1996:AGI

[2332] J. X. Su, D. L. Dill, and C. W. Barrett. Automatic generation of invariants in processor verification. Lecture Notes in Computer Science, 1166:377–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sentovich:1996:BSB

[2333] E. M. Sentovich. A brief study of BDD package performance. Lecture Notes in Computer Science, 1166:389–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinel:1996:LET

[2334] C. Meinel and T. Theobald. Local encoding transformations for optimizing OBDD-representations of finite state machines. Lecture Notes in Computer Science, 1166:404–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1996:DTE

[2335] J. Jain, A. Narayan, C. Coelho, and S. P. Khatri. Decomposition techniques for efficient ROBDD construction. Lecture Notes in Computer Science, 1166:419–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoneda:1996:BVZ

[2336] T. Yoneda, H. Hatori, A. Takahara, and S.-I. Minato. BDDs vs. zerosuppressed BDDs: For CTL symbolic model checking of Petri nets. Lecture Notes in Computer Science, 1166:435– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borrione:1996:HBI

[2337] D. Borrione, H. Bouamama, D. Deharbe, and C. Le Faou. HDL-Based integration of formal methods and CAD tools in the prevail environment. Lecture Notes in Computer Science, 1166: 450–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeller:1996:SOS

[2338] A. Zeller. Smooth operations with square operators — the version set model in ICE. Lecture Notes in Computer Science, 1167:8–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magnusson:1996:FGV

[2339] B. Magnusson and U. Asklund. Fine grained version control of configurations in COOP/ orm. Lecture Notes in Computer Science, 1167:31–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunt:1996:ESD

[2340] J. J. Hunt, K.-P. Vo, and W. F. Tichy. An empirical study of delta algorithms. Lecture Notes in Computer Science, 1167:49–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seiwald:1996:IFB

[2341] C. Seiwald. Inter-file branching — A practical method for representing variants. Lecture Notes in Computer Science, 1167:67–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tibrook:1996:ACS

[2342] D. Tibrook. An architecture for a construction system. Lecture Notes in Computer Science, 1167:76–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1996:CVS

[2343] R. Conradi and B. Westfechtel. Configuring versioned software products. Lecture Notes in Computer Science, 1167:88–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munch:1996:HMV

[2344] B. P. Munch. HiCoV: Managing the version space. Lecture Notes in Computer Science, 1167:110–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estublier:1996:WSM

[2345] J. Estublier. Work space management in software engineering environments. Lecture Notes in Computer Science, 1167:127–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perry:1996:SCS

[2346] D. E. Perry. System compositions and shared dependencies. Lecture Notes in Computer Science, 1167:139–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belanger:1996:IWA

[2347] D. Belanger, D. Korn, and H. Rao. Infrastructure for wide-area software development. Lecture Notes in Computer Science, 1167:154–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reuter:1996:DRC

[2348] J. Reuter, S. U. Haenssgen, J. J. Hunt, and W. F. Tichy. Distributed revision control via the WWW. Lecture Notes in Computer Science, 1167:166– ??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Micallef:1996:ASA

[2349] J. Micallef and G. M. Clemm. The Asgard system: Activity-based configuration management. Lecture Notes in Computer Science, 1167:175–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auer:1996:ERM

[2350] A. Auer and J. Taramaa. Experience report on the maturity of configuration management of embedded software. Lecture Notes in Computer Science, 1167:187–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gulla:1996:EUC

[2351] B. Gulla and J. Gorman. Experiences with the use of a configuration language. Lecture Notes in Computer Science, 1167:198–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolvik:1996:ICM

[2352] S. Kolvik. Introducing configuration management in an organisation. Lecture Notes in Computer Science, 1167:220–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:OTB

[2353] T. Lee, P. Thomas, and V. Lowen. An odyssey towards best SCM practices: The big picture. Lecture Notes in Computer Science, 1167:231–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dart:1996:BPC

[2354] S. Dart. Best practice for a CM solution. Lecture Notes in Computer Science, 1167:239–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bendix:1996:FSR

[2355] L. Bendix. Fully supported recursive workspaces. Lecture Notes in Computer Science, 1167:256–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crnkovic:1996:EUS

[2356] I. Crnkovic. Experience of using a simple SCM tool in a complex development environment. Lecture Notes in Computer Science, 1167:262–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eilfield:1996:CMG

[2357] P. Eilfield. Configuration management as “glueware” for development of client/server-applications in heterogeneous and distributed environments. Lecture Notes in Computer Science, 1167:264–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heimbigner:1996:PDC

[2358] D. Heimbigner and A. L. Wolf. Postdeployment configuration management. Lecture Notes in Computer Science, 1167:272–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacKay:1996:CSR

[2359] S. A. MacKay. Change sets revisited and CM of complex documents. Lecture Notes in Computer Science, 1167:277–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dix:1996:MSV

[2360] A. Dix, T. Rodden, and I. Sommerville. Modeling the sharing of versions. Lecture Notes in Computer Science, 1167:282–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aleven:1996:HDD

[2361] V. Aleven and K. D. Ashley. How different is different? Arguing about the significance of similarities and differences. Lecture Notes in Computer Science, 1168:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aarts:1996:TCB

[2362] R. J. Aarts and J. Rousu. Towards CBR for bioprocess planning. Lecture Notes in Computer Science, 1168:16– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergmann:1996:RAC

[2363] R. Bergmann and W. Wilke. On the role of abstraction in case-based reasoning. Lecture Notes in Computer Science, 1168:28–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonzano:1996:ICS

[2364] A. Bonzano, P. Cunningham, and C. Meckiff. ISAC: A CBR system for decision support in air traffic control. Lecture Notes in Computer Science, 1168:44–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boerner:1996:SSA

[2365] K. Boerner, E. Pippig, E.-C. Tammer, and C.-H. Coulon. Structural similarity and adaptation. Lecture Notes in Computer Science, 1168:58–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branting:1996:JSD

[2366] L. K. Branting and J. C. Lester.

Justification structures for document reuse. Lecture Notes in Computer Science, 1168:76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collins:1996:AGR

[2367] B. Collins and P. Cunningham. Adaptation guided retrieval in EBMT: a case-based approach to machine translation. Lecture Notes in Computer Science, 1168:91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daengdej:1996:DCI

[2368] J. Daengdej, D. Lukose, E. Tsui, and P. Beinat. Dynamically creating indices for two million cases: A real world problem. Lecture Notes in Computer Science, 1168:105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubitzky:1996:CMB

[2369] W. Dubitzky, J. G. Hughes, and D. A.

Bell. Case memory and the behaviouristic model of concepts and cognition. Lecture Notes in Computer Science, 1168:120–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Chamizo:1996:SOR

[2370] C. Fernandez-Chamizo, P. A. Gonzalez-Calero, M. Gomez-Albarran, and L. Hernandez-Yanez. Supporting object reuse through case-based reasoning. Lecture Notes in Computer Science, 1168:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goel:1996:MCE

[2371] A. K. Goel and J. W. Murdock. Metacases: Explaining case-based reasoning. Lecture Notes in Computer Science, 1168:150–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grimnes:1996:TLC

[2372] M. Grimnes and A. Aamodt. A two layer case-based reasoning architecture for medical image understanding. Lecture Notes in Computer Science, 1168:164–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanney:1996:LAR

[2373] K. Hanney and M. T. Keane. Learning adaptation rules from a casebase. Lecture Notes in Computer Science, 1168:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ye:1996:EAM

[2374] H. Ye. An evolutionary agent model of case-based classification. Lecture Notes in Computer Science, 1168:193– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamp:1996:UDL

[2375] G. Kamp. Using description logics for knowledge intensive case-based reasoning. Lecture Notes in Computer Science, 1168:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenz:1996:ACR

[2376] M. Lenz, H.-D. Burkhard, and S. Brueckner. Applying case retrieval nets to diagnostic tasks in technical domains. Lecture Notes in Computer Science, 1168:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macedo:1996:PSN

[2377] L. Macedo, F. C. Pereira, C. Grilo, and A. Cardoso. Plans as structured networks of hierarchically and temporally related case pieces. Lecture Notes in Computer Science, 1168:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maiden:1996:TPA

[2378] N. A. M. Maiden. The transfer problem in analogical reuse. Lecture Notes in Computer Science, 1168:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Micarelli:1996:CBS

[2379] A. Micarelli and F. Sciarrone. A casebased system for adaptive hypermedia navigation. Lecture Notes in Computer Science, 1168:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munoz-Avila:1996:FWE

[2380] H. Muñoz-Avila and J. Huellen. Feature weighting by explaining case-based planning episodes. Lecture Notes in Computer Science, 1168:280–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Napoli:1996:CBP

[2381] A. Napoli, J. Lieber, and R. Curien. Classification-based problem-solving in case-based reasoning. Lecture Notes in Computer Science, 1168:295–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osborne:1996:CBS

[2382] H. R. Osborne and D. G. Bridge. A case base similarity framework. Lecture Notes in Computer Science, 1168: 309–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaza:1996:ISE

[2383] E. Plaza, R. Lopez de Mantaras, and E. Armengol. On the importance of similitude: An entropy-based assessment. Lecture Notes in Computer Science, 1168:324–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahmer:1996:CEC

[2384] J. Rahmer and A. Voss. Case-enhanced configuration by resource balancing. Lecture Notes in Computer Science, 1168:339–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rousu:1996:ACC

[2385] J. Rousu and R. J. Aarts. Adaptation cost as a criterion for solution evaluation. Lecture Notes in Computer Science, 1168:354–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaaf:1996:FSN

[2386] J. W. Schaaf. Fish and shrink. A next step towards efficient case retrieval in large-scale case bases. Lecture Notes in Computer Science, 1168:362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1996:ADT

[2387] R. Schmidt, B. Heindl, B. Pollwein, and L. Gierl. Abstractions of data and time for multiparametric time course prognoses. Lecture Notes in Computer Science, 1168:377–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smyth:1996:UPA

[2388] B. Smyth and P. Cunningham. The utility problem analysed: a case-based reasoning perspective. Lecture Notes in Computer Science, 1168:392–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Surma:1996:RSF

[2389] J. Surma and B. Braunschweig. RE-PRO: Supporting flowsheet design by case-base retrieval. Lecture Notes in Computer Science, 1168:400–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tirri:1996:BFC

[2390] H. Tirri, P. Kontkanen, and P. Myllymaeki. A Bayesian framework for case-based reasoning. Lecture Notes in Computer Science, 1168:413–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voss:1996:PCR

[2391] A. Voss. Principles of case reusing systems. Lecture Notes in Computer Science, 1168:428–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber-Lee:1996:UTT

[2392] R. Weber-Lee, R. M. Barcia, A. Martins, and R. C. Pacheco. Using typicality theory to select the best match. Lecture Notes in Computer Science, 1168:445–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilke:1996:CDC

[2393] W. Wilke and R. Bergmann. Considering decision cost during learning of feature weights. Lecture Notes in Computer Science, 1168:460–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zdrahal:1996:CBP

[2394] Z. Zdrahal and E. Motta. Case-based problem solving methods for parametric design tasks. Lecture Notes in Computer Science, 1168:473–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barr:1996:CKM

[2395] J. M. Barr and R. V. Magaldi. Corporate knowledge management for the Millennium. Lecture Notes in Computer Science, 1168:487–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouchet:1996:CBR

[2396] J.-L. Bouchet and C. Eichenbaum-Voline. Case-based reasoning techniques applied to operation experience feedback in nuclear power plants. Lecture Notes in Computer Science, 1168: 497–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heider:1996:TCE

[2397] R. Heider. Troubleshooting CFM 56-3 engines for the Boeing 737 — using CBR and data-mining. Lecture Notes in Computer Science, 1168:512– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klahr:1996:GCB

[2398] P. Klahr. Global case-base development and deployment. Lecture Notes in Computer Science, 1168:519–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1996:RMS

[2399] M. Broy and L. Lamport. The RPC-Memory specification problem — problem statement. Lecture Notes in Computer Science, 1169:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1996:RMC

[2400] M. Broy, S. Merz, and K. Spies. The RPC-Memory case study: a synopsis. Lecture Notes in Computer Science, 1169:5–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1996:TSR

[2401] M. Abadi, L. Lamport, and S. Merz. A TLA solution to the RPC-Memory specification problem. Lecture Notes in Computer Science, 1169:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astesiano:1996:DSR

[2402] E. Astesiano and G. Reggio. A dynamic specification of the RPC-Memory problem. Lecture Notes in Computer Science, 1169:67–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1996:MMS

[2403] E. Best. A memory module specification using composable high-level Petri nets. Lecture Notes in Computer Science, 1169:109–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blom:1996:COT

[2404] J. Blom and B. Jonsson. Constraint oriented temporal logic specification. Lecture Notes in Computer Science, 1169:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1996:FSR

[2405] M. Broy. A functional solution to the RPC-Memory specification problem. Lecture Notes in Computer Science, 1169:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuellar:1996:SRM

[2406] J. Cuellar, D. Barnard, and M. Huber. A solution relying on the model checking of Boolean transition systems. Lecture Notes in Computer Science, 1169:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gotzhein:1996:ATL

[2407] R. Gotzhein. Applying a temporal logic to the RPC-Memory specification problem. Lecture Notes in Computer Science, 1169:253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooman:1996:UPA

[2408] J. Hooman. Using PVS for an assertional verification of the RPC-Memory specification problem. Lecture Notes in Computer Science, 1169:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hungar:1996:SVU

[2409] H. Hungar. Specification and verification using a visual formalism on top of temporal logic. Lecture Notes in Computer Science, 1169:305–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klarlund:1996:CSV

[2410] N. Klarlund, M. Nielsen, and K. Sunesen. A case study in verification based on trace abstractions. Lecture Notes in Computer Science, 1169:341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurki-Suonio:1996:ISJ

[2411] R. Kurki-Suonio. Incremental specification with joint actions: The RPC-Memory specification problem. Lecture Notes in Computer Science, 1169:375– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1996:MMC

[2412] K. G. Larsen, B. Steffen, and C. Weise. The methodology of modal constraints. Lecture Notes in Computer Science, 1169:405–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Romijn:1996:TRM

[2413] J. Romijn. Tackling the RPC-Memory specification problem with I/O automata. Lecture Notes in Computer Science, 1169:437–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoelen:1996:URS

[2414] K. Stoelen. Using relations on streams to solve the RPC-Memory specification problem. Lecture Notes in Computer Science, 1169:477–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Udink:1996:RMS

[2415] R. T. Udink and J. N. Kok. The RPC-Memory specification problem: UNITY + refinement calculus. Lecture Notes in Computer Science, 1169: 521–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlowski:1996:CFI

[2416] S. Orlowski. The changing face of information technology security. Lecture Notes in Computer Science, 1172:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardjono:1996:RKA

[2417] T. Hardjono and J. Seberry. Replicating the Kuperee authentication server for increased security and reliability. Lecture Notes in Computer Science, 1172:14–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taylor:1996:NRP

[2418] R. Taylor. Non-repudiation without public-key. Lecture Notes in Computer Science, 1172:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hah:1996:INR

[2419] Y. Hah. Investigation of nonrepudiation protocols. Lecture Notes in Computer Science, 1172:38–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1996:DSS

[2420] S.-J. Hwang and C.-C Chang. A dynamic secret sharing scheme with cheater detection. Lecture Notes in Computer Science, 1172:48–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Renvall:1996:NSS

[2421] A. Renvall and C. Ding. A nonlinear secret sharing scheme. Lecture Notes in Computer Science, 1172:56–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Renvall:1996:ASS

[2422] A. Renvall and C. Ding. The access structure of some secret-sharing schemes. Lecture Notes in Computer Science, 1172:67–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1996:CRF

[2423] C.-K. Wu and E. Dawson. On construction of resilient functions. Lecture Notes in Computer Science, 1172:79– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawson:1996:AAS

[2424] E. Dawson and J. He. Another approach to software key escrow encryption. Lecture Notes in Computer Science, 1172:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pieprzyk:1996:CBT

[2425] J. Pieprzyk, H. Ghodosi, C. Charnes, and R. Safavi-Naini. Cryptography based on transcendental numbers. Lecture Notes in Computer Science, 1172: 96–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Troy:1996:BSI

[2426] E. Troy. Breakthroughs in standardization of IT security criteria. Lecture Notes in Computer Science, 1172:108– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:TAP

[2427] L. Chen, D. Gollmann, and C. Mitchell.

Tailoring authentication protocols to match underlying mechanisms. Lecture Notes in Computer Science, 1172:121– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mu:1996:DSP

[2428] Y. Mu and V. Varadharajan. On the design of security protocols for mobile communications. Lecture Notes in Computer Science, 1172:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyd:1996:FDK

[2429] C. Boyd. A framework for design of key establishment protocols. Lecture Notes in Computer Science, 1172:146– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:1996:PMS

[2430] J. Golic. On period of multipleted sequences. Lecture Notes in Computer Science, 1172:158–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:1996:EDC

[2431] J. Golic. Edit distance correlation attacks on clock-controlled combiners with memory. Lecture Notes in Computer Science, 1172:169–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mihaljevic:1996:FCS

[2432] M. Mihaljevic. A faster cryptanalysis of the self-shrinking generator. Lecture Notes in Computer Science, 1172:182– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraani-Dastjerdi:1996:MML

[2433] A. Baraani-Dastjerdi, J. Pieprzyk, and R. Safavi-Naini. Modeling a multi-level secure object-oriented database using views. Lecture Notes in Computer Science, 1172:190–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varadharajan:1996:SJA

[2434] V. Varadharajan and P. Allen. Support for joint action based security policies. Lecture Notes in Computer Science, 1172:207–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandhu:1996:ACN

[2435] R. Sandhu. Access control: The neglected frontier. Lecture Notes in Computer Science, 1172:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1996:QAB

[2436] E. Jonsson and M. Andersson. On the quantitative assessment of behavioural security. Lecture Notes in Computer Science, 1172:228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustafson:1996:MPS

[2437] U. Gustafson, E. Jonsson, and T. Olovsson. On the modelling of preventive security based on a PC network intrusion experiment. Lecture Notes in Computer Science, 1172:242–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esmaili:1996:ERN

[2438] M. Esmaili, R. Safavi-Naini, and J. Pieprzyk. Evidential reasoning in network intrusion detection systems. Lecture Notes in Computer Science, 1172:253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1996:GOU

[2439] C.-H. Lin, C.-T. Wang, and C.-C.

Chang. A group-oriented (t, n) undeniable signature scheme without trusted center. Lecture Notes in Computer Science, 1172:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghodosi:1996:CHG

[2440] H. Ghodosi, J. Pieprzyk, C. Charnes, and R. Safavi-Naini. Cryptosystems for hierarchical groups. Lecture Notes in Computer Science, 1172:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bakhtiari:1996:SCH

[2441] S. Bakhtiari, R. Safavi-Naini, and J. Pieprzyk. On selectable collisionful hash functions. Lecture Notes in Computer Science, 1172:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bakhtiari:1996:PBA

[2442] S. Bakhtiari, R. Safavi-Naini, and J. Pieprzyk. On password-based authenticated key exchange using collisionful hash functions. Lecture Notes in Computer Science, 1172:299–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadeghiyan:1996:NUT

[2443] B. Sadeghiyan and J. Mohajeri. A new universal test for bit strings. Lecture Notes in Computer Science, 1172:311– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mu:1996:AMQ

[2444] Y. Mu and Y. Zheng. An alternative model of quantum key agreement via photon coupling. Lecture Notes in Computer Science, 1172:320–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rucklidge:1996:EVR

[2445] W. Rucklidge. Efficient visual recognition using the Hausdorff distance. Lecture Notes in Computer Science, 1173:xiii + 178, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahn:1996:HS

[2446] D. Kahn. The history of stenography. Lecture Notes in Computer Science, 1174:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franz:1996:CBS

[2447] E. Franz, A. Jerichow, S. Moeller, and A. Pfitzmann. Computer based steganography. Lecture Notes in Computer Science, 1174:7–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Handel:1996:HDO

[2448] T. Handel and M. Sandford. Hiding data in the OSI network model. Lecture Notes in Computer Science, 1174: 23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1996:SLS

[2449] R. Anderson. Stretching the limits of steganography. Lecture Notes in Computer Science, 1174:39–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfitzmann:1996:TTT

[2450] B. Pfitzmann. Trials of traced traitors.

Lecture Notes in Computer Science, 1174:49–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desmedt:1996:EBB

[2451] Y. Desmedt. Establishing big brother using covert channels and other covert techniques. Lecture Notes in Computer Science, 1174:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meadows:1996:CCC

[2452] C. Meadows and I. Moskowitz. Covert channels — A context-based view. Lecture Notes in Computer Science, 1174: 73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1996:CCA

[2453] M. Anderson and M. Ozols. Covert channel analysis for stubs. Lecture Notes in Computer Science, 1174:95– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1996:AAC

[2454] I. Jackson. Anonymous addresses and confidentiality of location. Lecture Notes in Computer Science, 1174:115– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Federrath:1996:MMC

[2455] H. Federrath, A. Jerichow, and A. Pfitzmann. MIXes in mobile communication systems: Location management with privacy. Lecture Notes in Computer Science, 1174:121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldschlag:1996:HRI

[2456] D. Goldschlag, M. Reed, and P. Syverson. Hiding routing information. Lecture Notes in Computer Science, 1174: 137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1996:NC

[2457] R. Anderson, S. Vaudenay, B. Preneel, and K. Nyberg. The Newton channel. Lecture Notes in Computer Science, 1174:151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burmester:1996:PRS

[2458] M. Burmester, Y. Desmedt, T. Itoh, and K. Sakurai. A progress report on subliminal-free channels. Lecture Notes in Computer Science, 1174:157– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Low:1996:MCP

[2459] S. Low and N. Maxemchuk. Modeling cryptographic protocols and their collusion analysis. Lecture Notes in Computer Science, 1174:169–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cox:1996:SRW

[2460] I. Cox, J. Kilian, T. Leighton, and T. Shamoon. A secure, robust watermark for multimedia. Lecture Notes in Computer Science, 1174:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:MIH

[2461] J. Smith and B. Comiskey. Modulation and information hiding in images. Lecture Notes in Computer Science, 1174:207–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brassil:1996:WDI

[2462] J. Brassil and L. O’Gorman. Watermarking document images with bounding box expansion. Lecture Notes in Computer Science, 1174:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simmons:1996:HSC

[2463] G. Simmons. The history of subliminal channels. Lecture Notes in Computer Science, 1174:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakurai:1996:BDB

[2464] K. Sakurai and Y. Yamane. Blind decoding, blind undeniable signatures, and their applications to privacy protection. Lecture Notes in Computer Science, 1174:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aura:1996:PID

[2465] T. Aura. Practical invisibility in digital communication. Lecture Notes in Computer Science, 1174:265–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davern:1996:FBI

[2466] P. Davern and M. Scott. Fractal based image steganography. Lecture Notes in Computer Science, 1174:279–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruhl:1996:EH

[2467] D. Gruhl, A. Lu, and W. Bender. Echo hiding. Lecture Notes in Computer Science, 1174:295–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aucsmith:1996:TRS

[2468] D. Aucsmith. Tamper resistant software. Lecture Notes in Computer Science, 1174:317–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaze:1996:OKE

[2469] M. Blaze. Oblivious key escrow. Lecture Notes in Computer Science, 1174: 335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berson:1996:HMO

[2470] T. Berson. Her Majesty’s orthography service. Lecture Notes in Computer Science, 1174:345–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfitzmann:1996:IHT

[2471] B. Pfitzmann. Information hiding terminology. Lecture Notes in Computer Science, 1174:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeza-Yates:1996:UVS

[2472] R. Baeza-Yates. A unified view to string matching algorithms. Lecture Notes in Computer Science, 1175:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MeyeraufderHeide:1996:CPS

[2473] F. Meyer auf der Heide and C. Scheideler. Communication in parallel systems. Lecture Notes in Computer Science, 1175:16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cahill:1996:OTO

[2474] V. Cahill. An overview of the Tigger object-support operating system framework. Lecture Notes in Computer Science, 1175:34–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dechamboux:1996:ADS

[2475] P. Dechamboux, D. Hagimont, J. Mossiere, and X. Rousset de Pina. The Arias distributed shared memory: An overview. Lecture Notes in Computer Science, 1175:56–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleindienst:1996:COS

[2476] J. Kleindienst, F. Plasil, and P. Tuma. CORBA and object services. Lecture Notes in Computer Science, 1175:74– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horvitz:1996:DTR

[2477] E. Horvitz. Decision-theoretic reasoning and the human-computer interface: Advances in embedded intelligent agents. Lecture Notes in Computer Science, 1175:94–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Overmars:1996:PAM

[2478] M. H. Overmars and P. Svestka. Probabilistic approaches to motion planning. Lecture Notes in Computer Science, 1175:95–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roth:1996:LOR

[2479] D. Roth. Learning in order to reason: The approach. Lecture Notes in Computer Science, 1175:113–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pokorny:1996:DSH

[2480] J. Pokorny. Database semantics in heterogeneous environment. Lecture Notes in Computer Science, 1175:125– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1996:BCP

[2481] K. Robinson and D. Mac Randal. Business case processing — rationale, survey and trends. Lecture Notes in Computer Science, 1175:143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siebes:1996:DMK

[2482] A. Siebes. Data mining and the KESO project. Lecture Notes in Computer Science, 1175:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domik:1996:CVC

[2483] G. Domik. Computer visualization — concepts, trends and current research. Lecture Notes in Computer Science, 1175:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matyska:1996:HPC

[2484] L. Matyska. High performance computing — the computational chemistry perspective. Lecture Notes in Computer Science, 1175:197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderVorst:1996:IMU

[2485] H. A. Van der Vorst. Iterative methods for unsymmetric linear systems. Lecture Notes in Computer Science, 1175: 217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenBrand:1996:CTS

[2486] M. Van den Brand, P. Klint, and C. Verhoef. Core technologies for system renovation. Lecture Notes in Computer Science, 1175:235–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBruin:1996:TGT

[2487] A. De Bruin and W. Pijls. Trends in game tree search. Lecture Notes in Computer Science, 1175:255–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dick:1996:VAV

[2488] J. Dick and J. Loubersac. A visual approach to VDM. Lecture Notes in Computer Science, 1175:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1996:SPP

[2489] D. Harel. Statecharts: Past, present and future. Lecture Notes in Computer Science, 1175:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kutten:1996:SFT

[2490] S. Kutten. Scalable fault tolerance.

Lecture Notes in Computer Science, 1175:286–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asirelli:1996:DDS

[2491] P. Asirelli, S. Gnesi, and M. C.

Rossi. A deductive database support to the specification of concurrent systems. Lecture Notes in Computer Science, 1175:309–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bednarek:1996:ICT

[2492] D. Bednarek, P. Merta, D. Obdrzalek, and J. Yaghob. Interprocess communication in the T4 system. Lecture Notes in Computer Science, 1175:317– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertino:1996:FRS

[2493] E. Bertino, D. Castelli, and F. Vitale. A formal representation for state diagrams in the OMT methodology. Lecture Notes in Computer Science, 1175:327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brattka:1996:FRR

[2494] V. Brattka and P. Hertling. Feasible real random access machines. Lecture Notes in Computer Science, 1175:335– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCastro:1996:TCF

[2495] C. De Castro. Temporal conversion functions for multitemporal relational databases. Lecture Notes in Computer Science, 1175:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaibisso:1996:OSS

[2496] C. Gaibisso, E. Nardelli, and G. Proietti. An output sensitive solution to the set union and intersection problem. Lecture Notes in Computer Science, 1175:351–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gambin:1996:SMI

[2497] A. Gambin and S. Lasota. On the semantics of multistage interconnection networks. Lecture Notes in Computer Science, 1175:359–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1996:DTK

[2498] J. L. Han. Decision trees, knowledge rules and some related data mining algorithms. Lecture Notes in Computer Science, 1175:369–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kucera:1996:RDN

[2499] A. Kucera. Regularity is decidable for normed BPA and normed BPP processes in polynomial time. Lecture Notes in Computer Science, 1175:377– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucas:1996:NPC

[2500] S. Lucas and J. Oliver. A new proposal of concurrent process calculus. Lecture Notes in Computer Science, 1175:385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchiori:1996:SVL

[2501] M. Marchiori. On the specification of visual languages: Multisets as first-class citizens. Lecture Notes in Computer Science, 1175:393–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mraz:1996:RAR

[2502] F. Mraz, M. Platek, and J. Vogel. Restarting automata with rewriting. Lecture Notes in Computer Science, 1175:401–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rozloznik:1996:ROG

[2503] M. Rozloznik, Z. Strakos, and M. Tuma. On the role of orthogonality in the GMRES method. Lecture Notes in Computer Science, 1175:409– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruzicka:1996:ETP

[2504] P. Ruzicka. Efficient tree pattern unification. Lecture Notes in Computer Science, 1175:417–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholz:1996:LWF

[2505] P. Scholz. A light-weight formalism for the specification of reactive systems. Lecture Notes in Computer Science, 1175:425–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shao:1996:KDA

[2506] J. Shao. Knowledge discovery in alarm data analysis. Lecture Notes in Computer Science, 1175:433–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sima:1996:APT

[2507] J. Sima. Aunt’s problem: Table rounding. Lecture Notes in Computer Science, 1175:441–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vovsova:1996:SHD

[2508] I. Vovsova. A seven-head dragon for optimal pattern searching. Lecture Notes in Computer Science, 1175:451– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werner:1996:CRV

[2509] T. Werner, V. Hlavac, A. Leonardis, and T. Pajdla. Choosing reference views for image-based representation. Lecture Notes in Computer Science, 1175:459–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zerovnik:1996:RGP

[2510] J. Zerovnik. Recognizing graph products and bundles. Lecture Notes in Computer Science, 1175:467–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zizka:1996:GAB

[2511] J. Zizka and Z. Kratochvil. Genetic algorithms-based creation of fuzzy-rule bases for the car control. Lecture Notes in Computer Science, 1175:475– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zukunft:1996:TMM

[2512] O. Zukunft. A transaction model for mobile and active database systems. Lecture Notes in Computer Science, 1175:483–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Francon:1996:RTD

[2513] J. Francon. On recent trends in discrete geometry in computer science. Lecture Notes in Computer Science, 1176:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zunic:1996:LSF

[2514] J. Zunic and D. M. Acketa. Least squares fitting of digital polynomial segments. Lecture Notes in Computer Science, 1176:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vialard:1996:GPE

[2515] A. Vialard. Geometrical parameters extraction from discrete paths. Lecture Notes in Computer Science, 1176:24– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aktouf:1996:TDH

[2516] Z. Aktouf, G. Bertrand, and L. Perroton. A three-dimensional holes closing algorithm. Lecture Notes in Computer Science, 1176:36–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgefors:1996:MRS

[2517] G. Borgefors, G. Ramella, and G. Sanniti di Baja. Multiresolution representation of shape in binary images. Lecture Notes in Computer Science, 1176:51–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruckstein:1996:DE

[2518] A. M. Bruckstein, R. J. Holt, and A. N. Netravali. Discrete elastica. Lecture Notes in Computer Science, 1176:59– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skarbek:1996:FRP

[2519] W. Skarbek, K. Ignasiak, and M. Ghuwar. Fractal representation of planar shapes. Lecture Notes in Computer Science, 1176:73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufman:1996:VS

[2520] A. E. Kaufman. Volume synthesis. Lecture Notes in Computer Science, 1176: 87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenoir:1996:FCN

[2521] A. Lenoir, R. Malgouyres, and M. Revenu. Fast computation of the normal vector field of the surface of a 3-D discrete object. Lecture Notes in Computer Science, 1176:101–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frey:1996:FEM

[2522] P. Frey and H. Borouchaki. Finite element meshes by means of voxels. Lecture Notes in Computer Science, 1176:115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kenmochi:1996:PGL

[2523] Y. Kenmochi, A. Imiya, and N. Ezquerra. Polyhedra generation from lattice points. Lecture Notes in Computer Science, 1176:127–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Francon:1996:RAS

[2524] J. Francon, J.-M. Schramm, and M. Tajine. Recognizing arithmetic straight lines and planes. Lecture Notes in Computer Science, 1176:141– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiorio:1996:TCR

[2525] C. Fiorio. A topologically consistent representation for image analysis: the topological graph of frontiers. Lecture Notes in Computer Science, 1176:151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayala:1996:DCC

[2526] R. Ayala, E. Dominguez, A. R. Frances, and A. Quintero. Determining the components of the complement of a digital (n − 1)-manifold in Z0n. Lecture Notes in Computer Science, 1176:163– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibathez:1996:DDS

[2527] L. Ibathez, C. Hamitouche, and C. Roux. Determination of discrete sampling grids with optimal topological and spectral properties. Lecture Notes in Computer Science, 1176:181– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braquelaire:1996:IPE

[2528] J.-P. Braquelaire, L. Brun, and A. Vialard. Inter-pixel Euclidean paths for image analysis. Lecture Notes in Computer Science, 1176:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mathieu:1996:FDC

[2529] G. Mathieu. Filling driven by contour marching. Lecture Notes in Computer Science, 1176:205–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bretto:1996:HMD

[2530] A. Bretto and S. Ubeda. Hypergraph model of digital topology for grey level images. Lecture Notes in Computer Science, 1176:217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scellier:1996:ASA

[2531] D. Scellier, J.-Y. Boire, C. Thouly, and J. Maublant. Application of skeletonization algorithms for myocardial spect quantification. Lecture Notes in Computer Science, 1176:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andres:1996:SP

[2532] E. Andres, C. Sibata, and R. Acharya.

Supercover 3D polygons. Lecture Notes in Computer Science, 1176:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lachaud:1996:TDI

[2533] J.-O Lachaud. Topologically defined iso-surfaces. Lecture Notes in Computer Science, 1176:245–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fidrich:1996:ISE

[2534] M. Fidrich. Iso-surface extraction in 4D with applications related to scale space. Lecture Notes in Computer Science, 1176:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boumghar:1996:CDS

[2535] F. Boumghar, S. Miguet, and J.- M. Nicod. Complexity of discrete surfaces in the dividing-cubes algorithm. Lecture Notes in Computer Science, 1176:269–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tougne:1996:CDC

[2536] L. Tougne. Circle digitization and cellular automata. Lecture Notes in Computer Science, 1176:283–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barcucci:1996:RCP

[2537] E. Barcucci, A. Del Lungo, M. Nivat, and R. Pinzani. Reconstruction convex polyominoes from horizontal and vertical projections II. Lecture Notes in Computer Science, 1176:295–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andres:1996:QSR

[2538] E. Andres. The quasi-shear rotation. Lecture Notes in Computer Science, 1176:307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devillers:1996:CGD

[2539] O. Devillers. Computational geometry and discrete computations. Lecture Notes in Computer Science, 1176:317– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertrand:1996:STP

[2540] G. Bertrand and R. Malgouyres. Some topological properties of discrete surfaces. Lecture Notes in Computer Science, 1176:325–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sharp:1996:SPS

[2541] N. G. Sharp and E. R. Hancock. A statistical process for surface tracking. Lecture Notes in Computer Science, 1176:337–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atallah:1996:ANS

[2542] M. J. Atallah and D. Z. Chen. Applications of a numbering scheme for polygonal obstacles in the plane. Lecture Notes in Computer Science, 1178:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turner:1996:MCH

[2543] J. S. Turner. Multicast communication in high speed networks. Lecture Notes in Computer Science, 1178:25– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bremner:1996:ICH

[2544] D. Bremner. Incremental convex hull algorithms are not output sensitive. Lecture Notes in Computer Science, 1178:26–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Efrat:1996:SSL

[2545] A. Efrat and O. Schwarzkopf. Separating and shattering long line segments. Lecture Notes in Computer Science, 1178:36–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devillers:1996:OLB

[2546] O. Devillers and M. J. Katz. Optimal line bipartitions of point sets. Lecture Notes in Computer Science, 1178:45– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fukuda:1996:IFA

[2547] T. Fukuda, Y. Morimoto, S. Morishita, and T. Tokuyama. Interval finding and its application to data mining. Lecture Notes in Computer Science, 1178:55–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Baca:1996:AST

[2548] D. Fernandez-Baca and J. Lagergren. On the approximability of the Steiner tree problem in phylogeny. Lecture Notes in Computer Science, 1178:65– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halldorsson:1996:ASC

[2549] M. M. Halldorsson and K. Tanaka. Approximation and special cases of common subtrees and editing distance. Lecture Notes in Computer Science, 1178: 75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1996:TDD

[2550] Y. Choi and Tak Wah Lam. Twodimensional dynamic dictionary matching. Lecture Notes in Computer Science, 1178:85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braezma:1996:DUU

[2551] A. Braezma, E. Ukkonen, and J. Vilo. Discovering unbounded unions of regular pattern languages from positive examples. Lecture Notes in Computer Science, 1178:95–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dey:1996:EPG

[2552] T. K. Dey and J. Pach. Extremal problems for geometric hypergraphs. Lecture Notes in Computer Science, 1178: 105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Efrat:1996:CFB

[2553] A. Efrat and M. J. Katz. Computing fair and bottleneck matchings in geometric graphs. Lecture Notes in Computer Science, 1178:115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBerg:1996:CMO

[2554] M. De Berg, O. Devillers, M. Van Kreveld, and O. Schwarzkopf. Computing the maximum overlap of two convex polygons under translations. Lecture Notes in Computer Science, 1178:126– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayase:1996:OMF

[2555] K. Hayase and H. Imai. OBDDs of a monotone function and of its prime implicants. Lecture Notes in Computer Science, 1178:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaug:1996:AMM

[2556] M.-S. Chaug. Algorithms for maximum matching and minimum fill-in on chordal bipartite graphs. Lecture Notes in Computer Science, 1178:146– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perng:1996:GSC

[2557] S.-L. Perng, M.-T. Ko, C.-W. Ho, and T.-S. Hsu. Graph searching on chordal graphs. Lecture Notes in Computer Science, 1178:156–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uno:1996:AEA

[2558] T. Uno. An algorithm for enumerating all directed spanning trees in a directed graph. Lecture Notes in Computer Science, 1178:166–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1996:VRA

[2559] T. Kloks, H. Mueller, and C. K. Wong. Vertex ranking of asteroidal triple-free graphs. Lecture Notes in Computer Science, 1178:174–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niedermeier:1996:RDP

[2560] R. Niedermeier. Recursively divisible problems. Lecture Notes in Computer Science, 1178:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allender:1996:SCW

[2561] E. Allender and K.-J. Lange. StUSPACE(log n ) contained within DSPACE(log_{2} n/ log log n ). Lecture Notes in Computer Science, 1178:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1996:FED

[2562] X. Zhou, S. Tamura, and T. Nishizeki.

Finding edge-disjoint paths in partial k-trees. Lecture Notes in Computer Science, 1178:203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1996:OAB

[2563] T.-S. Hsu and M.-Y. Kao. Optimal augmentation for bipartite componentwise biconnectivity in linear time. Lecture Notes in Computer Science, 1178: 213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chong:1996:TMP

[2564] Ka Wong Chong and Tak Wah Lam.

Towards more precise parallel biconnectivity approximation. Lecture Notes in Computer Science, 1178:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1996:CPV

[2565] A. Ambainis. The complexity of probabilistic versus deterministic finite automata. Lecture Notes in Computer Science, 1178:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Litow:1996:BLU

[2566] B. Litow. Bounded length UCFG equivalence. Lecture Notes in Computer Science, 1178:239–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:SMT

[2567] D. T. Lee and C. F. Shen. The Steiner minimal tree problem in the lambdageometry plane. Lecture Notes in Computer Science, 1178:247–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1996:SLS

[2568] S.-W. Cheng, N. Katoh, and M. Sugai. A study of the LMT-skeleton. Lecture Notes in Computer Science, 1178:256– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:NSM

[2569] Cao An Wang, F. Chin, and Y.-F Xu. A new subgraph of minimum weight triangulations. Lecture Notes in Computer Science, 1178:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pantziou:1996:DTR

[2570] G. E. Pantziou, A. Roberts, and A. Symvonis. Dynamic tree routing under the “matching with consumption” model. Lecture Notes in Computer Science, 1178:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Houle:1996:DET

[2571] M. E. Houle and G. Turner. Dimensionexchange token distribution on the mesh and the torus. Lecture Notes in Computer Science, 1178:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:DHP

[2572] J.-H. Lee, C.-S. Shin, and K.-Y. Chwa.

Directed Hamiltonian packing in ddimensional meshes and its application. Lecture Notes in Computer Science, 1178:295–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulou:1996:PNC

[2573] E. Papadopoulou. k-pairs non-crossing shortest paths in a simple polygon. Lecture Notes in Computer Science, 1178: 305–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lingas:1996:MCP

[2574] A. Lingas and V. Soltan. Minimum convex partition of a polygon with holes by cuts in given directions. Lecture Notes in Computer Science, 1178:315–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayashi:1996:ELR

[2575] T. Hayashi, K. Nakano, and S. Olariu.

Efficient list ranking on the reconfigurable mesh, with applications. Lecture Notes in Computer Science, 1178: 326–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kutytowski:1996:PMN

[2576] M. Kutytowski, K. Lorys, and B. Oesterdiekhoff. Periodic merging networks. Lecture Notes in Computer Science, 1178:336–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilfong:1996:MWA

[2577] G. Wilfong. Minimizing wavelengths in an all-optical ring network. Lecture Notes in Computer Science, 1178:346– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeh:1996:CAL

[2578] T.-H. Yeh, C.-M. Kuo, C.-L. Lei, and H.-C Yen. Competitive analysis of online disk scheduling. Lecture Notes in Computer Science, 1178:356–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verriet:1996:SIO

[2579] J. Verriet. Scheduling interval ordered tasks with non-uniform deadlines. Lecture Notes in Computer Science, 1178:366–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakurai:1996:CWR

[2580] K. Sakurai and Y. Zheng. Cryptographic weaknesses in the round transformation used in a block cipher with provable immunity against linear cryptanalysis. Lecture Notes in Computer Science, 1178:376–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takagi:1996:MVM

[2581] T. Takagi and S. Naito. The multivariable modular polynomial and its applications to cryptography. Lecture Notes in Computer Science, 1178:386– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1996:BAP

[2582] T.-S. Hsu and D. R. Lopez. Bounds and algorithms for a practical task allocation model. Lecture Notes in Computer Science, 1178:397–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fatouron:1996:SAS

[2583] P. Fatouron and P. Spirakis. Scheduling algorithms for strict multithreaded computations. Lecture Notes in Computer Science, 1178:407–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feuerstein:1996:MTP

[2584] E. Feuerstein and A. S. De Loma.

On multi-threaded paging. Lecture Notes in Computer Science, 1178:417– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryabko:1996:FEH

[2585] B. Ryabko and A. Fionov. A fast and efficient homophonic coding algorithm. Lecture Notes in Computer Science, 1178:427–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watanabe:1996:IDC

[2586] O. Watanabe and O. Yamashita. An improvement of the digital cash protocol of Okamoto and Ohta. Lecture Notes in Computer Science, 1178:436– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehne:1996:RPL

[2587] F. Dehne and S. W. Song. Randomized parallel list ranking for distributed memory multiprocessors. Lecture Notes in Computer Science, 1179:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1996:SAM

[2588] F. Shi. Suffix arrays for multiple strings: a method for on-line multiple string searches. Lecture Notes in Computer Science, 1179:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vollmer:1996:RAP

[2589] H. Vollmer. Relations among parallel and sequential computation models. Lecture Notes in Computer Science, 1179:23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greco:1996:EPP

[2590] S. Greco and D. Sacca. The expressive power of “possible-is-certain” semantics. Lecture Notes in Computer Science, 1179:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:USM

[2591] J. H. M. Lee, H. F. Leung, P. J. Stuckey, and V. W. L. Tam. Using stochastic methods to guide search in CLP: a preliminary report. Lecture Notes in Computer Science, 1179:43– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarau:1996:BSL

[2592] P. Tarau, V. Dahl, and A. Fall. Backtrackable state with linear arline implication and assumption grammars. Lecture Notes in Computer Science, 1179: 53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tessier:1996:DDC

[2593] A. Tessier. Declarative debugging in constraint logic programming. Lecture Notes in Computer Science, 1179:64– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1996:ILC

[2594] T. Park, S. B. Cho, and H. Y. Yeom.

An improved logging and checkpointing scheme for recoverable distributed shared memory. Lecture Notes in Computer Science, 1179:74–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sun:1996:DCC

[2595] C. Sun, Y. Yang, Y. Zhang, and D. Chen. Distributed concurrency control in real-time cooperative editing systems. Lecture Notes in Computer Science, 1179:84–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1996:VTB

[2596] H. Wu and W.-N. Chin. Variable timestamp-based distributed deadlock detection and resolution. Lecture Notes in Computer Science, 1179:96– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Byun:1996:LDT

[2597] S. Byun, R. Kennaway, and R. Sleep. Lambda-definable term rewriting systems. Lecture Notes in Computer Science, 1179:106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunter:1996:SAC

[2598] C. A. Gunter, T.-H. Ngair, and D. Subramanian. Sets as anti-chains. Lecture Notes in Computer Science, 1179: 116–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liquori:1996:TLC

[2599] L. Liquori and G. Castagna. A typed lamda calculus of objects. Lecture Notes in Computer Science, 1179:129– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valarcher:1996:IVE

[2600] P. Valarcher. Intensionality versus extensionality and primitive recursion. Lecture Notes in Computer Science, 1179:142–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vorobyov:1996:BTF

[2601] S. Vorobyov. On the bounded theories of finite trees. Lecture Notes in Computer Science, 1179:152–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamada:1996:BSG

[2602] Y. Hamada, A. Mei, F. Bao, and Y. Igarashi. Broadcasting in star graphs with Byzantine failures. Lecture Notes in Computer Science, 1179: 162–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juang:1996:BTS

[2603] W.-S. Juang and C.-L. Lei. Blind threshold signatures based on discrete logarithm. Lecture Notes in Computer Science, 1179:172–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaida:1996:CEL

[2604] T. Kaida, S. Uehara, and K. Imamura.

Computation of the k-error linear complexity of binary sequences with period 20n. Lecture Notes in Computer Science, 1179:182–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicomette:1996:SRV

[2605] V. Nicomette and Y. Deswarte. Symbolic rights and vouchers for access control in distributed object systems. Lecture Notes in Computer Science, 1179: 192–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1996:HCM

[2606] Z.-B. Xu and B.-T Yang. The heuristics for the constrained multicast routing. Lecture Notes in Computer Science, 1179:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1996:DAN

[2607] X. Yi. On design and analysis of a new block cipher. Lecture Notes in Computer Science, 1179:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elrad:1996:CSO

[2608] T. Elrad, B. Sheen, and N. V. Nastasic. CHESSBOARD: a synergy of objectoriented concurrent programming and program layering. Lecture Notes in Computer Science, 1179:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1996:CAF

[2609] G. Keller and M. Simons. A calculational approach to flattening nested data parallelism in functional languages. Lecture Notes in Computer Science, 1179:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramirez:1996:COO

[2610] R. Ramirez. Concurrent objectoriented programming in Tempo++. Lecture Notes in Computer Science, 1179:244–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shivers:1996:USF

[2611] O. Shivers. A universal scripting framework or lamda: The ultimate “little language”. Lecture Notes in Computer Science, 1179:254–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Theissinger:1996:DAL

[2612] M. Theissinger. Design and application of a light-weight persistency mechanism. Lecture Notes in Computer Science, 1179:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Browne:1996:HVU

[2613] I. A. Browne, Z. Manna, and H. B.

Sipma. Hierarchical verification using verification diagrams. Lecture Notes in Computer Science, 1179:276–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamid:1996:DES

[2614] I. A. Hamid and F. Khendek. A dynamic evolution for the specification of distributed systems. Lecture Notes in Computer Science, 1179:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Igarashi:1996:FVE

[2615] S. Igarashi, M. Shio, T. Shirogane, and T. Mizutani. Formal verification and evaluation of execution time in the envelope theory. Lecture Notes in Computer Science, 1179:299–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Majumdar:1996:DCL

[2616] R. Majumdar and R. K. Shyamasundar. Design of controllers for linear hybrid systems. Lecture Notes in Computer Science, 1179:309–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:CLD

[2617] X. Li and D. V. Hung. Checking linear duration invariants by linear programming. Lecture Notes in Computer Science, 1179:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carl:1996:CVR

[2618] M. Carl. Conjunctive vector representations for set valued feature descriptions. Lecture Notes in Computer Science, 1179:333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandru:1996:CND

[2619] V. Chandru, S. Roy, and R. Subrahmanyam. Constructive negation in definite constraint logic programs. Lecture Notes in Computer Science, 1179: 335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:MMS

[2620] H.-C. Chen, A. Lim, and N. A. Warsi. Multilevel master-slave parallel programming models. Lecture Notes in Computer Science, 1179:337–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chuang:1996:IXP

[2621] S.-C Chuang. Improving the X.509 protocol for ATM networks. Lecture Notes in Computer Science, 1179:339– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1996:ICI

[2622] F. K. Chan, J. Seng, L.-S. Ooi, and J.-K. Tan. InterMUD communications (IMC) protocol with extensions to VRMLise MultiUser domains (MUDs). Lecture Notes in Computer Science, 1179:341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goshi:1996:AHA

[2623] K. Goshi, J. Cheng, and K. Ushijima.

ALPS-HI: An annotated logic programming system with hypothetical implications. Lecture Notes in Computer Science, 1179:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1996:CMR

[2624] L. He, Y. Chao, Y. Simajiri, and H. Seki. Constructing more relevant models. Lecture Notes in Computer Science, 1179:345–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Igarashi:1996:FAC

[2625] S. Igarashi, T. Mizutani, T. Shirogane, and M. Shio. Formal analysis for continuous systems controlled by programs. Lecture Notes in Computer Science, 1179:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kusakabe:1996:ITD

[2626] S. Kusakabe, K. Inenaga, K. Nishimura, and M. Amamiya. Implementation tuning of a dataflow language on a stock parallel machine with visualization tools. Lecture Notes in Computer Science, 1179:349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:OSS

[2627] A. Li and O. Bukhres. Overall system for secure wireless mobile networks. Lecture Notes in Computer Science, 1179:351–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohania:1996:MVA

[2628] M. Mohania and G. Dong. Materialized view adaptation in distributed databases. Lecture Notes in Computer Science, 1179:353–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravindran:1996:COO

[2629] E. Ravindran and S. C. Hui. Coupling object-orientation and deduction — A modelling perspective. Lecture Notes in Computer Science, 1179:355–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thornton:1996:IPB

[2630] J. Thornton and A. Sattar. An integer programming-based nurse rostering system. Lecture Notes in Computer Science, 1179:357–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voicu:1996:SPD

[2631] R. Voicu. Synthesizing parallel divideand-conquer algorithms using the list interleave operator. Lecture Notes in Computer Science, 1179:359–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1996:STS

[2632] D. Harel. Some thoughts on statecharts, 13 years later. Lecture Notes in Computer Science, 1179:361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolot:1996:EEB

[2633] J.-C. Bolot. End-to-end behavior of the Internet: Measurements, analysis, and applications. Lecture Notes in Computer Science, 1179:362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolka:1996:OPL

[2634] G. Smolka. The Oz programming language and system. Lecture Notes in Computer Science, 1179:377–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tygar:1996:CPI

[2635] J. D. Tygar, B. S. Yee, and N. Heintze.

Cryptographic postage indicia. Lecture Notes in Computer Science, 1179:378– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjoerner:1996:SLM

[2636] D. Bjoerner. Semantics and laws of man-made infrastucture systems. Lecture Notes in Computer Science, 1181: 1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirth:1996:LLF

[2637] N. Wirth. The language lola, FPGAs, and PLDs in teaching digital circuit design. Lecture Notes in Computer Science, 1181:2–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dosch:1996:CDC

[2638] W. Dosch. Calculating digital counters. Lecture Notes in Computer Science, 1181:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stepanov:1996:GP

[2639] A. A. Stepanov. Generic programming. Lecture Notes in Computer Science, 1181:40–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zamulin:1996:ASO

[2640] A. V. Zamulin. Algebraic semantics of the Oberon target machine. Lecture Notes in Computer Science, 1181: 41–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koster:1996:MA

[2641] C. H. A. Koster. The making of Algol 68. Lecture Notes in Computer Science, 1181:55–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayoh:1996:ALP

[2642] B. Mayoh. Artificial life and pollution control: Explorations of a genetic algorithm system on the highly parallel connection machine. Lecture Notes in Computer Science, 1181:68–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yakhno:1996:ASM

[2643] T. M. Yakhno and E. S. Petrov. Application of subdefinite models for solving constraint satisfaction problems. Lecture Notes in Computer Science, 1181: 80–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Telerman:1996:PSO

[2644] V. Telerman, V. Sidorov, and D. Ushakov. Problem solving in the object-oriented technological environment NeMo+. Lecture Notes in Computer Science, 1181:91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seuren:1996:WUS

[2645] P. A. M. Seuren. What a universal semantic interlingua can and cannot do. Lecture Notes in Computer Science, 1181:101–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kononenko:1996:UST

[2646] I. Kononenko and S. Sharoff. Understanding short texts with integration of knowledge representation methods. Lecture Notes in Computer Science, 1181:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1996:CLR

[2647] R. Freivalds and T. Zeugmann. Colearning of recursive languages from positive data. Lecture Notes in Computer Science, 1181:122–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boutsinas:1996:AAV

[2648] B. Boutsinas, S. Papadimitriou, and G. Pavlides. Automatic analysis, verification and synthesis of rule-based real-time decision making systems with machine learning assistance. Lecture Notes in Computer Science, 1181:134– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czaja:1996:PCE

[2649] L. Czaja. Processes in cause-effect structures. Lecture Notes in Computer Science, 1181:146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmann:1996:PRD

[2650] P. Bachmann. On the power of recursion in dataflow schemes. Lecture Notes in Computer Science, 1181:166– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trichina:1996:DEP

[2651] E. Trichina. Derivation of explicitly parallel code from declarative program by transformations. Lecture Notes in Computer Science, 1181:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raja:1996:ACM

[2652] N. Raja and R. K. Shyamasundar. Actors as a coordinating model of computation. Lecture Notes in Computer Science, 1181:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nepomniaschaya:1996:AVP

[2653] A. S. Nepomniaschaya. An associative version of the Prim-Dijkstra algorithm and its application to some graph problems. Lecture Notes in Computer Science, 1181:203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frias:1996:ADS

[2654] M. F. Frias, G. A. Baum, and A. M.

Haeberer. Adding design strategies to fork algebras. Lecture Notes in Computer Science, 1181:214–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turchin:1996:STR

[2655] V. F. Turchin. Supercompilation:

Techniques and results. Lecture Notes in Computer Science, 1181:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nemylykh:1996:PTM

[2656] A. P. Nemylykh and V. A. Pinchuk.

Program transformation with metasystem transitions: Experiments with a supercompiler. Lecture Notes in Computer Science, 1181:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glueck:1996:FBT

[2657] R. Glueck and J. Joergensen. Fast binding-time analysis for multi-level specialization. Lecture Notes in Computer Science, 1181:261–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1996:BAE

[2658] N. D. Jones and A. J. Glenstrup. BTA algorithms to ensure termination of offline partial evaluation. Lecture Notes in Computer Science, 1181:273–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thiemann:1996:PEC

[2659] P. Thiemann and M. Sperber. Polyvariant, expansion and compiler generators. Lecture Notes in Computer Science, 1181:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mandel:1996:DCL

[2660] L. Mandel and M. V. Cengarle. The disjunctive constrained lambda calculus. Lecture Notes in Computer Science, 1181:297–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ivannikov:1996:EMC

[2661] V. Ivannikov, V. Zadorozhny, R. Kossmann, and B. Novikon. Efficient metaobject control using mediators. Lecture Notes in Computer Science, 1181:310–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pree:1996:FCS

[2662] W. Pree, G. Pomberger, and F. Kapsner. Framework component systems: Concepts, design heuristics, and perspectives. Lecture Notes in Computer Science, 1181:330–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froehlich:1996:PRT

[2663] J. H. Froehlich. Prototype of a runtime adaptable, object-oriented system. Lecture Notes in Computer Science, 1181:341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sabelfeld:1996:TEP

[2664] V. K. Sabelfeld. The tree equivalence problem for linear recursion schemes. Lecture Notes in Computer Science, 1181:350–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1996:MAL

[2665] L. Lu. A mode analysis of logic: Programs by abstract interpretation. Lecture Notes in Computer Science, 1181: 362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mogensen:1996:SBD

[2666] T. A. Mogensen. A semantics-based determinacy analysis for Prolog with cut. Lecture Notes in Computer Science, 1181:374–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frauenstein:1996:CFA

[2667] T. Frauenstein, W. Grieskamp, P. Pepper, and M. Suedholt. Communicating functional agents and their application to graphical user interfaces. Lecture Notes in Computer Science, 1181: 386–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voronkov:1996:MRD

[2668] A. Voronkov. Merging relational database technology with constraint technology. Lecture Notes in Computer Science, 1181:409–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekhlyar:1996:HBD

[2669] M. I. Dekhlyar and A. J. Dikovsky. On homeostatic behavior of dynamic deductive data bases. Lecture Notes in Computer Science, 1181:420–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afrali:1996:TLD

[2670] F. Afrali, M. Gergatsoulis, and M. Katzouraki. On transformations into linear database logic programs. Lecture Notes in Computer Science, 1181:433– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hibbard:1996:SDM

[2671] W. Hibbard, D. T. Kao, and A. Wierse. Scientific data modeling. Lecture Notes in Computer Science, 1183:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keim:1996:SID

[2672] D. Keim, J. P. Lee, B. Thuraisingham, and C. Wittenbrink. Supporting interactive database exploration. Lecture Notes in Computer Science, 1183:12– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1996:VRM

[2673] U. Lang, G. Grinstein, and R. D. Bergeron. Visualization related metadata. Lecture Notes in Computer Science, 1183:26–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:VVD

[2674] J. P. Lee. Views, visualization and databases. Lecture Notes in Computer Science, 1183:37–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grinstein:1996:DMD

[2675] G. Grinstein and B. Thuraisingham.

Data mining and data visualization. Lecture Notes in Computer Science, 1183:54–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukherjea:1996:RAI

[2676] S. Mukherjea and J. D. Foley. Requirements and architecture of an information visualization tool. Lecture Notes in Computer Science, 1183:57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hibbard:1996:MSD

[2677] W. Hibbard. Mathematical structures of data and their implications for visualization. Lecture Notes in Computer Science, 1183:76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1996:SMS

[2678] D. T. Kao, R. D. Bergeron, M. J.

Cullinane, and T. M. Sparr. Semantics and mathematics of scientific data sampling. Lecture Notes in Computer Science, 1183:86–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keim:1996:EVC

[2679] D. A. Keim. Enhancing the visual clustering of query-dependent database visualization techniques using screenfilling curves. Lecture Notes in Computer Science, 1183:101–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wittenbrink:1996:RDS

[2680] C. M. Wittenbrink, E. Rosen, A. Pang, and S. K. Lodha. Realtime database support for environmental visualization. Lecture Notes in Computer Science, 1183:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neugebauer:1996:MDP

[2681] M. Neugebauer. Metalevel database programming and visualization with POETView. Lecture Notes in Computer Science, 1183:131–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flanery:1996:LSD

[2682] R. E. Flanery and J. M. Donato. Large scale data analysis using AVS5 and POSTGRES. Lecture Notes in Computer Science, 1183:145–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1996:FAO

[2683] H. Davies. FAN — an array-oriented query language. Lecture Notes in Computer Science, 1183:155–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmeing:1996:LLD

[2684] W. Schmeing. LadMan — A large data management system. Lecture Notes in Computer Science, 1183:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiken:1996:TDV

[2685] A. Aiken, J. Chen, M. Lin, and M. Spalding. The Tioga-2 database visualization environment. Lecture Notes in Computer Science, 1183:181–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wierse:1996:CVB

[2686] A. Wierse. Collaborative visualization based on distributed data objects. Lecture Notes in Computer Science, 1183:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1996:PTI

[2687] R. C. Agarwal, F. G. Gustavson, and M. Zubair. Performance tuning IBM RS/6000 POWER2 systems. Lecture Notes in Computer Science, 1184:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1996:DIP

[2688] J. D. Andersen. Digital image processing: a 1996 review. Lecture Notes in Computer Science, 1184:14–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artemiev:1996:UPC

[2689] S. S. Artemiev, S. A. Gusev, and O. G.

Monakhov. Use of parallel computations for account of options by a Monte Carlo method. Lecture Notes in Computer Science, 1184:24–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barker:1996:WDE

[2690] V. A. Barker. Wavelets and differential equations. Lecture Notes in Computer Science, 1184:30–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billing:1996:FPI

[2691] G. D. Billing. From first principles to industrial applications. Lecture Notes in Computer Science, 1184:40– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjoerstad:1996:PIS

[2692] P. E. Bjoerstad, M. Dryja, and E. Vainikko. Parallel implementation of a Schwarz domain decomposition algorithm. Lecture Notes in Computer Science, 1184:46–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blackford:1996:PED

[2693] L. S. Blackford, A. Cleary, J. Demmel, and I. Dhillon. Practical experience in the dangers of heterogeneous computing. Lecture Notes in Computer Science, 1184:57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandt:1996:CAC

[2694] J. Brandt, I. Dimov, K. Georgiev, and J. Wasniewski. Coupling the advection and the chemical parts of large air pollution models. Lecture Notes in Computer Science, 1184:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:1996:AOP

[2695] P. Brezany and M. Dang. Advanced optimizations for parallel irregular outof-core programs. Lecture Notes in Computer Science, 1184:77–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:1996:SAM

[2696] P. Brezany, T. A. Mueck, and E. Schikuta. A software architecture for massively parallel input-output. Lecture Notes in Computer Science, 1184: 85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broda:1996:SGA

[2697] A. Broda and W. Dzwinel. Spatial genetic algorithm and its parallel implementation. Lecture Notes in Computer Science, 1184:97–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1996:AAT

[2698] J. Brown. Addressing algebra as a tool for parallel program development. Lecture Notes in Computer Science, 1184:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1996:ONB

[2699] J. Brown and M. Zagha. Optimizing the NAS BT application for the POWER CHALLENGE array. Lecture Notes in Computer Science, 1184:117– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Browne:1996:PAH

[2700] S. Browne, H. Casanova, and J. Dongarra. Providing access to high performance computing technologies. Lecture Notes in Computer Science, 1184:123– 133, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1996:POO

[2701] M. Bubak, W. Ciesla, and K. Sowa. Parallel object-oriented library of genetic algorithms. Lecture Notes in Computer Science, 1184:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1996:MPP

[2702] M. Bubak, W. Funika, and J. Moscinski. Monitoring of performance of PVM applications on virtual network computer. Lecture Notes in Computer Science, 1184:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buecker:1996:PVQ

[2703] H. M. Buecker and M. Sauren. A parallel version of the quasi-minimal residual method, based on coupled twoterm recurrences. Lecture Notes in Computer Science, 1184:157–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerioni:1996:PPS

[2704] F. Cerioni, M. Colajanini, S. Filippone, and S. Maiolatesi. A proposal for parallel sparse BLAS. Lecture Notes in Computer Science, 1184:166–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clausen:1996:PSB

[2705] J. Clausen. Parallel search-based methods in optimization. Lecture Notes in Computer Science, 1184:176–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dackland:1996:HAP

[2706] K. Dackland and B. Kaagstroem. An hierarchical approach for performance analysis of ScaLAPACK-based routines using the Distributed Linear Algebra Machine. Lecture Notes in Computer Science, 1184:186–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dillon:1996:MAA

[2707] E. Dillon, J. Guyard, and G. Wantz.

MeDLey: An abstract approach to message passing. Lecture Notes in Computer Science, 1184:196–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:ST

[2708] J. Dongarra and L. S. Blackford.

ScaLAPACK tutorial. Lecture Notes in Computer Science, 1184:204–215, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dracopoulos:1996:BSP

[2709] D. C. Dracopoulos and S. Kent. Bulk synchronous parallelisation of genetic programming. Lecture Notes in Computer Science, 1184:216–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duff:1996:FSS

[2710] I. S. Duff and J. A. Scott. Frontal software for the solution of sparse linear equations. Lecture Notes in Computer Science, 1184:227–238, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esposito:1996:PHB

[2711] A. Esposito and L. Tarricone. Parallel heuristics for bandwidth reduction of sparse matrices with IBM SP2 and Cray T3D. Lecture Notes in Computer Science, 1184:239–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filippone:1996:PLD

[2712] S. Filippone. Parallel libraries on distributed memory architectures: The IBM Parallel ESSL. Lecture Notes in Computer Science, 1184:247–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foresti:1996:PLS

[2713] S. Foresti, L. Zeng, G. Gullberg, and R. Huesman. Parallel least squares estimates of 2-D SPECT image reconstructions on the SGI Power Challenge. Lecture Notes in Computer Science, 1184:256–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fornberg:1996:PPC

[2714] B. Fornberg. The prospect for parallel computing in the oil industry. Lecture Notes in Computer Science, 1184: 262–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fragopoulou:1996:PEM

[2715] P. Fragopoulou and O. H. Nielsen. Performance evaluation and modeling of reduction operations on the IBM RS/ 6000 SP parallel computer. Lecture Notes in Computer Science, 1184:272– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujino:1996:VTF

[2716] S. Fujino and R. Himeno. A vectorization technique for a family of finite difference formulae and its performance evaluation. Lecture Notes in Computer Science, 1184:284–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1996:PSF

[2717] D. Garcia, F. J. Suarez, J. Garcia, and J. M. Lopez. Parallel simulation of finishing hot strip millsi. Lecture Notes in Computer Science, 1184:292–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1996:PTD

[2718] J. M. Garcia, J. L. Sanchez, and P. Gonzalez. PEPE: a trace-driven simulator to evaluate reconfigurable multicomputer architectures. Lecture Notes in Computer Science, 1184:302–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geng:1996:DAM

[2719] X. Geng, J. Zhao, Y. Yan, and Y. Yu.

Data acquisition and management in BEPC. Lecture Notes in Computer Science, 1184:312–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1996:FGP

[2720] A. Gupta and F. Gustavson. Fast graph partitioning and its application in sparse matrix ordering. Lecture Notes in Computer Science, 1184:316– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1996:DIE

[2721] A. Gupta, F. G. Gustavson, M. Joshi, and S. Toledo. The design, implementation, and evaluation of a banded linear solver for distributed-memory parallel computers. Lecture Notes in Computer Science, 1184:328–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustavson:1996:NPA

[2722] F. G. Gustavson and A. Gupta. A new parallel algorithm for tridiagonal symmetric positive definite systems of equations. Lecture Notes in Computer Science, 1184:341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hackenberg:1996:EPS

[2723] M. G. Hackenberg, W. Joppich, and S. Mijalkovic. An environment for the parallel solution of coupled problems on structured grids. Lecture Notes in Computer Science, 1184:350–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1996:PPR

[2724] J. A. J. Hall and K. I. M. McKinnon. PARSMI, a parallel revised simplex algorithm incorporating minor iterations and devex pricing. Lecture Notes in Computer Science, 1184:359–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamma:1996:POI

[2725] B. Hamma. Parallel optimization of interplanetary trajectories. Lecture Notes in Computer Science, 1184:369– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansen:1996:FMR

[2726] P. C. Hansen and S. H. Jensen. Filter model of reduced-rank noise reduction. Lecture Notes in Computer Science, 1184:379–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hegarty:1996:CPD

[2727] D. F. Hegarty, M. T. Kechadi, and K. A. Dawson. On the crossover points for dynamic load balancing of longchained molecules. Lecture Notes in Computer Science, 1184:388–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heuveline:1996:PCS

[2728] V. Heuveline and M. Sadkane. Parallel computation of spectral portrait of large matrices. Lecture Notes in Computer Science, 1184:398–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inutin:1996:CPM

[2729] S. A. Inutin. Computer parallel modular algebra. Lecture Notes in Computer Science, 1184:408–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:POS

[2730] H. J. Kim, J. K. Lee, and K. W. Rim. Parallel operating system for MPP system: Design and implementations. Lecture Notes in Computer Science, 1184:413–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:NGM

[2731] Y. Kim and T. Feng. New graph model and its routing algorithm for rearrangeable networks. Lecture Notes in Computer Science, 1184:423–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiper:1996:MDA

[2732] A. Kiper. Modified Dorn’s algorithm with improved speed-up. Lecture Notes in Computer Science, 1184:433–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitowski:1996:CTS

[2733] J. Kitowski, K. Boryczko, and J. Moscinski. Comparison of two shortrange molecular dynamics algorithms for parallel computing. Lecture Notes in Computer Science, 1184:443–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1996:PLA

[2734] J. Larsen, T. Christensen, and L. Frellesen. Parallelising large applications. Lecture Notes in Computer Science, 1184:450–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lorenzo:1996:HPF

[2735] P. A. R. Lorenzo, A. Mueller, Y. Murakami, and B. J. N. Wylie. High Performance Fortran interfacing to ScaLA-PACK. Lecture Notes in Computer Science, 1184:457–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manne:1996:PAM

[2736] F. Manne and T. Soerevik. Partitioning an array onto a mesh of processors. Lecture Notes in Computer Science, 1184:467–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maschhoff:1996:PEP

[2737] K. J. Maschhoff and D. C. Sorensen.

PARPACK: An efficient portable large scale eigenvalue package for distributed memory parallel architectures. Lecture Notes in Computer Science, 1184:478– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:1996:PDR

[2738] M. May. Parallel distributed representation of sparse grids using process arrays. Lecture Notes in Computer Science, 1184:487–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKee:1996:OSD

[2739] S. McKee, P. Oliveira, C. Aldridge, and J. McDonald. Optimal scheduling, decomposition, and parallelisation. Lecture Notes in Computer Science, 1184:494–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meajil:1996:ITB

[2740] A. I. Meajil. An implementation of a tree-based N-body algorithm on message-passing architectures. Lecture Notes in Computer Science, 1184:505– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meajil:1996:QAA

[2741] A. I. Meajil, T. El-Ghazawi, and T. Sterling. A quantitative approach for architecture-invariant parallel workload characterization. Lecture Notes in Computer Science, 1184:515–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monakhov:1996:PMP

[2742] O. G. Monakhov and O. Y. Chunikhin. Parallel mapping of program graphs into parallel computers by selforganization algorithm. Lecture Notes in Computer Science, 1184:525–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naur:1996:PCD

[2743] P. Naur. PARA’96 conference dinner talk, 1996 August 20. Lecture Notes in Computer Science, 1184:529–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nelson:1996:COR

[2744] C. A. Nelson. Compiler optimizations for red-black HPF codes. Lecture Notes in Computer Science, 1184: 534–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Onesti:1996:PLA

[2745] L. Onesti and T. Soerevik. Parallelization of a local area ocean model. Lecture Notes in Computer Science, 1184: 543–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ostromsky:1996:PSS

[2746] T. Ostromsky, S. Salvini, J. Wasniewski, and Z. Zlatev. Parallel solution of sparse problemns by using a sequence of large dense blocks. Lecture Notes in Computer Science, 1184:555– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oezaktas:1996:PSC

[2747] H. Oezaktas, M. Akguel, and M. C.

Pinar. The parallel surrogate constraint approach to the linear feasibility problem. Lecture Notes in Computer Science, 1184:565–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pardalos:1996:PGM

[2748] P. M. Pardalos, L. Pitsoulis, and M. G. C. Resende. A parallel GRASP for MAX-SAT problems. Lecture Notes in Computer Science, 1184:575–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perram:1996:AHM

[2749] J. W. Perram. Applications of HPCN in manufacturing industry. Lecture Notes in Computer Science, 1184:586– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinar:1996:EMD

[2750] A. Pinar and C. Aykanat. An effective model to decompose linear programs for parallel solution. Lecture Notes in Computer Science, 1184:592–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salvini:1996:NLS

[2751] S. Salvini. Numerical libraries on shared memory computers. Lecture Notes in Computer Science, 1184:602– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salvini:1996:LAS

[2752] S. Salvini and J. Wasniewski. Linear algebra subprograms on shared memory computers: Beyond LAPACK. Lecture Notes in Computer Science, 1184: 614–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skelboe:1996:IPS

[2753] S. Skelboe. Integration of partitioned stiff systems of ordinary differential equations. Lecture Notes in Computer Science, 1184:621–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soerensen:1996:DLE

[2754] J. N. Soerensen, S. A. Kristensen, and L. K. Christensen. Direct and large eddy simulations of thermo-convective flows. Lecture Notes in Computer Science, 1184:631–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzaki:1996:MTM

[2755] K. Suzaki, H. Tanuma, S. Hirano, and Y. Ichisugi. Multi-tasking method on parallel computers which combines a contiguous and non-contiguous processor partitioning algorithm. Lecture Notes in Computer Science, 1184:641– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szularz:1996:HPE

[2756] M. Szularz, J. Weston, M. Clint, and K. Murphy. A highly parallel explicitly restarted Lanczos algorithm. Lecture Notes in Computer Science, 1184:651–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomsen:1996:DSS

[2757] P. G. Thomsen. Dynamic systems and software. Lecture Notes in Computer Science, 1184:661–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tingleff:1996:PIC

[2758] O. Tingleff. Parallel implementations of classical optimization methods. Lecture Notes in Computer Science, 1184:668–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toxvaerd:1996:CEC

[2759] S. Toxvaerd. Computer engineering of complex physical and chemical systems by parallel molecular dynamics computations. Lecture Notes in Computer Science, 1184:676–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderVorst:1996:PFA

[2760] H. A. Van der Vorst and G. L. G. Sleijpen. A parallelizable and fast algorithm for very large generalized eigenproblems. Lecture Notes in Computer Science, 1184:686–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wcislo:1996:PCA

[2761] R. Wcislo, J. Kitowski, and J. Moscinski. Parallelisation of a code for animation of multi-object system. Lecture Notes in Computer Science, 1184: 697–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1996:PIP

[2762] T. Yang. Parallel inner productfree algorithm for least squares problems. Lecture Notes in Computer Science, 1184:710–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilde:1996:GSM

[2763] E. Wilde, P. Freiburghaus, D. Koller, and B. Plattner. A group and session management system for distributed multimedia applications. Lecture Notes in Computer Science, 1185:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waters:1996:LCA

[2764] A. G. Waters and J. S. Crawford. Low-cost ATM multicast routing with constrained delays. Lecture Notes in Computer Science, 1185:23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1996:AST

[2765] M. Hofmann. Adding scalability to transport level multicast. Lecture Notes in Computer Science, 1185:41– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1996:RBK

[2766] M. C. Chan, J.-F. Huard, A. A.

Lazar, and K.-S. Lim. On realizing a broadband kernel for multimedia networks. Lecture Notes in Computer Science, 1185:56–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waddington:1996:SQM

[2767] D. G. Waddington, G. Coulson, and D. Hutchison. Specifying QoS for multimedia communications within distributed programming environments. Lecture Notes in Computer Science, 1185:75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfeifer:1996:GCC

[2768] T. Pfeifer and R. Popescu-Zeletin.

Generic conversion of communication media for supporting personal mobility. Lecture Notes in Computer Science, 1185:104–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Almeida:1996:FDN

[2769] A. Almeida. A framework for the deployment of new services using hypermedia distributed systems. Lecture Notes in Computer Science, 1185:130– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quemada:1996:ICA

[2770] J. Quemada, T. De Miguel, A. Azcorra, and S. Pavon. ISABEL: a CSCW application for the distribution of events. Lecture Notes in Computer Science, 1185:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leopold:1996:BPA

[2771] H. Leopold and R. Hirn. The bookshop project: An Austrian interactive multimedia application case study. Lecture Notes in Computer Science, 1185: 154–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degermark:1996:IDN

[2772] M. Degermark and S. Pink. Issues in the design of a new network protocol. Lecture Notes in Computer Science, 1185:169–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadka:1996:SCC

[2773] A. H. Sadka, F. Eryurtlu, and A. M. Kondoz. Source and channel coding for mobile multimedia communications. Lecture Notes in Computer Science, 1185:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

HuisintVeld:1996:DCA

[2774] R. J. Huis in’t Veld, A.-N. Ladhani, B. Van der Waaij, and I. A. Widya. Developing a conference application on top of an advanced signalling infrastructure. Lecture Notes in Computer Science, 1185:201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GarciaAdanez:1996:NNA

[2775] X. Garcia Adanez, O. Verscheure, and J.-P. Hubaux. New network and ATM adaptation layers for real-time multimedia applications: a performance study based on psychophysics. Lecture Notes in Computer Science, 1185:216– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waldegg:1996:MAU

[2776] D. Bourges Waldegg, N. Lagha, and J.- P. Le Narzul. Multimedia applications on a Unix SVR4 kernel: Performance study. Lecture Notes in Computer Science, 1185:232–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verscheure:1996:PVQ

[2777] O. Verscheure and J.-P Hubaux. Perceptual video quality and activity metrics: Optimization of video services based on MPEG-2 encoding. Lecture Notes in Computer Science, 1185:249– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadimitriou:1996:PTQ

[2778] C. H. Papadimitriou. Planar topological queries. Lecture Notes in Computer Science, 1191:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaede:1996:IIC

[2779] V. Gaede and M. Wallace. An informal introduction to constraint database systems. Lecture Notes in Computer Science, 1191:7–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miranker:1996:QEC

[2780] D. P. Miranker, R. J. Bayardo, and V. Samoladas. Query evaluation as constraint search; an overview of early results. Lecture Notes in Computer Science, 1191:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toman:1996:CWF

[2781] D. Toman. Computing the wellfounded semantics for constraint extensions of Datalog. Lecture Notes in Computer Science, 1191:64–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:DLJ

[2782] Y. Wang, B. Shi, and Y. Qu. Decomposition and lossless join in constraint databases. Lecture Notes in Computer Science, 1191:80–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hacid:1996:RBC

[2783] M.-S. Hacid, P. Marcel, and C. Rigotti. A rule-based CQL for 2-dimensional tables. Lecture Notes in Computer Science, 1191:92–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afrati:1996:EQL

[2784] F. Afrati, T. Andronikos, and T. Kavalieros. On the expressiveness of query languages with linear constraints; capturing desirable spatial properties. Lecture Notes in Computer Science, 1191: 105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuijpers:1996:ETC

[2785] B. Kuijpers and M. Smits. On expressing topological connectivity in Spatial Datalog. Lecture Notes in Computer Science, 1191:116–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodsky:1996:CCO

[2786] A. Brodsky and V. E. Segal. The C03 constraint object-oriented database system: An overview. Lecture Notes in Computer Science, 1191:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertino:1996:ICC

[2787] E. Bertino, B. Catania, and S. Bressan. Integrity constraint checking in Chimera. Lecture Notes in Computer Science, 1191:160–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibrahim:1996:STI

[2788] H. Ibrahim, W. A. Gray, and N. J. Fiddian. SICSDD: Techniques and implementation. Lecture Notes in Computer Science, 1191:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alagic:1996:TCS

[2789] S. Alagic. A temporal constraint system for object-oriented databases. Lecture Notes in Computer Science, 1191: 208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doucet:1996:UDV

[2790] A. Doucet, M.-C. Fauvet, S. Gancarski, and G. Jomier. Using database versions to implement temporal integrity constraints. Lecture Notes in Computer Science, 1191:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Revesz:1996:GDA

[2791] P. Z. Revesz. Genomic database applications in DISCO. Lecture Notes in Computer Science, 1191:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toman:1996:CDP

[2792] D. Toman. Constraint databases and program analysis using abstract interpretation. Lecture Notes in Computer Science, 1191:246–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sexton:1996:QIF

[2793] A. P. Sexton. Querying indexed files. Lecture Notes in Computer Science, 1191:263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freeston:1996:CBT

[2794] M. Freeston. On the complexity of BV-tree updates. Lecture Notes in Computer Science, 1191:282–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bressan:1996:IID

[2795] S. Bressan, A. P. Sexton, and M. Wallace. Implementing index data structures using constraint logic programming. Lecture Notes in Computer Science, 1191:294–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Revesz:1996:PSD

[2796] P. Z. Revesz. Problem solving in the DISCO constraint database system. Lecture Notes in Computer Science, 1191:302–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yakhno:1996:IDL

[2797] T. M. Yakhno and S. Zilberfaine. Interval domain library for ECL0iPS0e and its application. Lecture Notes in Computer Science, 1191:316–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nigam:1996:SQO

[2798] S. Nigam and K. C. Davis. A semantic query optimization algorithm for object-oriented databases. Lecture Notes in Computer Science, 1191:329– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagl:1996:BTI

[2799] Manfred Nagl, editor. Building tightly integrated software development environments: the IPSEN approach, volume 1170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-540-61985-2. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 709 pp. LCCN QA76.76.O63B838 1996.

Franz:1996:AAR

[2800] Alexander Franz. Automatic ambiguity resolution in natural language processing, volume 1171 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-540-62004-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 155 pp. LCCN QA76.9.N38F75 1996.

Muller:1996:DIA

[2801] J. P. (Jorg P.) Muller. The design of intelligent agents: a layered approach, volume 1177 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-540-62003-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 227 pp. LCCN TJ217.5.M84 1996.

Hasan:1996:OSQ

[2802] Waqar Hasan. Optimization of SQL queries for parallel machines, volume 1182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-540-62065-6. ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 133 pp. LCCN QA76.73.S67H38 1996.

Chandru:1996:FST

[2803] V. Chandru and V. Vinay, editors. Foundations of software technology and theoretical computer science: 16th conference, Hyderabad, India, December 18–20, 1996: proceedings, volume 1180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-62034-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751 .F692 1996. FST and TCS 16: The proceedings of the Sixteenth Annual Conference on Foundations of Software Technology and Theoretical Computer Science.